### v2

# Continuous Diffusion for Mixed-Type Tabular Data

Abstract
Score-based generative models (or diffusion models for short) have proven successful for generating text and image data. However, the adaption of this model family to tabular data of mixed-type has fallen short so far. In this paper, we propose CDTD, a Continuous Diffusion model for mixed-type Tabular Data. Specifically, we combine score matching and score interpolation to ensure a common continuous noise distribution for both continuous and categorical features alike. We counteract the high heterogeneity inherent to data of mixed-type with distinct, adaptive noise schedules per feature or per data type. The learnable noise schedules ensure optimally allocated model capacity and balanced generative capability. We homogenize the data types further with model-specific loss calibration and initialization schemes tailored to mixed-type tabular data. Our experimental results show that CDTD consistently outperforms state-of-the-art benchmark models, captures feature correlations exceptionally well, and that heterogeneity in the noise schedule design boosts the sample quality.
1 Introduction
Score-based generative models (Song et al., 2021), also termed diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020), have shown remarkable potential for the generation of images (Dhariwal & Nichol, 2021; Rombach et al., 2022), videos (Ho et al., 2022), text (Li et al., 2022; Dieleman et al., 2022; Wu et al., 2023), molecules (Hoogeboom et al., 2022), and many other highly complex data structures with continuous features. The framework has since been adapted to categorical data in various ways, including discrete diffusion processes (Austin et al., 2021; Hoogeboom et al., 2021), diffusion in continuous embedding space (Dieleman et al., 2022; Li et al., 2022; Regol & Coates, 2023; Strudel et al., 2022), and others (Campbell et al., 2022; Meng et al., 2022; Sun et al., 2023). Diffusion models which include both, continuous and categorical features alike, build directly on advances from the image domain (Kim et al., 2023; Kotelnikov et al., 2023; Lee et al., 2023; Jolicoeur-Martineau et al., 2024) and thus, are not designed to deal with challenges specific to mixed-type data: The different diffusion processes are not aligned, and do not scale to larger datasets and/or features with a greater number of categories.
A crucial component in score-based generative models is the noise schedule (Kingma et al., 2022; Chen et al., 2022; Chen, 2023; Jabri et al., 2022; Wu et al., 2023). Typical noise schedules for image and text data are designed to focus model capacity on the noise levels most important to sample quality (Nichol & Dhariwal, 2021; Karras et al., 2022), while others attempt to learn the optimal noise schedule (Dieleman et al., 2022; Kingma et al., 2022). For mixed-type tabular data, existing approaches often combine distinct diffusion processes for the continuous and discrete features to derive a joint model (Kotelnikov et al., 2023; Lee et al., 2023). However, noise schedules are not directly transferable from one data modality to another and therefore, using specifications from image or text domain models is not optimal: First, the inherently different diffusion processes make it difficult to balance the noise schedules across features and feature types, and negatively affect the allocation of model capacity across timesteps. Further, and most importantly, the domain, nature and marginal distribution can vary significantly across features (Xu et al., 2019). For instance, any two continuous features may be subject to different levels of discretization or different bounds, even after applying common data pre-processing techniques; and any two categorical features may differ in the number of categories, or the degree of imbalance. This high heterogeneity warrants a rethinking of fundamental parts of the diffusion framework, including the noise schedule and the effective combination of diffusion processes for different data types.
In this paper, we introduce Continuous Diffusion for mixed-type Tabular Data (CDTD) to address the aforementioned shortcomings. We combine score matching (Hyvärinen, 2005) with score interpolation (Dieleman et al., 2022) to derive a score-based model that pushes the diffusion process for categorical data into embedding space, and uses a Gaussian diffusion process for both continuous and categorical features. This way, the different noise processes become directly comparable, easier to balance, and enable the application of, for instance, classifier-free guidance (Ho & Salimans, 2022), accelerated sampling (Lu et al., 2022), and other advances, to mixed-type tabular data.
We counteract the high feature heterogeneity inherent to data of mixed-type with distinct feature- or type-specific adaptive noise schedules. The learnable noise schedules allow the model to directly take feature or type heterogeneity into account during both training and generation, and thus avoid the reliance on image or text-specific noise schedules designs. This ensures a better allocation of the model’s capacity across features, feature types and timesteps, and guarantees high quality samples of tabular data. CDTD outperforms state-of-the-art baseline models across a diverse set of sample quality metrics, and computation time for data sets with an arbitrary number of categories and data points. Our experiments show that CDTD captures feature correlations exceptionally well, and that explicitly allowing for data-type heterogeneity in the noise schedules benefits sample quality.
In sum, we make several contributions specific to diffusion probabilistic modeling of tabular data:
  • We propose a joint continuous diffusion model for both continuous and categorical features such that all noise distributions can be Gaussian.
  • We balance model capacity across continuous and categorical features with a novel and effective loss calibration.
  • We derive a functional form to learn adaptive noise schedules and to allow for an efficient, exact evaluation and thus, greatly extend the idea of timewarping.
  • We drastically improve the scalability of tabular data diffusion models to features with a high number of categories.
  • We boost the quality of the generated samples with adaptive, feature- or type-specific noise schedules.
  • Our novel combination of score estimation methods allows for the application of state-of-the-art techniques, such as classifier-free guidance and accelerated sampling, to mixed-type tabular data for the first time.
Figure 1: CDTD framework. Adaptive noise schedules are trained to fit the (possibly aggregated) MSE and CE losses and transform the uniform timestep $t$ to a potentially feature-specific noise level to diffuse (“noisify”) the scalar values (for continuous features) or the embeddings (for categorical features). Associated sampling processes are highlighted in orange. The approximated score functions are concatenated and passed to a blackbox ODE solver for sample generation.
2 Score-based Generative Framework
We start with a brief outline of the score-based frameworks for continuous and categorical features. Next, we combine these into a single diffusion model to learn the joint distribution of mixed-type data.
2.1 Continuous Features
We denote the $i$-th continuous feature as $x_{\text{cont}}^{(i)}\in{\mathbb{R}}$ and the stacked feature vector as ${\mathbf{x}}_{0}\equiv{\mathbf{x}}_{\text{cont}}\in{\mathbb{R}}^{K_{\text{cont}}}$. Further, let $\{{\mathbf{x}}_{t}\}_{t=0}^{t=1}$ be a diffusion process that gradually adds noise in continuous time $t\in[0,1]$ to ${\mathbf{x}}_{0}$, and let $p_{t}({\mathbf{x}})$ denote the density function of the data at time $t$. Then, this process transforms the real data distribution $p_{0}({\mathbf{x}})$ into a terminal distribution of pure noise $p_{1}({\mathbf{x}})$ from which we can sample. Our goal is to learn the reverse process that allows us to go from noise ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}})$ to a new data sample ${\mathbf{x}}_{0}^{\ast}\sim p_{0}({\mathbf{x}})$.
The forward-pass of this continuous-time diffusion process is formulated as the solution to a stochastic differential equation (SDE):
$\mathrm{d}{\mathbf{x}}={\mathbf{f}}({\mathbf{x}},t)\mathrm{d}t+g(t)\mathrm{d}{\mathbf{w}},$ (1)
where ${\mathbf{f}}(\cdot,t):{\mathbb{R}}^{K_{\text{cont}}}\rightarrow{\mathbb{R}}^{K_{\text{cont}}}$ is the drift coefficient, $g(\cdot):{\mathbb{R}}\rightarrow{\mathbb{R}}$ is the diffusion coefficient, and ${\mathbf{w}}$ is a Brownian motion (Song et al., 2021). The reversion yields the trajectory of ${\mathbf{x}}$ as $t$ goes backwards in time from $1$ to 00, and is formulated as a probability flow ordinary differential equation (ODE):
$\mathrm{d}{\mathbf{x}}=\bigl{[}{\mathbf{f}}({\mathbf{x}},t)-\frac{1}{2}g(t)^{2}\nabla_{{\mathbf{x}}}\log p_{t}({\mathbf{x}})\bigr{]}\mathrm{d}t.$ (2)
We approximate the score function $\nabla_{{\mathbf{x}}}\log p_{t}({\mathbf{x}})$, the only unknown in Equation 2, by training a time-dependent score-based model ${\mathbf{s}}_{\bm{\theta}}({\mathbf{x}},t)$ via score matching (Hyvärinen, 2005). The parameters ${\bm{\theta}}$ are trained to minimize the denoising score matching objective:
$\mathbb{E}_{t}\bigl{[}\lambda_{t}\mathbb{E}_{{\mathbf{x}}_{0}}\mathbb{E}_{{\mathbf{x}}_{t}|{\mathbf{x}}_{0}}\left\lVert s_{\bm{\theta}}({\mathbf{x}}_{t},t)-\nabla_{{\mathbf{x}}_{t}}\log p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})\right\rVert_{2}^{2}\bigr{]},$ (3)
where $\lambda_{t}:[0,1]\rightarrow{\mathbb{R}}_{+}$ is a positive weighting function for timesteps $t\sim\mathcal{U}_{[0,1]}$, and $p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})$ is the density of the noisy ${\mathbf{x}}_{t}$ given the ground-truth data ${\mathbf{x}}_{0}$ (Vincent, 2011).
In this paper, we use the EDM formulation (Karras et al., 2022), that is, ${\mathbf{f}}(\cdot,t)={\bm{0}}$ and $g(t)=\sqrt{2[\frac{d}{dt}\sigma(t)]\sigma(t)}$ such that $p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})=\mathcal{N}({\mathbf{x}}_{t}|{\mathbf{x}}_{0},\sigma^{2}(t)I_{K_{\text{cont}}})$. We start the reverse process with sampling ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}})=\mathcal{N}({\bm{0}},\sigma^{2}(1)I_{K_{\text{cont}}})$ for $\sigma^{2}(1)$ being sufficiently large and $\mathbb{E}[{\mathbf{x}}_{0}]={\bm{0}}$. We then gradually guide ${\mathbf{x}}_{1}$ towards high density regions in the data space with ${\mathbf{s}}_{\bm{\theta}}({\mathbf{x}},t)$ replacing the unknown, true score function in Equation 2. In practice, ODE or predictor-corrector samplers can be used for this iterative denoising process (Song et al., 2021).
2.2 Categorical Features
For categorical features, the score function is undefined and the score-based generative framework cannot be directly applied. Therefore, we use score interpolation (Dieleman et al., 2022) and push the diffusion of categorical data into Euclidean embedding space. This allows us to impose a Gaussian noise distribution on categorical and continuous features.
Let $x_{\text{cat}}^{(j)}\in\{1,\dots,C_{j}\}$ be the $j$-th categorical feature with $C_{j}$ possible classes. The $i$-th class is associated to a distinct, trainable $d$-dimensional embedding vector ${\mathbf{e}}_{i}^{(j)}\in{\mathbb{R}}^{d}$. The embedding vector corresponding to the ground truth class is ${\mathbf{x}}_{0}^{(j)}\in\{{\mathbf{e}}_{1}^{(j)},\dots,{\mathbf{e}}_{C_{j}}^{(j)}\}$. We omit the superscript $j$ in what follows. The noisy variant at time $t$ is ${\mathbf{x}}_{t}\sim p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})=\mathcal{N}({\mathbf{x}}_{t}|{\mathbf{x}}_{0},\sigma^{2}(t)I_{d})$ similar to before. Given ${\mathbf{x}}_{t}$ and $t$, $\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}[\nabla_{{\mathbf{x}}_{t}}\log p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})]$ minimizes Equation 3, and we obtain
$\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}\nabla_{{\mathbf{x}}_{t}}\log p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})=\frac{1}{\sigma_{t}^{2}}[\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}[{\mathbf{x}}_{0}]-{\mathbf{x}}_{t}]\ .$ (4)
We compute $\hat{{\mathbf{x}}}_{0}=\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}[{\mathbf{x}}_{0}]$ as an average over the $C_{j}$ possible embedding vectors weighted by their probability. This yields a score function estimate for a single categorical feature. Since $p({\mathbf{x}}_{0}={\mathbf{e}}_{i}|{\mathbf{x}}_{t},t)=p(x_{\text{cat}}^{(j)}=i|{\mathbf{x}}_{t},t)$, an estimate of $p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)$ is obtained via a classifier that predicts $C_{j}$ class probabilities for the $j$-th feature and is trained via cross-entropy (CE) loss.
Likewise, we start the generative process for a categorical feature from an embedding vector ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}})=\mathcal{N}({\bm{0}},\sigma^{2}(1)I_{d})$. We then use score interpolation to iteratively denoise the embedding vector. After the last timestep in the reverse process, that is, $t\approx 0$, we directly infer the generated class from the feature-specific class probabilities, $p({\mathbf{x}}_{0}={\mathbf{e}}_{i}|{\mathbf{x}}_{t},t)$.
This framework can easily be extended to multiple categorical features. All embeddings are trained alongside the model. The model fully takes feature-specific uncertainty at intermediate timesteps into account (for each feature it only has to commit to a category at the final step of generation) which improves the consistency of the generated samples (Dieleman et al., 2022). Therefore, CDTD allows to capture subtle dependencies both within and across data types more accurately.
3 Method
In short, we combine score matching (Equation 3) with score interpolation (Equation 4) to model the joint distribution of mixed-type data. Next, we discuss the important components of our method. In particular, the combination of the different losses for score interpolation and score matching, initialization and loss weighting concerns, and the adaptive type- or feature-specific noise schedule designs.
3.1 General Framework
Figure 1 gives an overview of our Continuous Diffusion for mixed-type Tabular Data (CDTD) framework. The score model is conditioned on (1) all noisy continuous features, (2) the noisy embeddings of all categorical features in Euclidean space, and (3) the timestep $t$ which reflects potentially feature-specific, adaptive noise levels $\sigma_{\text{cont},i}$ and $\sigma_{\text{cat},j}$ for all $i$ and $j$. Additional conditioning information, such as the target feature for classification tasks, are straightforward to add. Note that while the Gaussian noise process acts directly on the continuous features, it acts on the embeddings of the categorical features. This way, we ensure a common continuous noise process for both data types.
During training, the model predicts the ground-truth value for continuous features and the class-specific probabilities for categorical features. During generation, we concatenate the score estimates, $\hat{s}_{\text{cont}}^{(i)}$ and $\hat{s}_{\text{cat}}^{(j)}$, for all features $i$ and $j$, and pass them to an ODE solver together with $\sigma_{\text{cont},i}$ and $\sigma_{\text{cat},j}$, the noise levels retrieved by transforming linearly spaced timesteps with the learned adaptive noise schedules. Further implementation details are provided in Appendix J.
3.2 Homogenization of Data Types
Let $\mathcal{L}_{\text{MSE}}(x_{\text{cont}}^{(i)},t)$ denote the time-weighted MSE (i.e., score matching in Equation 3) loss of the $i$-th continuous feature at a single timestep $t$, and $\mathcal{L}_{\text{CE}}(x_{\text{cat}}^{(j)},t)$ the CE (i.e. score interpolation) loss of the $j$-th categorical feature. Naturally, the two losses are defined on different scales. This leads to an unintended importance weighting of features in the generative process. We assume that an unconditional model should a priori, i.e., without having any information, be indifferent between all features. This reflects the state of the model at the terminal timestep $t=1$ in the diffusion process.
Formally, we aim to find calibrated losses, $\mathcal{L}^{\ast}_{\text{MSE}}$ and $\mathcal{L}^{\ast}_{\text{CE}}$ for all continuous features $i$ and categorical features $j$, such that
$\mathbb{E}[\mathcal{L}^{\ast}_{\text{MSE}}(x_{\text{cont}}^{(i)},1)]=\mathbb{E}[\mathcal{L}^{\ast}_{\text{CE}}(x_{\text{cat}}^{(j)},1)]=1.$ (5)
For continuous features, $\mathbb{E}[\mathcal{L}^{\ast}_{\text{MSE}}(x_{\text{cont}}^{(i)},1)]=1$ follows from standardizing $x_{\text{cont}}^{(i)}$ to zero mean and unit variance. For categorical features, we compute the normalization constant $\mathbb{E}[\mathcal{L}_{\text{CE}}(x_{\text{cat}}^{(j)},1)]$ directly as the CE of each predicted class in proportion to its empirical distribution in the train set (see Appendix A). We then average the calibrated losses to derive the joint loss function at a given timestep:
$\mathcal{L}(t)=\frac{1}{K}\Bigl{[}\sum_{i=1}^{K_{\text{cont}}}\mathcal{L}^{\ast}_{\text{MSE}}(x_{\text{cont}}^{(i)},t)+\sum_{j=1}^{K_{\text{cat}}}\mathcal{L}_{\text{CE}}^{\ast}(x_{\text{cat}}^{(j)},t)\Bigr{]},$ (6)
where $K=K_{\text{cont}}+K_{\text{cat}}$.
The loss calibration and the multiple data modalities have implications for the optimal initialization of the score model. We aim to initialize all feature-specific losses at one. We therefore initialize the output layer weights to zero (like in image diffusion models) and the output biases for continuous features to zero, and rely on the timestep weights of the EDM parameterization (Karras et al., 2022) to achieve a unit loss for all $t$. For the categorical features, we initialize the biases to match the category’s empirical probability in the training set (see Appendix B).
The initial equal importance across all timesteps will naturally change over the course of training. We employ an normalization scheme for the average diffusion loss (Karras et al., 2023; Kingma & Gao, 2023) to allow for changes in relative importance among features but ensure equal importance of all timesteps throughout training. To do so, we learn the time-specific normalization term $Z(t)$ such that $\mathcal{L}(t)/Z(t)\approx 1$. This ensures a consistent gradient signal and can be implemented by training a neural network to predict $\mathcal{L}(t)$ alongside our diffusion model (for details see Appendix C).
3.3 Noise Schedules
Since the optimal noise schedule of one feature impacts the noise schedules of other features, and the different data types have different sensitivities to additive noise, we introduce feature-specific or type-specific noise schedules. For instance, given the same embedding dimension, more noise is needed to remove the same amount of signal from embeddings of features with fewer classes. Likewise, a delayed noise schedule for one feature might improve sample quality as the model can rely on other correlated features that have been (partially) generated first. We make the noise schedules learnable, and therewith adaptive to avoid the reliance on designs for other data modalities.
We investigate the following noise schedule variants: (1) a single adaptive noise schedule, (2) adaptive noise schedules differentiated per data type and (3) feature-specific adaptive noise schedules. We only introduce the feature-specific noise schedules explicitly. The other noise schedule types are easily derived from our argument by appropriately aggregating terms across features.
Feature-specific Noise Schedules.
According to Equation 1, and following the EDM parameterization (Karras et al., 2022), we define the diffusion process of the $i$-th continuous feature as
$\mathrm{d}x_{{}_{\text{cont}}}^{(i)}=\sqrt{2\Bigl{[}\frac{d}{dt}h_{\text{cont},i}(t)\Bigr{]}h_{\text{cont},i}(t)}\mathrm{d}w_{t}^{(i)},$ (7)
and likewise the trajectory of the $j$-th categorical feature as
$\mathrm{d}{\mathbf{x}}_{\text{cat}}^{(j)}=\sqrt{2\Bigl{[}\frac{d}{dt}h_{\text{cat},j}(t)\Bigr{]}h_{\text{cat},j}(t)}\mathrm{d}{\mathbf{w}}_{t}^{(j)},$ (8)
where ${\mathbf{x}}_{\text{cat}}^{(j)}$ is the $d$-dimensional embedding of $x_{\text{cat}}^{(j)}$ in Euclidean space. The feature-specific noise schedules $h_{\text{cont},i}(t)$ and $h_{\text{cat},j}(t)$ represent the standard deviations of the added Gaussian noise such that $\sigma_{\text{cont},i}(t)=h_{\text{cont},i}(t)$ and $\sigma_{\text{cat},j}(t)=h_{\text{cat},j}(t)$. Thus, each continuous feature and each embedded categorical feature is affected by a distinct noise schedule.
Adaptive Noise Schedules.
We aim to learn a noise schedule $h_{k}:t\mapsto\sigma$ for all $K=K_{\text{cont}}+K_{\text{cat}}$ features. Note that $t\in[0,1]$, and with pre-specified minimum and maximum noise levels, we can scale $\sigma_{k}$ to lie in $[0,1]$ as well, without loss of generality. We will learn the feature-specific loss given the noise level, $F_{k}:\sigma_{k}\mapsto\ell_{k}$, alongside the score model, with $\ell_{k}$ the relevant (not explicitly weighted) training loss for the $k$-th feature. Then, our mapping of interest is $h_{k}=\tilde{F}_{k}^{-1}$, that is, the normalized and inverted function $F_{k}$. This encourages the relation between $t$ and $\ell_{k}$ to be linear.
Higher noise levels imply a lower signal-to-noise ratio, and therefore a larger incurred loss for the score model. Accordingly, $F_{k}$ must be a monotonically increasing and S-shaped function. We let $F_{k}=\gamma_{k}F_{\mathrm{d.a.log},k}(\sigma_{k})$ where $\gamma_{k}>0$ is a scaling factor that at $t=1$ enables fitting a loss $\ell_{k}>1$ early on in the training process, and a loss $\ell_{k}<1$ in case conditioning information is included. Further, we use the cdf of the domain-adapted Logistic distribution $F_{\mathrm{d.a.log},k}(\sigma_{k})$, where the input is pre-processed via a Logit function, with parameters $10<\mu_{k}<$ 1 (the location of the inflection point) and $\nu_{k}\geq 1$ (the steepness of the curve). Figure 2 illustrates the effect of the location parameter. The implicit importance of the noise levels is conveniently represented by the corresponding pdf $f_{\mathrm{d.a.log},k}$. To normalize and invert $F_{k}$, we set $\gamma_{k}=1$ and and directly utilize the quantile function $F_{\mathrm{d.a.log},k}^{-1}$. The detailed derivation of all relevant functions is given in Appendix D.
Our functional choice has several advantages. First, each noise schedule can be evaluated exactly without the need for approximations and only requires three parameters. Second, these parameters are well interpretable in the diffusion context and provide information on the inner workings of the model. For instance, for $\mu_{1}<\mu_{2}$, the model starts generating feature 2 before feature 1 in the reverse process. Third, the proposed functional form is less flexible than the original piece-wise linear function (Dieleman et al., 2022) such that an exponential moving average on the parameters is not necessary, and the fit is more robust to “outliers” encountered during training.
We use the adaptive noise schedules during both training and generation. We derive importance weights from $f_{\mathrm{d.a.log},k}$ to fit $h_{k}$ to avoid biasing the noise schedule to timesteps that are frequently sampled during training. Type-specific noise schedules refer to learning two functions $F_{1}$ and $F_{2}$ that predict the respective average loss over all features of a data type. Examples of learned noise schedules are given in Appendix K.
3.4 Additional Customization to Tabular Data
In the diffusion process, we add noise directly to the continuous features but to the embeddings of categorical features. We generally need more noise to remove all signal from the categorical representations. We therefore define type-specific minimum and maximum noise levels: For categorical features, we let $\sigma_{\text{cat},\min}=0.1$ and $\sigma_{\text{cat},\max}=100$; for continuous features, we set $\sigma_{\text{cont},\min}=0.002$ and $\sigma_{\text{cont},\max}=80$ (see Karras et al., 2022).
Figure 2: (Left) pdf ($f_{\mathrm{d.a.log},k}$) and cdf ($F_{\mathrm{d.a.log},k}$) of the domain-adapted Logistic distribution for five different values of the location parameter $\mu_{k}$ and for a given curve steepness $\nu_{k}=3$. (Right) impact of uniform vs. adjusted timewarping initialization on the pdf ($f_{\mathrm{d.a.log},k}$) and the cdf ($F_{\mathrm{d.a.log},k}$).
Lastly, an uninformative initialization of the adaptive noise schedules requires to set $\mu_{k}=0.5,\nu_{k}\approx 1$ and $\gamma_{k}=1$ such that $F_{\mathrm{d.a.log},k}$ corresponds approximately to the cdf of a uniform distribution. We can improve this with a more informative prior: In the image domain, diffusion models allocate substantial capacity towards generating the high level structure before generating details at lower noise levels. In tabular data, the location of features in the data matrix, and therefore the high level structure, is fixed. Instead, we are interested in generating details as accurately as possible, as these influence, for instance, subtle correlations among features. Note that the inflection point, $\mu_{k}$, of our adaptive noise schedule corresponds to the proportion of high (normalized) noise levels (i.e., $\sigma_{k}\geq 0.5$) in the distribution. Therefore, we adjust the initial noise schedules such that low noise levels ($\sigma_{k}<0.5$) are weighted by a factor of 3 relative to high noise levels ($\sigma_{k}\geq 0.5$) (see Figure 2). The proportion of high noise levels is decreased to $\mu_{k}=1/4$. We let $\nu_{k}\approx 1$ for a dispersed initial probability mass and initialize the scaling factor to $\gamma_{k}=1$.
4 Experiments
We benchmark our model against several generative models across multiple datasets. Additionally, we investigate three different noise schedule specifications: (1) a single adaptive noise schedule for both data types (single), (2) continuous and categorical data type-specific adaptive noise schedules (per type), and (3) feature-specific adaptive noise schedules (per feature).
Baseline models.
We use a diverse benchmark set of state-of-the-art generative models for mixed-type tabular data. This includes ARF (Watson et al., 2023), CTGAN (Xu et al., 2019), TVAE (Xu et al., 2019), TabDDPM (Kotelnikov et al., 2023), and CoDi (Lee et al., 2023). Each model follows a different design and/or modeling philosophy. Note that CoDi is an extension of STaSy (Kim et al., 2023, the same group of authors) that has shown to be superior in performance. For scaling reasons, the recently proposed Diffusion Forest (Jolicoeur-Martineau et al., 2024) is not an applicable benchmark. Further details on the respective benchmark models and their implementations are provided in Appendix F and Appendix G. To keep the comparison fair, we use the same architecture for CDTD as TabDDPM, with minor changes to accommodate the different inputs (see Appendix J).
Datasets.
We systematically investigate our model on eleven publicly available datasets. The datasets vary in size, prediction task (regression vs. binary classification), number of continuous and categorical features and their distributions. The number of categories for categorical features varies significantly across datasets (for more details, see Appendix E). We remove observations with missings in the target or any of the continuous features and encode missings in the categorical features as a separate category. All datasets are split in train (60%), validation (20%) and test (20%) partitions, hereinafter denoted $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{valid}}$ and $\mathcal{D}_{\text{test}}$, respectively. For classification tasks, we use stratification with respect to the outcome, we condition the model on the outcome during training and generation, and use the train set proportions for generation. In a last post-processing step, we round the integer-valued continuous features after generation for all models.
4.1 Evaluation Metrics
We evaluate all generative models on four sample quality criteria after generating synthetic data $\mathcal{D}_{\text{gen}}$ of size $\min(\lvert\mathcal{D}_{\text{train}}\rvert,50\,000)$. All metrics are averaged over five random seeds that affect the generative process.
Machine learning efficiency.
We follow the conventional train-synthetic-test-real strategy (see, Borisov et al., 2023; Liu et al., 2023; Kotelnikov et al., 2023; Kim et al., 2023; Xu et al., 2019; Watson et al., 2023). Hence, we train a group of models, consisting of a (logistic/ridge) regression, a random forest and a catboost model, on the data-specific prediction task (the corresponding hyperparameter settings are reported in Appendix I). We compare the model-averaged real test performance, $\text{Perf}(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})$, to the performance trained on the synthetic data, $\text{Perf}(\mathcal{D}_{\text{gen}},\mathcal{D}_{\text{test}})$. We subsample $\mathcal{D}_{\text{train}}$ in case of more than 50 000 observations to upper-bound the computational load. The results are averaged over ten different model seeds (in addition to the five random seeds that impact the sampling process). For regression tasks, we consider the RMSE and for classification tasks, the macro-averaged F1 and AUC scores. We only report $\lvert\text{Perf}(\mathcal{D}_{\text{gen}},\mathcal{D}_{\text{test}})-\text{Perf}(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})\rvert$ in the main part of this paper. An absolute difference close to zero, that is, synthetic and real data induce the same performance, indicates that the generative model performs well.
Detection score.
For each generative model, we report the accuracy of a catboost model that is trained to distinguish between real and generated (fake) samples (Borisov et al., 2023; Liu et al., 2023). First, we subsample the real data subsets, $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{valid}}$ and $\mathcal{D}_{\text{test}}$, to a maximum of 25 000 data samples to limit evaluation time. Then, we construct $\mathcal{D}_{\text{train}}^{\text{detect}},\mathcal{D}_{\text{valid}}^{\text{detect}}$ and $\mathcal{D}_{\text{test}}^{\text{detect}}$ with equal proportions of real and fake samples. We tune each catboost model on $\mathcal{D}_{\text{valid}}^{\text{detect}}$ and report the accuracy of the best-fitting model on $\mathcal{D}_{\text{test}}^{\text{detect}}$ (see Appendix H for details). A (perfect) detection score of 0.5 indicates the model is unable distinguish fake from real samples.
Statistical similarity.
We evaluate the statistical similarity between real and generated training data based on: (1) the Jensen-Shannon divergence (JSD; Lin, 1991) to quantify the difference in categorical distributions, (2) the Wasserstein distance (WD; Ramdas et al., 2017) to quantify the difference in continuous distributions, and (3) the L2 distance of the pair-wise correlation matrices. We use the Pearson correlation coefficient for the correlations between two continuous features, the Theil uncertainty coefficient for the correlations between two categorical features, and the correlation ratio across types (Kotelnikov et al., 2023; Zhao et al., 2021).
Distance to closest record.
That is, the minimum Euclidean distance of a generated data point to any observation in $\mathcal{D}_{\text{train}}$ (Borisov et al., 2023; Zhao et al., 2021). We one-hot encode categorical features and standardize all features to zero mean and unit variance to ensure each feature contributes equally to the distance. We compute the average distance to closest record (DCR) as a robust estimate. For brevity, we report the absolute difference of the DCR of the synthetic data and the DCR of the real test set. A good DCR value, that indicates realistic and sufficiently private data, should be close to zero.
4.2 Results
Table 1 shows the ranks of all generative averaged across all datasets for all considered metrics. The ranks in terms of the F1 and AUC scores are averaged over the classification task datasets. Likewise, the RMSE rank averages include the regression task datasets only. We assign the maximum possible rank when a model was not trainable on a given dataset. This includes TabDDPM, which outputs NaNs for acsincome and diabetes and CoDi, which we consider to be prohibitively expensive to train on diabetes (estimated 14.5 hours) and lending (estimated 60 hours). The dataset-specific results (including standard errors) and average metrics over all datasets are detailed in Appendix N. We provide visualizations of the captured correlations in the synthetic sample compared to the real training set in Appendix M and distribution plots for a qualitative comparison in Appendix L.
Sample quality.
CDTD consistently outperforms the considered benchmark models in most sample quality metrics. We see a major performance edge in terms of the detection score, the L2 distance of the correlation matrices and the regression-based metrics. Only for the Jensen-Shannon divergence, CDTD is dominated by ARF, a tree-based method that is expected to model categorical features particularly well. Interestingly, CDTD performs similar to TabDDPM on F1 scores but outperforms it dramatically for regression tasks. TabDDPM appears to favor modeling categorical features accurately, thereby sacrificing continuous features, as visualized in Appendix M. By utilizing score interpolation, CDTD is able to model intricate correlation structure more accurate than other (multinomial diffusion-based) frameworks. Most importantly, type-specific noise schedules always outperform the feature-specific noise schedule variant and often also the single noise schedule version. This illustrates the importance of accounting for the high heterogeneity in tabular data on the feature type level. Imposing different noise schedules per feature, however, appears to force too many constrains on the model and hence decreases sample quality. Per-feature noise schedules would requires more training steps to converge, as can be seen in Appendix K.
Table 1: Average performance rank of each generative model across eleven datasets. Per performance metric, bold indicates the best, underline the second best result. CoDi is prohibitively expensive to train on lending and diabetes, TabDDPM produces NaNs for acsincome and diabetes, and therefore get assigned rank 8 for these datasets.
ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
RMSE $3.000_{\pm 1.673}$ $6.400_{\pm 1.356}$ $6.200_{\pm 1.833}$ $6.400_{\pm 1.200}$ $5.600_{\pm 1.356}$ $\mathbf{2.400}_{\pm 0.490}$ $\underline{2.600}_{\pm 1.855}$ $3.400_{\pm 2.059}$
F1 $5.167_{\pm 1.951}$ $6.667_{\pm 1.374}$ $6.667_{\pm 0.745}$ $\underline{3.333}_{\pm 2.285}$ $5.333_{\pm 2.494}$ $\underline{3.333}_{\pm 1.247}$ $\mathbf{2.667}_{\pm 1.106}$ $3.000_{\pm 1.915}$
AUC $4.500_{\pm 1.607}$ $7.000_{\pm 1.155}$ $6.500_{\pm 0.764}$ $4.167_{\pm 2.115}$ $6.000_{\pm 2.160}$ $\underline{2.333}_{\pm 1.599}$ $\mathbf{2.167}_{\pm 0.898}$ $3.500_{\pm 1.500}$
L2 dist. of corr. $4.455_{\pm 1.305}$ $6.727_{\pm 1.213}$ $6.455_{\pm 1.157}$ $5.636_{\pm 2.346}$ $5.545_{\pm 1.876}$ $\underline{2.545}_{\pm 1.305}$ $\mathbf{2.091}_{\pm 0.793}$ $2.636_{\pm 1.553}$
Detection score $4.818_{\pm 1.336}$ $7.182_{\pm 1.113}$ $5.636_{\pm 0.881}$ $4.000_{\pm 2.412}$ $6.455_{\pm 1.777}$ $3.273_{\pm 0.862}$ $\mathbf{1.818}_{\pm 1.696}$ $\underline{2.909}_{\pm 1.505}$
JSD $\mathbf{1.273}_{\pm 0.617}$ $6.818_{\pm 1.113}$ $7.182_{\pm 0.716}$ $5.818_{\pm 1.641}$ $5.909_{\pm 1.083}$ $\underline{2.364}_{\pm 0.979}$ $3.000_{\pm 0.603}$ $3.727_{\pm 1.052}$
WD $4.273_{\pm 1.483}$ $6.000_{\pm 1.758}$ $6.455_{\pm 1.157}$ $5.000_{\pm 2.449}$ $6.727_{\pm 1.286}$ $3.000_{\pm 1.348}$ $\mathbf{2.273}_{\pm 1.052}$ $\underline{2.364}_{\pm 1.494}$
DCR $4.909_{\pm 1.564}$ $6.818_{\pm 1.466}$ $5.182_{\pm 2.443}$ $3.636_{\pm 2.837}$ $5.364_{\pm 1.920}$ $3.727_{\pm 2.049}$ $\underline{3.364}_{\pm 1.611}$ $\mathbf{3.091}_{\pm 1.443}$
Figure 3: Average wall-clock time for training and for sampling 1000 samples (excluding acsincome, diabetes and lending).
Training and sampling time.
Figure 3 shows the average wall-clock time over all datasets (for which training was feasible for all models) for training as well as the time for sampling 1000 data points of each baseline model and the most complex CDTD variant. Details are given in Appendix P. The use of embeddings (instead of one-hot encoding) for categorical features drastically reduces the training time. CDTD trains the fastest on average and is able to reach a maximum training time of less than 30 min across all datasets. Thus, CDTD exhibits improved scaling to an increase in the number of categories. The ODE definition of the diffusion process also implies competitive sampling speeds compared to TVAE and CTGAN and in particular, to the diffusion-based benchmarks CoDi and TabDDPM. In light of the vastly improved scalability, the CDTD results in Table 1 could in practice trivially be improved further by simply scaling up the architecture.
Table 2: Ablation study for five CDTD configurations with progressive addition of model components. We report the median performance metrics over acsincome, adult, beijing and churn.
Config. A B C D CDTD (per type)
RMSE (abs. diff.; $\downarrow$) 0.041 0.042 0.043 0.037 0.033
F1 (abs. diff.; $\downarrow$) 0.012 0.013 0.012 0.016 0.015
AUC (abs. diff.; $\downarrow$) 0.004 0.005 0.005 0.004 0.004
L2 distance of corr.  ($\downarrow$) 0.131 0.124 0.146 0.118 0.127
Detection score ($\downarrow$) 0.577 0.583 0.590 0.561 0.560
JSD ($\downarrow$) 0.011 0.011 0.011 0.012 0.013
WD ($\downarrow$) 0.004 0.005 0.005 0.003 0.003
DCR (abs. diff. to test; $\downarrow$) 0.405 0.361 0.386 0.299 0.372
Ablation study.
We conduct an ablation study over four datasets to investigate the separate components of our CDTD framework. The results are given in Table 2 (detailed results are in Appendix O). The baseline model Config. A includes a single noise schedule with the original piece-wise linear formulation (Dieleman et al., 2022) without loss normalization, improved model initialization or adaptive normalization, and the CE and MSE losses are naively averaged. Note that this configuration still is a novel contribution to the literature. Config. B adds our feature homogenization (i.e., loss normalization, improved initialization and adaptive normalization schemes), Config. C adds our proposed functional form for a single noise schedule with uniform initialization, and Config. D imposes per-type noise schedules. Lastly, we add the suggested (low noise level) overweighting timewarping initialization to arrive at the full CDTD (per type) model. We see the switch from the piece-wise linear functional form to our more robust noise schedule variant slightly harms sample quality. However, the per-type variant and the more informed initialization scheme compensate for this difference. Main differences are in the RMSE and detection score as well as training efficiency (the loss calibration and improved initialization facilitate model convergence). The final model especially works well on the larger datasets compared to the baseline (see Appendix O), as smaller datasets are relatively easy to fit with 3 million parameters, even without any model improvements.
5 Conclusion and Discussion
We propose a Continuous Diffusion model for mixed-type Tabular Data (CDTD) that combines score matching and score interpolation and imposes Gaussian diffusion processes on both continuous and embedded categorical features. We compared to various benchmark models and to a single noise schedule as typically used in image diffusion models. Our results indicate that acknowledging the high feature heterogeneity in tabular data on the feature type level and aligning type-specific diffusion elements, such as the noise schedules or losses, substantially benefits sample quality. CDTD also shows vastly improved scalability and can accommodate an arbitrary number of categories.
Our paper serves as an important step to adopting the diffusion probabilistic framework customized to tabular data. In particular, the common type of noise schedules allows for a framework that is easy to extend and may accelerate progress on diffusion models for tabular data. Crucially, CDTD allows the direct application of diffusion-related advances from the image domain, like classifier-free guidance, to tabular data. We leave further extensions, e.g., the exploration of accelerated sampling, efficient score model architectures, or different forms of adaptive noise schedules, to future work.
Finally, we want to highlight that synthetic data could be misused to support unwarranted claims. In any case, any generated data should not be blindly trusted and synthetic data based inferences should be compared to results derived from the real data. On the other hand, the correct usage of generative models can allow for better privacy preservation, facilitate data sharing and open science practices.
Limitations
The main limitation of the CDTD model is the addition of another hyperparameter, i.e., the type of noise schedule, and tuning a hyperparameter of a generative model can be a costly endeavor. However, our results also show that a per type schedule is most often optimal. Dieleman et al. (2022) show that the results of score interpolation can be sensitive to the initialization of the embeddings. We have not encountered any problems with the same standard deviation of 0.001 for all datasets. However, this may not apply in general. While the DCR indicates no privacy issues for the benchmark datasets used, additional caution must be taken when generating synthetic data from privacy sensitive sources. Lastly, our model may be outperformed by other generative models specialized for different types of tabular data, such as time-series.
Acknowledgements
This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-7437. We would also like to thank Sander Dieleman for helpful discussions.
Appendix A Loss Calibration
A priori, we let the model be indifferent between features. Therefore, we aim to scale the loss of each feature such that they attain the same loss at the terminal timestep $t=1$. In that case, the signal-to-noise ratio is sufficiently low to approximate a situation in which the model has no information about the data. We are looking for calibrated losses $\mathcal{L}_{\text{MSE}}^{\ast}(x_{\text{cont}}^{(i)},1)$ and $\mathcal{L}_{\text{CE}}^{\ast}(x_{\text{cat}}^{(j)},1)$ which in expectation achieve a unit loss at $t=1$. This way, the model treats all features equally before having seen any information.
For a single scalar feature and given timestep $t$, we can write the empirical denoising score matching loss (see Equation 3) in the EDM parameterization (Karras et al., 2022) as:
$\mathcal{L}_{\text{MSE}}(x_{\text{cont}}^{(i)},t)=\lambda(t)\Bigl{(}\underbrace{c_{\text{skip}}(t)x_{t}+c_{\text{out}}(t)F_{{\bm{\theta}}}^{(i)}}_{s_{\bm{\theta}}(x_{t},t)}-x_{\text{cont}}^{(i)}\Bigr{)}^{2},\$
where $F_{\bm{\theta}}^{(i)}$ denotes the neural network output for feature $i$ that parameterizes the score model $s_{\bm{\theta}}$. The parameters $c_{\text{skip}}(t)=\sigma_{\text{data}}^{2}/(\sigma^{2}(t)+\sigma_{\text{data}}^{2})$ and $c_{\text{out}}(t)=\sigma(t)\cdot\sigma_{\text{data}}/(\sqrt{\sigma^{2}(t)+\sigma_{\text{data}}^{2}})$ depend on $\sigma(t)$ (and $\sigma_{\text{data}}$) and therefore on timestep $t$. For $t\rightarrow 1$, $\sigma(t)$ approaches the maximum noise level $\sigma_{\text{cont},\max}$ and $c_{\text{skip}}(t)\rightarrow 0$ and $c_{\text{out}}(t)\rightarrow 1$ such that the score model directly predicts the data at high noise levels. For $t\rightarrow 0$, the model shifts increasingly towards predicting the error that has been added to the true data. In the EDM parameterization, the explicit timestep weight (used to achieve a unit loss across timesteps at initialization (see Appendix B) is $\lambda(t)=1/c_{\text{out}}(t)^{2}\approx 1$ for $t=1$.
At the terminal timestep $t=1$, we now have:
$\mathbb{E}_{p(x_{\text{cont}}^{(i)})}[\mathcal{L}_{\text{MSE}}(x_{\text{cont}}^{(i)},1)]$ $=\lambda(1)\,\mathbb{E}_{p(x_{\text{cont}}^{(i)})}\Bigl{(}c_{\text{skip}}(1)x_{1}+c_{\text{out}}(1)F_{\bm{\theta}}^{(i)}-x_{\text{cont}}^{(i)}\Bigr{)}^{2},$
$\approx\mathbb{E}_{p(x_{\text{cont}}^{(i)})}\Bigl{(}0\cdot x_{1}+1\cdot F_{\bm{\theta}}^{(i)}-x_{\text{cont}}^{(i)}\Bigr{)}^{2},$
$=\mathbb{E}_{p(x_{\text{cont}}^{(i)})}\Bigl{(}F_{\bm{\theta}}^{(i)}-x_{\text{cont}}^{(i)}\Bigr{)}^{2}.$
Without information, it is optimal to always predict the average value $\mathbb{E}_{p(x_{\text{cont}}^{(i)})}[x_{\text{cont}}^{(i)}]$ and thus, the minimum expected loss becomes:
$\mathbb{E}_{p(x_{\text{cont}}^{(i)})}[\mathcal{L}_{\text{MSE}}(x_{\text{cont}}^{(i)},1)]$ $=\mathbb{E}_{p(x_{\text{cont}}^{(i)})}\Bigl{(}\mathbb{E}_{p(x_{\text{cont}}^{(i)})}[x_{\text{cont}}^{(i)}]-x_{\text{cont}}^{(i)}\Bigr{)}^{2}=\mathrm{Var}[x_{\text{cont}}^{(i)}]\ .$
Therefore, we have $\mathcal{L}_{\text{MSE}}^{\ast}(x_{\text{cont}}^{(i)},1)=\mathcal{L}_{\text{MSE}}(x_{\text{cont}}^{(i)},1)$ as long as we standardize $x_{\text{cont}}^{(i)}$ to unit variance.
For a single categorical feature, $x^{(j)}_{\text{cat}}$ is distributed according to the proportions $p_{c}$ (for categories $c=1,\dots,C$). The denoising model for score interpolation is trained with the CE loss:
$\mathcal{L}_{\text{CE}}(x^{(j)}_{\text{cat}},t)=-\sum_{c=1}^{C}I(x^{(j)}_{\text{cat}}=c)\log F_{{\bm{\theta}},c}^{(j)}\ ,$
where $F_{{\bm{\theta}},c}^{(j)}$ denotes the score model’s prediction of the class probability at timestep $t$. Without information, it is optimal to assign the $c$-th category the same proportion as in the training set. At $t=1$, we thus let $F_{{\bm{\theta}},c}^{(j)}=p_{c}$ such that the minimum loss equals:
$\mathbb{E}_{p(x^{(j)}_{\text{cat}})}[\mathcal{L}_{\text{CE}}(x^{(j)}_{\text{cat}},1)]$ $=-\mathbb{E}_{p(x^{(j)}_{\text{cat}})}\sum_{c=1}^{C}I(x^{(j)}_{\text{cat}}=c)\log F_{{\bm{\theta}},c}^{(j)}\ ,$ (9)
$=-\sum_{c=1}^{C}\mathbb{E}_{p(x^{(j)}_{\text{cat}})}[I(x^{(j)}_{\text{cat}}=c)\log p_{c}]\ ,$ (10)
$=-\sum_{c=1}^{C}p_{c}\log p_{c}.$ (11)
We use the training set proportions to compute the normalization constant $Z_{j}=\leavevmode\nobreak\ -\sum_{c=1}^{C}p_{c}\log p_{c}$ to calibrate the loss for categorical features. Then,
$\mathbb{E}_{p(x^{(j)}_{\text{cat}})}[\mathcal{L}^{\ast}_{\text{CE}}(x^{(j)}_{\text{cat}},1)]=\mathbb{E}_{p(x^{(j)}_{\text{cat}})}[\mathcal{L}_{\text{CE}}(x^{(j)}_{\text{cat}},1)/Z_{j}]=1\ .$
We have thus achieved calibrated losses with respect to the terminal timestep $t=1$, that is, $\mathbb{E}_{p(x_{\text{cont}}^{(i)})}[\mathcal{L}_{\text{MSE}}^{\ast}(x_{\text{cont}}^{(i)},1)]=\mathbb{E}_{p(x^{(j)}_{\text{cat}})}[\mathcal{L}^{\ast}_{\text{CE}}(x^{(j)}_{\text{cat}},1)]=1$ for all continuous features $i$ and categorical features $j$.
Appendix B Output Layer Initialization
At initialization, we want the neural network to reflect the state of no information (see Appendix A). Likewise, our goal is a loss of one across all features and timesteps.
For continuous features $i$, we initialize the output layer weights (and biases) to zero such that the output of the score model for a single continuous feature, $F_{\bm{\theta}}^{(i)}$, is also zero. Since we use the EDM parameterization (Karras et al., 2022), we apply the associated explicit timestep weight $\lambda(t)=\frac{\sigma^{2}(t)+\sigma_{\text{data}}^{2}}{(\sigma(t)\cdot\sigma_{\text{data}})^{2}}$. This is explicitly designed to achieve a unit loss across timesteps at initialization and we show this analytically below. We denote the variances of the data $x_{\text{cont}}^{(i)}$ and of the Gaussian noise $\epsilon$ at time $t$ as $\sigma_{\text{data}}^{2}$ and $\sigma^{2}(t)$, respectively.
$\mathbb{E}_{p(x_{\text{cont}}^{(i)}),p(\epsilon)}[\mathcal{L}^{\ast}_{\text{MSE}}(x_{\text{cont}}^{(i)},t)]$ $=\lambda(t)\,\mathbb{E}_{p(x_{\text{cont}}^{(i)}),p(\epsilon)}\Bigl{(}c_{\text{skip}}(t)(x_{\text{cont}}^{(i)}+\epsilon)+c_{\text{out}}(t)F_{\bm{\theta}}^{(i)}-x_{\text{cont}}^{(i)}\Bigr{)}^{2}\ ,$
$=\lambda(t)\,\mathbb{E}_{p(x_{\text{cont}}^{(i)}),p(\epsilon)}\Bigl{(}c_{\text{skip}}(t)(x_{\text{cont}}^{(i)}+\epsilon)-x_{\text{cont}}^{(i)}\Bigr{)}^{2}\ ,$
$=\frac{\sigma^{2}(t)+\sigma_{\text{data}}^{2}}{(\sigma(t)\cdot\sigma_{\text{data}})^{2}}\mathbb{E}_{p(x_{\text{cont}}^{(i)}),p(\epsilon)}\Bigl{(}\frac{\sigma_{\text{data}}^{2}}{\sigma^{2}(t)+\sigma_{\text{data}}^{2}}(x_{\text{cont}}^{(i)}+\epsilon)-x_{\text{cont}}^{(i)}\Bigr{)}^{2}\ ,$
$=\frac{\sigma^{2}(t)+\sigma_{\text{data}}^{2}}{(\sigma(t)\cdot\sigma_{\text{data}})^{2}}\mathbb{E}_{p(x_{\text{cont}}^{(i)}),p(\epsilon)}\Bigl{(}\frac{\sigma_{\text{data}}^{2}\epsilon-\sigma^{2}(t)x_{\text{cont}}^{(i)}}{\sigma^{2}(t)+\sigma_{\text{data}}^{2}}\Bigr{)}^{2}\ ,$
$=\frac{1}{\sigma^{2}(t)+\sigma_{\text{data}}^{2}}\mathbb{E}_{p(x_{\text{cont}}^{(i)}),p(\epsilon)}\Bigl{(}\frac{\sigma_{\text{data}}}{\sigma(t)}\epsilon-\frac{\sigma(t)}{\sigma_{\text{data}}}x_{\text{cont}}^{(i)}\Bigr{)}^{2}\ ,$
$=\frac{1}{\sigma^{2}(t)+\sigma_{\text{data}}^{2}}\mathbb{E}_{p(x_{\text{cont}}^{(i)}),p(\epsilon)}\Bigl{(}\frac{\sigma_{\text{data}}^{2}}{\sigma^{2}(t)}\epsilon^{2}+\frac{\sigma^{2}(t)}{\sigma_{\text{data}}^{2}}(x_{\text{cont}}^{(i)})^{2}-2\epsilon x_{\text{cont}}^{(i)}\Bigr{)}\ ,$
$=\frac{1}{\sigma^{2}(t)+\sigma_{\text{data}}^{2}}\Bigl{(}\frac{\sigma_{\text{data}}^{2}}{\sigma^{2}(t)}\underbrace{\mathrm{Var}(\epsilon)}_{\sigma^{2}(t)}+\frac{\sigma^{2}(t)}{\sigma_{\text{data}}^{2}}\underbrace{\mathrm{Var}(x_{\text{cont}}^{(i)})}_{\sigma_{\text{data}}^{2}}-2\underbrace{\mathrm{Cov}(\epsilon,x_{\text{cont}}^{(i)})}_{0}\Bigr{)}\ ,$
$=\frac{1}{\sigma^{2}(t)+\sigma_{\text{data}}^{2}}\Bigl{(}\sigma_{\text{data}}^{2}+\sigma^{2}(t)\Bigr{)}=1.$
For categorical features $j$, we initialize the output layer such that the model achieves the respective losses under no information. Using the loss normalization constant $Z_{j}$ (see Appendix A) and dropping the expectation over $p(\epsilon)$, we have
$\mathbb{E}_{p(x_{\text{cat}}^{(j)})}[\mathcal{L}^{\ast}_{\text{CE}}(x_{\text{cat}}^{(j)},t)]=\mathbb{E}_{p(x_{\text{cat}}^{(j)})}[\mathcal{L}_{\text{CE}}(x_{\text{cat}}^{(j)},t)/Z_{j}]=\frac{1}{Z_{j}}\mathbb{E}_{p(x_{\text{cat}}^{(j)})}[\mathcal{L}_{\text{CE}}(x_{\text{cat}}^{(j)},t)].$
Hence, for $E_{p(x_{\text{cat}}^{(j)})}[\mathcal{L}_{\text{CE}}(x_{\text{cat}}^{(j)},t)]=Z_{j}$, we obtain an expected loss of one irrespective of $t$. The neural network outputs a vector of logits $F_{{\bm{\theta}}}^{(j)}$ that are transformed into probabilities with a softmax function for each categorical feature. We denote the $c$-th element of that vector $\mathrm{softmax}(\cdot)_{c}$. Since $Z_{j}$ is derived in Equation 11 by imposing probabilities equal to the training set proportions for that category, $p_{c}$, we have
$\log p_{c}=\log\mathrm{softmax}(F_{{\bm{\theta}}}^{(j)})_{c}=\log\frac{\exp(F_{{\bm{\theta}},c}^{(j)})}{\sum_{k=1}^{C}\exp(F_{{\bm{\theta}},k}^{(j)})}=F_{{\bm{\theta}},c}^{(j)}-\log\sum_{k=1}^{C}\exp(F_{{\bm{\theta}},k}^{(j)}).$
We initialize the neural network such that $F_{{\bm{\theta}},c}^{(j)}=\log p_{c}$ for all $c$. This can be achieved by initializing the output layer weights to zero and the output layer biases to the relevant training set log-proportions of the corresponding class. Hence, this initialization gives us
$F_{{\bm{\theta}},c}^{(j)}-\log\sum_{k=1}^{C}\exp(F_{{\bm{\theta}},k}^{(j)})=\log p_{c}-\log\sum_{k=1}^{C}p_{k}=\log p_{c},$
which in turn leads to an initial loss of $Z_{j}$ for all $t$ and therefore achieves a uniform, calibrated loss of one at initialization similar to the continuous feature case.
Appendix C Adaptive Normalization of the Average Diffusion Loss
Both the loss calibration (see Appendix A) and output layer initialization (see Appendix B) ensure that the losses across timesteps (and features) are equal at initialization. During training, the adaptive noise schedules allow the model to focus automatically on the noise levels that matter most, i.e., where the loss increase is steepest. However, the better the model becomes at a given timestep $t$, the lower the loss at that timestep, and the lower the gradient signal relative to the signal received for timesteps $\tilde{t}>t$. To counteract this, we additionally use an adaptive normalization of the average diffusion loss (averaged over the features) across timesteps. Specifically, we want to weight the average diffusion loss at timestep $t$, $\mathcal{L}(t)$ given in Equation 6, such that the normalized loss is the same (equal to one) for all $t$. Similar methods have been used by Karras et al. (2023) and Kingma & Gao (2023), we follow the latter in the setup of the corresponding network.
We train a neural network alongside our diffusion model to predict $\mathcal{L}(t)$ based on $t$ and using the MSE loss to learn this weighting. In particular, we first compute $c_{\text{noise}}(t)=\log(t)/4$ following the EDM parameterization (Karras et al., 2022). Then, we embed $c_{\text{noise}}$ in frequency space (1024-dimensional) using Fourier features. The result is then passed through a single linear layer to output a scalar value, which is in turn passed through an exponential function to ensure that the prediction ^$\hat{\mathcal{L}}(t)\geq 0$. We initialize the weights and biases to zero, to ensure that at model initialization we have a unit normalization.
Appendix D Derivation of the Functional Timewarping Form
Since higher noise levels, $\sigma$, imply a lower signal-to-noise ratio and, in turn, a larger loss, $\ell$, we know that the loss must be a monotonically increasing and S-shaped function of the noise level. Additionally, the function has to be easy to invert and differentiate. We incorporate this prior information in the functional timewarping form of $F:\sigma\mapsto\ell$. A convenient choice is the cdf of the logistic distribution:
$F_{\log}(y)=\bigl{[}1+\exp\bigl{(}-\nu(y-\mu^{\ast})\bigr{)}\bigr{]}^{-1},$ (12)
where $\mu^{\ast}$ describes the location of the inflection point of the S-shaped function and $\nu\geq 1$ indicates the steepness of the curve.
We let $y=\text{logit}(\sigma)=\log(\sigma/(1-\sigma))$ to change the domain of $F_{\log}$ from (,)($\infty$,$\infty$)( - ∞ , ∞ ) to $(0,1)$. The latter covers all possible values of the noise level $\sigma$ scaled to $[0,1]$ with the pre-specified minimum and maximum noise levels $\sigma_{\min}$ and $\sigma_{\max}$. To define the parameter $\mu$ in the same space and ensure that $10<\mu<$ 1, we also let $\mu^{\ast}=\text{logit}(\mu)$. Accordingly, we derive the cdf of the domain-adapted Logistic distribution:
$F_{\mathrm{d.a.log}}(\sigma)=\Biggl{[}1+\biggl{(}\frac{\sigma}{1-\sigma}\frac{1-\mu}{\mu}\biggr{)}^{-\nu}\Biggr{]}^{-1}.$ (13)
Since $\ell$ is not bounded, we introduce a multiplicative scale parameter, $\gamma>0$, such that for timewarping we predict the potentially feature-specific loss as ^$\hat{\ell}=F(\sigma)=\gamma F_{\mathrm{d.a.log}}(\sigma)$. $F_{\mathrm{d.a.log}}$ can also be initialized to the cdf of the uniform distribution with $\mu=0.5$, $\nu\approx 1$ and $\gamma=1$ such that all noise levels are initially equally weighted. However, an initial overweighting of lower noise levels is beneficial for tabular data (see also Section 3.4).
Likewise, we can derive the inverse cdf $F_{\mathrm{d.a.log}}^{-1}(t)$, that is our mapping of interest from timestep $t$ to noise level $\sigma$, in closed form:
$\sigma=F_{\mathrm{d.a.log}}^{-1}(t)=\text{sigmoid}(c),\text{ with }c=\ln\biggl{(}\frac{\mu}{1-\mu}\biggr{)}+\frac{1}{\nu}\ln\biggl{(}\frac{t}{1-t}\biggr{)}\ .$ (14)
When training the diffusion model, we learn the parameters of $F_{\mathrm{d.a.log}}$ as well as $\gamma$ by predicting the diffusion loss using $F(\sigma)$ and the noise levels scaled to $[0,1]$. At the beginning of each training step, we then use the current state of the parameters and $F_{\mathrm{d.a.log}}^{-1}$, with a sampled timestep $t\sim\mathcal{U}_{[0,1]}$ as input, to derive $\sigma$. To allow for feature-specific, adaptive noise schedules, we separately introduce $F_{k}(\sigma_{k})$ for each feature $k$, to predict the feature-specific loss $\ell_{k}$ based on the feature-specific scaled noise level $\sigma_{k}$.
Note that with timewarping we create a feedback loop in which we generate more and more $\sigma$s from the region of interest, decreasing the number of observations available to learn the parameters in different noise level regions. We thus weight the timewarping loss, ||^$\lvert\lvert\ell-\hat{\ell}\rvert\rvert^{2}_{2}$, when fitting $F(\sigma)$ to the data by the reciprocal of the pdf $f_{\mathrm{d.a.log}}(\sigma)$ to mitigate this adverse effect (see Dieleman et al., 2022). Again, this function is available to us in closed form. With $F_{\log}$ and $f_{\log}$ denoting the respective cdf and pdf of the Logistic distribution, we have
$f_{\mathrm{d.a.log}}(\sigma)$ $=\left.\frac{\partial}{\partial y}F_{\log}(y)\right|_{y=\text{logit}(\sigma)}\frac{\partial}{\partial\sigma}\ln\frac{\sigma}{1-\sigma}$
$=f_{\log}(\text{logit}(\sigma))\frac{1}{\sigma(1-\sigma)}$
$=\frac{\nu}{\sigma(1-\sigma)}\cdot\frac{Z(\sigma,\mu,\nu)}{\bigl{(}1+Z(\sigma,\mu,\nu)\bigr{)}^{2}},$
where we defined $Z(\sigma,\mu,\nu)=\bigl{(}\frac{\sigma}{1-\sigma}\frac{1-\mu}{\mu}\bigr{)}^{-\nu}$ and used the definitions of $f_{\log}$ and the parameter $\mu^{\ast}$.
Appendix E Datasets
Our selected benchmark datasets are highly diverse, particularly in the number of categories for a categorical feature. Almost all datasets are licensed under creative commons that permits their usage in our research. Only the nmes dataset license is unknown but the data is publicly accessible on the author’s website.
For the diabetes and covertype datasets, we transform the original multi-class classification problem into a binary classification task for easier comparison. For the covertype data, the task is converted into predicting whether a forest of type 2 is present in a given 30 $\times$ 30 meter area. In the diabetes data, we convert the task by predicting whether a patient was readmitted to a hospital.
Table 3: Overview of the selected experimental datasets. We count the outcome towards the respective features that remain after removing continuous features with an excessive number of missings. The minimum and maximum number of categories are taken over all categorical features.
Dataset License Prediction task Total no. No. of features No. of categories
observations categorical continuous    min. max.
acsincome (Ding et al., 2021) CC0 regression 1 664 500 8 3 2 529
adult (Becker & Kohavi, 1996) CC BY 4.0 binary classification 48 842 9 6 2 42
bank (Moro et al., 2012) CC BY 4.0 binary classification 41 188 11 10 2 12
beijing (Chen, 2017) CC BY 4.0 regression 41 757 1 10 4 4
churn (Keramati et al., 2014) CC BY 4.0 binary classification. 3 150 5 9 2 5
covertype (Blackard, 1998) CC BY 4.0 binary classification 581 012 44 10 2 2
default (Yeh, 2016) CC BY 4.0 binary classification 30 000 10 14 2 11
diabetes (Clore et al., 2014) CC BY 4.0 binary classification 101 766 28 9 2 716
lending (Club, 2015) DbCL 1.0 regression 9 182 10 34 2 3151
news (Fernandes et al., 2015) CC BY 4.0 regression 39 644 14 46 2 2
nmes (Deb & Trivedi, 1997) unknown regression 4 406 8 11 2 4
Appendix F Baseline Models
Below, we give a brief description of the used generative baseline models (including code sources).
ARF
(Watson et al., 2023) – a recent generative approach that is based on a random forest for density estimation. The implementation is available at https://github.com/bips-hb/arfpy and licensed under the MIT license. We use package version 0.1.1. For training, we utilize 16 CPU cores.
CTGAN
(Xu et al., 2019) – one of the most popular Generative-Adversarial-Network-based models for tabular data. The implementation is available as part of the Synthetic Data Vault (Patki et al., 2016) at https://github.com/sdv-dev/CTGAN and licensed under the Business Source License 1.1. We use package version 0.9.0.
TVAE
(Xu et al., 2019) – a Variational-Autoencoder-based model for tabular data. Similar to CTGAN. The implementation is available as part of the Synthetic Data Vault (Patki et al., 2016) at https://github.com/sdv-dev/CTGAN and licensed under the Business Source License 1.1. We use package version 0.9.0. Note that since we only use TVAE (and CTGAN) as benchmark, and do not provide a synthetic data creation service, the license permits the free usage.
TabDDPM
(Kotelnikov et al., 2023) – a diffusion-based generative model for tabular data that combines multinomial diffusion (Hoogeboom et al., 2021) and diffusion in continuous space. An implementation is available as part of the synthcity package (Qian et al., 2023) at https://github.com/vanderschaarlab/synthcity/ and licensed under the Apache 2.0 license. We use package version 0.2.7 with slightly adjusted code to allow for the manual specification of categorical features.
CoDi
(Lee et al., 2023) – a diffusion model trained using an additional contrastive loss, and which factorizes the joint distribution of mixed-type tabular data into a distribution for continuous data conditional on categorical features and a distribution for categorical data conditional on continuous features. Similar to TabDDPM, the authors utilize the multinomial diffusion framework (Hoogeboom et al., 2021) to model categorical data. An implementation is available at https://github.com/ChaejeongLee/CoDi under an unknown license.
Appendix G Implementation Details
Each of the selected benchmark models requires a rather different, more specialized neural network architecture. Imposing the same architecture across models is therefore not possible. The same inability holds for the comparison of CDTD to other diffusion-based models: Our model is the first to use a continuous noise distribution on both continuous and categorical features, and therefore the alignment of important design choices, like the noise schedule, across models is not possible. In particular, the Markov transition matrices based forward process of the multinomial diffusion framework (Hoogeboom et al., 2021) in TabDDPM and CoDi does not translate to our setting.
To ensure a fair comparison in terms of sampling steps, we set the steps for CDTD, TabDDPM and CoDi to $\max(200,\text{default})$. We therefore increase the default number of sampling steps for CoDi (from 50 steps) and TabDDPM (from 100 steps for classification datasets). For TabDDPM and regression datasets, we use the suggested default of 1000 sampling steps.
Table 4: Total number of trainable parameters per model on the adult dataset.
Model Trainable parameters
CTGAN 3 000 397
TVAE 2 996 408
TabDDPM 3 003 924
CoDi 2 998 043
CDTD (per type schedule, TabDDPM architecture) 2 999 721
We adjust each architecture to a total of ±$\pm 3$ million trainable parameters on the adult dataset to improve the comparability further (see Table 4) and use the same architectures for all considered datasets. Note that the total number of parameters may vary slightly across datasets due to different number of features and categories affecting the onehot encoding but is still comparable across models. We also align the embedding/bottleneck dimensions for CTGAN, TVAE, TabDDPM and CDTD to 256 to accommodate more complex datasets. If applicable, all models are trained for 30k steps on a single RTX 4090 instance, using PyTorch version 2.2.2.
Below, we briefly discuss our model-specific hyperparameter choices.
ARF
(Watson et al., 2023): We use the authors’s suggested default hyperparameters. In particular, we use 20 trees, $\delta=0$ and a minimum node size of 5. We follow the official package implementation and set the maximum number of iterations to 10 (see https://github.com/bips-hb/arfpy).
CTGAN
(Xu et al., 2019): We follow the popular implementation in the Synthetic Data Vault package (see https://github.com/sdv-dev/CTGAN). For this model to work, the batch size must be divisible by 10. Therefore, we adjust the batch size if necessary. We use a 258-dimensional embedding (instead of the default embedding dimension of 128) to better align the CTGAN architecture with TVAE, TabDDPM and CDCD.
TVAE
(Xu et al., 2019): We again follow the implementation in the Synthetic Data Vault. We use a 258-dimensional embedding to better align the architecture with CTGAN, TabDDPM and CDCD.
TabDDPM
(Kotelnikov et al., 2023): There are no general default hyperparameters provided. Hence, we mostly adapt the papers’ tuned hyperparameters for the adult dataset (one of the few used datasets that includes both continuous and categorical features). However, we decrease the learning rate from 0.002 to 0.001, since most of the tuned models in the paper used learning rates around 0.001. For regression task datasets, we use 1000 sampling steps in accordance with the author’s settings. For classification task datasets, we use 200 sampling steps (instead of the default 100 steps), to better align the model with CoDi and CDCD. Note also that for classification task datasets, TabDDPM models the conditional distribution $p(x|y)$, instead of the unconditional distribution $p(x)$ which is modeled for regression tasks. We adjust the dimension of the bottleneck to 256 (instead of the default 128) to also accommodate also larger datasets and align the model with CTGAN, TVAE,and CDCD.
CoDi
(Lee et al., 2023): We use the default hyperparameters from the official code (see https://github.com/ChaejeongLee/CoDi).
CDTD
(ours): To ensure comparability in particular to TabDDPM and CoDi, we use the same neural network architecture as TabDDPM. We only change the input layers to accommodate our embedding-based framework. In the input layer, we vectorize all embedded categorical features and concatenate them with the scalar valued continuous features. The adjusted output layer ensures that we predict a single value for each continuous features and set of class-specific probabilities for each categorical feature. Since our use of embeddings introduces additional parameters, we scale the hidden layers slightly down relative to the TabDDPM to ensure approximately 3 million trainable parameters (instead of 808 neurons per layer we use 806) on the adult dataset. More details on the CDTD implementation are given in Appendix J.
Appendix H Tuning of the Detection Model
We use a catboost model (Prokhorenkova et al., 2018) to test whether real and generated samples can be distinguished. We generate the same number of fake observations for each of the real train, validation and test sets. We cap the maximum size of the real data subsets to 25 000, and subsample them if necessary, to limit the computational load. We then combine real and fake observations per set to $\mathcal{D}_{\text{train}}^{\text{detect}},\mathcal{D}_{\text{valid}}^{\text{detect}}$, and $\mathcal{D}_{\text{test}}^{\text{detect}}$, respectively. The catboost model is trained on $\mathcal{D}_{\text{train}}^{\text{detect}}$ with the task of predicting whether an observation is real or fake. We tune the catboost model with optuna and for 50 trials to maximize the accuracy on $\mathcal{D}_{\text{valid}}^{\text{detect}}$. The catboost hyperparameter search space is given in Table 5. Afterwards, we repeat the sampling process and the creation of $\mathcal{D}_{\text{train}}^{\text{detect}},\mathcal{D}_{\text{valid}}^{\text{detect}}$ and $\mathcal{D}_{\text{test}}^{\text{detect}}$ for five different seeds. Each time, the model is trained on $\mathcal{D}_{\text{train}}^{\text{detect}}$ with the previously tuned hyperparameters, and evaluated on $\mathcal{D}_{\text{test}}^{\text{detect}}$. The average test set accuracy over the five seeds yields the estimated detection score.
Table 5: Catboost hyperparameter space settings. The model is tuned for 50 trials.
Parameter Distribution
no. iterations = 1000
learning rate Log Uniform [0.001, 1.0]
depth Cat([3,4,5,6,7,8])
L2 regularization Uniform [0.1, 10]
bagging temperature Uniform [0, 1]
leaf estimation iters Integer Uniform [1, 10]
Appendix I Machine Learning Efficiency Models
For the group of machine learning efficiency models, we use the scikit-learn and catboost package implementations including the default parameter settings, if not specified otherwise below:
Logistic or Ridge Regression:
max. iterations = 1000
Random Forest:
max. depth = 12, no. estimators = 100
Catboost:
no. iterations = 2000, early stopping rounds = 50, overfitting detector pval = 0.001
Appendix J CDTD Implementation Details
Figure 4: Overview of the CDTD architecture adapted from TabDDPM. The dimensions of the inputs and layer outputs are stated in the lower-left hand corner for a continuous features $x_{\text{cont}}$ and a categorical features $x_{\text{cat}}$. Note that each categorical features can have a different number of categories $\lvert C\rvert$, impacting the output dimension of the final layer. Scalars are colored orange, embeddings red and linear layers blue. The positional embedding highlighted in green refers to the positional sinusoidal embedding. CDTD only conditions on $y$, i.e., the target feature, for classification task datasets.
To enable a fair comparison to the other methods and in particular TabDDPM, the CDTD score model utilizes the exact same architecture and optimizer as Kotelnikov et al. (2023). An overview of the score model is given in Figure 4: First, the noisy data, including the noisy scalars for continuous features and the noisy embeddings for categorical features, are projected onto a 256-dimensional space. Similarly, timestep $t$ and possibly conditioning information $y$ are embedded in the same space. All 256-dimensional vectors are then added and the results is processed by a set of five fully-connected linear layers with ReLU activation functions. Lastly, a linear projection maps the output of the fully-connected layers to the required output dimensions, which depend on the number of features and number of categories per feature.
The only major difference to the TabDDPM setup are the inputs, as we need to embed the categorical features in Euclidean space. The output dimensions are the same, as we need to predict a single scalar for each $x_{\text{cont},i}$, and $\lvert C_{j}\rvert$ values for each $x_{\text{cat},j}$, with $C_{j}$ the set of categories of feature $j$. We change the initialization of the output layer as described in Appendix B: To handle our inputs, we embed the categorical features in 16-dimensional space and add a feature-specific bias of the same dimension, which captures feature-specific information common to all categories and is initialized to zero. We L2-normalize each embedding to prevent a degenerate embedding space in which embeddings are pushed further and further apart (see Dieleman et al., 2022). Also, Dieleman et al. (2022) argue that the standard deviation of the Normal distribution used to initialize the embeddings, denoted by $\sigma_{\text{cat emb}}$, is an important hyperparameter. In this paper, we set $\sigma_{\text{cat emb}}=0.001$ for all datasets and have not seen detrimental effects. Tuning this hyperparameter in practice could improve the results further.
Since we utilize embeddings, we have to scale the neurons per layer slightly down in the stack of the five fully-connected layers (from 808 for TabDDPM to 806). Also, since TabDDPM samples discrete steps from $[0,T]$, with $T\gg 1$, we adjust our timesteps $t\in[0,1]$ up by 1000. We use the same optimizer (Adam), learning rate (0.001), learning rate decay (linear), EMA decay (0.999), and training steps (30000). However, since we work with embeddings we add a linear warmup schedule over the first 100 steps.
Instead of using the vanilla uniform (time)step sampling as the TabDDPM, the CDTD model uses antithetic sampling Dieleman et al. (2022); Kingma et al. (2022). The timesteps are still uniformly distributed but spread out more evenly over the domain, which benefits the training of the adaptive noise schedules. For generation, we use an Euler sampler with 200 steps to minimize the discretization error.
Appendix K Examples of Learned Noise Schedules
Below we show the learned noise schedules for the smallest dataset (churn) and the largest dataset (acsincome). We also illustrate how well single, per type and per feature schedules fit the respective losses for each of the datasets.
Figure 5: (Left): Learned noise schedules for churn. This reflects $F_{\mathrm{d.a.log},k}^{-1}$. (Right): Implicit weighting of noise levels / timesteps. This visualizes $f_{\mathrm{d.a.log},k}$.
Figure 6: Illustration of the goodness of fit of the timewarping function $F_{k}$ for single (left), per type (middle) and per feature noise schedules (right) on the acsincome data.
Figure 7: (Left): Learned noise schedules for acsincome. This reflects $F_{\mathrm{d.a.log},k}^{-1}$. (Right): Implicit weighting of noise levels / timesteps. This visualizes $f_{\mathrm{d.a.log},k}$.
Figure 8: Illustration of the goodness of fit of the timewarping function $F_{k}$ for single (left), per type (middle) and per feature noise schedules (right) on the churn data.
Appendix L Qualitative Comparisons
Figure 9: Bivariate density for age and educational-num from the adult data.
Figure 10: Qualitative comparison of some univariate distributions for adult, bank, churn, nmes data.
Figure 11: Qualitative comparison of some univariate distributions for beijing, default, lending, news data. Note that CoDi is excluded from lending as it is prohibitively expensive to train on this dataset.
Appendix M Visualizations of Captured Correlations
Figure 12: Element-wise absolute differences of the correlation matrices between the real training set and the synthetic data for the acsincome dataset. TabDDPM generates NaNs for this dataset and is therefore excluded. Continuous (cont.) and categorical (cat.) features are indicated on the axes.
Figure 13: Element-wise absolute differences of the correlation matrices between the real training set and the synthetic data for the adult dataset. Continuous (cont.) and categorical (cat.) features are indicated on the axes.
Figure 14: Element-wise absolute differences of the correlation matrices between the real training set and the synthetic data for the bank dataset. Continuous (cont.) and categorical (cat.) features are indicated on the axes.
Figure 15: Element-wise absolute differences of the correlation matrices between the real training set and the synthetic data for the beijing dataset. Continuous (cont.) and categorical (cat.) features are indicated on the axes.
Figure 16: Element-wise absolute differences of the correlation matrices between the real training set and the synthetic data for the churn dataset. Continuous (cont.) and categorical (cat.) features are indicated on the axes.
Figure 17: Element-wise absolute differences of the correlation matrices between the real training set and the synthetic data for the covertype dataset. Continuous (cont.) and categorical (cat.) features are indicated on the axes.
Figure 18: Element-wise absolute differences of the correlation matrices between the real training set and the synthetic data for the default dataset. Continuous (cont.) and categorical (cat.) features are indicated on the axes.
Figure 19: Element-wise absolute differences of the correlation matrices between the real training set and the synthetic data for the diabetes dataset. TabDDPM generates NaNs for this dataset and is therefore excluded. CoDi is prohibitively expensive to train and therefore excluded. Continuous (cont.) and categorical (cat.) features are indicated on the axes.
Figure 20: Element-wise absolute differences of the correlation matrices between the real training set and the synthetic data for the lending dataset. CoDi is prohibitively expensive to train and therefore excluded. Continuous (cont.) and categorical (cat.) features are indicated on the axes.
Figure 21: Element-wise absolute differences of the correlation matrices between the real training set and the synthetic data for the news dataset. Continuous (cont.) and categorical (cat.) features are indicated on the axes.
Figure 22: Element-wise absolute differences of the correlation matrices between the real training set and the synthetic data for the nmes dataset. Continuous (cont.) and categorical (cat.) features are indicated on the axes.
Appendix N Detailed Results
CoDi is prohibitively expensive to train on lending and diabetes and TabDDPM produces NaNs for acsincome and diabetes. For both models, the performance metrics on these datasets are therefore not reported.
Table 6: Model evaluation results averaged over 11 datasets (skipping a dataset if the model was not trainable on it) for five benchmark models and for CDTD with three different noise schedules. Per performance metric, bold indicates the best, underline the second best result.
ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
RMSE (abs. diff.; $\downarrow$) 0.094 0.674 0.947 0.486 0.173 0.084 0.101 0.110
F1 (abs. diff.; $\downarrow$) 0.053 0.130 0.074 0.015 0.044 0.025 0.020 0.025
AUC (abs. diff.; $\downarrow$) 0.020 0.080 0.065 0.009 0.027 0.018 0.016 0.022
L2 distance of corr.  ($\downarrow$) 1.321 2.187 2.745 3.786 1.200 0.782 0.756 0.990
Detection score ($\downarrow$) 0.934 0.986 0.976 0.769 0.936 0.796 0.768 0.783
JSD ($\downarrow$) 0.011 0.114 0.152 0.051 0.038 0.015 0.016 0.018
WD ($\downarrow$) 0.011 0.023 0.025 0.061 0.022 0.010 0.007 0.009
DCR (abs. diff. to test; $\downarrow$) 1.588 3.336 1.621 0.568 1.000 0.796 0.806 0.758
Table 7: L2 norm (incl. standard errors in subscripts) of the correlation matrix differences of real and synthetic train sets for five benchmark models and for CDTD with three different noise schedules.
ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
acsincome $0.242_{\pm 0.002}$ $1.696_{\pm 0.008}$ $1.136_{\pm 0.004}$ $0.517_{\pm 0.006}$ $0.141_{\pm 0.003}$ $0.129_{\pm 0.003}$ $0.119_{\pm 0.002}$
adult $0.576_{\pm 0.006}$ $1.858_{\pm 0.010}$ $0.735_{\pm 0.012}$ $0.156_{\pm 0.006}$ $0.493_{\pm 0.009}$ $0.170_{\pm 0.007}$ $0.125_{\pm 0.009}$ $0.128_{\pm 0.010}$
bank $0.819_{\pm 0.024}$ $0.947_{\pm 0.019}$ $2.758_{\pm 0.049}$ $0.898_{\pm 0.025}$ $0.499_{\pm 0.021}$ $0.323_{\pm 0.008}$ $0.266_{\pm 0.011}$ $0.256_{\pm 0.015}$
beijing $0.133_{\pm 0.006}$ $1.445_{\pm 0.009}$ $1.642_{\pm 0.015}$ $1.133_{\pm 0.035}$ $0.363_{\pm 0.015}$ $0.075_{\pm 0.008}$ $0.073_{\pm 0.009}$ $0.071_{\pm 0.004}$
churn $0.635_{\pm 0.026}$ $1.355_{\pm 0.043}$ $1.301_{\pm 0.041}$ $0.327_{\pm 0.044}$ $0.746_{\pm 0.062}$ $0.302_{\pm 0.041}$ $0.289_{\pm 0.043}$ $0.282_{\pm 0.044}$
covertype $1.192_{\pm 0.017}$ $3.685_{\pm 0.005}$ $4.668_{\pm 0.003}$ $1.044_{\pm 0.001}$ $1.029_{\pm 0.032}$ $2.359_{\pm 0.011}$ $2.275_{\pm 0.009}$ $2.710_{\pm 0.009}$
default $1.228_{\pm 0.021}$ $2.697_{\pm 0.021}$ $1.564_{\pm 0.029}$ $3.408_{\pm 0.105}$ $1.672_{\pm 0.061}$ $0.627_{\pm 0.068}$ $0.652_{\pm 0.102}$ $0.737_{\pm 0.033}$
diabetes $1.189_{\pm 0.004}$ $1.654_{\pm 0.008}$ $5.351_{\pm 0.095}$ $1.201_{\pm 0.020}$ $0.803_{\pm 0.032}$ $1.345_{\pm 0.016}$
lending $3.473_{\pm 0.057}$ $2.420_{\pm 0.016}$ $5.895_{\pm 0.026}$ $10.675_{\pm 0.015}$ $1.042_{\pm 0.075}$ $1.189_{\pm 0.040}$ $1.363_{\pm 0.097}$
news $4.333_{\pm 0.128}$ $4.641_{\pm 0.028}$ $4.612_{\pm 0.016}$ $15.985_{\pm 0.081}$ $4.874_{\pm 0.148}$ $1.925_{\pm 0.527}$ $2.035_{\pm 0.475}$ $3.395_{\pm 0.950}$
nmes $0.717_{\pm 0.054}$ $1.663_{\pm 0.035}$ $0.532_{\pm 0.030}$ $0.447_{\pm 0.031}$ $0.609_{\pm 0.032}$ $0.433_{\pm 0.025}$ $0.478_{\pm 0.083}$ $0.481_{\pm 0.058}$
Table 8: Jensen-Shannon divergence (incl. standard errors in subscripts) for five benchmark models and for CDTD with three different noise schedules.
ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
acsincome $0.013_{\pm 0.001}$ $0.256_{\pm 0.000}$ $0.309_{\pm 0.000}$ $0.076_{\pm 0.001}$ $0.025_{\pm 0.001}$ $0.024_{\pm 0.000}$ $0.022_{\pm 0.001}$
adult $0.007_{\pm 0.001}$ $0.112_{\pm 0.001}$ $0.113_{\pm 0.001}$ $0.034_{\pm 0.001}$ $0.045_{\pm 0.001}$ $0.010_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.016_{\pm 0.000}$
bank $0.004_{\pm 0.000}$ $0.086_{\pm 0.001}$ $0.191_{\pm 0.001}$ $0.029_{\pm 0.001}$ $0.038_{\pm 0.001}$ $0.010_{\pm 0.000}$ $0.009_{\pm 0.001}$ $0.012_{\pm 0.001}$
beijing $0.005_{\pm 0.002}$ $0.147_{\pm 0.003}$ $0.257_{\pm 0.001}$ $0.035_{\pm 0.003}$ $0.018_{\pm 0.004}$ $0.003_{\pm 0.001}$ $0.005_{\pm 0.002}$ $0.005_{\pm 0.001}$
churn $0.011_{\pm 0.004}$ $0.095_{\pm 0.003}$ $0.048_{\pm 0.004}$ $0.014_{\pm 0.004}$ $0.043_{\pm 0.001}$ $0.012_{\pm 0.003}$ $0.012_{\pm 0.002}$ $0.011_{\pm 0.002}$
covertype $0.002_{\pm 0.000}$ $0.044_{\pm 0.000}$ $0.043_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.011_{\pm 0.000}$
default $0.008_{\pm 0.001}$ $0.194_{\pm 0.001}$ $0.177_{\pm 0.001}$ $0.027_{\pm 0.002}$ $0.073_{\pm 0.002}$ $0.013_{\pm 0.001}$ $0.015_{\pm 0.001}$ $0.015_{\pm 0.001}$
diabetes $0.009_{\pm 0.000}$ $0.093_{\pm 0.000}$ $0.187_{\pm 0.000}$ $0.022_{\pm 0.000}$ $0.023_{\pm 0.000}$ $0.026_{\pm 0.000}$
lending $0.049_{\pm 0.002}$ $0.092_{\pm 0.001}$ $0.188_{\pm 0.001}$ $0.243_{\pm 0.002}$ $0.055_{\pm 0.001}$ $0.056_{\pm 0.001}$ $0.064_{\pm 0.002}$
news $0.002_{\pm 0.001}$ $0.022_{\pm 0.001}$ $0.128_{\pm 0.001}$ $0.046_{\pm 0.000}$ $0.012_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$
nmes $0.008_{\pm 0.002}$ $0.117_{\pm 0.002}$ $0.029_{\pm 0.003}$ $0.028_{\pm 0.004}$ $0.027_{\pm 0.003}$ $0.008_{\pm 0.001}$ $0.009_{\pm 0.001}$ $0.013_{\pm 0.003}$
Table 9: Wasserstein distance (incl. standard errors in subscripts) for five benchmark models and for CDTD with three different noise schedules.
ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
acsincome $0.007_{\pm 0.000}$ $0.037_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.017_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.001_{\pm 0.000}$ $0.001_{\pm 0.000}$
adult $0.012_{\pm 0.000}$ $0.016_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$
bank $0.012_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.040_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.030_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$
beijing $0.009_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.058_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.019_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.002_{\pm 0.000}$
churn $0.013_{\pm 0.001}$ $0.027_{\pm 0.001}$ $0.032_{\pm 0.001}$ $0.008_{\pm 0.002}$ $0.048_{\pm 0.002}$ $0.008_{\pm 0.001}$ $0.007_{\pm 0.001}$ $0.006_{\pm 0.001}$
covertype $0.006_{\pm 0.000}$ $0.041_{\pm 0.000}$ $0.022_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.017_{\pm 0.000}$ $0.015_{\pm 0.000}$ $0.012_{\pm 0.000}$
default $0.005_{\pm 0.000}$ $0.011_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$
diabetes $0.012_{\pm 0.000}$ $0.020_{\pm 0.000}$ $0.038_{\pm 0.000}$ $0.038_{\pm 0.000}$ $0.020_{\pm 0.000}$ $0.042_{\pm 0.000}$
lending $0.013_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.016_{\pm 0.000}$ $0.425_{\pm 0.001}$ $0.009_{\pm 0.000}$ $0.010_{\pm 0.000}$ $0.011_{\pm 0.000}$
news $0.024_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.018_{\pm 0.000}$ $0.078_{\pm 0.001}$ $0.030_{\pm 0.000}$ $0.007_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.008_{\pm 0.000}$
nmes $0.012_{\pm 0.000}$ $0.036_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.007_{\pm 0.001}$ $0.016_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.000}$
Table 10: Detection score (incl. standard errors in subscripts) for five benchmark models and for CDTD with three different noise schedules.
ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
acsincome $0.808_{\pm 0.001}$ $0.989_{\pm 0.001}$ $0.985_{\pm 0.000}$ $0.825_{\pm 0.002}$ $0.540_{\pm 0.003}$ $0.532_{\pm 0.004}$ $0.526_{\pm 0.002}$
adult $0.889_{\pm 0.002}$ $0.997_{\pm 0.000}$ $0.967_{\pm 0.001}$ $0.590_{\pm 0.003}$ $0.992_{\pm 0.001}$ $0.604_{\pm 0.002}$ $0.588_{\pm 0.002}$ $0.591_{\pm 0.005}$
bank $0.955_{\pm 0.002}$ $1.000_{\pm 0.000}$ $0.988_{\pm 0.001}$ $0.783_{\pm 0.003}$ $1.000_{\pm 0.000}$ $0.795_{\pm 0.003}$ $0.739_{\pm 0.003}$ $0.694_{\pm 0.006}$
beijing $0.995_{\pm 0.000}$ $0.998_{\pm 0.000}$ $0.993_{\pm 0.001}$ $0.966_{\pm 0.002}$ $0.997_{\pm 0.001}$ $0.951_{\pm 0.002}$ $0.949_{\pm 0.001}$ $0.947_{\pm 0.002}$
churn $0.853_{\pm 0.002}$ $0.945_{\pm 0.006}$ $0.843_{\pm 0.011}$ $0.561_{\pm 0.005}$ $0.730_{\pm 0.012}$ $0.621_{\pm 0.016}$ $0.533_{\pm 0.007}$ $0.544_{\pm 0.031}$
covertype $0.945_{\pm 0.002}$ $0.997_{\pm 0.000}$ $0.989_{\pm 0.001}$ $0.586_{\pm 0.002}$ $0.900_{\pm 0.002}$ $0.991_{\pm 0.001}$ $0.992_{\pm 0.001}$ $0.991_{\pm 0.001}$
default $0.991_{\pm 0.001}$ $0.998_{\pm 0.001}$ $0.997_{\pm 0.001}$ $0.821_{\pm 0.002}$ $0.995_{\pm 0.000}$ $0.827_{\pm 0.004}$ $0.802_{\pm 0.003}$ $0.871_{\pm 0.001}$
diabetes $0.854_{\pm 0.002}$ $0.935_{\pm 0.002}$ $0.997_{\pm 0.001}$ $0.858_{\pm 0.001}$ $0.780_{\pm 0.002}$ $0.866_{\pm 0.002}$
lending $0.997_{\pm 0.001}$ $0.995_{\pm 0.002}$ $0.995_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.955_{\pm 0.006}$ $0.954_{\pm 0.009}$ $0.961_{\pm 0.004}$
news $0.998_{\pm 0.000}$ $1.000_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.966_{\pm 0.002}$ $1.000_{\pm 0.000}$ $0.973_{\pm 0.001}$ $0.953_{\pm 0.001}$ $0.977_{\pm 0.001}$
nmes $0.987_{\pm 0.002}$ $0.992_{\pm 0.003}$ $0.988_{\pm 0.002}$ $0.650_{\pm 0.014}$ $0.988_{\pm 0.000}$ $0.636_{\pm 0.008}$ $0.623_{\pm 0.008}$ $0.642_{\pm 0.010}$
Table 11: Distance to closest record of the generated data (incl. standard errors in subscripts) for five benchmark models and for CDTD with three different noise schedules.
ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
acsincome $8.637_{\pm 0.027}$ $10.758_{\pm 0.054}$ $6.652_{\pm 0.032}$ $10.877_{\pm 0.092}$ $8.346_{\pm 0.056}$ $8.322_{\pm 0.047}$ $8.349_{\pm 0.033}$
adult $2.523_{\pm 0.012}$ $5.012_{\pm 0.028}$ $2.227_{\pm 0.013}$ $1.647_{\pm 0.009}$ $2.735_{\pm 0.028}$ $1.112_{\pm 0.019}$ $1.231_{\pm 0.011}$ $1.294_{\pm 0.009}$
bank $3.025_{\pm 0.017}$ $3.840_{\pm 0.014}$ $3.136_{\pm 0.007}$ $2.327_{\pm 0.010}$ $3.062_{\pm 0.012}$ $1.828_{\pm 0.008}$ $1.943_{\pm 0.007}$ $2.062_{\pm 0.008}$
beijing $0.735_{\pm 0.003}$ $1.004_{\pm 0.006}$ $0.926_{\pm 0.003}$ $0.739_{\pm 0.006}$ $0.610_{\pm 0.002}$ $0.490_{\pm 0.002}$ $0.489_{\pm 0.001}$ $0.477_{\pm 0.002}$
churn $1.136_{\pm 0.015}$ $1.804_{\pm 0.036}$ $1.146_{\pm 0.039}$ $0.342_{\pm 0.031}$ $0.852_{\pm 0.016}$ $0.332_{\pm 0.021}$ $0.274_{\pm 0.021}$ $0.276_{\pm 0.012}$
covertype $1.741_{\pm 0.011}$ $5.773_{\pm 0.017}$ $3.173_{\pm 0.013}$ $0.889_{\pm 0.007}$ $1.508_{\pm 0.020}$ $2.297_{\pm 0.026}$ $2.209_{\pm 0.022}$ $2.252_{\pm 0.013}$
default $3.095_{\pm 0.026}$ $5.880_{\pm 0.020}$ $3.216_{\pm 0.013}$ $1.422_{\pm 0.013}$ $2.593_{\pm 0.020}$ $1.127_{\pm 0.028}$ $1.269_{\pm 0.014}$ $1.253_{\pm 0.012}$
diabetes $17.736_{\pm 0.107}$ $21.935_{\pm 0.046}$ $8.214_{\pm 0.022}$ $15.279_{\pm 0.026}$ $15.126_{\pm 0.058}$ $15.350_{\pm 0.059}$
lending $17.776_{\pm 0.132}$ $20.239_{\pm 0.222}$ $10.688_{\pm 0.025}$ $12.537_{\pm 0.076}$ $13.775_{\pm 0.147}$ $14.162_{\pm 0.188}$ $13.966_{\pm 0.282}$
news $6.147_{\pm 0.010}$ $4.789_{\pm 0.005}$ $5.821_{\pm 0.003}$ $4.960_{\pm 0.006}$ $4.661_{\pm 0.023}$ $3.635_{\pm 0.004}$ $3.687_{\pm 0.006}$ $3.749_{\pm 0.048}$
nmes $2.203_{\pm 0.028}$ $2.971_{\pm 0.008}$ $1.710_{\pm 0.019}$ $0.891_{\pm 0.033}$ $1.231_{\pm 0.024}$ $0.664_{\pm 0.029}$ $0.710_{\pm 0.032}$ $0.771_{\pm 0.023}$
Table 12: Machine learning efficiency F1 score for five benchmark models, the real training data and for CDTD with three different noise schedules. The standard deviation takes into account five different sampling seeds and uses the average results of the four machine learning efficiency models computed across ten model seeds.
Real Data ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
adult $0.797_{\pm 0.000}$ $0.769_{\pm 0.002}$ $0.647_{\pm 0.015}$ $0.756_{\pm 0.002}$ $0.787_{\pm 0.001}$ $0.745_{\pm 0.004}$ $0.787_{\pm 0.001}$ $0.787_{\pm 0.001}$ $0.787_{\pm 0.001}$
bank $0.745_{\pm 0.002}$ $0.682_{\pm 0.006}$ $0.680_{\pm 0.006}$ $0.629_{\pm 0.006}$ $0.720_{\pm 0.006}$ $0.673_{\pm 0.006}$ $0.776_{\pm 0.003}$ $0.767_{\pm 0.004}$ $0.737_{\pm 0.004}$
churn $0.873_{\pm 0.003}$ $0.780_{\pm 0.015}$ $0.761_{\pm 0.009}$ $0.802_{\pm 0.017}$ $0.857_{\pm 0.007}$ $0.865_{\pm 0.008}$ $0.854_{\pm 0.011}$ $0.852_{\pm 0.006}$ $0.845_{\pm 0.011}$
covertype $0.817_{\pm 0.001}$ $0.783_{\pm 0.001}$ $0.442_{\pm 0.008}$ $0.711_{\pm 0.002}$ $0.799_{\pm 0.001}$ $0.767_{\pm 0.001}$ $0.734_{\pm 0.002}$ $0.754_{\pm 0.001}$ $0.722_{\pm 0.002}$
default $0.674_{\pm 0.001}$ $0.627_{\pm 0.003}$ $0.686_{\pm 0.002}$ $0.632_{\pm 0.007}$ $0.678_{\pm 0.002}$ $0.638_{\pm 0.008}$ $0.670_{\pm 0.002}$ $0.671_{\pm 0.001}$ $0.673_{\pm 0.003}$
diabetes $0.621_{\pm 0.002}$ $0.572_{\pm 0.005}$ $0.557_{\pm 0.004}$ $0.553_{\pm 0.003}$ $0.617_{\pm 0.002}$ $0.617_{\pm 0.002}$ $0.611_{\pm 0.002}$
Table 13: Machine learning efficiency AUC score for five benchmark models, the real training data and for CDTD with three different noise schedules. The standard deviation takes into account five different sampling seeds and uses the average results of the four machine learning efficiency models computed across ten model seeds.
Real Data ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
adult $0.915_{\pm 0.000}$ $0.901_{\pm 0.000}$ $0.836_{\pm 0.006}$ $0.889_{\pm 0.002}$ $0.908_{\pm 0.000}$ $0.880_{\pm 0.005}$ $0.910_{\pm 0.000}$ $0.910_{\pm 0.001}$ $0.909_{\pm 0.000}$
bank $0.947_{\pm 0.000}$ $0.938_{\pm 0.001}$ $0.934_{\pm 0.003}$ $0.830_{\pm 0.020}$ $0.940_{\pm 0.005}$ $0.929_{\pm 0.005}$ $0.945_{\pm 0.000}$ $0.945_{\pm 0.001}$ $0.943_{\pm 0.004}$
churn $0.964_{\pm 0.001}$ $0.939_{\pm 0.007}$ $0.882_{\pm 0.006}$ $0.948_{\pm 0.004}$ $0.957_{\pm 0.004}$ $0.961_{\pm 0.001}$ $0.962_{\pm 0.001}$ $0.962_{\pm 0.001}$ $0.959_{\pm 0.003}$
covertype $0.892_{\pm 0.000}$ $0.860_{\pm 0.001}$ $0.677_{\pm 0.007}$ $0.777_{\pm 0.001}$ $0.876_{\pm 0.000}$ $0.845_{\pm 0.001}$ $0.816_{\pm 0.002}$ $0.828_{\pm 0.001}$ $0.802_{\pm 0.002}$
default $0.768_{\pm 0.000}$ $0.754_{\pm 0.002}$ $0.744_{\pm 0.002}$ $0.751_{\pm 0.004}$ $0.763_{\pm 0.002}$ $0.739_{\pm 0.008}$ $0.762_{\pm 0.003}$ $0.765_{\pm 0.002}$ $0.765_{\pm 0.002}$
diabetes $0.693_{\pm 0.001}$ $0.669_{\pm 0.002}$ $0.626_{\pm 0.003}$ $0.592_{\pm 0.002}$ $0.675_{\pm 0.001}$ $0.673_{\pm 0.001}$ $0.667_{\pm 0.001}$
Table 14: Machine learning efficiency RMSE for five benchmark models, the real training data and for CDTD with three different noise schedules. The standard deviation takes into account five different sampling seeds and uses the average results of the four machine learning efficiency models computed across ten model seeds.
Real Data ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
acsincome $0.804_{\pm 0.012}$ $0.757_{\pm 0.007}$ $2.292_{\pm 0.013}$ $1.054_{\pm 0.011}$ $0.857_{\pm 0.010}$ $0.838_{\pm 0.015}$ $0.811_{\pm 0.014}$ $0.820_{\pm 0.011}$
beijing $0.712_{\pm 0.001}$ $0.792_{\pm 0.007}$ $1.246_{\pm 0.010}$ $1.690_{\pm 0.016}$ $0.606_{\pm 0.006}$ $0.912_{\pm 0.005}$ $0.774_{\pm 0.005}$ $0.770_{\pm 0.005}$ $0.762_{\pm 0.005}$
lending $0.030_{\pm 0.000}$ $0.274_{\pm 0.007}$ $0.137_{\pm 0.007}$ $0.404_{\pm 0.007}$ $0.795_{\pm 0.031}$ $0.061_{\pm 0.001}$ $0.060_{\pm 0.001}$ $0.066_{\pm 0.002}$
news $1.001_{\pm 0.002}$ $0.923_{\pm 0.052}$ $1.906_{\pm 0.019}$ $3.999_{\pm 0.175}$ $0.083_{\pm 0.001}$ $1.302_{\pm 0.074}$ $0.819_{\pm 0.103}$ $0.776_{\pm 0.091}$ $0.755_{\pm 0.066}$
nmes $1.001_{\pm 0.003}$ $0.972_{\pm 0.024}$ $1.331_{\pm 0.052}$ $1.127_{\pm 0.047}$ $1.154_{\pm 0.047}$ $1.137_{\pm 0.052}$ $1.108_{\pm 0.083}$ $1.184_{\pm 0.076}$ $1.203_{\pm 0.081}$
Appendix O Ablation Study Details
Table 15: L2 norm (incl. standard errors in subscripts) of the correlation matrix differences of real and synthetic train sets for five CDTD configurations with progressive addition of model components.
Configuration A B C D CDTD (per type)
acsincome $0.131_{\pm 0.003}$ $0.119_{\pm 0.004}$ $0.124_{\pm 0.006}$ $0.129_{\pm 0.004}$ $0.129_{\pm 0.003}$
adult $0.131_{\pm 0.007}$ $0.128_{\pm 0.008}$ $0.168_{\pm 0.017}$ $0.107_{\pm 0.011}$ $0.125_{\pm 0.009}$
beijing $0.065_{\pm 0.009}$ $0.066_{\pm 0.012}$ $0.067_{\pm 0.011}$ $0.067_{\pm 0.010}$ $0.073_{\pm 0.009}$
churn $0.244_{\pm 0.015}$ $0.272_{\pm 0.034}$ $0.299_{\pm 0.066}$ $0.264_{\pm 0.012}$ $0.289_{\pm 0.043}$
Table 16: Jensen-Shannon divergence (incl. standard errors in subscripts) for five CDTD configurations with progressive addition of model components.
Configuration A B C D CDTD (per type)
acsincome $0.025_{\pm 0.000}$ $0.025_{\pm 0.001}$ $0.025_{\pm 0.001}$ $0.024_{\pm 0.001}$ $0.024_{\pm 0.000}$
adult $0.012_{\pm 0.001}$ $0.013_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.014_{\pm 0.001}$ $0.013_{\pm 0.001}$
beijing $0.004_{\pm 0.001}$ $0.006_{\pm 0.002}$ $0.005_{\pm 0.003}$ $0.004_{\pm 0.002}$ $0.005_{\pm 0.002}$
churn $0.010_{\pm 0.002}$ $0.008_{\pm 0.002}$ $0.009_{\pm 0.004}$ $0.010_{\pm 0.002}$ $0.012_{\pm 0.002}$
Table 17: Wasserstein distance (incl. standard errors in subscripts) for five CDTD configurations with progressive addition of model components.
Configuration A B C D CDTD (per type)
acsincome $0.002_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.001_{\pm 0.000}$ $0.001_{\pm 0.000}$
adult $0.004_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.004_{\pm 0.000}$
beijing $0.003_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.003_{\pm 0.000}$
churn $0.006_{\pm 0.001}$ $0.006_{\pm 0.000}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.007_{\pm 0.001}$
Table 18: Detection score (incl. standard errors in subscripts) for five CDTD configurations with progressive addition of model components.
Configuration A B C D CDTD (per type)
acsincome $0.534_{\pm 0.002}$ $0.534_{\pm 0.001}$ $0.538_{\pm 0.003}$ $0.532_{\pm 0.002}$ $0.532_{\pm 0.004}$
adult $0.597_{\pm 0.002}$ $0.593_{\pm 0.001}$ $0.615_{\pm 0.003}$ $0.580_{\pm 0.003}$ $0.588_{\pm 0.002}$
beijing $0.953_{\pm 0.002}$ $0.959_{\pm 0.001}$ $0.952_{\pm 0.003}$ $0.953_{\pm 0.001}$ $0.949_{\pm 0.001}$
churn $0.557_{\pm 0.014}$ $0.573_{\pm 0.014}$ $0.564_{\pm 0.012}$ $0.541_{\pm 0.015}$ $0.533_{\pm 0.007}$
Table 19: Distance to closest record of the generated data (incl. standard errors in subscripts) for five CDTD configurations with progressive addition of model components.
Real Test Set A B C D CDTD (per type)
acsincome $7.673_{\pm 0.017}$ $8.335_{\pm 0.064}$ $8.222_{\pm 0.035}$ $8.305_{\pm 0.021}$ $8.352_{\pm 0.025}$ $8.322_{\pm 0.047}$
adult $1.870_{\pm 0.000}$ $1.221_{\pm 0.018}$ $1.294_{\pm 0.015}$ $1.252_{\pm 0.014}$ $1.427_{\pm 0.008}$ $1.231_{\pm 0.011}$
beijing $0.385_{\pm 0.000}$ $0.545_{\pm 0.001}$ $0.559_{\pm 0.002}$ $0.539_{\pm 0.003}$ $0.541_{\pm 0.002}$ $0.489_{\pm 0.001}$
churn $0.347_{\pm 0.000}$ $0.307_{\pm 0.016}$ $0.326_{\pm 0.009}$ $0.294_{\pm 0.022}$ $0.298_{\pm 0.013}$ $0.274_{\pm 0.021}$
Table 20: Machine learning efficiency F1 score for five CDTD configurations with progressive addition of model components. The standard deviation accounts for five different sampling seeds and uses the average results of the four machine learning efficiency models across ten model seeds.
Real Data A B C D CDTD (per type)
adult $0.797_{\pm 0.000}$ $0.788_{\pm 0.001}$ $0.788_{\pm 0.001}$ $0.787_{\pm 0.001}$ $0.788_{\pm 0.002}$ $0.787_{\pm 0.001}$
churn $0.873_{\pm 0.003}$ $0.856_{\pm 0.008}$ $0.856_{\pm 0.014}$ $0.857_{\pm 0.007}$ $0.849_{\pm 0.006}$ $0.852_{\pm 0.006}$
Table 21: Machine learning efficiency AUC score for five CDTD configurations with progressive addition of model components. The standard deviation accounts for five different sampling seeds and uses the average results of the four machine learning efficiency models across ten model seeds.
Real Data A B C D CDTD (per type)
adult $0.915_{\pm 0.000}$ $0.909_{\pm 0.000}$ $0.910_{\pm 0.000}$ $0.909_{\pm 0.000}$ $0.910_{\pm 0.000}$ $0.910_{\pm 0.001}$
churn $0.964_{\pm 0.001}$ $0.962_{\pm 0.002}$ $0.961_{\pm 0.003}$ $0.960_{\pm 0.002}$ $0.961_{\pm 0.001}$ $0.962_{\pm 0.001}$
Table 22: Machine learning efficiency RMSE for five CDTD configurations with progressive addition of model components. The standard deviation accounts for five different sampling seeds and uses the average results of the four machine learning efficiency models across ten model seeds.
Real Data A B C D CDTD (per type)
acsincome $0.804_{\pm 0.012}$ $0.815_{\pm 0.009}$ $0.813_{\pm 0.018}$ $0.823_{\pm 0.017}$ $0.814_{\pm 0.014}$ $0.811_{\pm 0.014}$
beijing $0.712_{\pm 0.001}$ $0.782_{\pm 0.004}$ $0.785_{\pm 0.004}$ $0.778_{\pm 0.005}$ $0.776_{\pm 0.004}$ $0.770_{\pm 0.005}$
Appendix P Training and Sampling Times Details
Table 23: Training times in minutes. TabDDPM produces NaNs during training on acsincome and diabetes, and is therefore excluded for these data. CoDi is considered prohibitively expensive to train on diabetes and lending and we report estimated (est.) training times.
ARF CTGAN TVAE TabDDPM CoDi CDTD (per feature)
acsincome 80.3 59.9 26.0 - 231.9 5.8
adult 7.4 36.2 23.7 38.3 48.3 6.9
bank 11.0 37.6 24.6 40.5 42.7 26.3
beijing 3.7 34.3 23.9 36.1 24.9 23.4
churn 0.3 27.1 13.7 18.2 25.7 6.1
covertype 130.2 58.0 36.5 44.9 69.2 28.2
default 12.0 38.3 24.8 38.9 45.9 26.4
diabetes 58.5 90.1 25.3 - 870 (est.) 26.9
lending 5.2 157.9 36.6 48.7 3000 (est.) 25.3
news 23.0 48.8 33.3 37.2 41.5 25.2
nmes 0.4 32.8 17.2 24.9 30.2 6.3
Table 24: Sample times in seconds per 1000 samples.
ARF CTGAN TVAE TabDDPM CoDi CDTD (per feature)
acsincome 4.20 0.23 0.07 - 10.26 0.59
adult 1.78 0.31 0.16 0.82 3.65 0.56
bank 2.24 0.44 0.44 0.87 3.38 0.64
beijing 0.34 0.41 0.32 2.09 2.45 0.26
churn 1.00 0.40 0.24 0.95 2.78 0.39
covertype 9.74 0.28 0.25 2.45 4.35 1.97
default 2.07 0.27 0.25 0.86 3.48 0.60
diabetes 5.87 0.53 0.15 - - 1.33
lending 2.49 0.45 0.54 4.33 - 0.69
news 3.89 0.43 0.30 5.13 2.93 0.85
nmes 1.54 0.31 0.17 4.17 2.91 0.55

### v3

@@
 The framework has since been adapted to categorical data in various ways, including discrete diffusion processes (Austin et al., 2021; Hoogeboom et al., 2021), diffusion in continuous embedding space (Dieleman et al., 2022; Li et al., 2022; Regol & Coates, 2023; Strudel et al., 2022), and others (Campbell et al., 2022; Meng et al., 2022; Sun et al., 2023).
-Diffusion models which include both, continuous and categorical features alike, build directly on advances from the image domain (Kim et al., 2023; Kotelnikov et al., 2023; Lee et al., 2023; Jolicoeur-Martineau et al., 2024) and thus, are not designed to deal with challenges specific to mixed-type data: The different diffusion processes are not aligned, and do not scale to larger datasets and/or features with a greater number of categories.
+Diffusion models which include both, continuous and categorical features alike, build directly on advances from the image domain (Kim et al., 2023; Kotelnikov et al., 2023; Lee et al., 2023; Jolicoeur-Martineau et al., 2024) and thus, are not designed to deal with challenges specific to mixed-type data: The different diffusion processes and their losses are neither aligned nor balanced across data types, and do not scale to larger datasets and/or features with a greater number of categories.
+Models that naively combine different losses to integrate distinct generative processes may suffer from implicitly favoring the sample quality of some features or data types over others (Ma et al., 2020).
 A crucial component in score-based generative models is the noise schedule (Kingma et al., 2022; Chen et al., 2022; Chen, 2023; Jabri et al., 2022; Wu et al., 2023).
@@
 For instance, any two continuous features may be subject to different levels of discretization or different bounds, even after applying common data pre-processing techniques; and any two categorical features may differ in the number of categories, or the degree of imbalance.
-This high heterogeneity warrants a rethinking of fundamental parts of the diffusion framework, including the noise schedule and the effective combination of diffusion processes for different data types.
+The high heterogeneity and lack of balancing warrants a rethinking of fundamental parts of the diffusion framework, including the noise schedule and the effective combination of diffusion processes for different data types.
 In this paper, we introduce Continuous Diffusion for mixed-type Tabular Data (CDTD) to address the aforementioned shortcomings.
@@
 This way, the different noise processes become directly comparable, easier to balance, and enable the application of, for instance, classifier-free guidance (Ho & Salimans, 2022), accelerated sampling (Lu et al., 2022), and other advances, to mixed-type tabular data.
-We counteract the high feature heterogeneity inherent to data of mixed-type with distinct feature- or type-specific adaptive noise schedules.
-The learnable noise schedules allow the model to directly take feature or type heterogeneity into account during both training and generation, and thus avoid the reliance on image or text-specific noise schedules designs.
-This ensures a better allocation of the model’s capacity across features, feature types and timesteps, and guarantees high quality samples of tabular data.
-CDTD outperforms state-of-the-art baseline models across a diverse set of sample quality metrics, and computation time for data sets with an arbitrary number of categories and data points.
+We counteract the high feature heterogeneity inherent to data of mixed-type with distinct feature or type-specific adaptive noise schedules.
+The learnable noise schedules allow the model to directly take feature or type heterogeneity into account during both training and generation, and thus avoid the reliance on image or text-specific noise schedule designs.
+Moreover, we propose a diffusion-specific loss normalization and initialization scheme to homogenize different data types and their losses effectively.
+Our improvements ensure a better allocation of the model’s capacity across features, feature types and timesteps, and yield high quality samples of tabular data.
+CDTD outperforms state-of-the-art baseline models across a diverse set of sample quality metrics as well as computation time for data sets with an arbitrary number of categories and data points.
 Our experiments show that CDTD captures feature correlations exceptionally well, and that explicitly allowing for data-type heterogeneity in the noise schedules benefits sample quality.
 In sum, we make several contributions specific to diffusion probabilistic modeling of tabular data:
-  • We propose a joint continuous diffusion model for both continuous and categorical features such that all noise distributions can be Gaussian.
-  • We balance model capacity across continuous and categorical features with a novel and effective loss calibration.
-  • We derive a functional form to learn adaptive noise schedules and to allow for an efficient, exact evaluation and thus, greatly extend the idea of timewarping.
+  • We propose a joint continuous diffusion model for both continuous and categorical features such that all noise distributions are Gaussian.
+  • We balance model capacity across continuous and categorical features with a novel and effective loss calibration, an adjusted score model initialization and type or feature-specific noise schedules.
+  • We extend the idea of timewarping and propose a functional form to efficiently learn adaptive noise schedules, and to allow for exact evaluation and easy incorporation of prior information on the relative importance of noise levels.
   • We drastically improve the scalability of tabular data diffusion models to features with a high number of categories.
-  • We boost the quality of the generated samples with adaptive, feature- or type-specific noise schedules.
-  • Our novel combination of score estimation methods allows for the application of state-of-the-art techniques, such as classifier-free guidance and accelerated sampling, to mixed-type tabular data for the first time.
+  • We boost the quality of the generated samples with adaptive, feature or type-specific noise schedules.
+  • Our CDTD model allows the first-ever use of advanced techniques, like classifier-free guidance, for mixed-type tabular data directly in data space.
 Figure 1: CDTD framework.
@@
 Associated sampling processes are highlighted in orange.
-The approximated score functions are concatenated and passed to a blackbox ODE solver for sample generation.
+The approximated score functions are concatenated and passed to an ODE solver for sample generation.
 2 Score-based Generative Framework
@@
 2.1 Continuous Features
-We denote the $i$-th continuous feature as $x_{\text{cont}}^{(i)}\in{\mathbb{R}}$ and the stacked feature vector as ${\mathbf{x}}_{0}\equiv{\mathbf{x}}_{\text{cont}}\in{\mathbb{R}}^{K_{\text{cont}}}$.
+We denote $x_{\text{cont}}^{(i)}\in{\mathbb{R}}$ as the $i$-th continuous feature and ${\mathbf{x}}_{0}\equiv{\mathbf{x}}_{\text{cont}}\in{\mathbb{R}}^{K_{\text{cont}}}$ as the stacked feature vector.
 Further, let $\{{\mathbf{x}}_{t}\}_{t=0}^{t=1}$ be a diffusion process that gradually adds noise in continuous time $t\in[0,1]$ to ${\mathbf{x}}_{0}$, and let $p_{t}({\mathbf{x}})$ denote the density function of the data at time $t$.
@@
 Given ${\mathbf{x}}_{t}$ and $t$, $\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}[\nabla_{{\mathbf{x}}_{t}}\log p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})]$ minimizes Equation 3, and we obtain
-$\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}\nabla_{{\mathbf{x}}_{t}}\log p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})=\frac{1}{\sigma_{t}^{2}}[\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}[{\mathbf{x}}_{0}]-{\mathbf{x}}_{t}]\ .$ (4)
+$\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}\nabla_{{\mathbf{x}}_{t}}\log p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})=\frac{1}{\sigma^{2}(t)}[\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}[{\mathbf{x}}_{0}]-{\mathbf{x}}_{t}]\ .$ (4)
 We compute $\hat{{\mathbf{x}}}_{0}=\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}[{\mathbf{x}}_{0}]$ as an average over the $C_{j}$ possible embedding vectors weighted by their probability.
@@
 Since $p({\mathbf{x}}_{0}={\mathbf{e}}_{i}|{\mathbf{x}}_{t},t)=p(x_{\text{cat}}^{(j)}=i|{\mathbf{x}}_{t},t)$, an estimate of $p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)$ is obtained via a classifier that predicts $C_{j}$ class probabilities for the $j$-th feature and is trained via cross-entropy (CE) loss.
-Likewise, we start the generative process for a categorical feature from an embedding vector ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}})=\mathcal{N}({\bm{0}},\sigma^{2}(1)I_{d})$.
-We then use score interpolation to iteratively denoise the embedding vector.
-After the last timestep in the reverse process, that is, $t\approx 0$, we directly infer the generated class from the feature-specific class probabilities, $p({\mathbf{x}}_{0}={\mathbf{e}}_{i}|{\mathbf{x}}_{t},t)$.
+We start generating a categorical feature from an embedding vector ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}})=\mathcal{N}({\bm{0}},\sigma^{2}(1)I_{d})$.
+We then use score interpolation (Equation 4) to iteratively denoise this embedding vector.
+After the final reverse step, i.e., when  $t\approx 0$, we can directly infer the generated class from the feature-specific class probabilities, $p({\mathbf{x}}_{0}={\mathbf{e}}_{i}|{\mathbf{x}}_{t},t)$.
 This framework can easily be extended to multiple categorical features.
 All embeddings are trained alongside the model.
-The model fully takes feature-specific uncertainty at intermediate timesteps into account (for each feature it only has to commit to a category at the final step of generation) which improves the consistency of the generated samples (Dieleman et al., 2022).
+The model fully takes feature-specific uncertainty at intermediate timesteps into account (as for each feature it only has to commit to a category at the final step of generation) which improves the consistency of the generated samples (Dieleman et al., 2022).
 Therefore, CDTD allows to capture subtle dependencies both within and across data types more accurately.
@@
 Next, we discuss the important components of our method.
-In particular, the combination of the different losses for score interpolation and score matching, initialization and loss weighting concerns, and the adaptive type- or feature-specific noise schedule designs.
+In particular, the combination of the different losses for score matching and score interpolation, initialization and loss weighting concerns, and the adaptive type- or feature-specific noise schedule designs.
 3.1 General Framework
@@
 During generation, we concatenate the score estimates, $\hat{s}_{\text{cont}}^{(i)}$ and $\hat{s}_{\text{cat}}^{(j)}$, for all features $i$ and $j$, and pass them to an ODE solver together with $\sigma_{\text{cont},i}$ and $\sigma_{\text{cat},j}$, the noise levels retrieved by transforming linearly spaced timesteps with the learned adaptive noise schedules.
-Further implementation details are provided in Appendix J.
+Further details on the implementation and sampling are provided in Appendix J and Appendix K, respectively.
 3.2 Homogenization of Data Types
-Let $\mathcal{L}_{\text{MSE}}(x_{\text{cont}}^{(i)},t)$ denote the time-weighted MSE (i.e., score matching in Equation 3) loss of the $i$-th continuous feature at a single timestep $t$, and $\mathcal{L}_{\text{CE}}(x_{\text{cat}}^{(j)},t)$ the CE (i.e.
-score interpolation) loss of the $j$-th categorical feature.
+Let $\mathcal{L}_{\text{MSE}}(x_{\text{cont}}^{(i)},t)$ denote the time-weighted MSE (i.e., score matching) loss of the $i$-th continuous feature at a single timestep $t$, and $\mathcal{L}_{\text{CE}}(x_{\text{cat}}^{(j)},t)$ the CE (i.e., score interpolation) loss of the $j$-th categorical feature.
 Naturally, the two losses are defined on different scales.
-This leads to an unintended importance weighting of features in the generative process.
+This leads to an unintended importance weighting of features in the generative process (Ma et al., 2020).
 We assume that an unconditional model should a priori, i.e., without having any information, be indifferent between all features.
@@
 The initial equal importance across all timesteps will naturally change over the course of training.
-We employ an normalization scheme for the average diffusion loss (Karras et al., 2023; Kingma & Gao, 2023) to allow for changes in relative importance among features but ensure equal importance of all timesteps throughout training.
+We employ a normalization scheme for the average diffusion loss (Karras et al., 2023; Kingma & Gao, 2023) to allow for changes in relative importance among features but ensure equal importance of all timesteps throughout training.
 To do so, we learn the time-specific normalization term $Z(t)$ such that $\mathcal{L}(t)/Z(t)\approx 1$.
@@
 3.3 Noise Schedules
-Since the optimal noise schedule of one feature impacts the noise schedules of other features, and the different data types have different sensitivities to additive noise, we introduce feature-specific or type-specific noise schedules.
+Since the optimal noise schedule of one feature impacts the noise schedules of other features, and different data types have different sensitivities to additive noise, we introduce feature-specific or type-specific noise schedules.
 For instance, given the same embedding dimension, more noise is needed to remove the same amount of signal from embeddings of features with fewer classes.
@@
 Adaptive Noise Schedules.
-We aim to learn a noise schedule $h_{k}:t\mapsto\sigma$ for all $K=K_{\text{cont}}+K_{\text{cat}}$ features.
+Based on Dieleman et al.
+(2022), we aim to learn a noise schedule $h_{k}:t\mapsto\sigma$ for all $K=K_{\text{cont}}+K_{\text{cat}}$ features.
 Note that $t\in[0,1]$, and with pre-specified minimum and maximum noise levels, we can scale $\sigma_{k}$ to lie in $[0,1]$ as well, without loss of generality.
@@
 Type-specific noise schedules refer to learning two functions $F_{1}$ and $F_{2}$ that predict the respective average loss over all features of a data type.
-Examples of learned noise schedules are given in Appendix K.
+Examples of learned noise schedules are given in Appendix O.
 3.4 Additional Customization to Tabular Data
@@
 We use a diverse benchmark set of state-of-the-art generative models for mixed-type tabular data.
-This includes ARF (Watson et al., 2023), CTGAN (Xu et al., 2019), TVAE (Xu et al., 2019), TabDDPM (Kotelnikov et al., 2023), and CoDi (Lee et al., 2023).
+This includes SMOTE (Chawla et al., 2002), ARF (Watson et al., 2023), CTGAN (Xu et al., 2019), TVAE (Xu et al., 2019), TabDDPM (Kotelnikov et al., 2023), CoDi (Lee et al., 2023), TabSyn (Zhang et al., 2024).
 Each model follows a different design and/or modeling philosophy.
 Note that CoDi is an extension of STaSy (Kim et al., 2023, the same group of authors) that has shown to be superior in performance.
-For scaling reasons, the recently proposed Diffusion Forest (Jolicoeur-Martineau et al., 2024) is not an applicable benchmark.
+For scaling reasons, ForestDiffusion (Jolicoeur-Martineau et al., 2024) is not an applicable benchmark.
+[^1] Jolicoeur-Martineau et al.
+(2024) report in the appendix that they used 10-20 CPUs with 64-256 GB of memory for datasets with a median number of 540 observations.
+With the suggested hyperparameters (for improved efficiency) and 64 CPUs, the model took approx.
+500 min of training on the relatively small nmes data.
+Note that the model estimates $KT$ separate models, with $K$ being the number of features and $T$ the noise levels.
+Therefore, we consider ForestDiffusion to be prohibitively expensive for higher-dimensional data generation.
 Further details on the respective benchmark models and their implementations are provided in Appendix F and Appendix G.
-To keep the comparison fair, we use the same architecture for CDTD as TabDDPM, with minor changes to accommodate the different inputs (see Appendix J).
+We provide an in-depth comparison of CDTD to the diffusion-based baselines in Appendix N.
+To keep the comparison fair, we use the same architecture for CDTD as TabDDPM (the latter has also been adopted by TabSyn), with minor changes to accommodate the different inputs (see Appendix J).
 Datasets.
@@
 The datasets vary in size, prediction task (regression vs.
-binary classification), number of continuous and categorical features and their distributions.
+binary classification [^2] For ease of presentation, we only analyze binary targets.
+However, CDTD trivially extends to targets with multiple classes.), number of continuous and categorical features and their distributions.
 The number of categories for categorical features varies significantly across datasets (for more details, see Appendix E).
@@
 4.1 Evaluation Metrics
-We evaluate all generative models on four sample quality criteria after generating synthetic data $\mathcal{D}_{\text{gen}}$ of size $\min(\lvert\mathcal{D}_{\text{train}}\rvert,50\,000)$.
-All metrics are averaged over five random seeds that affect the generative process.
+In our experiments, we follow conventions from previous papers and use four sample quality criteria, which we assess using a comprehensive set of measures.
+All metrics are averaged over five random seeds that affect the generative process, which samples synthetic data $\mathcal{D}_{\text{gen}}$ of size $\min(\lvert\mathcal{D}_{\text{train}}\rvert,50\,000)$.
 Machine learning efficiency.
@@
 Hence, we train a group of models, consisting of a (logistic/ridge) regression, a random forest and a catboost model, on the data-specific prediction task (the corresponding hyperparameter settings are reported in Appendix I).
-We compare the model-averaged real test performance, $\text{Perf}(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})$, to the performance trained on the synthetic data, $\text{Perf}(\mathcal{D}_{\text{gen}},\mathcal{D}_{\text{test}})$.
+We compare the model-averaged real test performance, $\text{Perf}(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})$, to the performance when trained on the synthetic data, $\text{Perf}(\mathcal{D}_{\text{gen}},\mathcal{D}_{\text{test}})$.
 We subsample $\mathcal{D}_{\text{train}}$ in case of more than 50 000 observations to upper-bound the computational load.
@@
 Detection score.
-For each generative model, we report the accuracy of a catboost model that is trained to distinguish between real and generated (fake) samples (Borisov et al., 2023; Liu et al., 2023).
+For each generative model, we report the accuracy of a catboost model that is trained to distinguish between real and generated (fake) samples (Borisov et al., 2023; Liu et al., 2023; Zhang et al., 2024).
 First, we subsample the real data subsets, $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{valid}}$ and $\mathcal{D}_{\text{test}}$, to a maximum of 25 000 data samples to limit evaluation time.
@@
 Statistical similarity.
-We evaluate the statistical similarity between real and generated training data based on: (1) the Jensen-Shannon divergence (JSD; Lin, 1991) to quantify the difference in categorical distributions, (2) the Wasserstein distance (WD; Ramdas et al., 2017) to quantify the difference in continuous distributions, and (3) the L2 distance of the pair-wise correlation matrices.
-We use the Pearson correlation coefficient for the correlations between two continuous features, the Theil uncertainty coefficient for the correlations between two categorical features, and the correlation ratio across types (Kotelnikov et al., 2023; Zhao et al., 2021).
+We aim to assess the statistical similarity between real and generated data at both the feature and sample levels.
+We largely follow Zhao et al.
+(2021) and compare: (1) the Jensen-Shannon divergence (JSD; Lin, 1991) to quantify the difference in categorical distributions, (2) the Wasserstein distance (WD; Ramdas et al., 2017) to quantify the difference in continuous distributions, and (3) the L2 distance between pair-wise correlation matrices.
+We use the Pearson correlation coefficient for two continuous features, the Theil uncertainty coefficient for two categorical features, and the correlation ratio for mixed types.
+Similar metrics for the evaluation of statistical similarity have been used by Zhang et al.
+(2024).
 Distance to closest record.
@@
 For brevity, we report the absolute difference of the DCR of the synthetic data and the DCR of the real test set.
-A good DCR value, that indicates realistic and sufficiently private data, should be close to zero.
+A good DCR value, indicating both realistic and sufficiently private data, should be close to zero.
 4.2 Results
-Table 1 shows the ranks of all generative averaged across all datasets for all considered metrics.
+Table 1: Average performance rank of each generative model across eleven datasets.
+Per metric, bold indicates the best, underline the second best result.
+We assigned the rank 10 for CoDi on lending and diabetes, TabDDPM on acsincome and diabetes, SMOTE on acsincome and covertype.
+SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
+RMSE $\underline{3.400}_{\pm 3.382}$ $3.800_{\pm 2.482}$ $7.800_{\pm 1.470}$ $7.800_{\pm 2.227}$ $8.200_{\pm 1.470}$ $6.800_{\pm 1.939}$ $6.800_{\pm 1.720}$ $\mathbf{3.000}_{\pm 0.632}$ $\underline{3.400}_{\pm 2.059}$ $4.200_{\pm 2.227}$
+F1 $3.667_{\pm 3.145}$ $6.333_{\pm 2.285}$ $8.333_{\pm 1.491}$ $8.000_{\pm 1.000}$ $4.167_{\pm 2.794}$ $6.667_{\pm 3.091}$ $7.500_{\pm 1.607}$ $3.833_{\pm 1.462}$ $\mathbf{3.167}_{\pm 1.344}$ $\underline{3.500}_{\pm 1.979}$
+AUC $4.667_{\pm 2.749}$ $5.667_{\pm 2.055}$ $8.667_{\pm 1.106}$ $7.833_{\pm 1.067}$ $4.833_{\pm 2.794}$ $7.500_{\pm 2.872}$ $7.333_{\pm 1.374}$ $\underline{2.500}_{\pm 1.500}$ $\mathbf{2.333}_{\pm 0.943}$ $3.833_{\pm 1.675}$
+L2 dist.
+of corr.
+$4.818_{\pm 2.918}$ $5.636_{\pm 1.872}$ $8.091_{\pm 1.781}$ $7.909_{\pm 1.564}$ $7.000_{\pm 3.191}$ $6.909_{\pm 2.429}$ $6.818_{\pm 1.402}$ $\underline{2.727}_{\pm 1.286}$ $\mathbf{2.273}_{\pm 0.862}$ $3.000_{\pm 1.595}$
+Detection score $3.909_{\pm 3.502}$ $6.182_{\pm 1.696}$ $8.818_{\pm 1.466}$ $7.273_{\pm 1.213}$ $5.000_{\pm 3.045}$ $8.091_{\pm 2.391}$ $6.000_{\pm 1.595}$ $3.909_{\pm 1.164}$ $\mathbf{2.455}_{\pm 1.827}$ $\underline{3.545}_{\pm 1.725}$
+JSD $7.182_{\pm 2.167}$ $\mathbf{1.273}_{\pm 0.617}$ $8.182_{\pm 1.641}$ $8.818_{\pm 1.029}$ $6.909_{\pm 2.314}$ $7.000_{\pm 1.651}$ $6.545_{\pm 1.305}$ $\underline{2.455}_{\pm 1.076}$ $3.091_{\pm 0.793}$ $3.727_{\pm 1.052}$
+WD $\underline{3.091}_{\pm 3.315}$ $5.636_{\pm 1.611}$ $7.545_{\pm 1.827}$ $8.000_{\pm 1.477}$ $6.455_{\pm 3.144}$ $8.364_{\pm 1.823}$ $5.727_{\pm 2.339}$ $4.182_{\pm 1.466}$ $3.182_{\pm 1.192}$ $\mathbf{3.000}_{\pm 1.954}$
+DCR $6.000_{\pm 2.558}$ $6.182_{\pm 2.328}$ $8.455_{\pm 1.725}$ $6.182_{\pm 3.186}$ $4.455_{\pm 3.726}$ $6.545_{\pm 2.426}$ $5.909_{\pm 1.676}$ $4.091_{\pm 2.678}$ $\underline{3.818}_{\pm 2.124}$ $\mathbf{3.545}_{\pm 1.924}$
+Table 1 shows the average rank of each generative model across all datasets for the considered metrics.
 The ranks in terms of the F1 and AUC scores are averaged over the classification task datasets.
-Likewise, the RMSE rank averages include the regression task datasets only.
-We assign the maximum possible rank when a model was not trainable on a given dataset.
+Likewise, the RMSE rank averages include the regression task datasets.
+We assign the maximum possible rank when a model could not be trained on a given dataset or could not be evaluated in reasonable time.
 This includes TabDDPM, which outputs NaNs for acsincome and diabetes and CoDi, which we consider to be prohibitively expensive to train on diabetes (estimated 14.5 hours) and lending (estimated 60 hours).
-The dataset-specific results (including standard errors) and average metrics over all datasets are detailed in Appendix N.
-We provide visualizations of the captured correlations in the synthetic sample compared to the real training set in Appendix M and distribution plots for a qualitative comparison in Appendix L.
+Similarly, SMOTE is very inefficient in sampling for large datasets (78 min for 1000 samples on acsincome and 182 min on covertype) and does not finish the evaluation within 12 hours.
+The dataset-specific results (including standard errors) and average metrics over all datasets are detailed in Appendix R.
+We provide visualizations of the captured correlations in the synthetic sample compared to the real training set in Appendix Q and distribution plots for a qualitative comparison in Appendix P.
 Sample quality.
 CDTD consistently outperforms the considered benchmark models in most sample quality metrics.
-We see a major performance edge in terms of the detection score, the L2 distance of the correlation matrices and the regression-based metrics.
-Only for the Jensen-Shannon divergence, CDTD is dominated by ARF, a tree-based method that is expected to model categorical features particularly well.
-Interestingly, CDTD performs similar to TabDDPM on F1 scores but outperforms it dramatically for regression tasks.
-TabDDPM appears to favor modeling categorical features accurately, thereby sacrificing continuous features, as visualized in Appendix M.
-By utilizing score interpolation, CDTD is able to model intricate correlation structure more accurate than other (multinomial diffusion-based) frameworks.
-Most importantly, type-specific noise schedules always outperform the feature-specific noise schedule variant and often also the single noise schedule version.
+Specifically, we see a major performance edge in terms of the detection score, the L2 distance of the correlation matrices and the regression-based metrics.
+Only for the Jensen-Shannon divergence ARF, a tree-based method that is expected to model categorical features particularly well, outperforms CDTD.
+Interestingly, CDTD performs similar to TabDDPM on F1 scores, but outperforms it dramatically for regression tasks.
+TabDDPM appears to favor modeling categorical features accurately, thereby sacrificing continuous features, as visualized in Appendix Q.
+TabSyn, a latent-space diffusion model, performs worse than CDTD and often TabDDPM, which define diffusion in data space.
+In Appendix M, we further compare CDTD and TabSyn and investigate the benefits of defining a diffusion model in data space.
+By utilizing score interpolation, CDTD is able to model intricate correlation structure more accurately than other frameworks.
+Most importantly, type-specific noise schedules mostly outperform the feature-specific and single noise schedule variants.
 This illustrates the importance of accounting for the high heterogeneity in tabular data on the feature type level.
-Imposing different noise schedules per feature, however, appears to force too many constrains on the model and hence decreases sample quality.
-Per-feature noise schedules would requires more training steps to converge, as can be seen in Appendix K.
-Table 1: Average performance rank of each generative model across eleven datasets.
-Per performance metric, bold indicates the best, underline the second best result.
-CoDi is prohibitively expensive to train on lending and diabetes, TabDDPM produces NaNs for acsincome and diabetes, and therefore get assigned rank 8 for these datasets.
-ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
-RMSE $3.000_{\pm 1.673}$ $6.400_{\pm 1.356}$ $6.200_{\pm 1.833}$ $6.400_{\pm 1.200}$ $5.600_{\pm 1.356}$ $\mathbf{2.400}_{\pm 0.490}$ $\underline{2.600}_{\pm 1.855}$ $3.400_{\pm 2.059}$
-F1 $5.167_{\pm 1.951}$ $6.667_{\pm 1.374}$ $6.667_{\pm 0.745}$ $\underline{3.333}_{\pm 2.285}$ $5.333_{\pm 2.494}$ $\underline{3.333}_{\pm 1.247}$ $\mathbf{2.667}_{\pm 1.106}$ $3.000_{\pm 1.915}$
-AUC $4.500_{\pm 1.607}$ $7.000_{\pm 1.155}$ $6.500_{\pm 0.764}$ $4.167_{\pm 2.115}$ $6.000_{\pm 2.160}$ $\underline{2.333}_{\pm 1.599}$ $\mathbf{2.167}_{\pm 0.898}$ $3.500_{\pm 1.500}$
-L2 dist. of corr. $4.455_{\pm 1.305}$ $6.727_{\pm 1.213}$ $6.455_{\pm 1.157}$ $5.636_{\pm 2.346}$ $5.545_{\pm 1.876}$ $\underline{2.545}_{\pm 1.305}$ $\mathbf{2.091}_{\pm 0.793}$ $2.636_{\pm 1.553}$
-Detection score $4.818_{\pm 1.336}$ $7.182_{\pm 1.113}$ $5.636_{\pm 0.881}$ $4.000_{\pm 2.412}$ $6.455_{\pm 1.777}$ $3.273_{\pm 0.862}$ $\mathbf{1.818}_{\pm 1.696}$ $\underline{2.909}_{\pm 1.505}$
-JSD $\mathbf{1.273}_{\pm 0.617}$ $6.818_{\pm 1.113}$ $7.182_{\pm 0.716}$ $5.818_{\pm 1.641}$ $5.909_{\pm 1.083}$ $\underline{2.364}_{\pm 0.979}$ $3.000_{\pm 0.603}$ $3.727_{\pm 1.052}$
-WD $4.273_{\pm 1.483}$ $6.000_{\pm 1.758}$ $6.455_{\pm 1.157}$ $5.000_{\pm 2.449}$ $6.727_{\pm 1.286}$ $3.000_{\pm 1.348}$ $\mathbf{2.273}_{\pm 1.052}$ $\underline{2.364}_{\pm 1.494}$
-DCR $4.909_{\pm 1.564}$ $6.818_{\pm 1.466}$ $5.182_{\pm 2.443}$ $3.636_{\pm 2.837}$ $5.364_{\pm 1.920}$ $3.727_{\pm 2.049}$ $\underline{3.364}_{\pm 1.611}$ $\mathbf{3.091}_{\pm 1.443}$
-Figure 3: Average wall-clock time for training and for sampling 1000 samples (excluding acsincome, diabetes and lending).
-Training and sampling time.
-Figure 3 shows the average wall-clock time over all datasets (for which training was feasible for all models) for training as well as the time for sampling 1000 data points of each baseline model and the most complex CDTD variant.
-Details are given in Appendix P.
-The use of embeddings (instead of one-hot encoding) for categorical features drastically reduces the training time.
-CDTD trains the fastest on average and is able to reach a maximum training time of less than 30 min across all datasets.
-Thus, CDTD exhibits improved scaling to an increase in the number of categories.
-The ODE definition of the diffusion process also implies competitive sampling speeds compared to TVAE and CTGAN and in particular, to the diffusion-based benchmarks CoDi and TabDDPM.
-In light of the vastly improved scalability, the CDTD results in Table 1 could in practice trivially be improved further by simply scaling up the architecture.
+The different noise schedules per feature, however, appear to force too many constraints on the model and thus, decrease sample quality.
+Per-feature noise schedules would require more training steps to converge, as can be seen in Appendix O.
 Table 2: Ablation study for five CDTD configurations with progressive addition of model components. We report the median performance metrics over acsincome, adult, beijing and churn.
@@
 DCR (abs. diff. to test; $\downarrow$) 0.405 0.361 0.386 0.299 0.372
+Figure 3: Average training and sampling wall-clock time for 1000 samples (excl. acsincome, diabetes, lending).
+Training and sampling time.
+Figure 3 shows the average wall-clock time over all (for all models feasible) datasets for training as well as the time for sampling 1000 data points for each baseline model and the per feature CDTD variant (see Appendix T for details).
+We exclude SMOTE due to its considerably longer sampling with an average of 1377 seconds for 1000 samples.
+CDTD’s use of embeddings (instead of one-hot encoding) for categorical features drastically reduces training times and thus, improves scaling to increasing number of categories.
+The ODE formulation of the diffusion process implies competitive sampling speeds, in particular compared to the diffusion-based benchmarks CoDi, TabDDPM and TabSyn.
+Despite TabSyn utilizing a separately trained encoder, this does not result in a lower-dimensional latent space and therefore, does not speed up sampling.
 Ablation study.
 We conduct an ablation study over four datasets to investigate the separate components of our CDTD framework.
-The results are given in Table 2 (detailed results are in Appendix O).
+The results are given in Table 2 (detailed results are in Appendix S).
 The baseline model Config.
@@
 Main differences are in the RMSE and detection score as well as training efficiency (the loss calibration and improved initialization facilitate model convergence).
-The final model especially works well on the larger datasets compared to the baseline (see Appendix O), as smaller datasets are relatively easy to fit with 3 million parameters, even without any model improvements.
+The final model especially works well on the larger datasets compared to the baseline (see Appendix S), as smaller datasets are relatively easy to fit with 3 million parameters, even without any model improvements.
+We investigate the sensitivity of CDTD to important hyperparameters in Appendix L.
 5 Conclusion and Discussion
 We propose a Continuous Diffusion model for mixed-type Tabular Data (CDTD) that combines score matching and score interpolation and imposes Gaussian diffusion processes on both continuous and embedded categorical features.
-We compared to various benchmark models and to a single noise schedule as typically used in image diffusion models.
-Our results indicate that acknowledging the high feature heterogeneity in tabular data on the feature type level and aligning type-specific diffusion elements, such as the noise schedules or losses, substantially benefits sample quality.
-CDTD also shows vastly improved scalability and can accommodate an arbitrary number of categories.
-Our paper serves as an important step to adopting the diffusion probabilistic framework customized to tabular data.
-In particular, the common type of noise schedules allows for a framework that is easy to extend and may accelerate progress on diffusion models for tabular data.
-Crucially, CDTD allows the direct application of diffusion-related advances from the image domain, like classifier-free guidance, to tabular data.
-We leave further extensions, e.g., the exploration of accelerated sampling, efficient score model architectures, or different forms of adaptive noise schedules, to future work.
-Finally, we want to highlight that synthetic data could be misused to support unwarranted claims.
-In any case, any generated data should not be blindly trusted and synthetic data based inferences should be compared to results derived from the real data.
-On the other hand, the correct usage of generative models can allow for better privacy preservation, facilitate data sharing and open science practices.
+We compared CDTD to various benchmark models and to a single noise schedule as typically used in image diffusion models.
+Our results indicate that addressing the high feature heterogeneity in tabular data on the feature type level and aligning type-specific diffusion elements, such as the noise schedules or losses, substantially benefits sample quality.
+Moreover, CDTD shows vastly improved scalability and can accommodate an arbitrary number of categories.
+Our paper serves as an important step to customizing the diffusion probabilistic framework to tabular data.
+In particular, the common type of noise schedules allows for an easy to extend framework that might accelerate progress on diffusion models for tabular data.
+Crucially, CDTD allows the direct application of diffusion-related advances from the image domain, like classifier-free guidance, to tabular data without the need for a latent encoding.
+We leave further extensions to the tabular data domain, e.g., the exploration of accelerated sampling, efficient score model architectures, different forms of adaptive noise schedules, or the adaption to the data imputation task for future work.
+Finally, we want to emphasize the potential misuse of synthetic data to support unwarranted claims.
+Any generated data should therefore not be blindly trusted, and synthetic data based inferences should always be compared to results from the real data.
+However, the correct use of generative models enables better privacy preservation and facilitates data sharing and open science practices.
 Limitations
-The main limitation of the CDTD model is the addition of another hyperparameter, i.e., the type of noise schedule, and tuning a hyperparameter of a generative model can be a costly endeavor.
-However, our results also show that a per type schedule is most often optimal.
+The main limitation of CDTD is the addition of hyperparameters, and tuning hyperparameters of a generative model can be a costly endeavor.
+However, our results also show that (1) a per type schedule is most often optimal and (2) our default hyperparameters perform well across a diverse set of datasets.
 Dieleman et al.
-(2022) show that the results of score interpolation can be sensitive to the initialization of the embeddings.
-We have not encountered any problems with the same standard deviation of 0.001 for all datasets.
-However, this may not apply in general.
+(2022) show that the results of score interpolation for text data can be sensitive to the initialization of the embeddings.
+We have not encountered similar problems on tabular datasets (see Table 7).
 While the DCR indicates no privacy issues for the benchmark datasets used, additional caution must be taken when generating synthetic data from privacy sensitive sources.
-Lastly, our model may be outperformed by other generative models specialized for different types of tabular data, such as time-series.
+Lastly, for specific types of tabular data, such as time-series, our model may be outperformed by other generative models specialized for that type.
+While CDTD could be directly used for imputation using RePaint (Lugmayr et al., 2022), a separate training process is required to achieve the best results (Liu et al., 2024).
+Therefore, we leave the adaption of CDTD to the imputation task for future work.
 Acknowledgements
@@
 Appendix A Loss Calibration
-A priori, we let the model be indifferent between features.
-Therefore, we aim to scale the loss of each feature such that they attain the same loss at the terminal timestep $t=1$.
-In that case, the signal-to-noise ratio is sufficiently low to approximate a situation in which the model has no information about the data.
-We are looking for calibrated losses $\mathcal{L}_{\text{MSE}}^{\ast}(x_{\text{cont}}^{(i)},1)$ and $\mathcal{L}_{\text{CE}}^{\ast}(x_{\text{cat}}^{(j)},1)$ which in expectation achieve a unit loss at $t=1$.
-This way, the model treats all features equally before having seen any information.
-For a single scalar feature and given timestep $t$, we can write the empirical denoising score matching loss (see Equation 3) in the EDM parameterization (Karras et al., 2022) as:
+A priori, we let the model be indifferent between features, that is, we scale the loss of each feature such that at the terminal timestep the same loss is attained.
+Here, the signal-to-noise ratio is sufficiently low to approximate a situation in which the model has no information about the data.
+Thus, we are looking for calibrated losses $\mathcal{L}_{\text{MSE}}^{\ast}(x_{\text{cont}}^{(i)},1)$ and $\mathcal{L}_{\text{CE}}^{\ast}(x_{\text{cat}}^{(j)},1)$ which at $t=1$ achieve unit loss in expectation.
+For a single scalar feature and a given timestep $t$, we can write the empirical denoising score matching loss (Equation 3) in the EDM parameterization (Karras et al., 2022) as:
 $\mathcal{L}_{\text{MSE}}(x_{\text{cont}}^{(i)},t)=\lambda(t)\Bigl{(}\underbrace{c_{\text{skip}}(t)x_{t}+c_{\text{out}}(t)F_{{\bm{\theta}}}^{(i)}}_{s_{\bm{\theta}}(x_{t},t)}-x_{\text{cont}}^{(i)}\Bigr{)}^{2},\$
@@
 For $t\rightarrow 0$, the model shifts increasingly towards predicting the error that has been added to the true data.
-In the EDM parameterization, the explicit timestep weight (used to achieve a unit loss across timesteps at initialization (see Appendix B) is $\lambda(t)=1/c_{\text{out}}(t)^{2}\approx 1$ for $t=1$.
+In the EDM parameterization, the explicit timestep weight (used to achieve a unit loss across timesteps at initialization, see Appendix B) is $\lambda(t)=1/c_{\text{out}}(t)^{2}\approx 1$ for $t=1$.
 At the terminal timestep $t=1$, we now have:
@@
 We initialize the neural network such that $F_{{\bm{\theta}},c}^{(j)}=\log p_{c}$ for all $c$.
-This can be achieved by initializing the output layer weights to zero and the output layer biases to the relevant training set log-proportions of the corresponding class.
+This is achieved by initializing the output layer weights to zero and the output layer biases to the relevant training set log-proportions of the corresponding class.
 Hence, this initialization gives us
@@
 During training, the adaptive noise schedules allow the model to focus automatically on the noise levels that matter most, i.e., where the loss increase is steepest.
-However, the better the model becomes at a given timestep $t$, the lower the loss at that timestep, and the lower the gradient signal relative to the signal received for timesteps $\tilde{t}>t$.
-To counteract this, we additionally use an adaptive normalization of the average diffusion loss (averaged over the features) across timesteps.
+However, the better the model becomes at a given timestep $t$, the lower the loss at the respective timestep, and the lower the gradient signal relative to the signal for timesteps $\tilde{t}>t$.
+We counteract this with adaptive normalization of the average diffusion loss (averaged over the features) across timesteps.
 Specifically, we want to weight the average diffusion loss at timestep $t$, $\mathcal{L}(t)$ given in Equation 6, such that the normalized loss is the same (equal to one) for all $t$.
@@
 (2023) and Kingma & Gao (2023), we follow the latter in the setup of the corresponding network.
-We train a neural network alongside our diffusion model to predict $\mathcal{L}(t)$ based on $t$ and using the MSE loss to learn this weighting.
-In particular, we first compute $c_{\text{noise}}(t)=\log(t)/4$ following the EDM parameterization (Karras et al., 2022).
+We train a neural network alongside our diffusion model to predict $\mathcal{L}(t)$ based on $t$ and use the MSE loss to learn this weighting.
+First, we compute $c_{\text{noise}}(t)=\log(t)/4$ following the EDM parameterization (Karras et al., 2022).
 Then, we embed $c_{\text{noise}}$ in frequency space (1024-dimensional) using Fourier features.
-The result is then passed through a single linear layer to output a scalar value, which is in turn passed through an exponential function to ensure that the prediction ^$\hat{\mathcal{L}}(t)\geq 0$.
+The result is passed through a single linear layer to output a scalar value, passed through an exponential function to ensure that the prediction ^$\hat{\mathcal{L}}(t)\geq 0$.
 We initialize the weights and biases to zero, to ensure that at model initialization we have a unit normalization.
 Appendix D Derivation of the Functional Timewarping Form
-Since higher noise levels, $\sigma$, imply a lower signal-to-noise ratio and, in turn, a larger loss, $\ell$, we know that the loss must be a monotonically increasing and S-shaped function of the noise level.
+Since higher noise levels, $\sigma$, imply a lower signal-to-noise ratio, and in turn a larger loss, $\ell$, we know that the loss must be a monotonically increasing and S-shaped function of the noise level.
 Additionally, the function has to be easy to invert and differentiate.
@@
 where we defined $Z(\sigma,\mu,\nu)=\bigl{(}\frac{\sigma}{1-\sigma}\frac{1-\mu}{\mu}\bigr{)}^{-\nu}$ and used the definitions of $f_{\log}$ and the parameter $\mu^{\ast}$.
-Appendix E Datasets
-Our selected benchmark datasets are highly diverse, particularly in the number of categories for a categorical feature.
-Almost all datasets are licensed under creative commons that permits their usage in our research.
-Only the nmes dataset license is unknown but the data is publicly accessible on the author’s website.
-For the diabetes and covertype datasets, we transform the original multi-class classification problem into a binary classification task for easier comparison.
+Appendix E Benchmark Datasets
+Our selected benchmark datasets are highly diverse, particularly in the number of categories for categorical features (see Table 3).
+For the diabetes and covertype datasets, we transform the original multi-class classification problem into a binary classification task for ease of presentation.
 For the covertype data, the task is converted into predicting whether a forest of type 2 is present in a given 30 $\times$ 30 meter area.
 In the diabetes data, we convert the task by predicting whether a patient was readmitted to a hospital.
+All datasets are publicly accessible and (except nmes) licensed under creative commons.
 Table 3: Overview of the selected experimental datasets.
@@
 Appendix F Baseline Models
-Below, we give a brief description of the used generative baseline models (including code sources).
+Below, we give a brief description of our selected generative baseline models (including code sources).
+SMOTE
+(Chawla et al., 2002) – a technique (not a generative model) typically used to oversample minority classes based on interpolation between ground-truth observations.
+We use SMOTENC for mixed-type data from the scikit-learn package and mostly adapt the code from the TabDDPM repository (Kotelnikov et al., 2023).
+For sampling, we utilize 16 CPU cores.
 ARF
@@
 CoDi
-(Lee et al., 2023) – a diffusion model trained using an additional contrastive loss, and which factorizes the joint distribution of mixed-type tabular data into a distribution for continuous data conditional on categorical features and a distribution for categorical data conditional on continuous features.
-Similar to TabDDPM, the authors utilize the multinomial diffusion framework (Hoogeboom et al., 2021) to model categorical data.
+(Lee et al., 2023) – a diffusion model trained with an additional contrastive loss, and which factorizes the joint distribution of mixed-type tabular data into a distribution for continuous data conditional on categorical features and a distribution for categorical data conditional on continuous features.
+Similarly, the authors utilize the multinomial diffusion framework (Hoogeboom et al., 2021) to model categorical data.
 An implementation is available at https://github.com/ChaejeongLee/CoDi under an unknown license.
+TabSyn
+(Zhang et al., 2024) – a diffusion-based model that first learns a tranformer-based VAE to map mixed-type data to a continuous latent space.
+Then, the diffusion model is trained on that latent space.
+We use the official code available at https://github.com/amazon-science/tabsyn under the Apache 2.0 license.
 Appendix G Implementation Details
@@
 The same inability holds for the comparison of CDTD to other diffusion-based models: Our model is the first to use a continuous noise distribution on both continuous and categorical features, and therefore the alignment of important design choices, like the noise schedule, across models is not possible.
-In particular, the Markov transition matrices based forward process of the multinomial diffusion framework (Hoogeboom et al., 2021) in TabDDPM and CoDi does not translate to our setting.
-To ensure a fair comparison in terms of sampling steps, we set the steps for CDTD, TabDDPM and CoDi to $\max(200,\text{default})$.
-We therefore increase the default number of sampling steps for CoDi (from 50 steps) and TabDDPM (from 100 steps for classification datasets).
+In particular, the forward process of the multinomial diffusion framework (Hoogeboom et al., 2021) used in TabDDPM and CoDi, which is based on Markov transition matrices, does not translate to our setting.
+To ensure a fair comparison in terms of sampling steps, we set the steps for CDTD, TabDDPM, CoDi and TabSyn to $\max(200,\text{default})$.
+We therefore increase the default number of sampling steps for CoDi and TabSyn (from 50 steps) and TabDDPM (from 100 steps for classification datasets).
 For TabDDPM and regression datasets, we use the suggested default of 1000 sampling steps.
@@
 CoDi 2 998 043
+TabSyn 3 001 646
 CDTD (per type schedule, TabDDPM architecture) 2 999 721
@@
 Note that the total number of parameters may vary slightly across datasets due to different number of features and categories affecting the onehot encoding but is still comparable across models.
-We also align the embedding/bottleneck dimensions for CTGAN, TVAE, TabDDPM and CDTD to 256 to accommodate more complex datasets.
+We also align the embedding/bottleneck dimensions for CTGAN, TVAE, TabDDPM, TabSyn and CDTD to 256.
+To align TabDDPM, TabSyn and CDTD further, we use the TabDDPM architecture for all models, with appropriate adjustments for different input types and dimensions.
 If applicable, all models are trained for 30k steps on a single RTX 4090 instance, using PyTorch version 2.2.2.
 Below, we briefly discuss our model-specific hyperparameter choices.
+SMOTE
+(Chawla et al., 2002): We use the default hyperparameters suggested for the SMOTENC scikit-learn implementation.
 ARF
@@
 Therefore, we adjust the batch size if necessary.
-We use a 258-dimensional embedding (instead of the default embedding dimension of 128) to better align the CTGAN architecture with TVAE, TabDDPM and CDCD.
+We use a 256-dimensional embedding (instead of the default embedding dimension of 128) to better align the CTGAN architecture with TVAE, TabDDPM, TabSyn and CDTD.
 TVAE
-(Xu et al., 2019): We again follow the implementation in the Synthetic Data Vault. We use a 258-dimensional embedding to better align the architecture with CTGAN, TabDDPM and CDCD.
+(Xu et al., 2019): We again follow the implementation in the Synthetic Data Vault. We use a 256-dimensional embedding to better align the architecture with CTGAN, TabDDPM, TabSyn and CDTD.
 TabDDPM
@@
 Note also that for classification task datasets, TabDDPM models the conditional distribution $p(x|y)$, instead of the unconditional distribution $p(x)$ which is modeled for regression tasks.
-We adjust the dimension of the bottleneck to 256 (instead of the default 128) to also accommodate also larger datasets and align the model with CTGAN, TVAE,and CDCD.
+We adjust the dimension of the bottleneck to 256 (instead of the default 128) to also accommodate also larger datasets and align the model with CTGAN, TVAE,and CDTD.
 CoDi
 (Lee et al., 2023): We use the default hyperparameters from the official code (see https://github.com/ChaejeongLee/CoDi).
+TabSyn
+(Zhang et al., 2024): We use the default hyperparameters as suggested by the authors.
+The training steps that go towards training the VAE and the denoising network follow the proportions given in the official code (see https://github.com/amazon-science/tabsyn).
+To improve comparability to TabDDPM, CoDi and CDTD, we use the same neural network architecture as TabDDPM, which only differs slightly from the original architecture.
+We leave the VAE untouched.
 CDTD
-(ours): To ensure comparability in particular to TabDDPM and CoDi, we use the same neural network architecture as TabDDPM.
+(ours): To ensure comparability in particular to TabDDPM, CoDi and TabSyn, we use the same neural network architecture as TabDDPM.
 We only change the input layers to accommodate our embedding-based framework.
@@
 We cap the maximum size of the real data subsets to 25 000, and subsample them if necessary, to limit the computational load.
-We then combine real and fake observations per set to $\mathcal{D}_{\text{train}}^{\text{detect}},\mathcal{D}_{\text{valid}}^{\text{detect}}$, and $\mathcal{D}_{\text{test}}^{\text{detect}}$, respectively.
+Per set, we combine real and fake observations to $\mathcal{D}_{\text{train}}^{\text{detect}},\mathcal{D}_{\text{valid}}^{\text{detect}}$, and $\mathcal{D}_{\text{test}}^{\text{detect}}$, respectively.
 The catboost model is trained on $\mathcal{D}_{\text{train}}^{\text{detect}}$ with the task of predicting whether an observation is real or fake.
@@
 CDTD only conditions on $y$, i.e., the target feature, for classification task datasets.
-To enable a fair comparison to the other methods and in particular TabDDPM, the CDTD score model utilizes the exact same architecture and optimizer as Kotelnikov et al.
-(2023).
-An overview of the score model is given in Figure 4: First, the noisy data, including the noisy scalars for continuous features and the noisy embeddings for categorical features, are projected onto a 256-dimensional space.
+To enable a fair comparison to the other methods, and to TabDDPM and TabSyn in particular, the CDTD score model utilizes the exact same architecture and optimizer as Kotelnikov et al.
+(2023), which was also adapted by TabSyn (Zhang et al., 2024).
+An overview of the score model is provided in Figure 4: First, the noisy data, i.e., the noisy scalars for continuous features and the noisy embeddings for categorical features, are projected onto a 256-dimensional space.
 Similarly, timestep $t$ and possibly conditioning information $y$ are embedded in the same space.
-All 256-dimensional vectors are then added and the results is processed by a set of five fully-connected linear layers with ReLU activation functions.
+Then, all 256-dimensional vectors are added and the results is processed by a set of five fully-connected linear layers with ReLU activation functions.
 Lastly, a linear projection maps the output of the fully-connected layers to the required output dimensions, which depend on the number of features and number of categories per feature.
@@
 Also, Dieleman et al.
-(2022) argue that the standard deviation of the Normal distribution used to initialize the embeddings, denoted by $\sigma_{\text{cat emb}}$, is an important hyperparameter.
-In this paper, we set $\sigma_{\text{cat emb}}=0.001$ for all datasets and have not seen detrimental effects.
-Tuning this hyperparameter in practice could improve the results further.
+(2022) argue that the standard deviation of the Normal distribution used to initialize the embeddings, denoted by $\sigma_{\text{init}}$, is an important hyperparameter.
+In this paper, we set $\sigma_{\text{init}}=0.001$ for all datasets and have not seen detrimental effects.
+Table 7 indicates that CDTD is not sensitive to the choice of $\sigma_{\text{init}}$.
 Since we utilize embeddings, we have to scale the neurons per layer slightly down in the stack of the five fully-connected layers (from 808 for TabDDPM to 806).
-Also, since TabDDPM samples discrete steps from $[0,T]$, with $T\gg 1$, we adjust our timesteps $t\in[0,1]$ up by 1000.
+Also, since TabDDPM samples discrete steps from $[0,T]$, with $T\gg 1$, we scale our timesteps $t\in[0,1]$ up by 1000.
 We use the same optimizer (Adam), learning rate (0.001), learning rate decay (linear), EMA decay (0.999), and training steps (30000).
 However, since we work with embeddings we add a linear warmup schedule over the first 100 steps.
-Instead of using the vanilla uniform (time)step sampling as the TabDDPM, the CDTD model uses antithetic sampling Dieleman et al.
-(2022); Kingma et al.
-(2022).
+Instead of using the vanilla uniform (time)step sampling as the TabDDPM, the CDTD model uses antithetic sampling (Dieleman et al., 2022; Kingma et al., 2022).
 The timesteps are still uniformly distributed but spread out more evenly over the domain, which benefits the training of the adaptive noise schedules.
 For generation, we use an Euler sampler with 200 steps to minimize the discretization error.
-Appendix K Examples of Learned Noise Schedules
-Below we show the learned noise schedules for the smallest dataset (churn) and the largest dataset (acsincome).
-We also illustrate how well single, per type and per feature schedules fit the respective losses for each of the datasets.
+Appendix K CDTD Sampling
+To sample from our learned distribution, we need to run the reverse process of the probability flow ODE (Equation 2).
+For example, for two different features $x_{1}$ and $x_{2}$, we deconstruct the ODE as:
+$d{\mathbf{x}}$ $=-\frac{1}{2}{\mathbf{G}}(t){\mathbf{G}}(t)^{\mathsf{T}}\nabla_{\mathbf{x}}\log p_{t}({\mathbf{x}})dt$
+$=-\begin{bmatrix}\dot{\sigma_{1}}(t)\sigma_{1}(t)&\\ &\dot{\sigma_{2}}(t)\sigma_{2}(t)\end{bmatrix}\begin{bmatrix}\frac{\hat{x_{1}}-x_{1}}{\sigma_{1}(t)^{2}}\\ \frac{\hat{x_{2}}-x_{2}}{\sigma_{2}(t)^{2}}\end{bmatrix}dt$
+$=-\begin{bmatrix}\dot{\sigma_{1}}(t)&\\ &\dot{\sigma_{2}}(t)\end{bmatrix}\begin{bmatrix}\frac{\hat{x_{1}}-x_{1}}{\sigma_{1}(t)}\\ \frac{\hat{x_{2}}-x_{2}}{\sigma_{2}(t)}\end{bmatrix}dt$
+In practice, we use an Euler sampler with 200 discrete timesteps $\Delta t=t_{i+1}-t_{i}<0$.
+The timesteps are generated as a linearly spaced grid on $[0,1]$ and transformed afterwards into noise levels $\sigma_{k}(t)$ via the described timewarping procedure.
+For the discretized and simplified the ODE above, this yields
+${\mathbf{x}}_{i+1}$ $={\mathbf{x}}_{i}-\begin{bmatrix}\frac{\Delta\sigma_{1}(t)}{\Delta t}&\\ &\frac{\Delta\sigma_{2}(t)}{\Delta t}\end{bmatrix}\begin{bmatrix}\frac{\hat{x_{1}}-x_{1}}{\sigma_{1}(t)}\\ \frac{\hat{x_{2}}-x_{2}}{\sigma_{2}(t)}\end{bmatrix}\Delta t={\mathbf{x}}_{i}+\begin{bmatrix}\frac{x_{1}-\hat{x_{1}}}{\sigma_{1}(t)}\\ \frac{x_{2}-\hat{x_{2}}}{\sigma_{2}(t)}\end{bmatrix}\odot\begin{bmatrix}\Delta\sigma_{1}(t)\\ \Delta\sigma_{2}(t)\end{bmatrix}.$
+where $\odot$ denotes the element-wise product.
+Hence, we are effectively taking feature-specific steps of length $\Delta\sigma_{k}(t)$.
+The adaptive noise schedules (timewarping) therefore not only affect the training process, but also focus most work in the reverse process on the noise levels that matter most for sample quality (i.e., where $\Delta\sigma_{i}(t)$ is small).
+We use finite differences to approximate $\dot{\sigma}_{i}$, instead of the available, analytical variant, since $\frac{d\sigma_{k}(t)}{dt}\rightarrow\infty$ as $t\rightarrow 1$.
+The step $\Delta t$ would therefore be required to decrease as $t\rightarrow 1$ to ensure $\Delta t\approx dt$ holds.
+For a large number of steps, this assumption does not hold in practice, and for $\frac{d\sigma_{k}(t)}{dt}$ the update of ${\mathbf{x}}$ overshoots the target drastically.
+Intuitively, $\sigma_{k}(t)$ becomes too steep near the terminal timestep $t=1$ such that the step size can not sufficiently compensate for the slope increase to turn $\frac{d\sigma_{k}(t)}{dt}$ into a good approximation of the actual change in $\sigma_{k}(t)$.
+Moreover, the analytical solution would approximate $d\sigma_{k}(t)=\dot{\sigma}_{k}(t)dt$, i.e., the change in the noise level caused by a change in $t$.
+Since we know exactly where $\sigma_{k}(t)$ will end up when changing $t$, we are better off using that exact value and let $d\sigma_{k}(t)=\Delta\sigma_{k}(t)$.
+Table 6 shows that the gains in sample quality are marginal to non-existent after more than 500 sampling steps.
+Table 6: Performance sensitivity of CDTD (per type) to increasing number of sampling steps.
+Each metric is averaged over five seeds.
+As a robust measure, we report the median over the ablation study datasets acsincome, adult, beijing and churn .
+Steps RMSE F1 AUC L2 distance of corr. Detection score JSD WD DCR
+200 (default) 0.033 0.015 0.004 0.127 0.560 0.013 0.003 0.372
+500 0.028 0.018 0.005 0.130 0.565 0.012 0.003 0.372
+1000 0.028 0.018 0.005 0.129 0.560 0.012 0.003 0.373
+1500 0.028 0.018 0.005 0.129 0.561 0.012 0.003 0.374
+Appendix L Sensitivity to Important Hyperparameters
+The training and sampling processes of CDTD are affected by various novel hyperparameters.
+Generally, a per-type noise schedule works best as we show in our main results in Table 1 for a diverse set of benchmark datasets.
+Here, we examine the sensitivity of CDTD to two additional important hyperparameters: (1) the standard deviation of the noise used to initialize the embeddings (and therefore specific to score interpolation), $\sigma_{\text{init}}$, and (2) the weight of the low noise levels used to initialize the $\mu_{k}$ in the adaptive noise schedule parameterization.
+The experiments in Dieleman et al.
+(2022) show that $\sigma_{\text{init}}$ is a crucial hyperparameter for score interpolation on text data.
+The same sensitivity does not translate to the tabular data domain, as shown in our results in Table 7.
+The much smaller embedding dimension (16 vs.
+256) and the feature-specific embeddings significantly decrease the number of distinguishable categories.
+Compared to a vocabulary size of 32000 for text data (Dieleman et al., 2022), we only face a maximum of 3151 categories in the lending dataset (see Table 3).
+Thus, unlike other generative (diffusion) models for tabular data, CDTD scales to a practically arbitrary number of categories.
+Our proposed functional form for the adaptive noise schedules (see Appendix D) is the first to allow for the incorporation of prior information about the importance of low vs.
+high (normalized) noise levels.
+For this, we adjust the weight of low noise levels which directly determines the location of the inflection point $\mu_{k}$ (see Section 3.3).
+The results in Table 8 indicate low sample quality sensitivity to weight changes for a per-type noise schedule.
+The initialization only impacts the time to convergence but not (much) the location of the optimum.
+In our experiments, the number of training steps (30000) appears to be high enough for all model variants to converge.
+Table 7: Performance sensitivity of CDTD (per type) to changes in the standard deviation $\sigma_{\text{init}}$ in the initialization of the embeddings of categorical features.
+Each metric is averaged over five seeds.
+As a robust measure, we report the median over the ablation study datasets acsincome, adult, beijing and churn .
+$\sigma_{\text{init}}$ RMSE F1 AUC L2 distance of corr. Detection score JSD WD DCR
+1 0.032 0.017 0.006 0.126 0.564 0.011 0.004 0.311
+0.1 0.035 0.016 0.004 0.128 0.570 0.012 0.004 0.358
+0.01 0.032 0.017 0.005 0.131 0.566 0.011 0.004 0.369
+0.001 (default) 0.033 0.015 0.004 0.127 0.560 0.013 0.003 0.372
+Table 8: Performance sensitivity of CDTD (per type) to changes in the prior weight of low noise levels in the initialization of the adaptive noise schedules.
+Each metric is averaged over five seeds.
+As a robust measure, we report the median over the ablation study datasets acsincome, adult, beijing and churn .
+Weight RMSE F1 AUC L2 distance of corr. Detection score JSD WD DCR
+1 0.036 0.015 0.004 0.143 0.651 0.015 0.003 0.313
+2 0.030 0.014 0.005 0.147 0.651 0.014 0.003 0.352
+3 (default) 0.033 0.015 0.004 0.154 0.651 0.013 0.004 0.366
+4 0.034 0.019 0.005 0.148 0.656 0.013 0.004 0.370
+Appendix M Advantages of Diffusion in Data Space
+These days, inspired from diffusion models in the image and video domains, much work relies on the idea of latent diffusion.
+Here, we want to briefly discuss and emphasize that for tabular data, diffusion in latent space (represented by TabSyn) has important drawbacks and how CDTD, a diffusion model defined in data space differs from that.
+Latent diffusion models first encode the data and map it into a latent space.
+The diffusion model itself is then trained in that latent space.
+Hence, the performance of the diffusion model directly depends on a second, separate model, with a separate training procedure.
+TabSyn uses a VAE model to encode mixed-type data into a common continuous space that is not lower-dimensional, so as to minimize reconstruction errors.
+Any reconstruction errors caused by the incapability of the VAE in turn reduce the sample quality of the eventually generated samples, no matter the capacity of the diffusion model.
+This suggests that we would want to train a high capable encoder/decoder, which adds additional training costs.
+Figure 3 shows that latent diffusion is not necessarily more efficient in the tabular data domain.
+In particular, if the latent space is not lower-dimensional to minimize reconstruction error, then sampling speed is not improved.
+We further hypothesize that much tabular data, due to the lack of redundancy and spatial or sequential correlation, is difficult to summarize efficiently in a joint latent space.
+Hence, compared to other domains, larger VAEs and higher-dimensional latent spaces are required, increasing the training time.
+Also, there is the risk of the VAE not picking up on subtle correlations within the data or distorting existing correlations by mapping into the latent space.
+Any correlations not properly encoded in the latent space, cannot be learned or exploited by the diffusion model.
+Since we optimize the VAE on an average loss, its reconstruction and encoding performance of, for instance, minority classes or extreme values in long-tailed distributions is likely lacking.
+This makes the job of the diffusion model more difficult, if not impossible.
+Lastly, we take great care in homogenizing categorical and continuous features throughout the training process (see Appendix A and B).
+This is a crucial part of modeling mixed-type data.
+Using a VAE to define a diffusion process in latent space only shifts the necessity for homogenization to the VAE training process.
+Not balancing different feature- or data-types and their losses induces an implicit importance weight for each feature.
+Thus, the VAE may sacrifice the reconstruction quality of some features in favor of others (Kendall et al., 2018; Ma et al., 2020).
+To empirically investigate the difference of diffusion in data space (CDTD) and latent diffusion (TabSyn), we examine the worst feature-specific sample quality and other metrics that directly benefit from the model generating all features well.
+Our results in Table 9 show that, latent diffusion comes with a considerable decrease in sample quality (while imposing a similar architecture and number of parameters as well as sampling steps, see Appendix G).
+In particular, the attained maximum metrics indicate that TabSyn has issues modeling all features and their correlations sufficiently well.
+This supports our argument that a homogenization of data types is of crucial importance to avoid having the model implicitly favor one feature over another.
+Table 9: A comparison of the CDTD model to latent diffusion (TabSyn).
+We average each metric over five sampling seeds and as a robust measure report the median over the ablation datasets acsincome, adult, beijing and churn.
+Abs.
+diff.
+in corr.
+matrices refers to the absolute differences in the correlation matrices between ground truth and synthetic data.
+The maximum, minimum and mean are taken across features.
+Detection score L2 dist. of corr. JSD WD Abs. diff. in corr. matrices
+min mean max min mean max min max
+TabSyn 0.772 0.479 0.005 0.018 0.046 0.003 0.006 0.017 0 0.133
+CDTD (per type) 0.566 0.131 0.001 0.012 0.022 0.001 0.003 0.007 0 0.052
+improvement over TabSyn 1.364 3.656 5.000 1.500 2.091 3.000 2.000 2.429 0 2.558
+Appendix N Comparison to Related Work
+Table 10 summarizes our comparison of CDTD to the selected diffusion-based benchmark models, that is, TabSyn, TabDDPM and CoDi.
+Of those models, only TabSyn applies diffusion in latent space, which comes with both advantages and costs (as discussed in Appendix M).
+TabSyn is the only other model besides CDTD that avoids one-hot encoding categorical features by using embeddings.
+This improves the scalability to a higher number of categories without blowing up the input dimensions.
+Although both models utilize embeddings, TabSyn’s generative capabilities are more constrained by jointly encoding all features in a latent space.
+As such, it is still less flexible than CDTD, in particular when modeling very unbalanced categorical data.
+Information on rare categories may easily be cut off in favor of attributing more capacity in the latent space to more prominent categories or features.
+It should also be noted that TabSyn is the only model that makes use of a Transformer architecture in its VAE, which means that it scales quadratically in the number of features and therefore may not be easily scaled to high-dimensional data.
+The CDTD model is the first to utilize adaptive and type- or feature-specific noise schedules to model tabular data.
+Further, we take great care in homogenizing categorical and continuous features throughout the training process, including the model initialization (see Appendix A and B).
+No other model attempts balancing the different features types.
+This is problematic as it suggests that other models may suffer from feature-specific induced implicit importance weights that impact both training and generation processes.
+Hence, the sample quality of some features may be unintentionally sacrificed in favor of increasing the sample quality of other features (Kendall et al., 2018; Ma et al., 2020).
+Note that this also applies to TabSyn: Even though their diffusion model avoids this issue by relying on a single type of loss due to the continuous latent space, the VAE training process does not account for any balancing issues between the two data types.
+Hence, the balancing issue is not eliminated but got only shifted to the encoder VAE.
+Lastly, CDTD and TabSyn are the only models that define the diffusion process in continuous space.
+As such, other advanced techniques, like classifier-free guidance or ODE/SDE samplers, can be directly applied.
+To accommodate categorical data, CoDi and TabDDPM make use of multinomial diffusion (Hoogeboom et al., 2021), which is an inherently discrete process and therefore prohibits such applications.
+Table 10: Comparison of CDTD to the diffusion-based generative models CoDi, TabDDPM and TabSyn.
+($\ast$) Note that the VAE trained as part of the TabSyn model does not balance type-specific losses, which induces an implicit weighting among features.
+This can worsen the sample quality of some features in favor of others.
+defined in feature space avoids one-hot encoding balances feature types adaptive noise schedule type- or feature- specific noise schedules diffusion in continuous space
+CoDi
+TabDDPM
+TabSyn $\ast$
+CDTD (ours)
+Appendix O Examples of Learned Noise Schedules
+Next, we show the learned noise schedules for the smallest (churn) and the largest (acsincome) datasets.
+Additionally, we illustrate the fit of single, per type and per feature schedules to the respective losses.
 Figure 5: (Left): Learned noise schedules for churn. This reflects $F_{\mathrm{d.a.log},k}^{-1}$. (Right): Implicit weighting of noise levels / timesteps. This visualizes $f_{\mathrm{d.a.log},k}$.
@@
 Figure 8: Illustration of the goodness of fit of the timewarping function $F_{k}$ for single (left), per type (middle) and per feature noise schedules (right) on the churn data.
-Appendix L Qualitative Comparisons
+Appendix P Qualitative Comparisons
 Figure 9: Bivariate density for age and educational-num from the adult data.
-Figure 10: Qualitative comparison of some univariate distributions for adult, bank, churn, nmes data.
-Figure 11: Qualitative comparison of some univariate distributions for beijing, default, lending, news data.
-Note that CoDi is excluded from lending as it is prohibitively expensive to train on this dataset.
-Appendix M Visualizations of Captured Correlations
+Figure 10: Comparison of some univariate distributions for adult, bank, churn, nmes.
+Figure 11: Comparison of some univariate distributions for beijing, default, lending, news. (Note that CoDi is prohibitively expensive to train on lending and therefore excluded.)
+Appendix Q Visualizations of Captured Correlations
 Figure 12: Element-wise absolute differences of the correlation matrices between the real training set and the synthetic data for the acsincome dataset.
 TabDDPM generates NaNs for this dataset and is therefore excluded.
+SMOTE takes too long for sampling.
 Continuous (cont.) and categorical (cat.) features are indicated on the axes.
@@
 Figure 17: Element-wise absolute differences of the correlation matrices between the real training set and the synthetic data for the covertype dataset.
+SMOTE takes too long for sampling.
 Continuous (cont.) and categorical (cat.) features are indicated on the axes.
@@
 Continuous (cont.) and categorical (cat.) features are indicated on the axes.
-Appendix N Detailed Results
+Appendix R Detailed Results
 CoDi is prohibitively expensive to train on lending and diabetes and TabDDPM produces NaNs for acsincome and diabetes.
-For both models, the performance metrics on these datasets are therefore not reported.
-Table 6: Model evaluation results averaged over 11 datasets (skipping a dataset if the model was not trainable on it) for five benchmark models and for CDTD with three different noise schedules.
+SMOTE takes too long to sample datasets of a sufficient size for acsincome and covertype (see Table 29).
+For those models, the performance metrics on these datasets are therefore not reported.
+They are assigned a rank of 10 in Table 1 and are not taken into account when forming the average metrics reported in Table 11.
+Table 11: Model evaluation results averaged over 11 datasets (skipping a dataset if the model was not trainable on it, which due to extensive sampling times for SMOTE includes two of the most complex datasets, acsincome and covertype) for seven benchmark models and for CDTD with three different noise schedules.
 Per performance metric, bold indicates the best, underline the second best result.
-ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
-RMSE (abs. diff.; $\downarrow$) 0.094 0.674 0.947 0.486 0.173 0.084 0.101 0.110
-F1 (abs. diff.; $\downarrow$) 0.053 0.130 0.074 0.015 0.044 0.025 0.020 0.025
-AUC (abs. diff.; $\downarrow$) 0.020 0.080 0.065 0.009 0.027 0.018 0.016 0.022
-L2 distance of corr.  ($\downarrow$) 1.321 2.187 2.745 3.786 1.200 0.782 0.756 0.990
-Detection score ($\downarrow$) 0.934 0.986 0.976 0.769 0.936 0.796 0.768 0.783
-JSD ($\downarrow$) 0.011 0.114 0.152 0.051 0.038 0.015 0.016 0.018
-WD ($\downarrow$) 0.011 0.023 0.025 0.061 0.022 0.010 0.007 0.009
-DCR (abs. diff. to test; $\downarrow$) 1.588 3.336 1.621 0.568 1.000 0.796 0.806 0.758
+SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
+RMSE (abs. diff.; $\downarrow$) 0.083 0.094 0.674 0.947 0.486 0.173 0.313 0.084 0.101 0.110
+F1 (abs. diff.; $\downarrow$) 0.007 0.053 0.130 0.074 0.015 0.044 0.099 0.025 0.020 0.025
+AUC (abs. diff.; $\downarrow$) 0.008 0.020 0.080 0.065 0.009 0.027 0.059 0.018 0.016 0.022
+L2 distance of corr.  ($\downarrow$) 0.866 1.321 2.187 2.745 3.786 1.200 2.025 0.782 0.756 0.990
+Detection score ($\downarrow$) 0.661 0.934 0.986 0.976 0.769 0.936 0.877 0.796 0.768 0.783
+JSD ($\downarrow$) 0.055 0.011 0.114 0.152 0.051 0.038 0.048 0.015 0.016 0.018
+WD ($\downarrow$) 0.004 0.011 0.023 0.025 0.061 0.022 0.016 0.010 0.007 0.009
+DCR (abs. diff. to test; $\downarrow$) 1.278 1.588 3.336 1.621 0.568 1.000 2.593 0.796 0.806 0.758
 Table 12: L2 norm (incl.
-standard errors in subscripts) of the correlation matrix differences of real and synthetic train sets for five benchmark models and for CDTD with three different noise schedules.
-ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
-acsincome $0.242_{\pm 0.002}$ $1.696_{\pm 0.008}$ $1.136_{\pm 0.004}$ $0.517_{\pm 0.006}$ $0.141_{\pm 0.003}$ $0.129_{\pm 0.003}$ $0.119_{\pm 0.002}$
-adult $0.576_{\pm 0.006}$ $1.858_{\pm 0.010}$ $0.735_{\pm 0.012}$ $0.156_{\pm 0.006}$ $0.493_{\pm 0.009}$ $0.170_{\pm 0.007}$ $0.125_{\pm 0.009}$ $0.128_{\pm 0.010}$
-bank $0.819_{\pm 0.024}$ $0.947_{\pm 0.019}$ $2.758_{\pm 0.049}$ $0.898_{\pm 0.025}$ $0.499_{\pm 0.021}$ $0.323_{\pm 0.008}$ $0.266_{\pm 0.011}$ $0.256_{\pm 0.015}$
-beijing $0.133_{\pm 0.006}$ $1.445_{\pm 0.009}$ $1.642_{\pm 0.015}$ $1.133_{\pm 0.035}$ $0.363_{\pm 0.015}$ $0.075_{\pm 0.008}$ $0.073_{\pm 0.009}$ $0.071_{\pm 0.004}$
-churn $0.635_{\pm 0.026}$ $1.355_{\pm 0.043}$ $1.301_{\pm 0.041}$ $0.327_{\pm 0.044}$ $0.746_{\pm 0.062}$ $0.302_{\pm 0.041}$ $0.289_{\pm 0.043}$ $0.282_{\pm 0.044}$
-covertype $1.192_{\pm 0.017}$ $3.685_{\pm 0.005}$ $4.668_{\pm 0.003}$ $1.044_{\pm 0.001}$ $1.029_{\pm 0.032}$ $2.359_{\pm 0.011}$ $2.275_{\pm 0.009}$ $2.710_{\pm 0.009}$
-default $1.228_{\pm 0.021}$ $2.697_{\pm 0.021}$ $1.564_{\pm 0.029}$ $3.408_{\pm 0.105}$ $1.672_{\pm 0.061}$ $0.627_{\pm 0.068}$ $0.652_{\pm 0.102}$ $0.737_{\pm 0.033}$
-diabetes $1.189_{\pm 0.004}$ $1.654_{\pm 0.008}$ $5.351_{\pm 0.095}$ $1.201_{\pm 0.020}$ $0.803_{\pm 0.032}$ $1.345_{\pm 0.016}$
-lending $3.473_{\pm 0.057}$ $2.420_{\pm 0.016}$ $5.895_{\pm 0.026}$ $10.675_{\pm 0.015}$ $1.042_{\pm 0.075}$ $1.189_{\pm 0.040}$ $1.363_{\pm 0.097}$
-news $4.333_{\pm 0.128}$ $4.641_{\pm 0.028}$ $4.612_{\pm 0.016}$ $15.985_{\pm 0.081}$ $4.874_{\pm 0.148}$ $1.925_{\pm 0.527}$ $2.035_{\pm 0.475}$ $3.395_{\pm 0.950}$
-nmes $0.717_{\pm 0.054}$ $1.663_{\pm 0.035}$ $0.532_{\pm 0.030}$ $0.447_{\pm 0.031}$ $0.609_{\pm 0.032}$ $0.433_{\pm 0.025}$ $0.478_{\pm 0.083}$ $0.481_{\pm 0.058}$
-Table 8: Jensen-Shannon divergence (incl. standard errors in subscripts) for five benchmark models and for CDTD with three different noise schedules.
-ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
-acsincome $0.013_{\pm 0.001}$ $0.256_{\pm 0.000}$ $0.309_{\pm 0.000}$ $0.076_{\pm 0.001}$ $0.025_{\pm 0.001}$ $0.024_{\pm 0.000}$ $0.022_{\pm 0.001}$
-adult $0.007_{\pm 0.001}$ $0.112_{\pm 0.001}$ $0.113_{\pm 0.001}$ $0.034_{\pm 0.001}$ $0.045_{\pm 0.001}$ $0.010_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.016_{\pm 0.000}$
-bank $0.004_{\pm 0.000}$ $0.086_{\pm 0.001}$ $0.191_{\pm 0.001}$ $0.029_{\pm 0.001}$ $0.038_{\pm 0.001}$ $0.010_{\pm 0.000}$ $0.009_{\pm 0.001}$ $0.012_{\pm 0.001}$
-beijing $0.005_{\pm 0.002}$ $0.147_{\pm 0.003}$ $0.257_{\pm 0.001}$ $0.035_{\pm 0.003}$ $0.018_{\pm 0.004}$ $0.003_{\pm 0.001}$ $0.005_{\pm 0.002}$ $0.005_{\pm 0.001}$
-churn $0.011_{\pm 0.004}$ $0.095_{\pm 0.003}$ $0.048_{\pm 0.004}$ $0.014_{\pm 0.004}$ $0.043_{\pm 0.001}$ $0.012_{\pm 0.003}$ $0.012_{\pm 0.002}$ $0.011_{\pm 0.002}$
-covertype $0.002_{\pm 0.000}$ $0.044_{\pm 0.000}$ $0.043_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.011_{\pm 0.000}$
-default $0.008_{\pm 0.001}$ $0.194_{\pm 0.001}$ $0.177_{\pm 0.001}$ $0.027_{\pm 0.002}$ $0.073_{\pm 0.002}$ $0.013_{\pm 0.001}$ $0.015_{\pm 0.001}$ $0.015_{\pm 0.001}$
-diabetes $0.009_{\pm 0.000}$ $0.093_{\pm 0.000}$ $0.187_{\pm 0.000}$ $0.022_{\pm 0.000}$ $0.023_{\pm 0.000}$ $0.026_{\pm 0.000}$
-lending $0.049_{\pm 0.002}$ $0.092_{\pm 0.001}$ $0.188_{\pm 0.001}$ $0.243_{\pm 0.002}$ $0.055_{\pm 0.001}$ $0.056_{\pm 0.001}$ $0.064_{\pm 0.002}$
-news $0.002_{\pm 0.001}$ $0.022_{\pm 0.001}$ $0.128_{\pm 0.001}$ $0.046_{\pm 0.000}$ $0.012_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$
-nmes $0.008_{\pm 0.002}$ $0.117_{\pm 0.002}$ $0.029_{\pm 0.003}$ $0.028_{\pm 0.004}$ $0.027_{\pm 0.003}$ $0.008_{\pm 0.001}$ $0.009_{\pm 0.001}$ $0.013_{\pm 0.003}$
-Table 9: Wasserstein distance (incl. standard errors in subscripts) for five benchmark models and for CDTD with three different noise schedules.
-ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
-acsincome $0.007_{\pm 0.000}$ $0.037_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.017_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.001_{\pm 0.000}$ $0.001_{\pm 0.000}$
-adult $0.012_{\pm 0.000}$ $0.016_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$
-bank $0.012_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.040_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.030_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$
-beijing $0.009_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.058_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.019_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.002_{\pm 0.000}$
-churn $0.013_{\pm 0.001}$ $0.027_{\pm 0.001}$ $0.032_{\pm 0.001}$ $0.008_{\pm 0.002}$ $0.048_{\pm 0.002}$ $0.008_{\pm 0.001}$ $0.007_{\pm 0.001}$ $0.006_{\pm 0.001}$
-covertype $0.006_{\pm 0.000}$ $0.041_{\pm 0.000}$ $0.022_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.017_{\pm 0.000}$ $0.015_{\pm 0.000}$ $0.012_{\pm 0.000}$
-default $0.005_{\pm 0.000}$ $0.011_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$
-diabetes $0.012_{\pm 0.000}$ $0.020_{\pm 0.000}$ $0.038_{\pm 0.000}$ $0.038_{\pm 0.000}$ $0.020_{\pm 0.000}$ $0.042_{\pm 0.000}$
-lending $0.013_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.016_{\pm 0.000}$ $0.425_{\pm 0.001}$ $0.009_{\pm 0.000}$ $0.010_{\pm 0.000}$ $0.011_{\pm 0.000}$
-news $0.024_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.018_{\pm 0.000}$ $0.078_{\pm 0.001}$ $0.030_{\pm 0.000}$ $0.007_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.008_{\pm 0.000}$
-nmes $0.012_{\pm 0.000}$ $0.036_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.007_{\pm 0.001}$ $0.016_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.000}$
-Table 10: Detection score (incl. standard errors in subscripts) for five benchmark models and for CDTD with three different noise schedules.
-ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
-acsincome $0.808_{\pm 0.001}$ $0.989_{\pm 0.001}$ $0.985_{\pm 0.000}$ $0.825_{\pm 0.002}$ $0.540_{\pm 0.003}$ $0.532_{\pm 0.004}$ $0.526_{\pm 0.002}$
-adult $0.889_{\pm 0.002}$ $0.997_{\pm 0.000}$ $0.967_{\pm 0.001}$ $0.590_{\pm 0.003}$ $0.992_{\pm 0.001}$ $0.604_{\pm 0.002}$ $0.588_{\pm 0.002}$ $0.591_{\pm 0.005}$
-bank $0.955_{\pm 0.002}$ $1.000_{\pm 0.000}$ $0.988_{\pm 0.001}$ $0.783_{\pm 0.003}$ $1.000_{\pm 0.000}$ $0.795_{\pm 0.003}$ $0.739_{\pm 0.003}$ $0.694_{\pm 0.006}$
-beijing $0.995_{\pm 0.000}$ $0.998_{\pm 0.000}$ $0.993_{\pm 0.001}$ $0.966_{\pm 0.002}$ $0.997_{\pm 0.001}$ $0.951_{\pm 0.002}$ $0.949_{\pm 0.001}$ $0.947_{\pm 0.002}$
-churn $0.853_{\pm 0.002}$ $0.945_{\pm 0.006}$ $0.843_{\pm 0.011}$ $0.561_{\pm 0.005}$ $0.730_{\pm 0.012}$ $0.621_{\pm 0.016}$ $0.533_{\pm 0.007}$ $0.544_{\pm 0.031}$
-covertype $0.945_{\pm 0.002}$ $0.997_{\pm 0.000}$ $0.989_{\pm 0.001}$ $0.586_{\pm 0.002}$ $0.900_{\pm 0.002}$ $0.991_{\pm 0.001}$ $0.992_{\pm 0.001}$ $0.991_{\pm 0.001}$
-default $0.991_{\pm 0.001}$ $0.998_{\pm 0.001}$ $0.997_{\pm 0.001}$ $0.821_{\pm 0.002}$ $0.995_{\pm 0.000}$ $0.827_{\pm 0.004}$ $0.802_{\pm 0.003}$ $0.871_{\pm 0.001}$
-diabetes $0.854_{\pm 0.002}$ $0.935_{\pm 0.002}$ $0.997_{\pm 0.001}$ $0.858_{\pm 0.001}$ $0.780_{\pm 0.002}$ $0.866_{\pm 0.002}$
-lending $0.997_{\pm 0.001}$ $0.995_{\pm 0.002}$ $0.995_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.955_{\pm 0.006}$ $0.954_{\pm 0.009}$ $0.961_{\pm 0.004}$
-news $0.998_{\pm 0.000}$ $1.000_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.966_{\pm 0.002}$ $1.000_{\pm 0.000}$ $0.973_{\pm 0.001}$ $0.953_{\pm 0.001}$ $0.977_{\pm 0.001}$
-nmes $0.987_{\pm 0.002}$ $0.992_{\pm 0.003}$ $0.988_{\pm 0.002}$ $0.650_{\pm 0.014}$ $0.988_{\pm 0.000}$ $0.636_{\pm 0.008}$ $0.623_{\pm 0.008}$ $0.642_{\pm 0.010}$
-Table 11: Distance to closest record of the generated data (incl. standard errors in subscripts) for five benchmark models and for CDTD with three different noise schedules.
-ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
-acsincome $8.637_{\pm 0.027}$ $10.758_{\pm 0.054}$ $6.652_{\pm 0.032}$ $10.877_{\pm 0.092}$ $8.346_{\pm 0.056}$ $8.322_{\pm 0.047}$ $8.349_{\pm 0.033}$
-adult $2.523_{\pm 0.012}$ $5.012_{\pm 0.028}$ $2.227_{\pm 0.013}$ $1.647_{\pm 0.009}$ $2.735_{\pm 0.028}$ $1.112_{\pm 0.019}$ $1.231_{\pm 0.011}$ $1.294_{\pm 0.009}$
-bank $3.025_{\pm 0.017}$ $3.840_{\pm 0.014}$ $3.136_{\pm 0.007}$ $2.327_{\pm 0.010}$ $3.062_{\pm 0.012}$ $1.828_{\pm 0.008}$ $1.943_{\pm 0.007}$ $2.062_{\pm 0.008}$
-beijing $0.735_{\pm 0.003}$ $1.004_{\pm 0.006}$ $0.926_{\pm 0.003}$ $0.739_{\pm 0.006}$ $0.610_{\pm 0.002}$ $0.490_{\pm 0.002}$ $0.489_{\pm 0.001}$ $0.477_{\pm 0.002}$
-churn $1.136_{\pm 0.015}$ $1.804_{\pm 0.036}$ $1.146_{\pm 0.039}$ $0.342_{\pm 0.031}$ $0.852_{\pm 0.016}$ $0.332_{\pm 0.021}$ $0.274_{\pm 0.021}$ $0.276_{\pm 0.012}$
-covertype $1.741_{\pm 0.011}$ $5.773_{\pm 0.017}$ $3.173_{\pm 0.013}$ $0.889_{\pm 0.007}$ $1.508_{\pm 0.020}$ $2.297_{\pm 0.026}$ $2.209_{\pm 0.022}$ $2.252_{\pm 0.013}$
-default $3.095_{\pm 0.026}$ $5.880_{\pm 0.020}$ $3.216_{\pm 0.013}$ $1.422_{\pm 0.013}$ $2.593_{\pm 0.020}$ $1.127_{\pm 0.028}$ $1.269_{\pm 0.014}$ $1.253_{\pm 0.012}$
-diabetes $17.736_{\pm 0.107}$ $21.935_{\pm 0.046}$ $8.214_{\pm 0.022}$ $15.279_{\pm 0.026}$ $15.126_{\pm 0.058}$ $15.350_{\pm 0.059}$
-lending $17.776_{\pm 0.132}$ $20.239_{\pm 0.222}$ $10.688_{\pm 0.025}$ $12.537_{\pm 0.076}$ $13.775_{\pm 0.147}$ $14.162_{\pm 0.188}$ $13.966_{\pm 0.282}$
-news $6.147_{\pm 0.010}$ $4.789_{\pm 0.005}$ $5.821_{\pm 0.003}$ $4.960_{\pm 0.006}$ $4.661_{\pm 0.023}$ $3.635_{\pm 0.004}$ $3.687_{\pm 0.006}$ $3.749_{\pm 0.048}$
-nmes $2.203_{\pm 0.028}$ $2.971_{\pm 0.008}$ $1.710_{\pm 0.019}$ $0.891_{\pm 0.033}$ $1.231_{\pm 0.024}$ $0.664_{\pm 0.029}$ $0.710_{\pm 0.032}$ $0.771_{\pm 0.023}$
-Table 12: Machine learning efficiency F1 score for five benchmark models, the real training data and for CDTD with three different noise schedules.
+standard errors in subscripts) of the correlation matrix differences of real and synthetic train sets for seven benchmark models and for CDTD with three different noise schedules.
+SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
+acsincome - $0.242_{\pm 0.002}$ $1.696_{\pm 0.008}$ $1.136_{\pm 0.004}$ $0.517_{\pm 0.006}$ $0.524_{\pm 0.010}$ $0.141_{\pm 0.003}$ $0.129_{\pm 0.003}$ $0.119_{\pm 0.002}$
+adult $0.414_{\pm 0.016}$ $0.576_{\pm 0.006}$ $1.858_{\pm 0.010}$ $0.735_{\pm 0.012}$ $0.156_{\pm 0.006}$ $0.493_{\pm 0.009}$ $0.449_{\pm 0.011}$ $0.170_{\pm 0.007}$ $0.125_{\pm 0.009}$ $0.128_{\pm 0.010}$
+bank $0.404_{\pm 0.015}$ $0.819_{\pm 0.024}$ $0.947_{\pm 0.019}$ $2.758_{\pm 0.049}$ $0.898_{\pm 0.025}$ $0.499_{\pm 0.021}$ $0.677_{\pm 0.015}$ $0.323_{\pm 0.008}$ $0.266_{\pm 0.011}$ $0.256_{\pm 0.015}$
+beijing $0.081_{\pm 0.007}$ $0.133_{\pm 0.006}$ $1.445_{\pm 0.009}$ $1.642_{\pm 0.015}$ $1.133_{\pm 0.035}$ $0.363_{\pm 0.015}$ $0.096_{\pm 0.008}$ $0.075_{\pm 0.008}$ $0.073_{\pm 0.009}$ $0.071_{\pm 0.004}$
+churn $0.264_{\pm 0.036}$ $0.635_{\pm 0.026}$ $1.355_{\pm 0.043}$ $1.301_{\pm 0.041}$ $0.327_{\pm 0.044}$ $0.746_{\pm 0.062}$ $0.509_{\pm 0.053}$ $0.302_{\pm 0.041}$ $0.289_{\pm 0.043}$ $0.282_{\pm 0.044}$
+covertype - $1.192_{\pm 0.017}$ $3.685_{\pm 0.005}$ $4.668_{\pm 0.003}$ $1.044_{\pm 0.001}$ $1.029_{\pm 0.032}$ $3.958_{\pm 0.243}$ $2.359_{\pm 0.011}$ $2.275_{\pm 0.009}$ $2.710_{\pm 0.009}$
+default $0.709_{\pm 0.048}$ $1.228_{\pm 0.021}$ $2.697_{\pm 0.021}$ $1.564_{\pm 0.029}$ $3.408_{\pm 0.105}$ $1.672_{\pm 0.061}$ $1.121_{\pm 0.042}$ $0.627_{\pm 0.068}$ $0.652_{\pm 0.102}$ $0.737_{\pm 0.033}$
+diabetes $2.355_{\pm 0.026}$ $1.189_{\pm 0.004}$ $1.654_{\pm 0.008}$ $5.351_{\pm 0.095}$ $2.381_{\pm 0.026}$ $1.201_{\pm 0.020}$ $0.803_{\pm 0.032}$ $1.345_{\pm 0.016}$
+lending $1.321_{\pm 0.063}$ $3.473_{\pm 0.057}$ $2.420_{\pm 0.016}$ $5.895_{\pm 0.026}$ $10.675_{\pm 0.015}$ $6.701_{\pm 0.034}$ $1.042_{\pm 0.075}$ $1.189_{\pm 0.040}$ $1.363_{\pm 0.097}$
+news $1.684_{\pm 1.466}$ $4.333_{\pm 0.128}$ $4.641_{\pm 0.028}$ $4.612_{\pm 0.016}$ $15.985_{\pm 0.081}$ $4.874_{\pm 0.148}$ $4.990_{\pm 0.024}$ $1.925_{\pm 0.527}$ $2.035_{\pm 0.475}$ $3.395_{\pm 0.950}$
+nmes $0.565_{\pm 0.047}$ $0.717_{\pm 0.054}$ $1.663_{\pm 0.035}$ $0.532_{\pm 0.030}$ $0.447_{\pm 0.031}$ $0.609_{\pm 0.032}$ $0.867_{\pm 0.046}$ $0.433_{\pm 0.025}$ $0.478_{\pm 0.083}$ $0.481_{\pm 0.058}$
+Table 13: Jensen-Shannon divergence (incl. standard errors in subscripts) for seven benchmark models and for CDTD with three different noise schedules.
+SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
+acsincome - $0.013_{\pm 0.001}$ $0.256_{\pm 0.000}$ $0.309_{\pm 0.000}$ $0.076_{\pm 0.001}$ $0.045_{\pm 0.001}$ $0.025_{\pm 0.001}$ $0.024_{\pm 0.000}$ $0.022_{\pm 0.001}$
+adult $0.064_{\pm 0.001}$ $0.007_{\pm 0.001}$ $0.112_{\pm 0.001}$ $0.113_{\pm 0.001}$ $0.034_{\pm 0.001}$ $0.045_{\pm 0.001}$ $0.020_{\pm 0.001}$ $0.010_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.016_{\pm 0.000}$
+bank $0.039_{\pm 0.001}$ $0.004_{\pm 0.000}$ $0.086_{\pm 0.001}$ $0.191_{\pm 0.001}$ $0.029_{\pm 0.001}$ $0.038_{\pm 0.001}$ $0.054_{\pm 0.001}$ $0.010_{\pm 0.000}$ $0.009_{\pm 0.001}$ $0.012_{\pm 0.001}$
+beijing $0.006_{\pm 0.002}$ $0.005_{\pm 0.002}$ $0.147_{\pm 0.003}$ $0.257_{\pm 0.001}$ $0.035_{\pm 0.003}$ $0.018_{\pm 0.004}$ $0.007_{\pm 0.002}$ $0.003_{\pm 0.001}$ $0.005_{\pm 0.002}$ $0.005_{\pm 0.001}$
+churn $0.012_{\pm 0.004}$ $0.011_{\pm 0.004}$ $0.095_{\pm 0.003}$ $0.048_{\pm 0.004}$ $0.014_{\pm 0.004}$ $0.043_{\pm 0.001}$ $0.017_{\pm 0.002}$ $0.012_{\pm 0.003}$ $0.012_{\pm 0.002}$ $0.011_{\pm 0.002}$
+covertype - $0.002_{\pm 0.000}$ $0.044_{\pm 0.000}$ $0.043_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.049_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.011_{\pm 0.000}$
+default $0.042_{\pm 0.001}$ $0.008_{\pm 0.001}$ $0.194_{\pm 0.001}$ $0.177_{\pm 0.001}$ $0.027_{\pm 0.002}$ $0.073_{\pm 0.002}$ $0.082_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.015_{\pm 0.001}$ $0.015_{\pm 0.001}$
+diabetes $0.067_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.093_{\pm 0.000}$ $0.187_{\pm 0.000}$ $0.095_{\pm 0.000}$ $0.022_{\pm 0.000}$ $0.023_{\pm 0.000}$ $0.026_{\pm 0.000}$
+lending $0.143_{\pm 0.001}$ $0.049_{\pm 0.002}$ $0.092_{\pm 0.001}$ $0.188_{\pm 0.001}$ $0.243_{\pm 0.002}$ $0.114_{\pm 0.002}$ $0.055_{\pm 0.001}$ $0.056_{\pm 0.001}$ $0.064_{\pm 0.002}$
+news $0.063_{\pm 0.001}$ $0.002_{\pm 0.001}$ $0.022_{\pm 0.001}$ $0.128_{\pm 0.001}$ $0.046_{\pm 0.000}$ $0.012_{\pm 0.001}$ $0.016_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$
+nmes $0.060_{\pm 0.001}$ $0.008_{\pm 0.002}$ $0.117_{\pm 0.002}$ $0.029_{\pm 0.003}$ $0.028_{\pm 0.004}$ $0.027_{\pm 0.003}$ $0.026_{\pm 0.001}$ $0.008_{\pm 0.001}$ $0.009_{\pm 0.001}$ $0.013_{\pm 0.003}$
+Table 14: Wasserstein distance (incl. standard errors in subscripts) for seven benchmark models and for CDTD with three different noise schedules.
+SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
+acsincome - $0.007_{\pm 0.000}$ $0.037_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.017_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.001_{\pm 0.000}$ $0.001_{\pm 0.000}$
+adult $0.003_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.016_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$
+bank $0.002_{\pm 0.001}$ $0.012_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.040_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.030_{\pm 0.001}$ $0.005_{\pm 0.000}$ $0.006_{\pm 0.001}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$
+beijing $0.002_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.058_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.019_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.002_{\pm 0.000}$
+churn $0.006_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.027_{\pm 0.001}$ $0.032_{\pm 0.001}$ $0.008_{\pm 0.002}$ $0.048_{\pm 0.002}$ $0.013_{\pm 0.002}$ $0.008_{\pm 0.001}$ $0.007_{\pm 0.001}$ $0.006_{\pm 0.001}$
+covertype - $0.006_{\pm 0.000}$ $0.041_{\pm 0.000}$ $0.022_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.017_{\pm 0.000}$ $0.017_{\pm 0.000}$ $0.015_{\pm 0.000}$ $0.012_{\pm 0.000}$
+default $0.002_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.011_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$
+diabetes $0.004_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.020_{\pm 0.000}$ $0.038_{\pm 0.000}$ $0.011_{\pm 0.000}$ $0.038_{\pm 0.000}$ $0.020_{\pm 0.000}$ $0.042_{\pm 0.000}$
+lending $0.006_{\pm 0.000}$ $0.013_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.016_{\pm 0.000}$ $0.425_{\pm 0.001}$ $0.050_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.010_{\pm 0.000}$ $0.011_{\pm 0.000}$
+news $0.007_{\pm 0.000}$ $0.024_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.018_{\pm 0.000}$ $0.078_{\pm 0.001}$ $0.030_{\pm 0.000}$ $0.025_{\pm 0.000}$ $0.007_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.008_{\pm 0.000}$
+nmes $0.005_{\pm 0.001}$ $0.012_{\pm 0.000}$ $0.036_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.007_{\pm 0.001}$ $0.016_{\pm 0.001}$ $0.038_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.000}$
+Table 15: Detection score (incl. standard errors in subscripts) for seven benchmark models and for CDTD with three different noise schedules.
+SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
+acsincome - $0.808_{\pm 0.001}$ $0.989_{\pm 0.001}$ $0.985_{\pm 0.000}$ $0.825_{\pm 0.002}$ $0.680_{\pm 0.002}$ $0.540_{\pm 0.003}$ $0.532_{\pm 0.004}$ $0.526_{\pm 0.002}$
+adult $0.320_{\pm 0.006}$ $0.889_{\pm 0.002}$ $0.997_{\pm 0.000}$ $0.967_{\pm 0.001}$ $0.590_{\pm 0.003}$ $0.992_{\pm 0.001}$ $0.630_{\pm 0.003}$ $0.604_{\pm 0.002}$ $0.588_{\pm 0.002}$ $0.591_{\pm 0.005}$
+bank $0.633_{\pm 0.008}$ $0.955_{\pm 0.002}$ $1.000_{\pm 0.000}$ $0.988_{\pm 0.001}$ $0.783_{\pm 0.003}$ $1.000_{\pm 0.000}$ $0.843_{\pm 0.002}$ $0.795_{\pm 0.003}$ $0.739_{\pm 0.003}$ $0.694_{\pm 0.006}$
+beijing $0.976_{\pm 0.001}$ $0.995_{\pm 0.000}$ $0.998_{\pm 0.000}$ $0.993_{\pm 0.001}$ $0.966_{\pm 0.002}$ $0.997_{\pm 0.001}$ $0.966_{\pm 0.001}$ $0.951_{\pm 0.002}$ $0.949_{\pm 0.001}$ $0.947_{\pm 0.002}$
+churn $0.339_{\pm 0.020}$ $0.853_{\pm 0.002}$ $0.945_{\pm 0.006}$ $0.843_{\pm 0.011}$ $0.561_{\pm 0.005}$ $0.730_{\pm 0.012}$ $0.865_{\pm 0.012}$ $0.621_{\pm 0.016}$ $0.533_{\pm 0.007}$ $0.544_{\pm 0.031}$
+covertype - $0.945_{\pm 0.002}$ $0.997_{\pm 0.000}$ $0.989_{\pm 0.001}$ $0.586_{\pm 0.002}$ $0.900_{\pm 0.002}$ $0.979_{\pm 0.001}$ $0.991_{\pm 0.001}$ $0.992_{\pm 0.001}$ $0.991_{\pm 0.001}$
+default $0.493_{\pm 0.009}$ $0.991_{\pm 0.001}$ $0.998_{\pm 0.001}$ $0.997_{\pm 0.001}$ $0.821_{\pm 0.002}$ $0.995_{\pm 0.000}$ $0.902_{\pm 0.001}$ $0.827_{\pm 0.004}$ $0.802_{\pm 0.003}$ $0.871_{\pm 0.001}$
+diabetes $0.367_{\pm 0.001}$ $0.854_{\pm 0.002}$ $0.935_{\pm 0.002}$ $0.997_{\pm 0.001}$ $0.940_{\pm 0.001}$ $0.858_{\pm 0.001}$ $0.780_{\pm 0.002}$ $0.866_{\pm 0.002}$
+lending $0.926_{\pm 0.004}$ $0.997_{\pm 0.001}$ $0.995_{\pm 0.002}$ $0.995_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.998_{\pm 0.001}$ $0.955_{\pm 0.006}$ $0.954_{\pm 0.009}$ $0.961_{\pm 0.004}$
+news $0.993_{\pm 0.001}$ $0.998_{\pm 0.000}$ $1.000_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.966_{\pm 0.002}$ $1.000_{\pm 0.000}$ $0.999_{\pm 0.000}$ $0.973_{\pm 0.001}$ $0.953_{\pm 0.001}$ $0.977_{\pm 0.001}$
+nmes $0.905_{\pm 0.007}$ $0.987_{\pm 0.002}$ $0.992_{\pm 0.003}$ $0.988_{\pm 0.002}$ $0.650_{\pm 0.014}$ $0.988_{\pm 0.000}$ $0.841_{\pm 0.008}$ $0.636_{\pm 0.008}$ $0.623_{\pm 0.008}$ $0.642_{\pm 0.010}$
+Table 16: Distance to closest record of the generated data (incl. standard errors in subscripts) for seven benchmark models and for CDTD with three different noise schedules.
+Test Set SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
+acsincome $7.673_{\pm 0.017}$ $8.637_{\pm 0.027}$ $10.758_{\pm 0.054}$ $6.652_{\pm 0.032}$ $10.877_{\pm 0.092}$ $10.305_{\pm 0.073}$ $8.346_{\pm 0.056}$ $8.322_{\pm 0.047}$ $8.349_{\pm 0.033}$
+adult $1.870_{\pm 0.000}$ $1.371_{\pm 0.018}$ $2.523_{\pm 0.012}$ $5.012_{\pm 0.028}$ $2.227_{\pm 0.013}$ $1.647_{\pm 0.009}$ $2.735_{\pm 0.028}$ $2.341_{\pm 0.013}$ $1.112_{\pm 0.019}$ $1.231_{\pm 0.011}$ $1.294_{\pm 0.009}$
+bank $2.369_{\pm 0.000}$ $1.369_{\pm 0.011}$ $3.025_{\pm 0.017}$ $3.840_{\pm 0.014}$ $3.136_{\pm 0.007}$ $2.327_{\pm 0.010}$ $3.062_{\pm 0.012}$ $2.973_{\pm 0.012}$ $1.828_{\pm 0.008}$ $1.943_{\pm 0.007}$ $2.062_{\pm 0.008}$
+beijing $0.385_{\pm 0.000}$ $0.139_{\pm 0.003}$ $0.735_{\pm 0.003}$ $1.004_{\pm 0.006}$ $0.926_{\pm 0.003}$ $0.739_{\pm 0.006}$ $0.610_{\pm 0.002}$ $0.626_{\pm 0.001}$ $0.490_{\pm 0.002}$ $0.489_{\pm 0.001}$ $0.477_{\pm 0.002}$
+churn $0.347_{\pm 0.000}$ $0.232_{\pm 0.028}$ $1.136_{\pm 0.015}$ $1.804_{\pm 0.036}$ $1.146_{\pm 0.039}$ $0.342_{\pm 0.031}$ $0.852_{\pm 0.016}$ $1.130_{\pm 0.018}$ $0.332_{\pm 0.021}$ $0.274_{\pm 0.021}$ $0.276_{\pm 0.012}$
+covertype $0.529_{\pm 0.001}$ $1.741_{\pm 0.011}$ $5.773_{\pm 0.017}$ $3.173_{\pm 0.013}$ $0.889_{\pm 0.007}$ $1.508_{\pm 0.020}$ $3.086_{\pm 0.009}$ $2.297_{\pm 0.026}$ $2.209_{\pm 0.022}$ $2.252_{\pm 0.013}$
+default $1.812_{\pm 0.000}$ $1.032_{\pm 0.010}$ $3.095_{\pm 0.026}$ $5.880_{\pm 0.020}$ $3.216_{\pm 0.013}$ $1.422_{\pm 0.013}$ $2.593_{\pm 0.020}$ $2.603_{\pm 0.018}$ $1.127_{\pm 0.028}$ $1.269_{\pm 0.014}$ $1.253_{\pm 0.012}$
+diabetes $15.608_{\pm 0.055}$ $13.909_{\pm 0.050}$ $17.736_{\pm 0.107}$ $21.935_{\pm 0.046}$ $8.214_{\pm 0.022}$ $28.955_{\pm 0.060}$ $15.279_{\pm 0.026}$ $15.126_{\pm 0.058}$ $15.350_{\pm 0.059}$
+lending $11.184_{\pm 0.000}$ $17.752_{\pm 0.143}$ $17.776_{\pm 0.132}$ $20.239_{\pm 0.222}$ $10.688_{\pm 0.025}$ $12.537_{\pm 0.076}$ $16.222_{\pm 0.092}$ $13.775_{\pm 0.147}$ $14.162_{\pm 0.188}$ $13.966_{\pm 0.282}$
+news $3.615_{\pm 0.000}$ $3.553_{\pm 0.134}$ $6.147_{\pm 0.010}$ $4.789_{\pm 0.005}$ $5.821_{\pm 0.003}$ $4.960_{\pm 0.006}$ $4.661_{\pm 0.023}$ $5.351_{\pm 0.008}$ $3.635_{\pm 0.004}$ $3.687_{\pm 0.006}$ $3.749_{\pm 0.048}$
+nmes $1.931_{\pm 0.000}$ $1.394_{\pm 0.019}$ $2.203_{\pm 0.028}$ $2.971_{\pm 0.008}$ $1.710_{\pm 0.019}$ $0.891_{\pm 0.033}$ $1.231_{\pm 0.024}$ $2.260_{\pm 0.034}$ $0.664_{\pm 0.029}$ $0.710_{\pm 0.032}$ $0.771_{\pm 0.023}$
+Table 17: Machine learning efficiency F1 score for seven benchmark models, the real training data and for CDTD with three different noise schedules.
 The standard deviation takes into account five different sampling seeds and uses the average results of the four machine learning efficiency models computed across ten model seeds.
-Real Data ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
-adult $0.797_{\pm 0.000}$ $0.769_{\pm 0.002}$ $0.647_{\pm 0.015}$ $0.756_{\pm 0.002}$ $0.787_{\pm 0.001}$ $0.745_{\pm 0.004}$ $0.787_{\pm 0.001}$ $0.787_{\pm 0.001}$ $0.787_{\pm 0.001}$
-bank $0.745_{\pm 0.002}$ $0.682_{\pm 0.006}$ $0.680_{\pm 0.006}$ $0.629_{\pm 0.006}$ $0.720_{\pm 0.006}$ $0.673_{\pm 0.006}$ $0.776_{\pm 0.003}$ $0.767_{\pm 0.004}$ $0.737_{\pm 0.004}$
-churn $0.873_{\pm 0.003}$ $0.780_{\pm 0.015}$ $0.761_{\pm 0.009}$ $0.802_{\pm 0.017}$ $0.857_{\pm 0.007}$ $0.865_{\pm 0.008}$ $0.854_{\pm 0.011}$ $0.852_{\pm 0.006}$ $0.845_{\pm 0.011}$
-covertype $0.817_{\pm 0.001}$ $0.783_{\pm 0.001}$ $0.442_{\pm 0.008}$ $0.711_{\pm 0.002}$ $0.799_{\pm 0.001}$ $0.767_{\pm 0.001}$ $0.734_{\pm 0.002}$ $0.754_{\pm 0.001}$ $0.722_{\pm 0.002}$
-default $0.674_{\pm 0.001}$ $0.627_{\pm 0.003}$ $0.686_{\pm 0.002}$ $0.632_{\pm 0.007}$ $0.678_{\pm 0.002}$ $0.638_{\pm 0.008}$ $0.670_{\pm 0.002}$ $0.671_{\pm 0.001}$ $0.673_{\pm 0.003}$
-diabetes $0.621_{\pm 0.002}$ $0.572_{\pm 0.005}$ $0.557_{\pm 0.004}$ $0.553_{\pm 0.003}$ $0.617_{\pm 0.002}$ $0.617_{\pm 0.002}$ $0.611_{\pm 0.002}$
-Table 13: Machine learning efficiency AUC score for five benchmark models, the real training data and for CDTD with three different noise schedules.
+Real Data SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
+adult $0.797_{\pm 0.000}$ $0.784_{\pm 0.001}$ $0.769_{\pm 0.002}$ $0.647_{\pm 0.015}$ $0.756_{\pm 0.002}$ $0.787_{\pm 0.001}$ $0.745_{\pm 0.004}$ $0.782_{\pm 0.001}$ $0.787_{\pm 0.001}$ $0.787_{\pm 0.001}$ $0.787_{\pm 0.001}$
+bank $0.745_{\pm 0.002}$ $0.740_{\pm 0.004}$ $0.682_{\pm 0.006}$ $0.680_{\pm 0.006}$ $0.629_{\pm 0.006}$ $0.720_{\pm 0.006}$ $0.673_{\pm 0.006}$ $0.711_{\pm 0.007}$ $0.776_{\pm 0.003}$ $0.767_{\pm 0.004}$ $0.737_{\pm 0.004}$
+churn $0.873_{\pm 0.003}$ $0.865_{\pm 0.008}$ $0.780_{\pm 0.015}$ $0.761_{\pm 0.009}$ $0.802_{\pm 0.017}$ $0.857_{\pm 0.007}$ $0.865_{\pm 0.008}$ $0.771_{\pm 0.014}$ $0.854_{\pm 0.011}$ $0.852_{\pm 0.006}$ $0.845_{\pm 0.011}$
+covertype $0.817_{\pm 0.001}$ $0.783_{\pm 0.001}$ $0.442_{\pm 0.008}$ $0.711_{\pm 0.002}$ $0.799_{\pm 0.001}$ $0.767_{\pm 0.001}$ $0.614_{\pm 0.015}$ $0.734_{\pm 0.002}$ $0.754_{\pm 0.001}$ $0.722_{\pm 0.002}$
+default $0.674_{\pm 0.001}$ $0.677_{\pm 0.001}$ $0.627_{\pm 0.003}$ $0.686_{\pm 0.002}$ $0.632_{\pm 0.007}$ $0.678_{\pm 0.002}$ $0.638_{\pm 0.008}$ $0.496_{\pm 0.009}$ $0.670_{\pm 0.002}$ $0.671_{\pm 0.001}$ $0.673_{\pm 0.003}$
+diabetes $0.621_{\pm 0.002}$ $0.615_{\pm 0.002}$ $0.572_{\pm 0.005}$ $0.557_{\pm 0.004}$ $0.553_{\pm 0.003}$ $0.560_{\pm 0.006}$ $0.617_{\pm 0.002}$ $0.617_{\pm 0.002}$ $0.611_{\pm 0.002}$
+Table 18: Machine learning efficiency AUC score for seven benchmark models, the real training data and for CDTD with three different noise schedules.
 The standard deviation takes into account five different sampling seeds and uses the average results of the four machine learning efficiency models computed across ten model seeds.
-Real Data ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
-adult $0.915_{\pm 0.000}$ $0.901_{\pm 0.000}$ $0.836_{\pm 0.006}$ $0.889_{\pm 0.002}$ $0.908_{\pm 0.000}$ $0.880_{\pm 0.005}$ $0.910_{\pm 0.000}$ $0.910_{\pm 0.001}$ $0.909_{\pm 0.000}$
-bank $0.947_{\pm 0.000}$ $0.938_{\pm 0.001}$ $0.934_{\pm 0.003}$ $0.830_{\pm 0.020}$ $0.940_{\pm 0.005}$ $0.929_{\pm 0.005}$ $0.945_{\pm 0.000}$ $0.945_{\pm 0.001}$ $0.943_{\pm 0.004}$
-churn $0.964_{\pm 0.001}$ $0.939_{\pm 0.007}$ $0.882_{\pm 0.006}$ $0.948_{\pm 0.004}$ $0.957_{\pm 0.004}$ $0.961_{\pm 0.001}$ $0.962_{\pm 0.001}$ $0.962_{\pm 0.001}$ $0.959_{\pm 0.003}$
-covertype $0.892_{\pm 0.000}$ $0.860_{\pm 0.001}$ $0.677_{\pm 0.007}$ $0.777_{\pm 0.001}$ $0.876_{\pm 0.000}$ $0.845_{\pm 0.001}$ $0.816_{\pm 0.002}$ $0.828_{\pm 0.001}$ $0.802_{\pm 0.002}$
-default $0.768_{\pm 0.000}$ $0.754_{\pm 0.002}$ $0.744_{\pm 0.002}$ $0.751_{\pm 0.004}$ $0.763_{\pm 0.002}$ $0.739_{\pm 0.008}$ $0.762_{\pm 0.003}$ $0.765_{\pm 0.002}$ $0.765_{\pm 0.002}$
-diabetes $0.693_{\pm 0.001}$ $0.669_{\pm 0.002}$ $0.626_{\pm 0.003}$ $0.592_{\pm 0.002}$ $0.675_{\pm 0.001}$ $0.673_{\pm 0.001}$ $0.667_{\pm 0.001}$
-Table 14: Machine learning efficiency RMSE for five benchmark models, the real training data and for CDTD with three different noise schedules.
+Real Data SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
+adult $0.915_{\pm 0.000}$ $0.906_{\pm 0.001}$ $0.901_{\pm 0.000}$ $0.836_{\pm 0.006}$ $0.889_{\pm 0.002}$ $0.908_{\pm 0.000}$ $0.880_{\pm 0.005}$ $0.906_{\pm 0.001}$ $0.910_{\pm 0.000}$ $0.910_{\pm 0.001}$ $0.909_{\pm 0.000}$
+bank $0.947_{\pm 0.000}$ $0.943_{\pm 0.001}$ $0.938_{\pm 0.001}$ $0.934_{\pm 0.003}$ $0.830_{\pm 0.020}$ $0.940_{\pm 0.005}$ $0.929_{\pm 0.005}$ $0.939_{\pm 0.003}$ $0.945_{\pm 0.000}$ $0.945_{\pm 0.001}$ $0.943_{\pm 0.004}$
+churn $0.964_{\pm 0.001}$ $0.961_{\pm 0.002}$ $0.939_{\pm 0.007}$ $0.882_{\pm 0.006}$ $0.948_{\pm 0.004}$ $0.957_{\pm 0.004}$ $0.961_{\pm 0.001}$ $0.919_{\pm 0.006}$ $0.962_{\pm 0.001}$ $0.962_{\pm 0.001}$ $0.959_{\pm 0.003}$
+covertype $0.892_{\pm 0.000}$ $0.860_{\pm 0.001}$ $0.677_{\pm 0.007}$ $0.777_{\pm 0.001}$ $0.876_{\pm 0.000}$ $0.845_{\pm 0.001}$ $0.671_{\pm 0.013}$ $0.816_{\pm 0.002}$ $0.828_{\pm 0.001}$ $0.802_{\pm 0.002}$
+default $0.768_{\pm 0.000}$ $0.759_{\pm 0.003}$ $0.754_{\pm 0.002}$ $0.744_{\pm 0.002}$ $0.751_{\pm 0.004}$ $0.763_{\pm 0.002}$ $0.739_{\pm 0.008}$ $0.746_{\pm 0.011}$ $0.762_{\pm 0.003}$ $0.765_{\pm 0.002}$ $0.765_{\pm 0.002}$
+diabetes $0.693_{\pm 0.001}$ $0.679_{\pm 0.001}$ $0.669_{\pm 0.002}$ $0.626_{\pm 0.003}$ $0.592_{\pm 0.002}$ $0.645_{\pm 0.002}$ $0.675_{\pm 0.001}$ $0.673_{\pm 0.001}$ $0.667_{\pm 0.001}$
+Table 19: Machine learning efficiency RMSE for seven benchmark models, the real training data and for CDTD with three different noise schedules.
 The standard deviation takes into account five different sampling seeds and uses the average results of the four machine learning efficiency models computed across ten model seeds.
-Real Data ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
-acsincome $0.804_{\pm 0.012}$ $0.757_{\pm 0.007}$ $2.292_{\pm 0.013}$ $1.054_{\pm 0.011}$ $0.857_{\pm 0.010}$ $0.838_{\pm 0.015}$ $0.811_{\pm 0.014}$ $0.820_{\pm 0.011}$
-beijing $0.712_{\pm 0.001}$ $0.792_{\pm 0.007}$ $1.246_{\pm 0.010}$ $1.690_{\pm 0.016}$ $0.606_{\pm 0.006}$ $0.912_{\pm 0.005}$ $0.774_{\pm 0.005}$ $0.770_{\pm 0.005}$ $0.762_{\pm 0.005}$
-lending $0.030_{\pm 0.000}$ $0.274_{\pm 0.007}$ $0.137_{\pm 0.007}$ $0.404_{\pm 0.007}$ $0.795_{\pm 0.031}$ $0.061_{\pm 0.001}$ $0.060_{\pm 0.001}$ $0.066_{\pm 0.002}$
-news $1.001_{\pm 0.002}$ $0.923_{\pm 0.052}$ $1.906_{\pm 0.019}$ $3.999_{\pm 0.175}$ $0.083_{\pm 0.001}$ $1.302_{\pm 0.074}$ $0.819_{\pm 0.103}$ $0.776_{\pm 0.091}$ $0.755_{\pm 0.066}$
-nmes $1.001_{\pm 0.003}$ $0.972_{\pm 0.024}$ $1.331_{\pm 0.052}$ $1.127_{\pm 0.047}$ $1.154_{\pm 0.047}$ $1.137_{\pm 0.052}$ $1.108_{\pm 0.083}$ $1.184_{\pm 0.076}$ $1.203_{\pm 0.081}$
-Appendix O Ablation Study Details
+Real Data SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
+acsincome $0.804_{\pm 0.012}$ $0.757_{\pm 0.007}$ $2.292_{\pm 0.013}$ $1.054_{\pm 0.011}$ $0.857_{\pm 0.010}$ $0.959_{\pm 0.012}$ $0.838_{\pm 0.015}$ $0.811_{\pm 0.014}$ $0.820_{\pm 0.011}$
+beijing $0.712_{\pm 0.001}$ $0.739_{\pm 0.002}$ $0.792_{\pm 0.007}$ $1.246_{\pm 0.010}$ $1.690_{\pm 0.016}$ $0.606_{\pm 0.006}$ $0.912_{\pm 0.005}$ $0.788_{\pm 0.011}$ $0.774_{\pm 0.005}$ $0.770_{\pm 0.005}$ $0.762_{\pm 0.005}$
+lending $0.030_{\pm 0.000}$ $0.042_{\pm 0.001}$ $0.274_{\pm 0.007}$ $0.137_{\pm 0.007}$ $0.404_{\pm 0.007}$ $0.795_{\pm 0.031}$ $0.268_{\pm 0.004}$ $0.061_{\pm 0.001}$ $0.060_{\pm 0.001}$ $0.066_{\pm 0.002}$
+news $1.001_{\pm 0.002}$ $1.180_{\pm 0.107}$ $0.923_{\pm 0.052}$ $1.906_{\pm 0.019}$ $3.999_{\pm 0.175}$ $0.083_{\pm 0.001}$ $1.302_{\pm 0.074}$ $0.374_{\pm 0.028}$ $0.819_{\pm 0.103}$ $0.776_{\pm 0.091}$ $0.755_{\pm 0.066}$
+nmes $1.001_{\pm 0.003}$ $1.112_{\pm 0.044}$ $0.972_{\pm 0.024}$ $1.331_{\pm 0.052}$ $1.127_{\pm 0.047}$ $1.154_{\pm 0.047}$ $1.137_{\pm 0.052}$ $0.535_{\pm 0.013}$ $1.108_{\pm 0.083}$ $1.184_{\pm 0.076}$ $1.203_{\pm 0.081}$
+Appendix S Ablation Study Details
 Table 20: L2 norm (incl.
@@
 beijing $0.712_{\pm 0.001}$ $0.782_{\pm 0.004}$ $0.785_{\pm 0.004}$ $0.778_{\pm 0.005}$ $0.776_{\pm 0.004}$ $0.770_{\pm 0.005}$
-Appendix P Training and Sampling Times Details
+Appendix T Training and Sampling Times Details
 Table 28: Training times in minutes.
@@
 CoDi is considered prohibitively expensive to train on diabetes and lending and we report estimated (est.) training times.
-ARF CTGAN TVAE TabDDPM CoDi CDTD (per feature)
-acsincome 80.3 59.9 26.0 - 231.9 5.8
-adult 7.4 36.2 23.7 38.3 48.3 6.9
-bank 11.0 37.6 24.6 40.5 42.7 26.3
-beijing 3.7 34.3 23.9 36.1 24.9 23.4
-churn 0.3 27.1 13.7 18.2 25.7 6.1
-covertype 130.2 58.0 36.5 44.9 69.2 28.2
-default 12.0 38.3 24.8 38.9 45.9 26.4
-diabetes 58.5 90.1 25.3 - 870 (est.) 26.9
-lending 5.2 157.9 36.6 48.7 3000 (est.) 25.3
-news 23.0 48.8 33.3 37.2 41.5 25.2
-nmes 0.4 32.8 17.2 24.9 30.2 6.3
+SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (per feature)
+acsincome - 80.3 59.9 26.0 - 231.9 13.4 5.8
+adult - 7.4 36.2 23.7 38.3 48.3 32.7 6.9
+bank - 11.0 37.6 24.6 40.5 42.7 48.5 26.3
+beijing - 3.7 34.3 23.9 36.1 24.9 25.8 23.4
+churn - 0.3 27.1 13.7 18.2 25.7 21.5 6.1
+covertype - 130.2 58.0 36.5 44.9 69.2 30.7 28.2
+default - 12.0 38.3 24.8 38.9 45.9 40.1 26.4
+diabetes - 58.5 90.1 25.3 - 870 (est.) 34.6 26.9
+lending - 5.2 157.9 36.6 48.7 3000 (est.) 42.1 25.3
+news - 23.0 48.8 33.3 37.2 41.5 57.9 25.2
+nmes - 0.4 32.8 17.2 24.9 30.2 31.0 6.3
 Table 29: Sample times in seconds per 1000 samples.
-ARF CTGAN TVAE TabDDPM CoDi CDTD (per feature)
-acsincome 4.20 0.23 0.07 - 10.26 0.59
-adult 1.78 0.31 0.16 0.82 3.65 0.56
-bank 2.24 0.44 0.44 0.87 3.38 0.64
-beijing 0.34 0.41 0.32 2.09 2.45 0.26
-churn 1.00 0.40 0.24 0.95 2.78 0.39
-covertype 9.74 0.28 0.25 2.45 4.35 1.97
-default 2.07 0.27 0.25 0.86 3.48 0.60
-diabetes 5.87 0.53 0.15 - - 1.33
-lending 2.49 0.45 0.54 4.33 - 0.69
-news 3.89 0.43 0.30 5.13 2.93 0.85
-nmes 1.54 0.31 0.17 4.17 2.91 0.55
+SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (per feature)
+acsincome 4674.45 4.20 0.23 0.07 - 10.26 3.53 0.59
+adult 10.71 1.78 0.31 0.16 0.82 3.65 0.88 0.56
+bank 16.19 2.24 0.44 0.44 0.87 3.38 0.80 0.64
+beijing 3.98 0.34 0.41 0.32 2.09 2.45 0.99 0.26
+churn 0.52 1.00 0.40 0.24 0.95 2.78 0.80 0.39
+covertype 10913.34 9.74 0.28 0.25 2.45 4.35 0.85 1.97
+default 10.00 2.07 0.27 0.25 0.86 3.48 0.82 0.60
+diabetes 166.75 5.87 0.53 0.15 - - 0.83 1.33
+lending 4.06 2.49 0.45 0.54 4.33 - 0.85 0.69
+news 66.49 3.89 0.43 0.30 5.13 2.93 0.86 0.85
+nmes 0.69 1.54 0.31 0.17 4.17 2.91 0.82 0.55

### v4

@@
 Abstract
-Score-based generative models (or diffusion models for short) have proven successful for generating text and image data.
-However, the adaption of this model family to tabular data of mixed-type has fallen short so far.
-In this paper, we propose CDTD, a Continuous Diffusion model for mixed-type Tabular Data.
-Specifically, we combine score matching and score interpolation to ensure a common continuous noise distribution for both continuous and categorical features alike.
-We counteract the high heterogeneity inherent to data of mixed-type with distinct, adaptive noise schedules per feature or per data type.
-The learnable noise schedules ensure optimally allocated model capacity and balanced generative capability.
-We homogenize the data types further with model-specific loss calibration and initialization schemes tailored to mixed-type tabular data.
-Our experimental results show that CDTD consistently outperforms state-of-the-art benchmark models, captures feature correlations exceptionally well, and that heterogeneity in the noise schedule design boosts the sample quality.
+Score-based generative models, commonly referred to as diffusion models, have proven to be successful at generating text and image data.
+However, their adaptation to mixed-type tabular data remains underexplored.
+In this work, we propose CDTD, a Continuous Diffusion model for mixed-type Tabular Data.
+CDTD is based on a novel combination of score matching and score interpolation to enforce a unified continuous noise distribution for both continuous and categorical features.
+We explicitly acknowledge the necessity of homogenizing distinct data types by relying on model-specific loss calibration and initialization schemes.
+To further address the high heterogeneity in mixed-type tabular data, we introduce adaptive feature- or type-specific noise schedules.
+These ensure balanced generative performance across features and optimize the allocation of model capacity across features and diffusion time.
+Our experimental results show that CDTD consistently outperforms state-of-the-art benchmark models, captures feature correlations exceptionally well, and that heterogeneity in the noise schedule design boosts sample quality.
+Replication code is available at https://github.com/muellermarkus/cdtd.
 1 Introduction
-Score-based generative models (Song et al., 2021), also termed diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020), have shown remarkable potential for the generation of images (Dhariwal & Nichol, 2021; Rombach et al., 2022), videos (Ho et al., 2022), text (Li et al., 2022; Dieleman et al., 2022; Wu et al., 2023), molecules (Hoogeboom et al., 2022), and many other highly complex data structures with continuous features.
+Score-based generative models (Song et al., 2021), also termed diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020), have shown remarkable potential for the generation of images (Dhariwal & Nichol, 2021; Rombach et al., 2022), videos (Ho et al., 2022), text (Li et al., 2022; Dieleman et al., 2022; Wu et al., 2023), molecules (Hoogeboom et al., 2022), and other highly complex data structures with continuous features.
 The framework has since been adapted to categorical data in various ways, including discrete diffusion processes (Austin et al., 2021; Hoogeboom et al., 2021), diffusion in continuous embedding space (Dieleman et al., 2022; Li et al., 2022; Regol & Coates, 2023; Strudel et al., 2022), and others (Campbell et al., 2022; Meng et al., 2022; Sun et al., 2023).
-Diffusion models which include both, continuous and categorical features alike, build directly on advances from the image domain (Kim et al., 2023; Kotelnikov et al., 2023; Lee et al., 2023; Jolicoeur-Martineau et al., 2024) and thus, are not designed to deal with challenges specific to mixed-type data: The different diffusion processes and their losses are neither aligned nor balanced across data types, and do not scale to larger datasets and/or features with a greater number of categories.
-Models that naively combine different losses to integrate distinct generative processes may suffer from implicitly favoring the sample quality of some features or data types over others (Ma et al., 2020).
-A crucial component in score-based generative models is the noise schedule (Kingma et al., 2022; Chen et al., 2022; Chen, 2023; Jabri et al., 2022; Wu et al., 2023).
-Typical noise schedules for image and text data are designed to focus model capacity on the noise levels most important to sample quality (Nichol & Dhariwal, 2021; Karras et al., 2022), while others attempt to learn the optimal noise schedule (Dieleman et al., 2022; Kingma et al., 2022).
-For mixed-type tabular data, existing approaches often combine distinct diffusion processes for the continuous and discrete features to derive a joint model (Kotelnikov et al., 2023; Lee et al., 2023).
-However, noise schedules are not directly transferable from one data modality to another and therefore, using specifications from image or text domain models is not optimal: First, the inherently different diffusion processes make it difficult to balance the noise schedules across features and feature types, and negatively affect the allocation of model capacity across timesteps.
-Further, and most importantly, the domain, nature and marginal distribution can vary significantly across features (Xu et al., 2019).
-For instance, any two continuous features may be subject to different levels of discretization or different bounds, even after applying common data pre-processing techniques; and any two categorical features may differ in the number of categories, or the degree of imbalance.
-The high heterogeneity and lack of balancing warrants a rethinking of fundamental parts of the diffusion framework, including the noise schedule and the effective combination of diffusion processes for different data types.
+However, their adaptation to mixed-type tabular data, which includes both continuous and categorical features, remains under explored.
+Existing models build directly on advances from the image domain (Kim et al., 2023; Kotelnikov et al., 2023; Lee et al., 2023; Jolicoeur-Martineau et al., 2024) and, therefore, are not designed to deal with challenges specific to mixed-type tabular data: The type-specific diffusion processes and their losses are neither aligned nor balanced.
+A naive combination of different losses can cause the generative model to favor the sample quality of some features or data types over others (Ma et al., 2020).
+Furthermore, tabular data often includes categorical features of high cardinality.
+However, existing diffusion models for tabular data typically rely on a discrete diffusion framework to model one-hot-encoded categorical features (e.g., Kotelnikov et al., 2023; Lee et al., 2023).
+As a consequence, these models do not scale well and fail to capture the full uncertainty during the denoising process, as a data sample can never be ‘in-between’ categories.
+Noise schedules determine the amount of noise at each diffusion timestep and are typically defined manually (Nichol & Dhariwal, 2021; Karras et al., 2022), but can also be learned (Dieleman et al., 2022; Kingma et al., 2021).
+They are a crucial component in score-based generative models (Kingma et al., 2021; Chen et al., 2023; Chen, 2023; Jabri et al., 2023; Wu et al., 2023) as they aim to focus the model capacity on the noise levels most important to sample quality.
+However, despite their importance, previous work on diffusion models for tabular data simply adopted noise schedules from the image domain, which is not optimal.
+First, diffusion models for mixed-type tabular data inherently rely on combining type-specific diffusion processes.
+This makes it important, but difficult, to balance noise schedules across feature types.
+Unbalanced noise schedules negatively affect the allocation of model capacity.
+For instance, both TabDDPM (Kotelnikov et al., 2023) and CoDi (Lee et al., 2023) use the discrete multinomial diffusion framework (Hoogeboom et al., 2021) to model categorical features.
+This induces different types of noise for continuous and categorical features, making alignment or comparison of noise schedules impossible.
+Second, the domain, nature, and marginal distribution can vary significantly across features (Xu et al., 2019).
+For instance, any two continuous features may be subject to different levels of discretization or skewness, even after applying common data preprocessing techniques; and any two categorical features may differ in the number of categories, or the degree of imbalance.
+The high heterogeneity and lack of balance warrant a rethinking of fundamental parts of the diffusion framework, including the noise schedule and the effective combination of diffusion processes for different data types, rather than simply copying what has been working for images.
 In this paper, we introduce Continuous Diffusion for mixed-type Tabular Data (CDTD) to address the aforementioned shortcomings.
-We combine score matching (Hyvärinen, 2005) with score interpolation (Dieleman et al., 2022) to derive a score-based model that pushes the diffusion process for categorical data into embedding space, and uses a Gaussian diffusion process for both continuous and categorical features.
-This way, the different noise processes become directly comparable, easier to balance, and enable the application of, for instance, classifier-free guidance (Ho & Salimans, 2022), accelerated sampling (Lu et al., 2022), and other advances, to mixed-type tabular data.
-We counteract the high feature heterogeneity inherent to data of mixed-type with distinct feature or type-specific adaptive noise schedules.
-The learnable noise schedules allow the model to directly take feature or type heterogeneity into account during both training and generation, and thus avoid the reliance on image or text-specific noise schedule designs.
-Moreover, we propose a diffusion-specific loss normalization and initialization scheme to homogenize different data types and their losses effectively.
-Our improvements ensure a better allocation of the model’s capacity across features, feature types and timesteps, and yield high quality samples of tabular data.
-CDTD outperforms state-of-the-art baseline models across a diverse set of sample quality metrics as well as computation time for data sets with an arbitrary number of categories and data points.
+We combine score matching (Hyvärinen, 2005) with score interpolation (Dieleman et al., 2022) to derive a score-based model that pushes the diffusion process for categorical data into embedding space, and thus enables a Gaussian diffusion process for both continuous and categorical features.
+This way, the different noise processes become directly comparable, easier to balance, and enable the application of, for instance, classifier-free guidance (Ho & Salimans, 2022), accelerated sampling (Lu et al., 2022), and other advances to mixed-type tabular data.
+We counteract the high feature heterogeneity inherent to mixed-type tabular data with distinct feature- or type-specific adaptive noise schedules.
+The learnable noise schedules allow the model to directly take feature or type heterogeneity into account during both training and generation, and thus avoid the reliance on image-specific noise schedule designs.
+Likewise, Shi et al.
+(2024) also propose the use of feature-specific noise schedules for tabular data but do not account for the necessity of homogenizing different data types.
+In contrast, we propose a diffusion-specific loss calibration and initialization scheme for effective data type homogenization.
+These improvements ensure a better allocation of our model’s capacity across features, feature types and timesteps, and yield high quality samples.
+CDTD outperforms state-of-the-art baseline models across a diverse set of sample quality metrics and datasets as well as computation time.
 Our experiments show that CDTD captures feature correlations exceptionally well, and that explicitly allowing for data-type heterogeneity in the noise schedules benefits sample quality.
-In sum, we make several contributions specific to diffusion probabilistic modeling of tabular data:
-  • We propose a joint continuous diffusion model for both continuous and categorical features such that all noise distributions are Gaussian.
-  • We balance model capacity across continuous and categorical features with a novel and effective loss calibration, an adjusted score model initialization and type or feature-specific noise schedules.
-  • We extend the idea of timewarping and propose a functional form to efficiently learn adaptive noise schedules, and to allow for exact evaluation and easy incorporation of prior information on the relative importance of noise levels.
-  • We drastically improve the scalability of tabular data diffusion models to features with a high number of categories.
-  • We boost the quality of the generated samples with adaptive, feature or type-specific noise schedules.
-  • Our CDTD model allows the first-ever use of advanced techniques, like classifier-free guidance, for mixed-type tabular data directly in data space.
+In sum, we make several contributions specific to score-based modeling of tabular data:
+  • We propose a unified continuous diffusion model for both continuous and categorical features such that all noise distributions are Gaussian.
+  • We balance model capacity across continuous and categorical features with a novel loss calibration, an adjusted score model initialization and adaptive type- or feature-specific noise schedules.
+  • We suggest a novel functional form to efficiently learn adaptive noise schedules, and to allow for an exact evaluation and incorporation of prior information on the relative importance of noise levels.
+  • We drastically improve the scalability of tabular data diffusion models to high-cardinality features.
+  • We are the first to enable the use of advanced techniques, like classifier-free guidance, for mixed-type tabular data directly in data space as opposed to a latent space.
 Figure 1: CDTD framework.
@@
 2 Score-based Generative Framework
-We start with a brief outline of the score-based frameworks for continuous and categorical features.
-Next, we combine these into a single diffusion model to learn the joint distribution of mixed-type data.
+We start with outlining the score-based frameworks for continuous and categorical features.
+Next, we combine these into a single, unified model to learn the joint distribution of mixed-type tabular data.
 2.1 Continuous Features
@@
 where $\lambda_{t}:[0,1]\rightarrow{\mathbb{R}}_{+}$ is a positive weighting function for timesteps $t\sim\mathcal{U}_{[0,1]}$, and $p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})$ is the density of the noisy ${\mathbf{x}}_{t}$ given the ground-truth data ${\mathbf{x}}_{0}$ (Vincent, 2011).
-In this paper, we use the EDM formulation (Karras et al., 2022), that is, ${\mathbf{f}}(\cdot,t)={\bm{0}}$ and $g(t)=\sqrt{2[\frac{d}{dt}\sigma(t)]\sigma(t)}$ such that $p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})=\mathcal{N}({\mathbf{x}}_{t}|{\mathbf{x}}_{0},\sigma^{2}(t)I_{K_{\text{cont}}})$.
-We start the reverse process with sampling ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}})=\mathcal{N}({\bm{0}},\sigma^{2}(1)I_{K_{\text{cont}}})$ for $\sigma^{2}(1)$ being sufficiently large and $\mathbb{E}[{\mathbf{x}}_{0}]={\bm{0}}$.
+In this paper, we use the EDM formulation (Karras et al., 2022), that is, ${\mathbf{f}}(\cdot,t)={\bm{0}}$ and $g(t)=\sqrt{2[\frac{\mathrm{d}}{\mathrm{d}t}\sigma(t)]\sigma(t)}$ such that $p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})=\mathcal{N}({\mathbf{x}}_{t}|{\mathbf{x}}_{0},\sigma^{2}(t){\mathbf{I}}_{K_{\text{cont}}})$.
+We standardize ${\mathbf{x}}_{0}$ to zero mean and unit variance.
+Then, for a sufficiently large $\sigma^{2}(1)$, we can start the reverse process with sampling ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}})=\mathcal{N}({\bm{0}},\sigma^{2}(1){\mathbf{I}}_{K_{\text{cont}}})$.
 We then gradually guide ${\mathbf{x}}_{1}$ towards high density regions in the data space with ${\mathbf{s}}_{\bm{\theta}}({\mathbf{x}},t)$ replacing the unknown, true score function in Equation 2.
-In practice, ODE or predictor-corrector samplers can be used for this iterative denoising process (Song et al., 2021).
 2.2 Categorical Features
-For categorical features, the score function is undefined and the score-based generative framework cannot be directly applied.
-Therefore, we use score interpolation (Dieleman et al., 2022) and push the diffusion of categorical data into Euclidean embedding space.
-This allows us to impose a Gaussian noise distribution on categorical and continuous features.
-Let $x_{\text{cat}}^{(j)}\in\{1,\dots,C_{j}\}$ be the $j$-th categorical feature with $C_{j}$ possible classes.
-The $i$-th class is associated to a distinct, trainable $d$-dimensional embedding vector ${\mathbf{e}}_{i}^{(j)}\in{\mathbb{R}}^{d}$.
-The embedding vector corresponding to the ground truth class is ${\mathbf{x}}_{0}^{(j)}\in\{{\mathbf{e}}_{1}^{(j)},\dots,{\mathbf{e}}_{C_{j}}^{(j)}\}$.
-We omit the superscript $j$ in what follows.
-The noisy variant at time $t$ is ${\mathbf{x}}_{t}\sim p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})=\mathcal{N}({\mathbf{x}}_{t}|{\mathbf{x}}_{0},\sigma^{2}(t)I_{d})$ similar to before.
-Given ${\mathbf{x}}_{t}$ and $t$, $\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}[\nabla_{{\mathbf{x}}_{t}}\log p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})]$ minimizes Equation 3, and we obtain
-$\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}\nabla_{{\mathbf{x}}_{t}}\log p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})=\frac{1}{\sigma^{2}(t)}[\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}[{\mathbf{x}}_{0}]-{\mathbf{x}}_{t}]\ .$ (4)
-We compute $\hat{{\mathbf{x}}}_{0}=\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}[{\mathbf{x}}_{0}]$ as an average over the $C_{j}$ possible embedding vectors weighted by their probability.
-This yields a score function estimate for a single categorical feature.
-Since $p({\mathbf{x}}_{0}={\mathbf{e}}_{i}|{\mathbf{x}}_{t},t)=p(x_{\text{cat}}^{(j)}=i|{\mathbf{x}}_{t},t)$, an estimate of $p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)$ is obtained via a classifier that predicts $C_{j}$ class probabilities for the $j$-th feature and is trained via cross-entropy (CE) loss.
-We start generating a categorical feature from an embedding vector ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}})=\mathcal{N}({\bm{0}},\sigma^{2}(1)I_{d})$.
-We then use score interpolation (Equation 4) to iteratively denoise this embedding vector.
-After the final reverse step, i.e., when  $t\approx 0$, we can directly infer the generated class from the feature-specific class probabilities, $p({\mathbf{x}}_{0}={\mathbf{e}}_{i}|{\mathbf{x}}_{t},t)$.
-This framework can easily be extended to multiple categorical features.
-All embeddings are trained alongside the model.
-The model fully takes feature-specific uncertainty at intermediate timesteps into account (as for each feature it only has to commit to a category at the final step of generation) which improves the consistency of the generated samples (Dieleman et al., 2022).
-Therefore, CDTD allows to capture subtle dependencies both within and across data types more accurately.
+Let $x_{\text{cat}}^{(j)}\in\{1,\dots,C_{j}\}$ be the $j$-th categorical feature with $C_{j}$ distinct classes.
+We learn a feature-specific encoder to represent each category $c$ as a $d$-dimensional vector ${\mathbf{e}}_{x_{\text{cat}}}^{(j)}=\text{Enc}_{j}(x_{\text{cat}}^{(j)})$.
+Further, let ${\mathbf{x}}_{0}^{(j)}\in\{{\mathbf{e}}_{1}^{(j)},\dots,{\mathbf{e}}_{C_{j}}^{(j)}\}$ be the noiseless embedding at $t=0$.
+To unify the diffusion frameworks for categorical and continuous data as much as possible, we base both on the same Gaussian-type noise.
+Thus, instead of adding noise to the categorical variable directly, we add noise to the embedding ${\mathbf{x}}^{(j)}_{t}\sim p_{0t}({\mathbf{x}}_{t}^{(j)}|{\mathbf{x}}^{(j)}_{0})=\mathcal{N}({\mathbf{x}}^{(j)}_{t}|{\mathbf{x}}^{(j)}_{0},\sigma^{2}(t)I_{d})$ such that ${\mathbf{x}}^{(j)}_{1}\sim p_{1}({\mathbf{x}}^{(j)})=\mathcal{N}({\bm{0}},\sigma^{2}(1)I_{d})$, analogous to score matching.
+For categorical data, denoising score matching (Equation 3) is not directly applicable to learn $\nabla_{{\mathbf{x}}^{(j)}_{t}}\log p_{0t}({\mathbf{x}}^{(j)}_{t}|{\mathbf{x}}^{(j)}_{0})$, since the score can only take on $C_{j}$ distinct values.
+To proceed, we transform the score matching approach into a discrete choice problem.
+Note that for a given $t$ and ${\mathbf{x}}^{(j)}_{t}$ it is sufficient to find $\mathbb{E}_{\mathrm{P}({\mathbf{x}}^{(j)}_{0}|{\mathbf{x}}^{(j)}_{t},t)}[\nabla_{{\mathbf{x}}^{(j)}_{t}}\log p_{0t}({\mathbf{x}}^{(j)}_{t}|{\mathbf{x}}^{(j)}_{0})]$ as it minimizes Equation 3.
+Presuming Gaussian noise, we have
+$\mathbb{E}_{\mathrm{P}({\mathbf{x}}^{(j)}_{0}|{\mathbf{x}}^{(j)}_{t},t)}\Bigl{[}\nabla_{{\mathbf{x}}^{(j)}_{t}}\log p_{0t}({\mathbf{x}}^{(j)}_{t}|{\mathbf{x}}^{(j)}_{0})\Bigr{]}=\frac{1}{\sigma^{2}(t)}\Bigl{[}\mathbb{E}_{\mathrm{P}({\mathbf{x}}^{(j)}_{0}|{\mathbf{x}}^{(j)}_{t},t)}[{\mathbf{x}}^{(j)}_{0}]-{\mathbf{x}}^{(j)}_{t}\Bigr{]}\ .$ (4)
+We can thus approximate the score by computing the probability-weighted average of the $C_{j}$ possible embedding vectors, that is, $\hat{{\mathbf{x}}}^{(j)}_{0}=\mathbb{E}_{\mathrm{P}({\mathbf{x}}^{(j)}_{0}|{\mathbf{x}}^{(j)}_{t},t)}[{\mathbf{x}}^{(j)}_{0}]$.
+Since $\mathrm{P}({\mathbf{x}}^{(j)}_{0}={\mathbf{e}}^{(j)}_{x_{\text{cat}}}|{\mathbf{x}}^{(j)}_{t},t)=\mathrm{P}(x_{\text{cat}}^{(j)}=c|{\mathbf{x}}^{(j)}_{t},t)$, we can estimate $\mathrm{P}({\mathbf{x}}^{(j)}_{0}|{\mathbf{x}}^{(j)}_{t},t)$ via a classifier that predicts the $C_{j}$ class probabilities and is trained to minimize the cross-entropy (CE).
+This procedure effectively interpolates between the $C_{j}$ ground-truth embeddings ${\mathbf{x}}_{0}^{(j)}$ and is therefore known as score interpolation (Dieleman et al., 2022).
+This framework easily extends to multiple features.
+Most importantly, $\text{Enc}_{j}$ is trained alongside the model such that ${\mathbf{x}}_{0}^{(j)}$ is directly optimized for denoising the data.
+During sampling, the model only has to commit to a category at the final step of generation, i.e., we allow for a smooth, continuous transition between states.
+This is unlike multinomial diffusion (Hoogeboom et al., 2021), which imposes discrete transitioning steps and is the framework used by several existing diffusion models for tabular data (Kotelnikov et al., 2023; Lee et al., 2023).
+By defining diffusion for categorical data in embedding space, we allow our model to take uncertainty at intermediate timesteps fully into account, which improves the consistency of the generated samples (Dieleman et al., 2022).
+Thus, CDTD can more accurately capture subtile dependencies both within and across data types.
 3 Method
-In short, we combine score matching (Equation 3) with score interpolation (Equation 4) to model the joint distribution of mixed-type data.
+To model the joint distribution of mixed-type data, we combine score matching (Equation 3) with score interpolation (Equation 4).
 Next, we discuss the important components of our method.
@@
 Figure 1 gives an overview of our Continuous Diffusion for mixed-type Tabular Data (CDTD) framework.
-The score model is conditioned on (1) all noisy continuous features, (2) the noisy embeddings of all categorical features in Euclidean space, and (3) the timestep $t$ which reflects potentially feature-specific, adaptive noise levels $\sigma_{\text{cont},i}$ and $\sigma_{\text{cat},j}$ for all $i$ and $j$.
-Additional conditioning information, such as the target feature for classification tasks, are straightforward to add.
-Note that while the Gaussian noise process acts directly on the continuous features, it acts on the embeddings of the categorical features.
-This way, we ensure a common continuous noise process for both data types.
-During training, the model predicts the ground-truth value for continuous features and the class-specific probabilities for categorical features.
-During generation, we concatenate the score estimates, $\hat{s}_{\text{cont}}^{(i)}$ and $\hat{s}_{\text{cat}}^{(j)}$, for all features $i$ and $j$, and pass them to an ODE solver together with $\sigma_{\text{cont},i}$ and $\sigma_{\text{cat},j}$, the noise levels retrieved by transforming linearly spaced timesteps with the learned adaptive noise schedules.
-Further details on the implementation and sampling are provided in Appendix J and Appendix K, respectively.
+The score model is conditioned on (1) the noisy continuous features, (2) the noisy embeddings of categorical features, and (3) the timestep $t$.
+It predicts the ground-truth value $x_{\text{cont}}^{(i)}$ for continuous features and the class-specific probabilities $\mathrm{P}(x_{\text{cat}}^{(j)}=c)$ for categorical features.
+Additional conditioning information is straightforward to add.
+Note that while the Gaussian noise process acts directly on continuous features, it acts on the embedded categorical features ${\mathbf{x}}_{0}^{(j)}$.
+This way, we ensure a unified continuous noise process for both data types.
+Further details on the implementation are provided in Appendix K.
+During generation, we concatenate the score estimates, $\hat{s}_{\text{cont}}^{(i)}$ and $\hat{s}_{\text{cat}}^{(j)}$, for all features $i$ and $j$ before passing them to an ODE solver.
+Our model is the first to utilize feature- or type-specific noise schedules also during the generation of tabular data, which allows the sampler to take feature- or type-specific steps.
+Details on the sampling process and the algorithm are given in Appendix L.
 3.2 Homogenization of Data Types
-Let $\mathcal{L}_{\text{MSE}}(x_{\text{cont}}^{(i)},t)$ denote the time-weighted MSE (i.e., score matching) loss of the $i$-th continuous feature at a single timestep $t$, and $\mathcal{L}_{\text{CE}}(x_{\text{cat}}^{(j)},t)$ the CE (i.e., score interpolation) loss of the $j$-th categorical feature.
+Let $\mathcal{L}_{\text{MSE}}(x_{\text{cont}}^{(i)},t)$ be the time-weighted MSE (i.e., score matching) loss of the $i$-th continuous feature at a single timestep $t$, and $\mathcal{L}_{\text{CE}}(x_{\text{cat}}^{(j)},t)$ the CE (i.e., score interpolation) loss of the $j$-th categorical feature.
 Naturally, the two losses are defined on different scales.
 This leads to an unintended importance weighting of features in the generative process (Ma et al., 2020).
-We assume that an unconditional model should a priori, i.e., without having any information, be indifferent between all features.
-This reflects the state of the model at the terminal timestep $t=1$ in the diffusion process.
-Formally, we aim to find calibrated losses, $\mathcal{L}^{\ast}_{\text{MSE}}$ and $\mathcal{L}^{\ast}_{\text{CE}}$ for all continuous features $i$ and categorical features $j$, such that
-$\mathbb{E}[\mathcal{L}^{\ast}_{\text{MSE}}(x_{\text{cont}}^{(i)},1)]=\mathbb{E}[\mathcal{L}^{\ast}_{\text{CE}}(x_{\text{cat}}^{(j)},1)]=1.$ (5)
-For continuous features, $\mathbb{E}[\mathcal{L}^{\ast}_{\text{MSE}}(x_{\text{cont}}^{(i)},1)]=1$ follows from standardizing $x_{\text{cont}}^{(i)}$ to zero mean and unit variance.
-For categorical features, we compute the normalization constant $\mathbb{E}[\mathcal{L}_{\text{CE}}(x_{\text{cat}}^{(j)},1)]$ directly as the CE of each predicted class in proportion to its empirical distribution in the train set (see Appendix A).
-We then average the calibrated losses to derive the joint loss function at a given timestep:
-$\mathcal{L}(t)=\frac{1}{K}\Bigl{[}\sum_{i=1}^{K_{\text{cont}}}\mathcal{L}^{\ast}_{\text{MSE}}(x_{\text{cont}}^{(i)},t)+\sum_{j=1}^{K_{\text{cat}}}\mathcal{L}_{\text{CE}}^{\ast}(x_{\text{cat}}^{(j)},t)\Bigr{]},$ (6)
-where $K=K_{\text{cont}}+K_{\text{cat}}$.
-The loss calibration and the multiple data modalities have implications for the optimal initialization of the score model.
-We aim to initialize all feature-specific losses at one.
-We therefore initialize the output layer weights to zero (like in image diffusion models) and the output biases for continuous features to zero, and rely on the timestep weights of the EDM parameterization (Karras et al., 2022) to achieve a unit loss for all $t$.
-For the categorical features, we initialize the biases to match the category’s empirical probability in the training set (see Appendix B).
-The initial equal importance across all timesteps will naturally change over the course of training.
-We employ a normalization scheme for the average diffusion loss (Karras et al., 2023; Kingma & Gao, 2023) to allow for changes in relative importance among features but ensure equal importance of all timesteps throughout training.
-To do so, we learn the time-specific normalization term $Z(t)$ such that $\mathcal{L}(t)/Z(t)\approx 1$.
-This ensures a consistent gradient signal and can be implemented by training a neural network to predict $\mathcal{L}(t)$ alongside our diffusion model (for details see Appendix C).
+To solve this, we observe that an unconditional generative model should a priori, i.e., without having any information, be indifferent between all features.
+For diffusion models this reflects the state of the model at the terminal timestep $t=1$.
+Definition 1 (Calibrated losses).
+Scaled losses $\mathcal{L}^{\ast}_{\text{MSE}}$ and $\mathcal{L}^{\ast}_{\text{CE}}$ are called calibrated if
+$\mathbb{E}[\mathcal{L}^{\ast}_{\text{MSE}}(x_{\text{cont}}^{(i)},1)]=\mathbb{E}[\mathcal{L}^{\ast}_{\text{CE}}(x_{\text{cat}}^{(j)},1)]=1,$
+for all continuous features $i$ and categorical features $j$.
+Assumption 1 (No information at $t=1$).
+The noise level $\sigma(1)$ is high enough such that the input to the score model cannot be distinguished from noise, i.e., the model has no information.
+Proposition 1 (Homogenization of feature-specific losses).
+Under Assumption 1, if each $x_{\text{cont}}^{(i)}$ has unit variance and $Z_{j}$ is the feature-specific entropy of categorical feature $j$, then the losses $\mathcal{L}^{\ast}_{\text{MSE}}=\mathcal{L}_{\text{MSE}}$ and $\mathcal{L}^{\ast}_{\text{CE}}=\mathcal{L}_{\text{CE}}(x_{\text{cat}}^{(j)},1)/Z_{j}$ are calibrated for all continuous features $i$ and categorical features $j$.
+See Appendix B for a detailed proof.
+Given Assumption 1 and Proposition 1, we can derive a joint loss function without unintended importance weighting by averaging the $K=K_{\text{cont}}+K_{\text{cat}}$ calibrated losses at a given $t$:
+$\mathcal{L}(t)=\frac{1}{K}\Bigl{[}\sum_{i=1}^{K_{\text{cont}}}\mathcal{L}^{\ast}_{\text{MSE}}(x_{\text{cont}}^{(i)},t)+\sum_{j=1}^{K_{\text{cat}}}\mathcal{L}_{\text{CE}}^{\ast}(x_{\text{cat}}^{(j)},t)\Bigr{]}.$ (5)
+Implications for the score model initialization.
+The loss calibration and the multi-modality of the data have implications for the optimal initialization of the score model.
+To match the loss calibration, we aim to initialize the model such that all feature-specific losses are one.
+We therefore initialize the output layer weights and biases for continuous features to zero and rely on the timestep weights of the EDM parameterization (Karras et al., 2022) to achieve a unit loss for all $t$.
+For the categorical features, we initialize the biases to match each category’s empirical log probability in the training set (see Appendix C for details).
+Weighting across time.
+The initial equal importance across all $t$ will change over the course of training.
+To allow for changes in the relative importance among features but ensure equal importance of all timesteps throughout training, we employ a normalization scheme for the average diffusion loss (Karras et al., 2024; Kingma & Gao, 2023).
+Specifically, we learn the time-dependent normalization $Z(t)$ such that $\mathcal{L}(t)/Z(t)\approx 1$.
+This ensures a consistent gradient signal and can be implemented by training a neural network to predict $\mathcal{L}(t)$ alongside our diffusion model (see Appendix D for details).
 3.3 Noise Schedules
-Since the optimal noise schedule of one feature impacts the noise schedules of other features, and different data types have different sensitivities to additive noise, we introduce feature-specific or type-specific noise schedules.
-For instance, given the same embedding dimension, more noise is needed to remove the same amount of signal from embeddings of features with fewer classes.
+Evidently, the noise schedule of one feature impacts the optimality of noise schedules for other features, and different data types have different sensitivities to additive noise.
+For instance, given the same embedding dimension, more noise may be needed to remove the same amount of signal from embeddings of features with fewer classes (see Appendix A).
 Likewise, a delayed noise schedule for one feature might improve sample quality as the model can rely on other correlated features that have been (partially) generated first.
+Therefore, we introduce feature-specific or type-specific noise schedules.
 We make the noise schedules learnable, and therewith adaptive to avoid the reliance on designs for other data modalities.
@@
 Feature-specific noise schedules.
-According to Equation 1, and following the EDM parameterization (Karras et al., 2022), we define the diffusion process of the $i$-th continuous feature as
-$\mathrm{d}x_{{}_{\text{cont}}}^{(i)}=\sqrt{2\Bigl{[}\frac{d}{dt}h_{\text{cont},i}(t)\Bigr{]}h_{\text{cont},i}(t)}\mathrm{d}w_{t}^{(i)},$ (7)
-and likewise the trajectory of the $j$-th categorical feature as
-$\mathrm{d}{\mathbf{x}}_{\text{cat}}^{(j)}=\sqrt{2\Bigl{[}\frac{d}{dt}h_{\text{cat},j}(t)\Bigr{]}h_{\text{cat},j}(t)}\mathrm{d}{\mathbf{w}}_{t}^{(j)},$ (8)
-where ${\mathbf{x}}_{\text{cat}}^{(j)}$ is the $d$-dimensional embedding of $x_{\text{cat}}^{(j)}$ in Euclidean space.
-The feature-specific noise schedules $h_{\text{cont},i}(t)$ and $h_{\text{cat},j}(t)$ represent the standard deviations of the added Gaussian noise such that $\sigma_{\text{cont},i}(t)=h_{\text{cont},i}(t)$ and $\sigma_{\text{cat},j}(t)=h_{\text{cat},j}(t)$.
-Thus, each continuous feature and each embedded categorical feature is affected by a distinct noise schedule.
+Following Equation 1, we define the diffusion process of $x_{{}_{\text{cont}}}^{(i)}$ as
+$\mathrm{d}x_{{}_{\text{cont}}}^{(i)}=\sqrt{2\Bigl{[}\frac{\mathrm{d}}{\mathrm{d}t}h_{\text{cont},i}(t)\Bigr{]}h_{\text{cont},i}(t)}\mathrm{d}w_{t}^{(i)},$ (6)
+and likewise the trajectory of ${\mathbf{x}}_{\text{cat}}^{(j)}$ as
+$\mathrm{d}{\mathbf{x}}_{\text{cat}}^{(j)}=\sqrt{2\Bigl{[}\frac{\mathrm{d}}{\mathrm{d}t}h_{\text{cat},j}(t)\Bigr{]}h_{\text{cat},j}(t)}\mathrm{d}{\mathbf{w}}_{t}^{(j)},$ (7)
+where ${\mathbf{x}}_{\text{cat}}^{(j)}\in{\mathbb{R}}^{d}$ is the embedding of $x_{\text{cat}}^{(j)}$ in Euclidean space.
+The noise schedules $h_{\text{cont},i}(t)$ and $h_{\text{cat},j}(t)$ represent the feature-specific standard deviations $\sigma_{\text{cont},i}(t)$ and $\sigma_{\text{cat},j}(t)$ of the added Gaussian noise, respectively.
+Thus, each feature is affected by a distinct adaptive noise schedule.
+On the other hand, type-specific noise schedules involve only two functions, $h_{\text{cat}}(t)$ and $h_{\text{cont}}(t)$, such that all features of the same type are affected by the same noise schedule.
+Figure 2: (Left) pdf ($f_{\mathrm{d.a.log},k}$) and cdf ($F_{\mathrm{d.a.log},k}$) of the domain-adapted Logistic distribution for five different values of the location parameter $\mu_{k}$ and for a given curve steepness $\nu_{k}=3$.
+(Right) impact of uniform vs.
+adjusted timewarping initialization on the pdf ($f_{\mathrm{d.a.log},k}$) and the cdf ($F_{\mathrm{d.a.log},k}$).
 Adaptive noise schedules.
-Based on Dieleman et al.
-(2022), we aim to learn a noise schedule $h_{k}:t\mapsto\sigma$ for all $K=K_{\text{cont}}+K_{\text{cat}}$ features.
-Note that $t\in[0,1]$, and with pre-specified minimum and maximum noise levels, we can scale $\sigma_{k}$ to lie in $[0,1]$ as well, without loss of generality.
-We will learn the feature-specific loss given the noise level, $F_{k}:\sigma_{k}\mapsto\ell_{k}$, alongside the score model, with $\ell_{k}$ the relevant (not explicitly weighted) training loss for the $k$-th feature.
-Then, our mapping of interest is $h_{k}=\tilde{F}_{k}^{-1}$, that is, the normalized and inverted function $F_{k}$.
+We aim to learn a noise schedule $h_{k}:t\mapsto\sigma_{k}$ for each feature $k$.
+Inspired by Dieleman et al.
+(2022), we learn a function $F_{k}$ that predicts the feature-specific (not explicitly weighted) loss $\ell_{k}$ given the noise level $\sigma_{k}$.
+By normalizing and inverting $F_{k}$, we achieve the mapping of interest $h_{k}=\tilde{F}_{k}^{-1}$.
 This encourages the relation between $t$ and $\ell_{k}$ to be linear.
-Higher noise levels imply a lower signal-to-noise ratio, and therefore a larger incurred loss for the score model.
+Higher noise levels imply a lower signal-to-noise ratio and therefore a larger incurred loss for the score model.
 Accordingly, $F_{k}$ must be a monotonically increasing and S-shaped function.
-We let $F_{k}=\gamma_{k}F_{\mathrm{d.a.log},k}(\sigma_{k})$ where $\gamma_{k}>0$ is a scaling factor that at $t=1$ enables fitting a loss $\ell_{k}>1$ early on in the training process, and a loss $\ell_{k}<1$ in case conditioning information is included.
+We let $F_{k}=\gamma_{k}F_{\mathrm{d.a.log},k}(\sigma_{k})$ where $\gamma_{k}>0$ is a scaling factor that enables fitting a loss $\ell_{k}>1$ at $t=1$ near the start of training, and a loss $\ell_{k}<1$ in case conditioning information is included.
 Further, we use the cdf of the domain-adapted Logistic distribution $F_{\mathrm{d.a.log},k}(\sigma_{k})$, where the input is pre-processed via a Logit function, with parameters $10<\mu_{k}<$ 1 (the location of the inflection point) and $\nu_{k}\geq 1$ (the steepness of the curve).
+Note that with pre-specified minimum and maximum noise levels, we can always scale $\sigma_{k}$ to lie in $[0,1]$.
 Figure 2 illustrates the effect of the location parameter.
-The implicit importance of the noise levels is conveniently represented by the corresponding pdf $f_{\mathrm{d.a.log},k}$.
-To normalize and invert $F_{k}$, we set $\gamma_{k}=1$ and and directly utilize the quantile function $F_{\mathrm{d.a.log},k}^{-1}$.
-The detailed derivation of all relevant functions is given in Appendix D.
+The implicit importance of the noise levels is conveniently represented by the pdf $f_{\mathrm{d.a.log},k}$.
+To normalize and invert $F_{k}$, we set $\gamma_{k}=1$ and directly use the closed-form quantile function $F_{\mathrm{d.a.log},k}^{-1}$.
+The detailed derivation of all relevant functions is given in Appendix E.
+To avoid biasing the noise schedule to frequently sampled timesteps during training, we derive importance weights from $f_{\mathrm{d.a.log},k}$ when fitting $h_{k}$.
+We use the adaptive noise schedules during both training and generation, and give examples of learned noise schedules in Appendix P.
 Our functional choice has several advantages.
@@
 For instance, for $\mu_{1}<\mu_{2}$, the model starts generating feature 2 before feature 1 in the reverse process.
-Third, the proposed functional form is less flexible than the original piece-wise linear function (Dieleman et al., 2022) such that an exponential moving average on the parameters is not necessary, and the fit is more robust to “outliers” encountered during training.
-We use the adaptive noise schedules during both training and generation.
-We derive importance weights from $f_{\mathrm{d.a.log},k}$ to fit $h_{k}$ to avoid biasing the noise schedule to timesteps that are frequently sampled during training.
-Type-specific noise schedules refer to learning two functions $F_{1}$ and $F_{2}$ that predict the respective average loss over all features of a data type.
-Examples of learned noise schedules are given in Appendix O.
+Third, the proposed functional form is less flexible than the original piece-wise linear function (Dieleman et al., 2022) such that an EMA on the parameters is not necessary, and the fit is more robust to “outliers” encountered during training.
+This is crucial when using a feature-specific specification as opposed to a single noise schedule.
 3.4 Additional Customization to Tabular Data
 In the diffusion process, we add noise directly to the continuous features but to the embeddings of categorical features.
-We generally need more noise to remove all signal from the categorical representations.
-We therefore define type-specific minimum and maximum noise levels: For categorical features, we let $\sigma_{\text{cat},\min}=0.1$ and $\sigma_{\text{cat},\max}=100$; for continuous features, we set $\sigma_{\text{cont},\min}=0.002$ and $\sigma_{\text{cont},\max}=80$ (see Karras et al., 2022).
-Figure 2: (Left) pdf ($f_{\mathrm{d.a.log},k}$) and cdf ($F_{\mathrm{d.a.log},k}$) of the domain-adapted Logistic distribution for five different values of the location parameter $\mu_{k}$ and for a given curve steepness $\nu_{k}=3$.
-(Right) impact of uniform vs.
-adjusted timewarping initialization on the pdf ($f_{\mathrm{d.a.log},k}$) and the cdf ($F_{\mathrm{d.a.log},k}$).
-Lastly, an uninformative initialization of the adaptive noise schedules requires to set $\mu_{k}=0.5,\nu_{k}\approx 1$ and $\gamma_{k}=1$ such that $F_{\mathrm{d.a.log},k}$ corresponds approximately to the cdf of a uniform distribution.
-We can improve this with a more informative prior: In the image domain, diffusion models allocate substantial capacity towards generating the high level structure before generating details at lower noise levels.
-In tabular data, the location of features in the data matrix, and therefore the high level structure, is fixed.
-Instead, we are interested in generating details as accurately as possible, as these influence, for instance, subtle correlations among features.
-Note that the inflection point, $\mu_{k}$, of our adaptive noise schedule corresponds to the proportion of high (normalized) noise levels (i.e., $\sigma_{k}\geq 0.5$) in the distribution.
-Therefore, we adjust the initial noise schedules such that low noise levels ($\sigma_{k}<0.5$) are weighted by a factor of 3 relative to high noise levels ($\sigma_{k}\geq 0.5$) (see Figure 2).
-The proportion of high noise levels is decreased to $\mu_{k}=1/4$.
+We generally need more noise to remove all the signal from the categorical representations.
+We therefore define type-specific minimum and maximum noise levels: For categorical features, we let $\sigma_{\text{cat},\min}=0$ and $\sigma_{\text{cat},\max}=100$; for continuous features, we set $\sigma_{\text{cont},\min}=0$ and $\sigma_{\text{cont},\max}=80$.
+Lastly, an uninformative initialization of the adaptive noise schedules requires $\mu_{k}=0.5,\nu_{k}\approx 1$ and $\gamma_{k}=1$ such that $F_{\mathrm{d.a.log},k}$ corresponds approximately to the cdf of a uniform distribution.
+We can improve upon this with a more informative prior: As opposed to images, in tabular data the location of features in the data matrix, and therefore the high-level structure, is fixed.
+Instead, we are interested in generating details as accurately as possible.
+Note that the inflection point, $\mu_{k}$, corresponds to the proportion of high noise levels (i.e., $\sigma_{k}\geq 0.5\cdot(\sigma_{\max}+\sigma_{\min})$) in the distribution.
+Therefore, we empirically choose $\mu_{k}=1/4$ such that low noise levels (i.e., $\sigma_{k}<0.5\cdot(\sigma_{\max}+\sigma_{\min})$) are initially presumed to be three times more important than high noise levels (see Figure 2).
 We let $\nu_{k}\approx 1$ for a dispersed initial probability mass and initialize the scaling factor to $\gamma_{k}=1$.
@@
 We use a diverse benchmark set of state-of-the-art generative models for mixed-type tabular data.
-This includes SMOTE (Chawla et al., 2002), ARF (Watson et al., 2023), CTGAN (Xu et al., 2019), TVAE (Xu et al., 2019), TabDDPM (Kotelnikov et al., 2023), CoDi (Lee et al., 2023), TabSyn (Zhang et al., 2024).
+This includes SMOTE (Chawla et al., 2002), ARF (Watson et al., 2023), CTGAN (Xu et al., 2019), TVAE (Xu et al., 2019), TabDDPM (Kotelnikov et al., 2023), CoDi (Lee et al., 2023) and TabSyn (Zhang et al., 2024).
 Each model follows a different design and/or modeling philosophy.
@@
 [^1] Jolicoeur-Martineau et al.
-(2024) report in the appendix that they used 10-20 CPUs with 64-256 GB of memory for datasets with a median number of 540 observations.
+(2024) report that they used 10-20 CPUs with 64-256 GB of memory for datasets with a median number of 540 observations.
 With the suggested hyperparameters (for improved efficiency) and 64 CPUs, the model took approx.
-500 min of training on the relatively small nmes data.
+500 min of training on the small nmes data.
 Note that the model estimates $KT$ separate models, with $K$ being the number of features and $T$ the noise levels.
 Therefore, we consider ForestDiffusion to be prohibitively expensive for higher-dimensional data generation.
-Further details on the respective benchmark models and their implementations are provided in Appendix F and Appendix G.
-We provide an in-depth comparison of CDTD to the diffusion-based baselines in Appendix N.
-To keep the comparison fair, we use the same architecture for CDTD as TabDDPM (the latter has also been adopted by TabSyn), with minor changes to accommodate the different inputs (see Appendix J).
+Further details on the respective benchmark models and their implementations are provided in Appendix G and Appendix H, respectively.
+We provide an in-depth comparison of CDTD to the diffusion-based baselines in Appendix O.
+To keep the comparison fair, we use the same internal architecture for CDTD as TabDDPM (which was also adopted by TabSyn), with minor changes to accommodate the different inputs (see Appendix K).
 Datasets.
-We systematically investigate our model on eleven publicly available datasets.
+We systematically investigate our model on 10 publicly available datasets (see Appendix F for details).
 The datasets vary in size, prediction task (regression vs.
@@
 However, CDTD trivially extends to targets with multiple classes.), number of continuous and categorical features and their distributions.
-The number of categories for categorical features varies significantly across datasets (for more details, see Appendix E).
+The number of categories for categorical features varies significantly across datasets.
 We remove observations with missings in the target or any of the continuous features and encode missings in the categorical features as a separate category.
 All datasets are split in train (60%), validation (20%) and test (20%) partitions, hereinafter denoted $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{valid}}$ and $\mathcal{D}_{\text{test}}$, respectively.
-For classification tasks, we use stratification with respect to the outcome, we condition the model on the outcome during training and generation, and use the train set proportions for generation.
-In a last post-processing step, we round the integer-valued continuous features after generation for all models.
+For classification tasks, we use stratification with respect to the outcome.
+We round the integer-valued continuous features after generation.
 4.1 Evaluation Metrics
 In our experiments, we follow conventions from previous papers and use four sample quality criteria, which we assess using a comprehensive set of measures.
-All metrics are averaged over five random seeds that affect the generative process, which samples synthetic data $\mathcal{D}_{\text{gen}}$ of size $\min(\lvert\mathcal{D}_{\text{train}}\rvert,50\,000)$.
+All metrics are averaged over five random seeds that affect the sampling process of synthetic data $\mathcal{D}_{\text{gen}}$ of size $\min(\lvert\mathcal{D}_{\text{train}}\rvert,50\,000)$.
 Machine learning efficiency.
 We follow the conventional train-synthetic-test-real strategy (see, Borisov et al., 2023; Liu et al., 2023; Kotelnikov et al., 2023; Kim et al., 2023; Xu et al., 2019; Watson et al., 2023).
-Hence, we train a group of models, consisting of a (logistic/ridge) regression, a random forest and a catboost model, on the data-specific prediction task (the corresponding hyperparameter settings are reported in Appendix I).
-We compare the model-averaged real test performance, $\text{Perf}(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})$, to the performance when trained on the synthetic data, $\text{Perf}(\mathcal{D}_{\text{gen}},\mathcal{D}_{\text{test}})$.
-We subsample $\mathcal{D}_{\text{train}}$ in case of more than 50 000 observations to upper-bound the computational load.
-The results are averaged over ten different model seeds (in addition to the five random seeds that impact the sampling process).
+We train a logistic/ridge regression, a random forest and a catboost model, on the data-specific prediction task (see Appendix J for details) and then compare the model-averaged real test performance, $\text{Perf}(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})$, to the performance when trained on the synthetic data, $\text{Perf}(\mathcal{D}_{\text{gen}},\mathcal{D}_{\text{test}})$.
+The results are averaged over ten different model seeds (in addition to the five seeds for the sampling).
 For regression tasks, we consider the RMSE and for classification tasks, the macro-averaged F1 and AUC scores.
-We only report $\lvert\text{Perf}(\mathcal{D}_{\text{gen}},\mathcal{D}_{\text{test}})-\text{Perf}(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})\rvert$ in the main part of this paper.
-An absolute difference close to zero, that is, synthetic and real data induce the same performance, indicates that the generative model performs well.
+We report $\lvert\text{Perf}(\mathcal{D}_{\text{gen}},\mathcal{D}_{\text{test}})-\text{Perf}(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})\rvert$.
+An absolute difference close to zero is preferable since then synthetic and real data induce the same performance.
 Detection score.
@@
 Then, we construct $\mathcal{D}_{\text{train}}^{\text{detect}},\mathcal{D}_{\text{valid}}^{\text{detect}}$ and $\mathcal{D}_{\text{test}}^{\text{detect}}$ with equal proportions of real and fake samples.
-We tune each catboost model on $\mathcal{D}_{\text{valid}}^{\text{detect}}$ and report the accuracy of the best-fitting model on $\mathcal{D}_{\text{test}}^{\text{detect}}$ (see Appendix H for details).
-A (perfect) detection score of 0.5 indicates the model is unable distinguish fake from real samples.
+We tune each catboost model on $\mathcal{D}_{\text{valid}}^{\text{detect}}$ and report the accuracy of the best-fitting model on $\mathcal{D}_{\text{test}}^{\text{detect}}$ (see Appendix I for details).
+A (perfect) detection score of 0.5 indicates that the model is unable to distinguish fake from real samples.
 Statistical similarity.
-We aim to assess the statistical similarity between real and generated data at both the feature and sample levels.
+Similar to Zhang et al.
+(2024), we assess the statistical similarity between real and generated data at both the feature and sample levels.
 We largely follow Zhao et al.
-(2021) and compare: (1) the Jensen-Shannon divergence (JSD; Lin, 1991) to quantify the difference in categorical distributions, (2) the Wasserstein distance (WD; Ramdas et al., 2017) to quantify the difference in continuous distributions, and (3) the L2 distance between pair-wise correlation matrices.
+(2021) and compare: (1) the Jensen-Shannon divergence (JSD; Lin, 1991) to quantify the difference in categorical distributions, (2) the Wasserstein distance (WD; Ramdas et al., 2017) to quantify the difference in continuous distributions, and (3) the L2 distance between pairwise correlation matrices.
 We use the Pearson correlation coefficient for two continuous features, the Theil uncertainty coefficient for two categorical features, and the correlation ratio for mixed types.
-Similar metrics for the evaluation of statistical similarity have been used by Zhang et al.
-(2024).
-Distance to closest record.
-That is, the minimum Euclidean distance of a generated data point to any observation in $\mathcal{D}_{\text{train}}$ (Borisov et al., 2023; Zhao et al., 2021).
+Privacy.
+We compute the distance to closest record (DCR) as the minimum Euclidean distance of a generated data point to any observation in $\mathcal{D}_{\text{train}}$ (Borisov et al., 2023; Zhao et al., 2021).
 We one-hot encode categorical features and standardize all features to zero mean and unit variance to ensure each feature contributes equally to the distance.
-We compute the average distance to closest record (DCR) as a robust estimate.
+We compute the average DCR as a robust estimate.
 For brevity, we report the absolute difference of the DCR of the synthetic data and the DCR of the real test set.
@@
 We assigned the rank 10 for CoDi on lending and diabetes, TabDDPM on acsincome and diabetes, SMOTE on acsincome and covertype.
+RMSE, F1, AUC and DCR are measured in abs.
+differences to the real test set.
 SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
-RMSE $\underline{3.400}_{\pm 3.382}$ $3.800_{\pm 2.482}$ $7.800_{\pm 1.470}$ $7.800_{\pm 2.227}$ $8.200_{\pm 1.470}$ $6.800_{\pm 1.939}$ $6.800_{\pm 1.720}$ $\mathbf{3.000}_{\pm 0.632}$ $\underline{3.400}_{\pm 2.059}$ $4.200_{\pm 2.227}$
-F1 $3.667_{\pm 3.145}$ $6.333_{\pm 2.285}$ $8.333_{\pm 1.491}$ $8.000_{\pm 1.000}$ $4.167_{\pm 2.794}$ $6.667_{\pm 3.091}$ $7.500_{\pm 1.607}$ $3.833_{\pm 1.462}$ $\mathbf{3.167}_{\pm 1.344}$ $\underline{3.500}_{\pm 1.979}$
-AUC $4.667_{\pm 2.749}$ $5.667_{\pm 2.055}$ $8.667_{\pm 1.106}$ $7.833_{\pm 1.067}$ $4.833_{\pm 2.794}$ $7.500_{\pm 2.872}$ $7.333_{\pm 1.374}$ $\underline{2.500}_{\pm 1.500}$ $\mathbf{2.333}_{\pm 0.943}$ $3.833_{\pm 1.675}$
-L2 dist.
-of corr.
-$4.818_{\pm 2.918}$ $5.636_{\pm 1.872}$ $8.091_{\pm 1.781}$ $7.909_{\pm 1.564}$ $7.000_{\pm 3.191}$ $6.909_{\pm 2.429}$ $6.818_{\pm 1.402}$ $\underline{2.727}_{\pm 1.286}$ $\mathbf{2.273}_{\pm 0.862}$ $3.000_{\pm 1.595}$
-Detection score $3.909_{\pm 3.502}$ $6.182_{\pm 1.696}$ $8.818_{\pm 1.466}$ $7.273_{\pm 1.213}$ $5.000_{\pm 3.045}$ $8.091_{\pm 2.391}$ $6.000_{\pm 1.595}$ $3.909_{\pm 1.164}$ $\mathbf{2.455}_{\pm 1.827}$ $\underline{3.545}_{\pm 1.725}$
-JSD $7.182_{\pm 2.167}$ $\mathbf{1.273}_{\pm 0.617}$ $8.182_{\pm 1.641}$ $8.818_{\pm 1.029}$ $6.909_{\pm 2.314}$ $7.000_{\pm 1.651}$ $6.545_{\pm 1.305}$ $\underline{2.455}_{\pm 1.076}$ $3.091_{\pm 0.793}$ $3.727_{\pm 1.052}$
-WD $\underline{3.091}_{\pm 3.315}$ $5.636_{\pm 1.611}$ $7.545_{\pm 1.827}$ $8.000_{\pm 1.477}$ $6.455_{\pm 3.144}$ $8.364_{\pm 1.823}$ $5.727_{\pm 2.339}$ $4.182_{\pm 1.466}$ $3.182_{\pm 1.192}$ $\mathbf{3.000}_{\pm 1.954}$
-DCR $6.000_{\pm 2.558}$ $6.182_{\pm 2.328}$ $8.455_{\pm 1.725}$ $6.182_{\pm 3.186}$ $4.455_{\pm 3.726}$ $6.545_{\pm 2.426}$ $5.909_{\pm 1.676}$ $4.091_{\pm 2.678}$ $\underline{3.818}_{\pm 2.124}$ $\mathbf{3.545}_{\pm 1.924}$
+RMSE $3.6_{\pm 3.3}$ $3.4_{\pm 2.1}$ $8.0_{\pm 1.6}$ $7.8_{\pm 2.2}$ $8.4_{\pm 1.0}$ $7.0_{\pm 1.8}$ $7.2_{\pm 1.5}$ $4.0_{\pm 1.7}$ $\mathbf{2.6}_{\pm 1.0}$ $\underline{3.2}_{\pm 1.5}$
+F1 $4.2_{\pm 2.9}$ $6.2_{\pm 2.3}$ $8.0_{\pm 1.4}$ $8.0_{\pm 1.0}$ $4.3_{\pm 3.1}$ $6.5_{\pm 3.0}$ $8.2_{\pm 1.7}$ $4.0_{\pm 1.1}$ $\mathbf{2.3}_{\pm 1.4}$ $\underline{3.5}_{\pm 1.4}$
+AUC $4.3_{\pm 3.0}$ $5.7_{\pm 1.9}$ $8.3_{\pm 1.2}$ $7.8_{\pm 1.1}$ $4.3_{\pm 3.1}$ $6.8_{\pm 3.1}$ $8.2_{\pm 1.6}$ $\underline{3.3}_{\pm 1.1}$ $\mathbf{2.7}_{\pm 1.5}$ $3.7_{\pm 1.2}$
+L2 dist. of corr. $5.0_{\pm 2.8}$ $5.7_{\pm 2.2}$ $8.3_{\pm 1.9}$ $7.9_{\pm 1.4}$ $6.0_{\pm 3.3}$ $7.0_{\pm 2.5}$ $7.1_{\pm 1.2}$ $3.4_{\pm 1.1}$ $\mathbf{2.0}_{\pm 0.8}$ $\underline{2.8}_{\pm 1.5}$
+Detection score $5.7_{\pm 2.5}$ $6.1_{\pm 1.7}$ $8.8_{\pm 1.5}$ $7.6_{\pm 1.1}$ $4.8_{\pm 3.2}$ $7.9_{\pm 2.4}$ $6.2_{\pm 1.6}$ $\underline{3.1}_{\pm 1.5}$ $\mathbf{1.6}_{\pm 1.0}$ $3.3_{\pm 1.1}$
+JSD $7.2_{\pm 2.2}$ $\mathbf{1.2}_{\pm 0.4}$ $7.6_{\pm 2.4}$ $8.8_{\pm 1.0}$ $6.9_{\pm 2.1}$ $7.2_{\pm 1.5}$ $6.8_{\pm 1.3}$ $\underline{2.5}_{\pm 0.7}$ $2.8_{\pm 1.1}$ $4.1_{\pm 0.9}$
+WD $\underline{3.2}_{\pm 3.3}$ $5.8_{\pm 1.8}$ $7.4_{\pm 2.4}$ $7.9_{\pm 1.6}$ $5.6_{\pm 3.5}$ $8.4_{\pm 1.8}$ $5.8_{\pm 2.0}$ $4.5_{\pm 1.6}$ $\mathbf{3.2}_{\pm 1.5}$ $3.4_{\pm 1.9}$
+DCR $5.8_{\pm 2.7}$ $6.5_{\pm 2.1}$ $8.4_{\pm 1.7}$ $6.1_{\pm 2.9}$ $4.5_{\pm 3.4}$ $6.6_{\pm 2.4}$ $6.3_{\pm 2.0}$ $4.1_{\pm 2.5}$ $\underline{3.9}_{\pm 2.3}$ $\mathbf{3.0}_{\pm 2.0}$
 Table 1 shows the average rank of each generative model across all datasets for the considered metrics.
+Detailed results (including standard errors) for each model and dataset are reported in Appendix S.
 The ranks in terms of the F1 and AUC scores are averaged over the classification task datasets.
@@
 Similarly, SMOTE is very inefficient in sampling for large datasets (78 min for 1000 samples on acsincome and 182 min on covertype) and does not finish the evaluation within 12 hours.
-The dataset-specific results (including standard errors) and average metrics over all datasets are detailed in Appendix R.
-We provide visualizations of the captured correlations in the synthetic sample compared to the real training set in Appendix Q and distribution plots for a qualitative comparison in Appendix P.
+We provide visualizations of the captured correlations in the synthetic sample compared to the real training set in Appendix R and distribution plots for a qualitative comparison in Appendix Q.
 Sample quality.
 CDTD consistently outperforms the considered benchmark models in most sample quality metrics.
-Specifically, we see a major performance edge in terms of the detection score, the L2 distance of the correlation matrices and the regression-based metrics.
-Only for the Jensen-Shannon divergence ARF, a tree-based method that is expected to model categorical features particularly well, outperforms CDTD.
-Interestingly, CDTD performs similar to TabDDPM on F1 scores, but outperforms it dramatically for regression tasks.
-TabDDPM appears to favor modeling categorical features accurately, thereby sacrificing continuous features, as visualized in Appendix Q.
-TabSyn, a latent-space diffusion model, performs worse than CDTD and often TabDDPM, which define diffusion in data space.
-In Appendix M, we further compare CDTD and TabSyn and investigate the benefits of defining a diffusion model in data space.
-By utilizing score interpolation, CDTD is able to model intricate correlation structure more accurately than other frameworks.
+Specifically, we see a major performance edge in terms of the detection score, the L2 distance of the correlation matrices and the ML efficiency metrics.
+Using score interpolation, CDTD is able to model the intricate correlation structure more accurately than other frameworks.
+Due to their inductive biases, SMOTE and ARF perform very well in terms of univariate distribution fit for continuous features (as measured by WD) and categorical (as measured by JSD), respectively.
+However, in both of those cases, CDTD performs competitively, in particular when compared to the diffusion-based models.
+Interestingly, TabSyn, a latent space diffusion model, performs considerably worse than CDTD and often TabDDPM, which define diffusion in data space.
+In Appendix N, we further compare CDTD and TabSyn and investigate the benefits of defining a diffusion model in data space.
 Most importantly, type-specific noise schedules mostly outperform the feature-specific and single noise schedule variants.
-This illustrates the importance of accounting for the high heterogeneity in tabular data on the feature type level.
-The different noise schedules per feature, however, appear to force too many constraints on the model and thus, decrease sample quality.
-Per-feature noise schedules would require more training steps to converge, as can be seen in Appendix O.
-Table 2: Ablation study for five CDTD configurations with progressive addition of model components. We report the median performance metrics over acsincome, adult, beijing and churn.
-Config. A B C D CDTD (per type)
-RMSE (abs. diff.; $\downarrow$) 0.041 0.042 0.043 0.037 0.033
-F1 (abs. diff.; $\downarrow$) 0.012 0.013 0.012 0.016 0.015
-AUC (abs. diff.; $\downarrow$) 0.004 0.005 0.005 0.004 0.004
-L2 distance of corr.  ($\downarrow$) 0.131 0.124 0.146 0.118 0.127
-Detection score ($\downarrow$) 0.577 0.583 0.590 0.561 0.560
-JSD ($\downarrow$) 0.011 0.011 0.011 0.012 0.013
-WD ($\downarrow$) 0.004 0.005 0.005 0.003 0.003
-DCR (abs. diff. to test; $\downarrow$) 0.405 0.361 0.386 0.299 0.372
-Figure 3: Average training and sampling wall-clock time for 1000 samples (excl. acsincome, diabetes, lending).
+This illustrates the necessity to account for the high heterogeneity in tabular data on the feature-type level.
+Having distinct noise schedules per feature instead appears to force too many constraints on the model and, thus, decreases sample quality.
+Per-feature noise schedules would also require more training steps to fully converge, as can be seen in Appendix P.
+We investigate the sensitivity of CDTD to important hyperparameters in Appendix M.
+Table 2: Ablation study for five CDTD configurations. We report the median performance. The grey column depicts the impact of our data type homogenization.
+{NiceTabular}
+[…]
+RMSE (abs. diff.; $\downarrow$) 0.090 0.069 0.060 0.058 0.055
+F1 (abs. diff.; $\downarrow$) 0.015 0.011 0.010 0.006 0.009
+AUC (abs. diff.; $\downarrow$) 0.008 0.005 0.005 0.005 0.006
+L2 distance of corr.  ($\downarrow$) 0.483 0.421 0.457 0.450 0.444
+Detection score ($\downarrow$) 0.762 0.739 0.774 0.662 0.701
+JSD ($\downarrow$) 0.010 0.013 0.011 0.011 0.011
+WD ($\downarrow$) 0.008 0.006 0.007 0.005 0.006
+DCR (abs. diff. to test; $\downarrow$) 0.639 0.552 0.574 0.350 0.544
+Figure 3: Average training and sampling wall-clock times (excl. SMOTE, acsincome, diabetes, lending).
 Training and sampling time.
-Figure 3 shows the average wall-clock time over all (for all models feasible) datasets for training as well as the time for sampling 1000 data points for each baseline model and the per feature CDTD variant (see Appendix T for details).
-We exclude SMOTE due to its considerably longer sampling with an average of 1377 seconds for 1000 samples.
-CDTD’s use of embeddings (instead of one-hot encoding) for categorical features drastically reduces training times and thus, improves scaling to increasing number of categories.
-The ODE formulation of the diffusion process implies competitive sampling speeds, in particular compared to the diffusion-based benchmarks CoDi, TabDDPM and TabSyn.
-Despite TabSyn utilizing a separately trained encoder, this does not result in a lower-dimensional latent space and therefore, does not speed up sampling.
+Figure 3 shows the average training and sampling wall-clock times (for all feasible models) over all datasets (see Appendix U for details on sample quality as a function of sampling time).
+We exclude SMOTE due to its considerably longer sampling times with an average of 1377 seconds for 1000 samples.
+CDTD’s use of embeddings (instead of one-hot encoding) for categorical features improves its scaling to increasing number of categories and thus, drastically reduces training times.
+The sampling speed of CDTD is competitive, in particular compared to the diffusion-based benchmarks CoDi, TabDDPM and TabSyn.
 Ablation study.
-We conduct an ablation study over four datasets to investigate the separate components of our CDTD framework.
-The results are given in Table 2 (detailed results are in Appendix S).
+We investigate the separate components of our CDTD framework.
+The summarized results are given in Table 4.2 and the detailed results in Appendix T.
 The baseline model Config.
-A includes a single noise schedule with the original piece-wise linear formulation (Dieleman et al., 2022) without loss normalization, improved model initialization or adaptive normalization, and the CE and MSE losses are naively averaged.
-Note that this configuration still is a novel contribution to the literature.
+A includes a single noise schedule with the original piece-wise linear functional form (Dieleman et al., 2022) and the CE and MSE losses are naively averaged.
+Note that this configuration is still a novel contribution.
 Config.
-B adds our feature homogenization (i.e., loss normalization, improved initialization and adaptive normalization schemes), Config.
+B adds our data type homogenization (i.e., loss calibration, improved initialization and time-dependent normalization schemes), Config.
 C adds our proposed functional form for a single noise schedule with uniform initialization, and Config.
 D imposes per-type noise schedules.
-Lastly, we add the suggested (low noise level) overweighting timewarping initialization to arrive at the full CDTD (per type) model.
-We see the switch from the piece-wise linear functional form to our more robust noise schedule variant slightly harms sample quality.
-However, the per-type variant and the more informed initialization scheme compensate for this difference.
-Main differences are in the RMSE and detection score as well as training efficiency (the loss calibration and improved initialization facilitate model convergence).
-The final model especially works well on the larger datasets compared to the baseline (see Appendix S), as smaller datasets are relatively easy to fit with 3 million parameters, even without any model improvements.
-We investigate the sensitivity of CDTD to important hyperparameters in Appendix L.
+Lastly, we increase the importance of low noise levels at initialization to arrive at the full CDTD (per type) model.
+The results show that data type homogenization benefits sample quality significantly.
+Metrics associated with continuous features, i.e., RMSE and WD, as well as those relying on all features being generated well, i.e., the detection score, L2 distance of corr.
+matrices and DCR, improve dramatically.
+Switching from the piecewise linear noise schedule to our more robust functional form slightly harms sample quality.
+However, the per-type variant and the improved initialization more than compensate for this.
+The latter appears to trade-off some sample quality for faster convergence during training.
 5 Conclusion and Discussion
-We propose a Continuous Diffusion model for mixed-type Tabular Data (CDTD) that combines score matching and score interpolation and imposes Gaussian diffusion processes on both continuous and embedded categorical features.
-We compared CDTD to various benchmark models and to a single noise schedule as typically used in image diffusion models.
-Our results indicate that addressing the high feature heterogeneity in tabular data on the feature type level and aligning type-specific diffusion elements, such as the noise schedules or losses, substantially benefits sample quality.
+In this paper, we introduce a Continuous Diffusion model for mixed-type Tabular Data (CDTD) that combines score matching and score interpolation and imposes Gaussian diffusion processes on both continuous and embedded categorical features.
+Our results indicate that addressing the high feature heterogeneity in tabular data on the feature type level by aligning type-specific diffusion elements, such as the noise schedules or losses, substantially benefits sample quality.
 Moreover, CDTD shows vastly improved scalability and can accommodate an arbitrary number of categories.
-Our paper serves as an important step to customizing the diffusion probabilistic framework to tabular data.
-In particular, the common type of noise schedules allows for an easy to extend framework that might accelerate progress on diffusion models for tabular data.
+Our paper serves as an important step to customizing score-based models to tabular data.
+The common type of noise schedules allows for an easy to extend framework.
 Crucially, CDTD allows the direct application of diffusion-related advances from the image domain, like classifier-free guidance, to tabular data without the need for a latent encoding.
-We leave further extensions to the tabular data domain, e.g., the exploration of accelerated sampling, efficient score model architectures, different forms of adaptive noise schedules, or the adaption to the data imputation task for future work.
-Finally, we want to emphasize the potential misuse of synthetic data to support unwarranted claims.
-Any generated data should therefore not be blindly trusted, and synthetic data based inferences should always be compared to results from the real data.
-However, the correct use of generative models enables better privacy preservation and facilitates data sharing and open science practices.
+We leave further extensions to the tabular data domain, e.g., the exploration of accelerated sampling, efficient score model architectures, or the adaptation to the data imputation task for future work.
+Finally, we want to warn against the potential misuse of synthetic data to support unwarranted claims.
+Any generated data should not be blindly trusted, and synthetic data based inferences should always be compared to results from the real data.
+However, the correct use of generative models for tabular data enables better privacy preservation and facilitates data sharing and open science practices.
 Limitations
@@
 While CDTD could be directly used for imputation using RePaint (Lugmayr et al., 2022), a separate training process is required to achieve the best results (Liu et al., 2024).
-Therefore, we leave the adaption of CDTD to the imputation task for future work.
+Therefore, we leave the adaptation of CDTD to the imputation task for future work.
 Acknowledgements
-This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-7437. We would also like to thank Sander Dieleman for helpful discussions.
-Appendix A Loss Calibration
-A priori, we let the model be indifferent between features, that is, we scale the loss of each feature such that at the terminal timestep the same loss is attained.
-Here, the signal-to-noise ratio is sufficiently low to approximate a situation in which the model has no information about the data.
-Thus, we are looking for calibrated losses $\mathcal{L}_{\text{MSE}}^{\ast}(x_{\text{cont}}^{(i)},1)$ and $\mathcal{L}_{\text{CE}}^{\ast}(x_{\text{cat}}^{(j)},1)$ which at $t=1$ achieve unit loss in expectation.
+This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-7437. We also thank Sander Dieleman for helpful discussions.
+Appendix A Noise implications of heterogeneous cardinalities
+One distinct characteristic of mixed-type tabular data is that each categorical feature may have a different cardinality, i.e., a different number of categories.
+Below, we briefly empirically investigate how this data characteristic warrants a rethinking of diffusion noise schedules.
+We show that, under some assumptions, categorical features of higher cardinality require more added noise than lower cardinality features to remove the same amount of signal.
+Let $x_{\text{cat}}^{(j)}$ be a single categorical feature with $C_{j}$ categories.
+We assume that each category $c$ has equal probability $\mathrm{P}(c)=1/C_{j}$ and train a CDTD model with a fixed, linear noise schedule to learn the distribution of a single $x_{\text{cat}}^{(j)}$.
+For each class $c\in\{1,\dots,C_{j}\}$, we extract the associated learned embedding ${\mathbf{e}}_{c}^{(j)}$.
+For 500 timesteps $t$, linearly spaced on $[0,1]$, we derive the associated noise levels $\sigma(t)$ from the learned noise schedules and for each $\sigma(t)$, we sample 1000 noisy embeddings ${\mathbf{e}}_{c}^{(j)}(t)={\mathbf{e}}_{c}^{(j)}+\sigma(t){\mathbf{\epsilon}}$ such that ${\mathbf{\epsilon}}\sim\mathcal{N}({\bm{0}},{\mathbf{I}})$.
+We input all ${\mathbf{e}}_{c}^{(j)}(t)$ into the score model to compute the average (calibrated) cross-entropy (CE) loss.
+We define the remaining signal of the feature at a given $t$ as $S_{c}=1-\text{CE}_{c}(t)$, where $\text{CE}_{c}$ reflects the CE loss when using the noisy embedding for class $c$ to predict the true class.
+Lastly, we compute the overall remaining signal over all classes as $\max_{c}S_{c}$.
+Figure 4 shows the result of repeating this procedure with varying $C_{j}$ and four-dimensional embeddings.
+As we can see, lower cardinality features, i.e., those with a low $C_{j}$, attain the same signal only at a higher noise levels $\sigma(t)$.
+This example of the effect of feature heterogeneity in tabular data encourages us to investigate feature-specific noise schedules.
+Figure 4: Comparison of signal remaining in embedded categorical features given by $\max_{c}S_{c}$, where $S_{c}=1-\text{CE}_{c}(t)$ and $\text{CE}_{c}$ reflects the CE loss when using the noisy embedding for class $c\in\{1,\dots,C_{j}\}$ to predict the true class.
+Lower cardinality features tend to require a higher noise level $\sigma(t)$ to achieve the same amount of remaining signal as higher cardinality features.
+Appendix B Loss Calibration
+Under Assumption 1, the signal-to-noise ratio at the terminal timestep is sufficiently low to approximate a situation in which the model has no information about the data.
+We want to let the model be indifferent between features, that is, we scale the loss of each feature such that at the terminal timestep the same expected loss is attained.
+Therefore, we are looking for scaled losses $\mathcal{L}_{\text{MSE}}^{\ast}(x_{\text{cont}}^{(i)},1)$ and $\mathcal{L}_{\text{CE}}^{\ast}(x_{\text{cat}}^{(j)},1)$ which at $t=1$ achieve unit loss in expectation.
 For a single scalar feature and a given timestep $t$, we can write the empirical denoising score matching loss (Equation 3) in the EDM parameterization (Karras et al., 2022) as:
@@
 For $t\rightarrow 0$, the model shifts increasingly towards predicting the error that has been added to the true data.
-In the EDM parameterization, the explicit timestep weight (used to achieve a unit loss across timesteps at initialization, see Appendix B) is $\lambda(t)=1/c_{\text{out}}(t)^{2}\approx 1$ for $t=1$.
-At the terminal timestep $t=1$, we now have:
+In the EDM parameterization, the explicit timestep weight (used to achieve a unit loss across timesteps at initialization, see Appendix C) is $\lambda(t)=1/c_{\text{out}}(t)^{2}\approx 1$ for $t=1$.
+At the terminal timestep $t=1$, we then have:
 $\mathbb{E}_{p(x_{\text{cont}}^{(i)})}[\mathcal{L}_{\text{MSE}}(x_{\text{cont}}^{(i)},1)]$ $=\lambda(1)\,\mathbb{E}_{p(x_{\text{cont}}^{(i)})}\Bigl{(}c_{\text{skip}}(1)x_{1}+c_{\text{out}}(1)F_{\bm{\theta}}^{(i)}-x_{\text{cont}}^{(i)}\Bigr{)}^{2},$
@@
 Therefore, we have $\mathcal{L}_{\text{MSE}}^{\ast}(x_{\text{cont}}^{(i)},1)=\mathcal{L}_{\text{MSE}}(x_{\text{cont}}^{(i)},1)$ as long as we standardize $x_{\text{cont}}^{(i)}$ to unit variance.
-For a single categorical feature, $x^{(j)}_{\text{cat}}$ is distributed according to the proportions $p_{c}$ (for categories $c=1,\dots,C$).
+For a single categorical feature, $x^{(j)}_{\text{cat}}$ is distributed according to the proportions $p_{c}$ (for categories $c=1,\dots,C_{j}$).
 The denoising model for score interpolation is trained with the CE loss:
-$\mathcal{L}_{\text{CE}}(x^{(j)}_{\text{cat}},t)=-\sum_{c=1}^{C}I(x^{(j)}_{\text{cat}}=c)\log F_{{\bm{\theta}},c}^{(j)}\ ,$
+$\mathcal{L}_{\text{CE}}(x^{(j)}_{\text{cat}},t)=-\sum_{c=1}^{C_{j}}I(x^{(j)}_{\text{cat}}=c)\log F_{{\bm{\theta}},c}^{(j)}\ ,$
 where $F_{{\bm{\theta}},c}^{(j)}$ denotes the score model’s prediction of the class probability at timestep $t$.
 Without information, it is optimal to assign the $c$-th category the same proportion as in the training set.
-At $t=1$, we thus let $F_{{\bm{\theta}},c}^{(j)}=p_{c}$ such that the minimum loss equals:
-$\mathbb{E}_{p(x^{(j)}_{\text{cat}})}[\mathcal{L}_{\text{CE}}(x^{(j)}_{\text{cat}},1)]$ $=-\mathbb{E}_{p(x^{(j)}_{\text{cat}})}\sum_{c=1}^{C}I(x^{(j)}_{\text{cat}}=c)\log F_{{\bm{\theta}},c}^{(j)}\ ,$ (9)
-$=-\sum_{c=1}^{C}\mathbb{E}_{p(x^{(j)}_{\text{cat}})}[I(x^{(j)}_{\text{cat}}=c)\log p_{c}]\ ,$ (10)
-$=-\sum_{c=1}^{C}p_{c}\log p_{c}.$ (11)
-We use the training set proportions to compute the normalization constant $Z_{j}=\leavevmode\nobreak\ -\sum_{c=1}^{C}p_{c}\log p_{c}$ to calibrate the loss for categorical features. Then,
+At $t=1$, we thus let $F_{{\bm{\theta}},c}^{(j)}=p_{c}$ such that the loss equals:
+$\mathbb{E}_{p(x^{(j)}_{\text{cat}})}[\mathcal{L}_{\text{CE}}(x^{(j)}_{\text{cat}},1)]$ $=-\mathbb{E}_{p(x^{(j)}_{\text{cat}})}\sum_{c=1}^{C_{j}}I(x^{(j)}_{\text{cat}}=c)\log F_{{\bm{\theta}},c}^{(j)}\ ,$ (8)
+$=-\sum_{c=1}^{C_{j}}\mathbb{E}_{p(x^{(j)}_{\text{cat}})}[I(x^{(j)}_{\text{cat}}=c)\log p_{c}]\ ,$ (9)
+$=-\sum_{c=1}^{C_{j}}p_{c}\log p_{c}.$ (10)
+Note that this is the feature-specific entropy and we can use the training set proportions to compute this normalization constant $Z_{j}=\leavevmode\nobreak\ -\sum_{c=1}^{C_{j}}p_{c}\log p_{c}$ to scale the loss for categorical features.
+Then,
 $\mathbb{E}_{p(x^{(j)}_{\text{cat}})}[\mathcal{L}^{\ast}_{\text{CE}}(x^{(j)}_{\text{cat}},1)]=\mathbb{E}_{p(x^{(j)}_{\text{cat}})}[\mathcal{L}_{\text{CE}}(x^{(j)}_{\text{cat}},1)/Z_{j}]=1\ .$
 We have thus achieved calibrated losses with respect to the terminal timestep $t=1$, that is, $\mathbb{E}_{p(x_{\text{cont}}^{(i)})}[\mathcal{L}_{\text{MSE}}^{\ast}(x_{\text{cont}}^{(i)},1)]=\mathbb{E}_{p(x^{(j)}_{\text{cat}})}[\mathcal{L}^{\ast}_{\text{CE}}(x^{(j)}_{\text{cat}},1)]=1$ for all continuous features $i$ and categorical features $j$.
-Appendix B Output Layer Initialization
-At initialization, we want the neural network to reflect the state of no information (see Appendix A). Likewise, our goal is a loss of one across all features and timesteps.
+Appendix C Output Layer Initialization
+At initialization, we want the neural network to reflect the state of no information (see Appendix B). Likewise, our goal is a loss of one across all features and timesteps.
 For continuous features $i$, we initialize the output layer weights (and biases) to zero such that the output of the score model for a single continuous feature, $F_{\bm{\theta}}^{(i)}$, is also zero.
@@
 We denote the variances of the data $x_{\text{cont}}^{(i)}$ and of the Gaussian noise $\epsilon$ at time $t$ as $\sigma_{\text{data}}^{2}$ and $\sigma^{2}(t)$, respectively.
+At initialization we have:
 $\mathbb{E}_{p(x_{\text{cont}}^{(i)}),p(\epsilon)}[\mathcal{L}^{\ast}_{\text{MSE}}(x_{\text{cont}}^{(i)},t)]$ $=\lambda(t)\,\mathbb{E}_{p(x_{\text{cont}}^{(i)}),p(\epsilon)}\Bigl{(}c_{\text{skip}}(t)(x_{\text{cont}}^{(i)}+\epsilon)+c_{\text{out}}(t)F_{\bm{\theta}}^{(i)}-x_{\text{cont}}^{(i)}\Bigr{)}^{2}\ ,$
@@
 For categorical features $j$, we initialize the output layer such that the model achieves the respective losses under no information.
-Using the loss normalization constant $Z_{j}$ (see Appendix A) and dropping the expectation over $p(\epsilon)$, we have
+Using the loss normalization constant $Z_{j}$ (see Appendix B) and dropping the expectation over $p(\epsilon)$, we have
 $\mathbb{E}_{p(x_{\text{cat}}^{(j)})}[\mathcal{L}^{\ast}_{\text{CE}}(x_{\text{cat}}^{(j)},t)]=\mathbb{E}_{p(x_{\text{cat}}^{(j)})}[\mathcal{L}_{\text{CE}}(x_{\text{cat}}^{(j)},t)/Z_{j}]=\frac{1}{Z_{j}}\mathbb{E}_{p(x_{\text{cat}}^{(j)})}[\mathcal{L}_{\text{CE}}(x_{\text{cat}}^{(j)},t)].$
@@
 which in turn leads to an initial loss of $Z_{j}$ for all $t$ and therefore achieves a uniform, calibrated loss of one at initialization similar to the continuous feature case.
-Appendix C Adaptive Normalization of the Average Diffusion Loss
-Both the loss calibration (see Appendix A) and output layer initialization (see Appendix B) ensure that the losses across timesteps (and features) are equal at initialization.
+Appendix D Adaptive Normalization of the Average Diffusion Loss
+Both the loss calibration (see Appendix B) and output layer initialization (see Appendix C) ensure that the losses across timesteps (and features) are equal at initialization.
 During training, the adaptive noise schedules allow the model to focus automatically on the noise levels that matter most, i.e., where the loss increase is steepest.
@@
 Then, we embed $c_{\text{noise}}$ in frequency space (1024-dimensional) using Fourier features.
-The result is passed through a single linear layer to output a scalar value, passed through an exponential function to ensure that the prediction ^$\hat{\mathcal{L}}(t)\geq 0$.
+The result is passed through a single linear layer to output a scalar value and through an exponential function to ensure that the prediction ^$\hat{\mathcal{L}}(t)\geq 0$.
 We initialize the weights and biases to zero, to ensure that at model initialization we have a unit normalization.
-Appendix D Derivation of the Functional Timewarping Form
+Appendix E Derivation of the Functional Timewarping Form
 Since higher noise levels, $\sigma$, imply a lower signal-to-noise ratio, and in turn a larger loss, $\ell$, we know that the loss must be a monotonically increasing and S-shaped function of the noise level.
@@
 where we defined $Z(\sigma,\mu,\nu)=\bigl{(}\frac{\sigma}{1-\sigma}\frac{1-\mu}{\mu}\bigr{)}^{-\nu}$ and used the definitions of $f_{\log}$ and the parameter $\mu^{\ast}$.
-Appendix E Benchmark Datasets
+Appendix F Benchmark Datasets
 Our selected benchmark datasets are highly diverse, particularly in the number of categories for categorical features (see Table 3).
 For the diabetes and covertype datasets, we transform the original multi-class classification problem into a binary classification task for ease of presentation.
+For the diabetes data, we convert the task to predicting whether a patient was readmitted to a hospital.
 For the covertype data, the task is converted into predicting whether a forest of type 2 is present in a given 30 $\times$ 30 meter area.
-In the diabetes data, we convert the task by predicting whether a patient was readmitted to a hospital.
 All datasets are publicly accessible and (except nmes) licensed under creative commons.
 Table 3: Overview of the selected experimental datasets.
-We count the outcome towards the respective features that remain after removing continuous features with an excessive number of missings.
+We count the target towards the respective features that remain after removing continuous features with an excessive number of missings.
 The minimum and maximum number of categories are taken over all categorical features.
 Dataset License Prediction task Total no. No. of features No. of categories
-observations categorical continuous    min. max.
+observations categorical continuous    min max
 acsincome (Ding et al., 2021) CC0 regression 1 664 500 8 3 2 529
-adult (Becker & Kohavi, 1996) CC BY 4.0 binary classification 48 842 9 6 2 42
-bank (Moro et al., 2012) CC BY 4.0 binary classification 41 188 11 10 2 12
-beijing (Chen, 2017) CC BY 4.0 regression 41 757 1 10 4 4
-churn (Keramati et al., 2014) CC BY 4.0 binary classification. 3 150 5 9 2 5
-covertype (Blackard, 1998) CC BY 4.0 binary classification 581 012 44 10 2 2
-default (Yeh, 2016) CC BY 4.0 binary classification 30 000 10 14 2 11
-diabetes (Clore et al., 2014) CC BY 4.0 binary classification 101 766 28 9 2 716
-lending (Club, 2015) DbCL 1.0 regression 9 182 10 34 2 3151
+adult (Becker & Kohavi, 1996) CC BY 4.0 binary class. 48 842 9 6 2 42
+bank (Moro et al., 2014) CC BY 4.0 binary class. 41 188 11 10 2 12
+beijing (Chen, 2015) CC BY 4.0 regression 41 757 1 10 4 4
+churn (Keramati et al., 2014) CC BY 4.0 binary class. 3 150 5 9 2 5
+covertype (Blackard, 1998) CC BY 4.0 binary class. 581 012 44 10 2 2
+default (Yeh, 2009) CC BY 4.0 binary class. 30 000 10 14 2 11
+diabetes (Clore et al., 2014) CC BY 4.0 binary class. 101 766 28 9 2 716
+lending (Lending Club, 2015) DbCL 1.0 regression 9 182 10 34 2 3151
 news (Fernandes et al., 2015) CC BY 4.0 regression 39 644 14 46 2 2
 nmes (Deb & Trivedi, 1997) unknown regression 4 406 8 11 2 4
-Appendix F Baseline Models
+Appendix G Baseline Models
 Below, we give a brief description of our selected generative baseline models (including code sources).
@@
 TabSyn
-(Zhang et al., 2024) – a diffusion-based model that first learns a tranformer-based VAE to map mixed-type data to a continuous latent space.
+(Zhang et al., 2024) – a diffusion-based model that first learns a transformer-based VAE to map mixed-type data to a continuous latent space.
 Then, the diffusion model is trained on that latent space.
+Note that despite TabSyn utilizing a separately trained encoder, this does not result in a lower-dimensional latent space and therefore, does not speed up sampling.
 We use the official code available at https://github.com/amazon-science/tabsyn under the Apache 2.0 license.
-Appendix G Implementation Details
+Appendix H Implementation Details
 Each of the selected benchmark models requires a rather different, more specialized neural network architecture.
@@
 TVAE 2 996 408
-TabDDPM 3 003 924
+TabDDPM 3 001 786
 CoDi 2 998 043
-TabSyn 3 001 646
-CDTD (per type schedule, TabDDPM architecture) 2 999 721
-We adjust each architecture to a total of ±$\pm 3$ million trainable parameters on the adult dataset to improve the comparability further (see Table 4) and use the same architectures for all considered datasets.
-Note that the total number of parameters may vary slightly across datasets due to different number of features and categories affecting the onehot encoding but is still comparable across models.
+TabSyn 2 997 765
+CDTD (per type) 3 002 969
+We adjust each architecture to a total of $\sim 3$ million trainable parameters on the adult dataset to improve the comparability further (see Table 4) and use the same architectures for all considered datasets.
+Note that the total number of parameters may vary slightly across datasets due to different number of features and categories affecting the one-hot encoding but is still comparable across models.
 We also align the embedding/bottleneck dimensions for CTGAN, TVAE, TabDDPM, TabSyn and CDTD to 256.
@@
 The adjusted output layer ensures that we predict a single value for each continuous features and set of class-specific probabilities for each categorical feature.
-Since our use of embeddings introduces additional parameters, we scale the hidden layers slightly down relative to the TabDDPM to ensure approximately 3 million trainable parameters (instead of 808 neurons per layer we use 806) on the adult dataset.
-More details on the CDTD implementation are given in Appendix J.
-Appendix H Tuning of the Detection Model
+Since our use of embeddings introduces additional parameters, we scale the hidden layers slightly down relative to the TabDDPM to ensure approximately 3 million trainable parameters (instead of 798 neurons per layer we use 796) on the adult dataset.
+More details on the CDTD implementation are given in Appendix K.
+Appendix I Tuning of the Detection Model
 We use a catboost model (Prokhorenkova et al., 2018) to test whether real and generated samples can be distinguished.
@@
 leaf estimation iters Integer Uniform [1, 10]
-Appendix I Machine Learning Efficiency Models
+Appendix J Machine Learning Efficiency Models
 For the group of machine learning efficiency models, we use the scikit-learn and catboost package implementations including the default parameter settings, if not specified otherwise below:
@@
 no. iterations = 2000, early stopping rounds = 50, overfitting detector pval = 0.001
-Appendix J CDTD Implementation Details
+We subsample $\mathcal{D}_{\text{train}}$ in case of more than 50 000 observations to upper-bound the computational load.
+Appendix K CDTD Implementation Details
 Figure 5: Overview of the CDTD architecture adapted from TabDDPM.
@@
 CDTD only conditions on $y$, i.e., the target feature, for classification task datasets.
-To enable a fair comparison to the other methods, and to TabDDPM and TabSyn in particular, the CDTD score model utilizes the exact same architecture and optimizer as Kotelnikov et al.
+To enable a fair comparison to the other methods, and to TabDDPM and TabSyn in particular, the CDTD score model utilizes the exact same architecture as Kotelnikov et al.
 (2023), which was also adapted by TabSyn (Zhang et al., 2024).
-An overview of the score model is provided in Figure 4: First, the noisy data, i.e., the noisy scalars for continuous features and the noisy embeddings for categorical features, are projected onto a 256-dimensional space.
-Similarly, timestep $t$ and possibly conditioning information $y$ are embedded in the same space.
-Then, all 256-dimensional vectors are added and the results is processed by a set of five fully-connected linear layers with ReLU activation functions.
+We use the QuantileTransformer to pre-process continuous features, followed by a standardization to zero mean and unit variance.
+An overview of the score model is provided in Figure 5: First, the noisy data, i.e., the noisy scalars for continuous features and the noisy embeddings for categorical features, and the timestep $t$, are projected onto a 256-dimensional space.
+Then, all 256-dimensional vectors are added and the result is processed by a set of five fully-connected linear layers with ReLU activation functions.
 Lastly, a linear projection maps the output of the fully-connected layers to the required output dimensions, which depend on the number of features and number of categories per feature.
 The only major difference to the TabDDPM setup are the inputs, as we need to embed the categorical features in Euclidean space.
-The output dimensions are the same, as we need to predict a single scalar for each $x_{\text{cont},i}$, and $\lvert C_{j}\rvert$ values for each $x_{\text{cat},j}$, with $C_{j}$ the set of categories of feature $j$.
-We change the initialization of the output layer as described in Appendix B: To handle our inputs, we embed the categorical features in 16-dimensional space and add a feature-specific bias of the same dimension, which captures feature-specific information common to all categories and is initialized to zero.
+The output dimensions are the same, as we need to predict a single scalar for each $x_{\text{cont}}^{(i)}$, and $C_{j}$ values for each $x_{\text{cat}}^{(j)}$, with $C_{j}$ being the number of categories of feature $j$.
+We change the initialization of the output layer as described in Appendix C.
+To handle our inputs, we embed the categorical features in 16-dimensional space and add a feature-specific bias of the same dimension, which captures feature-specific information common to all categories and is initialized to zero.
 We L2-normalize each embedding to prevent a degenerate embedding space in which embeddings are pushed further and further apart (see Dieleman et al., 2022).
@@
 Table 7 indicates that CDTD is not sensitive to the choice of $\sigma_{\text{init}}$.
-Since we utilize embeddings, we have to scale the neurons per layer slightly down in the stack of the five fully-connected layers (from 808 for TabDDPM to 806).
+Since we utilize embeddings, we have to scale the neurons per layer slightly down in the stack of the five fully-connected layers (from 798 for TabDDPM to 796).
 Also, since TabDDPM samples discrete steps from $[0,T]$, with $T\gg 1$, we scale our timesteps $t\in[0,1]$ up by 1000.
 We use the same optimizer (Adam), learning rate (0.001), learning rate decay (linear), EMA decay (0.999), and training steps (30000).
-However, since we work with embeddings we add a linear warmup schedule over the first 100 steps.
-Instead of using the vanilla uniform (time)step sampling as the TabDDPM, the CDTD model uses antithetic sampling (Dieleman et al., 2022; Kingma et al., 2022).
+However, since we work with embeddings we add a linear warmup schedule over the first 1000 steps.
+Lastly, instead of using uniform (time)step sampling as TabDDPM, the CDTD model uses antithetic sampling (Dieleman et al., 2022; Kingma et al., 2021).
 The timesteps are still uniformly distributed but spread out more evenly over the domain, which benefits the training of the adaptive noise schedules.
-For generation, we use an Euler sampler with 200 steps to minimize the discretization error.
-Appendix K CDTD Sampling
-To sample from our learned distribution, we need to run the reverse process of the probability flow ODE (Equation 2).
+For generation, we use a deterministic (Euler) sampler with 200 steps to minimize the discretization error (see Appendix L for details).
+Appendix L CDTD Sampling
+Algorithm 1 Deterministic Sampling
+0:  $N$ sampling steps Sample ${\mathbf{x}}_{\text{cont},t_{0}}^{(i)}\sim\mathcal{N}({\bm{0}},\sigma_{\text{cont},\max}^{2}{\mathbf{I}}_{K_{\text{cont}}})\,\forall i\in\{1,\dots,K_{\text{cont}}\}$ Sample ${\mathbf{x}}_{\text{cat},t_{0}}^{(j)}\sim\mathcal{N}({\bm{0}},\sigma_{\text{cat},\max}^{2}{\mathbf{I}}_{K_{\text{cat}}})\,\forall j\in\{1,\dots,K_{\text{cat}}\}$
+1:  for $s\in\{0,\dots,N-1\}$ do
+2:     $t_{s}=1-s/N$
+3:     $t_{s+1}=1-(s+1)/N$
+4:     ${\mathbf{x}}_{s}\leftarrow({\mathbf{x}}_{\text{cat},t_{s}}^{(1)},\dots,{\mathbf{x}}_{\text{cat},t_{s}}^{(K_{\text{cat}})},x_{\text{cont},t_{s}}^{(1)},\dots,x_{\text{cont},t_{s}}^{(K_{\text{cont}})})$
+5:     for all $i\in\{1,\dots,K_{\text{cont}}\}$ do
+6:        $\mathrm{d}x_{\text{cont}}^{(i)}=(x_{\text{cont},t_{s}}^{(i)}-D^{\text{cont},i}_{\bm{\theta}}({\mathbf{x}}_{s},t_{s}))/\sigma_{\text{cont},i}(t_{s})$
+7:        $x_{\text{cont},t_{s+1}}^{(i)}\leftarrow x_{\text{cont},t_{s}}^{(i)}+[\sigma_{\text{cont},i}(t_{s+1})-\sigma_{\text{cont},i}(t_{s})]\,\mathrm{d}x_{\text{cont}}^{(i)}$
+8:     end for
+9:     for all $j\in\{1,\dots,K_{\text{cat}}\}$ do
+10:        $\hat{\mathrm{P}}({\mathbf{x}}_{\text{cat},0}^{(j)}|{\mathbf{x}}_{\text{cat},t_{s}}^{(j)})=D^{\text{cat},j}_{\bm{\theta}}({\mathbf{x}}_{s},t_{s})$
+11:        $\mathrm{d}{\mathbf{x}}_{\text{cat}}^{(j)}=\Bigl{(}{\mathbf{x}}_{\text{cat},t_{s}}^{(j)}-\mathbb{E}_{\hat{\mathrm{P}}({\mathbf{x}}_{\text{cat},0}^{(j)}|{\mathbf{x}}_{\text{cat},t_{s}}^{(j)})}[{\mathbf{x}}_{\text{cat},0}^{(j)}]\Bigr{)}/\sigma_{\text{cat},j}(t_{s})$ from Equation 4
+12:        ${\mathbf{x}}_{\text{cat},t_{s+1}}^{(j)}\leftarrow{\mathbf{x}}_{\text{cat},t_{s}}^{(j)}+[\sigma_{\text{cat},j}(t_{s+1})-\sigma_{\text{cat},j}(t_{s})]\,\mathrm{d}{\mathbf{x}}_{\text{cat}}^{(j)}$
+13:     end for
+14:  end forRecover classes from embeddings with additional pass to score model
+15:  ${\mathbf{x}}_{N}\leftarrow({\mathbf{x}}_{\text{cat},t_{N}}^{(1)},\dots,{\mathbf{x}}_{\text{cat},t_{N}}^{(K_{\text{cat}})},x_{\text{cont},t_{N}}^{(1)},\dots,x_{\text{cont},t_{N}}^{(K_{\text{cont}})})$
+16:  for all $j\in\{1,\dots,K_{\text{cat}}\}$ do
+17:     $\hat{\mathrm{P}}({\mathbf{x}}_{\text{cat},0}^{(j)}|{\mathbf{x}}_{\text{cat},t_{N}}^{(j)})=D^{\text{cat},j}_{\bm{\theta}}({\mathbf{x}}_{N},t_{N-1})$
+18:     $x_{\text{cat}}^{(j)}\leftarrow\operatorname*{arg\,max}_{c\in C_{j}}\hat{\mathrm{P}}({\mathbf{x}}_{\text{cat},0}^{(j)}={\mathbf{e}}_{c}^{(j)}|{\mathbf{x}}_{\text{cat},t_{N}}^{(j)})$ (pick most likely category)
+19:  end for
+20:  return  $(x_{\text{cat}}^{(1)},\dots,x_{\text{cat}}^{(K_{\text{cat}})},x_{\text{cont},t_{N}}^{(1)},\dots,x_{\text{cont},t_{N}}^{(K_{\text{cont}})})$
+Algorithm 1 shows our deterministic sampling approach, where $D^{\text{cont},i}_{\bm{\theta}}$ and $D^{\text{cat},j}_{\bm{\theta}}$ represent the score model output for the $i$-th continuous and $j$-th categorical feature, respectively.
+We found results similar to Karras et al.
+(2022), i.e., that adding stochasticity to the sampling process does not benefit sample quality.
+We also experimented with second-order samplers (e.g., Heun) but found them to add little benefit over our first-order method while requiring double the NFEs.
+As a default, we use 200 sampling steps.
+Table 6 shows that the gains in sample quality are marginal to non-existent after more than 500 sampling steps.
+There is another subtlety in our sampler: Unlike in EDM (Karras et al., 2022), in the final step the sampler steps into $t=0$, which is not associated with $\sigma_{0}=0$ but $\sigma_{0}=\sigma_{\text{min}}$.
+This is a consequence of utilizing more than one noise schedule.
+We condition the score model not on $\sigma$ but on $t$, which indicates the “global” time common to all noise schedules.
+Moving from $\sigma_{0}=\sigma_{\text{min}}$ to $\sigma=0$ in the final step, would imply $t<0$.
+Therefore, we let $\sigma_{\text{min}}=0$ for all feature types.
+To sample from the learned distribution, we need to run the reverse process of the probability flow ODE (Equation 2).
 For example, for two different features $x_{1}$ and $x_{2}$, we deconstruct the ODE as:
-$d{\mathbf{x}}$ $=-\frac{1}{2}{\mathbf{G}}(t){\mathbf{G}}(t)^{\mathsf{T}}\nabla_{\mathbf{x}}\log p_{t}({\mathbf{x}})dt$
-$=-\begin{bmatrix}\dot{\sigma_{1}}(t)\sigma_{1}(t)&\\ &\dot{\sigma_{2}}(t)\sigma_{2}(t)\end{bmatrix}\begin{bmatrix}\frac{\hat{x_{1}}-x_{1}}{\sigma_{1}(t)^{2}}\\ \frac{\hat{x_{2}}-x_{2}}{\sigma_{2}(t)^{2}}\end{bmatrix}dt$
-$=-\begin{bmatrix}\dot{\sigma_{1}}(t)&\\ &\dot{\sigma_{2}}(t)\end{bmatrix}\begin{bmatrix}\frac{\hat{x_{1}}-x_{1}}{\sigma_{1}(t)}\\ \frac{\hat{x_{2}}-x_{2}}{\sigma_{2}(t)}\end{bmatrix}dt$
+$\mathrm{d}{\mathbf{x}}$ $=-\frac{1}{2}{\mathbf{G}}(t){\mathbf{G}}(t)^{\mathsf{T}}\nabla_{\mathbf{x}}\log p_{t}({\mathbf{x}})\mathrm{d}t$
+$=-\begin{bmatrix}\dot{\sigma_{1}}(t)\sigma_{1}(t)&\\ &\dot{\sigma_{2}}(t)\sigma_{2}(t)\end{bmatrix}\begin{bmatrix}\frac{\hat{x_{1}}-x_{1}}{\sigma_{1}(t)^{2}}\\ \frac{\hat{x_{2}}-x_{2}}{\sigma_{2}(t)^{2}}\end{bmatrix}\mathrm{d}t$
+$=-\begin{bmatrix}\dot{\sigma_{1}}(t)&\\ &\dot{\sigma_{2}}(t)\end{bmatrix}\begin{bmatrix}\frac{\hat{x_{1}}-x_{1}}{\sigma_{1}(t)}\\ \frac{\hat{x_{2}}-x_{2}}{\sigma_{2}(t)}\end{bmatrix}\mathrm{d}t$
 In practice, we use an Euler sampler with 200 discrete timesteps $\Delta t=t_{i+1}-t_{i}<0$.
 The timesteps are generated as a linearly spaced grid on $[0,1]$ and transformed afterwards into noise levels $\sigma_{k}(t)$ via the described timewarping procedure.
-For the discretized and simplified the ODE above, this yields
-${\mathbf{x}}_{i+1}$ $={\mathbf{x}}_{i}-\begin{bmatrix}\frac{\Delta\sigma_{1}(t)}{\Delta t}&\\ &\frac{\Delta\sigma_{2}(t)}{\Delta t}\end{bmatrix}\begin{bmatrix}\frac{\hat{x_{1}}-x_{1}}{\sigma_{1}(t)}\\ \frac{\hat{x_{2}}-x_{2}}{\sigma_{2}(t)}\end{bmatrix}\Delta t={\mathbf{x}}_{i}+\begin{bmatrix}\frac{x_{1}-\hat{x_{1}}}{\sigma_{1}(t)}\\ \frac{x_{2}-\hat{x_{2}}}{\sigma_{2}(t)}\end{bmatrix}\odot\begin{bmatrix}\Delta\sigma_{1}(t)\\ \Delta\sigma_{2}(t)\end{bmatrix}.$
+For the discretized and simplified ODE at step $i$, this yields
+${\mathbf{x}}_{i+1}$ $={\mathbf{x}}_{i}-\begin{bmatrix}\frac{\Delta\sigma_{1}(t)}{\Delta t}&\\ &\frac{\Delta\sigma_{2}(t)}{\Delta t}\end{bmatrix}\begin{bmatrix}\frac{\hat{x_{1}}-x_{1}}{\sigma_{1}(t_{i})}\\ \frac{\hat{x_{2}}-x_{2}}{\sigma_{2}(t_{i})}\end{bmatrix}\Delta t={\mathbf{x}}_{i}+\begin{bmatrix}\frac{x_{1}-\hat{x_{1}}}{\sigma_{1}(t_{i})}\\ \frac{x_{2}-\hat{x_{2}}}{\sigma_{2}(t_{i})}\end{bmatrix}\odot\begin{bmatrix}\Delta\sigma_{1}(t)\\ \Delta\sigma_{2}(t)\end{bmatrix},$
 where $\odot$ denotes the element-wise product.
 Hence, we are effectively taking feature-specific steps of length $\Delta\sigma_{k}(t)$.
-The adaptive noise schedules (timewarping) therefore not only affect the training process, but also focus most work in the reverse process on the noise levels that matter most for sample quality (i.e., where $\Delta\sigma_{i}(t)$ is small).
+The adaptive noise schedules (timewarping) therefore not only affect the training process, but also focus most work in the reverse process on the noise levels that matter most for sample quality (i.e., where $\Delta\sigma_{k}(t)$ is small).
 We use finite differences to approximate $\dot{\sigma}_{i}$, instead of the available, analytical variant, since $\frac{d\sigma_{k}(t)}{dt}\rightarrow\infty$ as $t\rightarrow 1$.
-The step $\Delta t$ would therefore be required to decrease as $t\rightarrow 1$ to ensure $\Delta t\approx dt$ holds.
-For a large number of steps, this assumption does not hold in practice, and for $\frac{d\sigma_{k}(t)}{dt}$ the update of ${\mathbf{x}}$ overshoots the target drastically.
-Intuitively, $\sigma_{k}(t)$ becomes too steep near the terminal timestep $t=1$ such that the step size can not sufficiently compensate for the slope increase to turn $\frac{d\sigma_{k}(t)}{dt}$ into a good approximation of the actual change in $\sigma_{k}(t)$.
-Moreover, the analytical solution would approximate $d\sigma_{k}(t)=\dot{\sigma}_{k}(t)dt$, i.e., the change in the noise level caused by a change in $t$.
-Since we know exactly where $\sigma_{k}(t)$ will end up when changing $t$, we are better off using that exact value and let $d\sigma_{k}(t)=\Delta\sigma_{k}(t)$.
-Table 6 shows that the gains in sample quality are marginal to non-existent after more than 500 sampling steps.
+The step $\Delta t$ would therefore be required to decrease as $t\rightarrow 1$ to ensure $\Delta t\approx\mathrm{d}t$ holds.
+For a large number of steps, this assumption does not hold in practice, and for $\frac{\mathrm{d}\sigma_{k}(t)}{\mathrm{d}t}$ the update of ${\mathbf{x}}$ overshoots the target drastically.
+Intuitively, $\sigma_{k}(t)$ becomes too steep near the terminal timestep $t=1$ such that the step size can not sufficiently compensate for the slope increase to turn $\frac{\mathrm{d}\sigma_{k}(t)}{\mathrm{d}t}$ into a good approximation of the actual change in $\sigma_{k}(t)$.
+Moreover, the analytical solution would approximate $\mathrm{d}\sigma_{k}(t)=\dot{\sigma}_{k}(t)\mathrm{d}t$, i.e., the change in the noise level caused by a change in $t$.
+Since we know exactly where $\sigma_{k}(t)$ will end up when changing $t$, we are better off using that exact value and let $\mathrm{d}\sigma_{k}(t)=\Delta\sigma_{k}(t)$.
 Table 6: Performance sensitivity of CDTD (per type) to increasing number of sampling steps.
@@
 Steps RMSE F1 AUC L2 distance of corr. Detection score JSD WD DCR
-200 (default) 0.033 0.015 0.004 0.127 0.560 0.013 0.003 0.372
-500 0.028 0.018 0.005 0.130 0.565 0.012 0.003 0.372
-1000 0.028 0.018 0.005 0.129 0.560 0.012 0.003 0.373
-1500 0.028 0.018 0.005 0.129 0.561 0.012 0.003 0.374
-Appendix L Sensitivity to Important Hyperparameters
+100 0.038 0.011 0.007 0.131 0.574 0.012 0.003 0.315
+200 (default) 0.030 0.009 0.006 0.130 0.565 0.012 0.003 0.316
+500 0.027 0.009 0.006 0.130 0.557 0.013 0.002 0.313
+1000 0.028 0.008 0.006 0.130 0.562 0.013 0.002 0.311
+Appendix M Sensitivity to Important Hyperparameters
 The training and sampling processes of CDTD are affected by various novel hyperparameters.
-Generally, a per-type noise schedule works best as we show in our main results in Table 1 for a diverse set of benchmark datasets.
+Generally, a per-type noise schedule works best, as we show in our main results in Table 1 for a diverse set of benchmark datasets.
 Here, we examine the sensitivity of CDTD to two additional important hyperparameters: (1) the standard deviation of the noise used to initialize the embeddings (and therefore specific to score interpolation), $\sigma_{\text{init}}$, and (2) the weight of the low noise levels used to initialize the $\mu_{k}$ in the adaptive noise schedule parameterization.
@@
 (2022) show that $\sigma_{\text{init}}$ is a crucial hyperparameter for score interpolation on text data.
-The same sensitivity does not translate to the tabular data domain, as shown in our results in Table 7.
-The much smaller embedding dimension (16 vs.
-256) and the feature-specific embeddings significantly decrease the number of distinguishable categories.
+Table 7 shows that this sensitivity does not translate to the tabular data domain.
+This may be explained by the much smaller embedding dimension (16 vs.
+256) or by our usage of feature-specific embeddings.
 Compared to a vocabulary size of 32000 for text data (Dieleman et al., 2022), we only face a maximum of 3151 categories in the lending dataset (see Table 3).
 Thus, unlike other generative (diffusion) models for tabular data, CDTD scales to a practically arbitrary number of categories.
-Our proposed functional form for the adaptive noise schedules (see Appendix D) is the first to allow for the incorporation of prior information about the importance of low vs.
+Our proposed functional form for the adaptive noise schedules (see Appendix E) is the first to allow for the incorporation of prior information about the importance of low vs.
 high (normalized) noise levels.
 For this, we adjust the weight of low noise levels which directly determines the location of the inflection point $\mu_{k}$ (see Section 3.3).
-The results in Table 8 indicate low sample quality sensitivity to weight changes for a per-type noise schedule.
+The results in Table 8 indicate low sensitivity of sample quality to weight changes for a per-type noise schedule.
 The initialization only impacts the time to convergence but not (much) the location of the optimum.
-In our experiments, the number of training steps (30000) appears to be high enough for all model variants to converge.
+In our experiments, the number of training steps (30000) appears to be high enough for all model variants to converge to similar states.
 Table 7: Performance sensitivity of CDTD (per type) to changes in the standard deviation $\sigma_{\text{init}}$ in the initialization of the embeddings of categorical features.
@@
 $\sigma_{\text{init}}$ RMSE F1 AUC L2 distance of corr. Detection score JSD WD DCR
-1 0.032 0.017 0.006 0.126 0.564 0.011 0.004 0.311
-0.1 0.035 0.016 0.004 0.128 0.570 0.012 0.004 0.358
-0.01 0.032 0.017 0.005 0.131 0.566 0.011 0.004 0.369
-0.001 (default) 0.033 0.015 0.004 0.127 0.560 0.013 0.003 0.372
+1 0.031 0.012 0.004 0.123 0.555 0.013 0.003 0.294
+0.1 0.027 0.012 0.004 0.125 0.556 0.013 0.003 0.323
+0.01 0.031 0.007 0.005 0.129 0.575 0.013 0.003 0.320
+0.001 (default) 0.030 0.009 0.006 0.130 0.565 0.012 0.003 0.316
 Table 8: Performance sensitivity of CDTD (per type) to changes in the prior weight of low noise levels in the initialization of the adaptive noise schedules.
@@
 Weight RMSE F1 AUC L2 distance of corr. Detection score JSD WD DCR
-1 0.036 0.015 0.004 0.143 0.651 0.015 0.003 0.313
-2 0.030 0.014 0.005 0.147 0.651 0.014 0.003 0.352
-3 (default) 0.033 0.015 0.004 0.154 0.651 0.013 0.004 0.366
-4 0.034 0.019 0.005 0.148 0.656 0.013 0.004 0.370
-Appendix M Advantages of Diffusion in Data Space
+1 0.031 0.010 0.005 0.126 0.570 0.012 0.002 0.238
+2 0.030 0.009 0.004 0.125 0.570 0.013 0.002 0.289
+3 (default) 0.030 0.009 0.006 0.130 0.565 0.012 0.003 0.316
+4 0.028 0.011 0.005 0.134 0.574 0.011 0.003 0.350
+Appendix N Advantages of Diffusion in Data Space
 These days, inspired from diffusion models in the image and video domains, much work relies on the idea of latent diffusion.
-Here, we want to briefly discuss and emphasize that for tabular data, diffusion in latent space (represented by TabSyn) has important drawbacks and how CDTD, a diffusion model defined in data space differs from that.
-Latent diffusion models first encode the data and map it into a latent space.
-The diffusion model itself is then trained in that latent space.
-Hence, the performance of the diffusion model directly depends on a second, separate model, with a separate training procedure.
+Here, we want to briefly discuss and emphasize that for tabular data, diffusion in latent space (represented by TabSyn) has important drawbacks and how CDTD, a diffusion model defined in data space alleviates those.
+Latent diffusion models first encode the data in a latent space.
+The diffusion model itself is then trained in that latent space instead of directly on the features.
+Hence, the performance of the diffusion model directly depends on a second model, with a separate training procedure.
 TabSyn uses a VAE model to encode mixed-type data into a common continuous space that is not lower-dimensional, so as to minimize reconstruction errors.
-Any reconstruction errors caused by the incapability of the VAE in turn reduce the sample quality of the eventually generated samples, no matter the capacity of the diffusion model.
-This suggests that we would want to train a high capable encoder/decoder, which adds additional training costs.
+Any reconstruction errors caused by the VAE reduce the sample quality of the eventually generated samples, no matter the capacity of the diffusion model.
+This suggests that we would want to train a highly capable encoder/decoder, which adds additional training costs.
 Figure 3 shows that latent diffusion is not necessarily more efficient in the tabular data domain.
-In particular, if the latent space is not lower-dimensional to minimize reconstruction error, then sampling speed is not improved.
+In particular, if the latent space is not lower-dimensional or the encoder/decoder is very complex, then sampling speed is not improved.
 We further hypothesize that much tabular data, due to the lack of redundancy and spatial or sequential correlation, is difficult to summarize efficiently in a joint latent space.
@@
 Also, there is the risk of the VAE not picking up on subtle correlations within the data or distorting existing correlations by mapping into the latent space.
-Any correlations not properly encoded in the latent space, cannot be learned or exploited by the diffusion model.
+Any correlations not properly encoded cannot be learned by the diffusion model.
 Since we optimize the VAE on an average loss, its reconstruction and encoding performance of, for instance, minority classes or extreme values in long-tailed distributions is likely lacking.
-This makes the job of the diffusion model more difficult, if not impossible.
-Lastly, we take great care in homogenizing categorical and continuous features throughout the training process (see Appendix A and B).
+This makes the job of the diffusion model more difficult, if not sometimes impossible.
+Lastly, we take great care in homogenizing categorical and continuous features throughout the training process (see Appendix B and C).
 This is a crucial part of modeling mixed-type data.
 Using a VAE to define a diffusion process in latent space only shifts the necessity for homogenization to the VAE training process.
-Not balancing different feature- or data-types and their losses induces an implicit importance weight for each feature.
+Not balancing different feature- or data-types and their losses induces implicit importance weights.
 Thus, the VAE may sacrifice the reconstruction quality of some features in favor of others (Kendall et al., 2018; Ma et al., 2020).
-To empirically investigate the difference of diffusion in data space (CDTD) and latent diffusion (TabSyn), we examine the worst feature-specific sample quality and other metrics that directly benefit from the model generating all features well.
-Our results in Table 9 show that, latent diffusion comes with a considerable decrease in sample quality (while imposing a similar architecture and number of parameters as well as sampling steps, see Appendix G).
-In particular, the attained maximum metrics indicate that TabSyn has issues modeling all features and their correlations sufficiently well.
-This supports our argument that a homogenization of data types is of crucial importance to avoid having the model implicitly favor one feature over another.
+To empirically investigate the difference of diffusion in data space (CDTD) and latent diffusion (TabSyn), we examine feature-specific sample quality metrics and those that directly benefit from all features being generated well.
+Our results in Table N show that latent diffusion comes with a considerable decrease in sample quality (while imposing a similar architecture and number of parameters as well as sampling steps, see Appendix H).
+In particular, the attained maximum univariate metrics as well as the detection score and the L2 distance of the correlation matrices indicate that TabSyn has issues modeling all features and their correlations sufficiently well.
+This supports our argument that a homogenization of data types is important and that an encoding in latent space may complicate the learning of joint data characteristics.
 Table 9: A comparison of the CDTD model to latent diffusion (TabSyn).
-We average each metric over five sampling seeds and as a robust measure report the median over the ablation datasets acsincome, adult, beijing and churn.
+We average each metric over five sampling seeds and as a robust measure report the median over the 11 datasets.
 Abs.
@@
 matrices refers to the absolute differences in the correlation matrices between ground truth and synthetic data.
-The maximum, minimum and mean are taken across features.
-Detection score L2 dist. of corr. JSD WD Abs. diff. in corr. matrices
+The max, min and mean are taken across features.
+{NiceTabular}
+[…]
 min mean max min mean max min max
-TabSyn 0.772 0.479 0.005 0.018 0.046 0.003 0.006 0.017 0 0.133
-CDTD (per type) 0.566 0.131 0.001 0.012 0.022 0.001 0.003 0.007 0 0.052
-improvement over TabSyn 1.364 3.656 5.000 1.500 2.091 3.000 2.000 2.429 0 2.558
-Appendix N Comparison to Related Work
-Table 10 summarizes our comparison of CDTD to the selected diffusion-based benchmark models, that is, TabSyn, TabDDPM and CoDi.
-Of those models, only TabSyn applies diffusion in latent space, which comes with both advantages and costs (as discussed in Appendix M).
+TabSyn 0.859 0.919 0.004 0.044 0.141 0.002 0.012 0.025 0.000 0.261
+CDTD (per type) 0.701 0.444 0.002 0.011 0.025 0.001 0.006 0.010 0.000 0.088
+Improv. over TabSyn 18.4% 51.7% 50.0% 75.0% 82.3% 50.0% 50.0% 60.0% - 66.3%
+Appendix O Comparison to Related Work
+Table O summarizes our comparison of CDTD to the diffusion-based benchmark models, that is, TabSyn, TabDDPM and CoDi.
+Of those models, only TabSyn applies diffusion in latent space, which comes with both advantages and costs (as discussed in Appendix N).
 TabSyn is the only other model besides CDTD that avoids one-hot encoding categorical features by using embeddings.
@@
 Although both models utilize embeddings, TabSyn’s generative capabilities are more constrained by jointly encoding all features in a latent space.
-As such, it is still less flexible than CDTD, in particular when modeling very unbalanced categorical data.
-Information on rare categories may easily be cut off in favor of attributing more capacity in the latent space to more prominent categories or features.
-It should also be noted that TabSyn is the only model that makes use of a Transformer architecture in its VAE, which means that it scales quadratically in the number of features and therefore may not be easily scaled to high-dimensional data.
-The CDTD model is the first to utilize adaptive and type- or feature-specific noise schedules to model tabular data.
-Further, we take great care in homogenizing categorical and continuous features throughout the training process, including the model initialization (see Appendix A and B).
+It should also be noted that TabSyn makes use of a Transformer architecture in its VAE, which means that it scales quadratically in the number of features and therefore may not be easily scaled to high-dimensional data.
+CDTD is the first model to utilize adaptive and type- or feature-specific noise schedules to model tabular data.
+Further, we take great care in homogenizing categorical and continuous features throughout the training process (see Appendix B and C).
 No other model attempts balancing the different features types.
-This is problematic as it suggests that other models may suffer from feature-specific induced implicit importance weights that impact both training and generation processes.
+This is problematic as it suggests that other models may suffer from feature-specific implicit importance weights that impact both training and generation.
 Hence, the sample quality of some features may be unintentionally sacrificed in favor of increasing the sample quality of other features (Kendall et al., 2018; Ma et al., 2020).
 Note that this also applies to TabSyn: Even though their diffusion model avoids this issue by relying on a single type of loss due to the continuous latent space, the VAE training process does not account for any balancing issues between the two data types.
-Hence, the balancing issue is not eliminated but got only shifted to the encoder VAE.
+Hence, the balancing issue is not eliminated but got only shifted to the VAE.
 Lastly, CDTD and TabSyn are the only models that define the diffusion process in continuous space.
@@
 This can worsen the sample quality of some features in favor of others.
-defined in feature space avoids one-hot encoding balances feature types adaptive noise schedule type- or feature- specific noise schedules diffusion in continuous space
-CoDi
-TabDDPM
-TabSyn $\ast$
-CDTD (ours)
-Appendix O Examples of Learned Noise Schedules
-Next, we show the learned noise schedules for the smallest (churn) and the largest (acsincome) datasets.
-Additionally, we illustrate the fit of single, per type and per feature schedules to the respective losses.
+{NiceTabular}
+[…]
+CoDi ✓
+TabDDPM ✓
+TabSyn ✓ $\ast$
+CDTD (ours) ✓ ✓ ✓ ✓ ✓ ✓
+Appendix P Examples of Learned Noise Schedules
+Next, we show the learned noise schedules for the largest (acsincome) and the smallest (churn) datasets.
+Additionally, we illustrate the fit of single, per type and per feature schedules to the respective diffusion losses they were trained to fit.
+Figure 6: (Left): Learned noise schedules for acsincome. This reflects $F_{\mathrm{d.a.log},k}^{-1}$. (Right): Implicit weighting of noise levels / timesteps. This visualizes $f_{\mathrm{d.a.log},k}$.
+Figure 7: Illustration of the goodness of fit of the timewarping function $F_{k}$ for single (left), per type (middle) and per feature noise schedules (right) on the acsincome data.
 Figure 8: (Left): Learned noise schedules for churn. This reflects $F_{\mathrm{d.a.log},k}^{-1}$. (Right): Implicit weighting of noise levels / timesteps. This visualizes $f_{\mathrm{d.a.log},k}$.
-Figure 6: Illustration of the goodness of fit of the timewarping function $F_{k}$ for single (left), per type (middle) and per feature noise schedules (right) on the acsincome data.
-Figure 7: (Left): Learned noise schedules for acsincome. This reflects $F_{\mathrm{d.a.log},k}^{-1}$. (Right): Implicit weighting of noise levels / timesteps. This visualizes $f_{\mathrm{d.a.log},k}$.
 Figure 9: Illustration of the goodness of fit of the timewarping function $F_{k}$ for single (left), per type (middle) and per feature noise schedules (right) on the churn data.
-Appendix P Qualitative Comparisons
+Appendix Q Qualitative Comparisons
 Figure 10: Bivariate density for age and educational-num from the adult data.
@@
 Figure 12: Comparison of some univariate distributions for beijing, default, lending, news. (Note that CoDi is prohibitively expensive to train on lending and therefore excluded.)
-Appendix Q Visualizations of Captured Correlations
+Appendix R Visualizations of Captured Correlations
 Figure 13: Element-wise absolute differences of the correlation matrices between the real training set and the synthetic data for the acsincome dataset.
@@
 Continuous (cont.) and categorical (cat.) features are indicated on the axes.
-Appendix R Detailed Results
+Appendix S Detailed Results
 CoDi is prohibitively expensive to train on lending and diabetes and TabDDPM produces NaNs for acsincome and diabetes.
@@
 For those models, the performance metrics on these datasets are therefore not reported.
-They are assigned a rank of 10 in Table 1 and are not taken into account when forming the average metrics reported in Table 11.
-Table 11: Model evaluation results averaged over 11 datasets (skipping a dataset if the model was not trainable on it, which due to extensive sampling times for SMOTE includes two of the most complex datasets, acsincome and covertype) for seven benchmark models and for CDTD with three different noise schedules.
+They are assigned a rank of 10 in Table 1 and the worst metric-specific performance across all models before computing the average metrics reported in Table 11.
+Table 11: Model evaluation results averaged over 11 datasets (if a model was not trainable on a given dataset, we assign the maximum, i.e., worst, value over all models for that dataset to this model) for seven benchmark models and for CDTD with three different noise schedules.
 Per performance metric, bold indicates the best, underline the second best result.
 SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
-RMSE (abs. diff.; $\downarrow$) 0.083 0.094 0.674 0.947 0.486 0.173 0.313 0.084 0.101 0.110
-F1 (abs. diff.; $\downarrow$) 0.007 0.053 0.130 0.074 0.015 0.044 0.099 0.025 0.020 0.025
-AUC (abs. diff.; $\downarrow$) 0.008 0.020 0.080 0.065 0.009 0.027 0.059 0.018 0.016 0.022
-L2 distance of corr.  ($\downarrow$) 0.866 1.321 2.187 2.745 3.786 1.200 2.025 0.782 0.756 0.990
-Detection score ($\downarrow$) 0.661 0.934 0.986 0.976 0.769 0.936 0.877 0.796 0.768 0.783
-JSD ($\downarrow$) 0.055 0.011 0.114 0.152 0.051 0.038 0.048 0.015 0.016 0.018
-WD ($\downarrow$) 0.004 0.011 0.023 0.025 0.061 0.022 0.016 0.010 0.007 0.009
-DCR (abs. diff. to test; $\downarrow$) 1.278 1.588 3.336 1.621 0.568 1.000 2.593 0.796 0.806 0.758
+RMSE (abs. diff.; $\downarrow$) 0.366 0.091 0.635 0.868 0.674 0.278 0.309 0.086 0.091 0.102
+F1 (abs. diff.; $\downarrow$) 0.068 0.053 0.130 0.074 0.020 0.048 0.111 0.020 0.015 0.014
+AUC (abs. diff.; $\downarrow$) 0.043 0.020 0.080 0.065 0.023 0.039 0.066 0.015 0.014 0.015
+L2 distance of corr.  ($\downarrow$) 1.287 1.321 2.190 2.707 3.001 2.383 2.097 0.792 0.684 0.920
+Detection score ($\downarrow$) 0.869 0.933 0.986 0.977 0.790 0.947 0.858 0.761 0.739 0.770
+JSD ($\downarrow$) 0.077 0.011 0.101 0.135 0.086 0.073 0.052 0.015 0.016 0.018
+WD ($\downarrow$) 0.011 0.011 0.024 0.023 0.050 0.059 0.017 0.010 0.008 0.009
+DCR (abs. diff. to test; $\downarrow$) 1.813 1.588 3.317 1.602 2.061 2.838 2.648 0.927 0.867 0.760
 Table 12: L2 norm (incl.
@@
 SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
-acsincome - $0.242_{\pm 0.002}$ $1.696_{\pm 0.008}$ $1.136_{\pm 0.004}$ $0.517_{\pm 0.006}$ $0.524_{\pm 0.010}$ $0.141_{\pm 0.003}$ $0.129_{\pm 0.003}$ $0.119_{\pm 0.002}$
-adult $0.414_{\pm 0.016}$ $0.576_{\pm 0.006}$ $1.858_{\pm 0.010}$ $0.735_{\pm 0.012}$ $0.156_{\pm 0.006}$ $0.493_{\pm 0.009}$ $0.449_{\pm 0.011}$ $0.170_{\pm 0.007}$ $0.125_{\pm 0.009}$ $0.128_{\pm 0.010}$
-bank $0.404_{\pm 0.015}$ $0.819_{\pm 0.024}$ $0.947_{\pm 0.019}$ $2.758_{\pm 0.049}$ $0.898_{\pm 0.025}$ $0.499_{\pm 0.021}$ $0.677_{\pm 0.015}$ $0.323_{\pm 0.008}$ $0.266_{\pm 0.011}$ $0.256_{\pm 0.015}$
-beijing $0.081_{\pm 0.007}$ $0.133_{\pm 0.006}$ $1.445_{\pm 0.009}$ $1.642_{\pm 0.015}$ $1.133_{\pm 0.035}$ $0.363_{\pm 0.015}$ $0.096_{\pm 0.008}$ $0.075_{\pm 0.008}$ $0.073_{\pm 0.009}$ $0.071_{\pm 0.004}$
-churn $0.264_{\pm 0.036}$ $0.635_{\pm 0.026}$ $1.355_{\pm 0.043}$ $1.301_{\pm 0.041}$ $0.327_{\pm 0.044}$ $0.746_{\pm 0.062}$ $0.509_{\pm 0.053}$ $0.302_{\pm 0.041}$ $0.289_{\pm 0.043}$ $0.282_{\pm 0.044}$
-covertype - $1.192_{\pm 0.017}$ $3.685_{\pm 0.005}$ $4.668_{\pm 0.003}$ $1.044_{\pm 0.001}$ $1.029_{\pm 0.032}$ $3.958_{\pm 0.243}$ $2.359_{\pm 0.011}$ $2.275_{\pm 0.009}$ $2.710_{\pm 0.009}$
-default $0.709_{\pm 0.048}$ $1.228_{\pm 0.021}$ $2.697_{\pm 0.021}$ $1.564_{\pm 0.029}$ $3.408_{\pm 0.105}$ $1.672_{\pm 0.061}$ $1.121_{\pm 0.042}$ $0.627_{\pm 0.068}$ $0.652_{\pm 0.102}$ $0.737_{\pm 0.033}$
-diabetes $2.355_{\pm 0.026}$ $1.189_{\pm 0.004}$ $1.654_{\pm 0.008}$ $5.351_{\pm 0.095}$ $2.381_{\pm 0.026}$ $1.201_{\pm 0.020}$ $0.803_{\pm 0.032}$ $1.345_{\pm 0.016}$
-lending $1.321_{\pm 0.063}$ $3.473_{\pm 0.057}$ $2.420_{\pm 0.016}$ $5.895_{\pm 0.026}$ $10.675_{\pm 0.015}$ $6.701_{\pm 0.034}$ $1.042_{\pm 0.075}$ $1.189_{\pm 0.040}$ $1.363_{\pm 0.097}$
-news $1.684_{\pm 1.466}$ $4.333_{\pm 0.128}$ $4.641_{\pm 0.028}$ $4.612_{\pm 0.016}$ $15.985_{\pm 0.081}$ $4.874_{\pm 0.148}$ $4.990_{\pm 0.024}$ $1.925_{\pm 0.527}$ $2.035_{\pm 0.475}$ $3.395_{\pm 0.950}$
-nmes $0.565_{\pm 0.047}$ $0.717_{\pm 0.054}$ $1.663_{\pm 0.035}$ $0.532_{\pm 0.030}$ $0.447_{\pm 0.031}$ $0.609_{\pm 0.032}$ $0.867_{\pm 0.046}$ $0.433_{\pm 0.025}$ $0.478_{\pm 0.083}$ $0.481_{\pm 0.058}$
+acsincome - $0.242_{\pm 0.002}$ $1.696_{\pm 0.008}$ $1.136_{\pm 0.004}$ $0.517_{\pm 0.006}$ $0.560_{\pm 0.005}$ $0.134_{\pm 0.002}$ $0.135_{\pm 0.002}$ $0.123_{\pm 0.002}$
+adult $0.414_{\pm 0.016}$ $0.576_{\pm 0.006}$ $1.858_{\pm 0.010}$ $0.735_{\pm 0.012}$ $0.160_{\pm 0.012}$ $0.493_{\pm 0.009}$ $0.514_{\pm 0.013}$ $0.175_{\pm 0.007}$ $0.125_{\pm 0.005}$ $0.123_{\pm 0.011}$
+bank $0.404_{\pm 0.015}$ $0.819_{\pm 0.024}$ $0.947_{\pm 0.019}$ $2.758_{\pm 0.049}$ $0.529_{\pm 0.050}$ $0.499_{\pm 0.021}$ $0.759_{\pm 0.013}$ $0.333_{\pm 0.018}$ $0.231_{\pm 0.009}$ $0.288_{\pm 0.013}$
+beijing $0.081_{\pm 0.007}$ $0.128_{\pm 0.009}$ $1.470_{\pm 0.007}$ $1.226_{\pm 0.008}$ $0.368_{\pm 0.085}$ $0.373_{\pm 0.008}$ $0.086_{\pm 0.008}$ $0.073_{\pm 0.009}$ $0.073_{\pm 0.007}$ $0.071_{\pm 0.009}$
+churn $0.264_{\pm 0.036}$ $0.635_{\pm 0.026}$ $1.355_{\pm 0.043}$ $1.301_{\pm 0.041}$ $0.273_{\pm 0.069}$ $0.746_{\pm 0.062}$ $0.613_{\pm 0.022}$ $0.269_{\pm 0.028}$ $0.255_{\pm 0.021}$ $0.271_{\pm 0.040}$
+covertype - $1.192_{\pm 0.017}$ $3.685_{\pm 0.005}$ $4.668_{\pm 0.003}$ $1.124_{\pm 0.183}$ $1.029_{\pm 0.032}$ $3.749_{\pm 0.181}$ $1.970_{\pm 0.010}$ $1.357_{\pm 0.155}$ $1.972_{\pm 0.011}$
+default $0.709_{\pm 0.048}$ $1.228_{\pm 0.021}$ $2.697_{\pm 0.021}$ $1.564_{\pm 0.029}$ $0.685_{\pm 0.131}$ $1.672_{\pm 0.061}$ $1.125_{\pm 0.049}$ $0.724_{\pm 0.116}$ $0.641_{\pm 0.130}$ $0.681_{\pm 0.037}$
+diabetes $2.355_{\pm 0.026}$ $1.189_{\pm 0.004}$ $1.654_{\pm 0.008}$ $5.351_{\pm 0.095}$ $2.796_{\pm 0.066}$ $1.381_{\pm 0.016}$ $1.213_{\pm 0.029}$ $1.351_{\pm 0.044}$
+lending $1.321_{\pm 0.063}$ $3.473_{\pm 0.057}$ $2.420_{\pm 0.016}$ $5.895_{\pm 0.026}$ $10.046_{\pm 0.007}$ $6.792_{\pm 0.034}$ $1.148_{\pm 0.087}$ $1.239_{\pm 0.090}$ $1.351_{\pm 0.050}$
+news $1.684_{\pm 1.466}$ $4.333_{\pm 0.128}$ $4.641_{\pm 0.028}$ $4.612_{\pm 0.016}$ $12.356_{\pm 0.097}$ $4.874_{\pm 0.148}$ $5.153_{\pm 0.014}$ $2.050_{\pm 0.594}$ $1.811_{\pm 0.295}$ $3.446_{\pm 1.111}$
+nmes $0.565_{\pm 0.047}$ $0.717_{\pm 0.054}$ $1.663_{\pm 0.035}$ $0.532_{\pm 0.030}$ $0.426_{\pm 0.041}$ $0.609_{\pm 0.032}$ $0.919_{\pm 0.067}$ $0.454_{\pm 0.043}$ $0.444_{\pm 0.075}$ $0.445_{\pm 0.071}$
 Table 13: Jensen-Shannon divergence (incl. standard errors in subscripts) for seven benchmark models and for CDTD with three different noise schedules.
 SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
-acsincome - $0.013_{\pm 0.001}$ $0.256_{\pm 0.000}$ $0.309_{\pm 0.000}$ $0.076_{\pm 0.001}$ $0.045_{\pm 0.001}$ $0.025_{\pm 0.001}$ $0.024_{\pm 0.000}$ $0.022_{\pm 0.001}$
-adult $0.064_{\pm 0.001}$ $0.007_{\pm 0.001}$ $0.112_{\pm 0.001}$ $0.113_{\pm 0.001}$ $0.034_{\pm 0.001}$ $0.045_{\pm 0.001}$ $0.020_{\pm 0.001}$ $0.010_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.016_{\pm 0.000}$
-bank $0.039_{\pm 0.001}$ $0.004_{\pm 0.000}$ $0.086_{\pm 0.001}$ $0.191_{\pm 0.001}$ $0.029_{\pm 0.001}$ $0.038_{\pm 0.001}$ $0.054_{\pm 0.001}$ $0.010_{\pm 0.000}$ $0.009_{\pm 0.001}$ $0.012_{\pm 0.001}$
-beijing $0.006_{\pm 0.002}$ $0.005_{\pm 0.002}$ $0.147_{\pm 0.003}$ $0.257_{\pm 0.001}$ $0.035_{\pm 0.003}$ $0.018_{\pm 0.004}$ $0.007_{\pm 0.002}$ $0.003_{\pm 0.001}$ $0.005_{\pm 0.002}$ $0.005_{\pm 0.001}$
-churn $0.012_{\pm 0.004}$ $0.011_{\pm 0.004}$ $0.095_{\pm 0.003}$ $0.048_{\pm 0.004}$ $0.014_{\pm 0.004}$ $0.043_{\pm 0.001}$ $0.017_{\pm 0.002}$ $0.012_{\pm 0.003}$ $0.012_{\pm 0.002}$ $0.011_{\pm 0.002}$
-covertype - $0.002_{\pm 0.000}$ $0.044_{\pm 0.000}$ $0.043_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.049_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.011_{\pm 0.000}$
-default $0.042_{\pm 0.001}$ $0.008_{\pm 0.001}$ $0.194_{\pm 0.001}$ $0.177_{\pm 0.001}$ $0.027_{\pm 0.002}$ $0.073_{\pm 0.002}$ $0.082_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.015_{\pm 0.001}$ $0.015_{\pm 0.001}$
-diabetes $0.067_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.093_{\pm 0.000}$ $0.187_{\pm 0.000}$ $0.095_{\pm 0.000}$ $0.022_{\pm 0.000}$ $0.023_{\pm 0.000}$ $0.026_{\pm 0.000}$
-lending $0.143_{\pm 0.001}$ $0.049_{\pm 0.002}$ $0.092_{\pm 0.001}$ $0.188_{\pm 0.001}$ $0.243_{\pm 0.002}$ $0.114_{\pm 0.002}$ $0.055_{\pm 0.001}$ $0.056_{\pm 0.001}$ $0.064_{\pm 0.002}$
-news $0.063_{\pm 0.001}$ $0.002_{\pm 0.001}$ $0.022_{\pm 0.001}$ $0.128_{\pm 0.001}$ $0.046_{\pm 0.000}$ $0.012_{\pm 0.001}$ $0.016_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$
-nmes $0.060_{\pm 0.001}$ $0.008_{\pm 0.002}$ $0.117_{\pm 0.002}$ $0.029_{\pm 0.003}$ $0.028_{\pm 0.004}$ $0.027_{\pm 0.003}$ $0.026_{\pm 0.001}$ $0.008_{\pm 0.001}$ $0.009_{\pm 0.001}$ $0.013_{\pm 0.003}$
+acsincome - $0.013_{\pm 0.001}$ $0.256_{\pm 0.000}$ $0.309_{\pm 0.000}$ $0.076_{\pm 0.001}$ $0.052_{\pm 0.001}$ $0.021_{\pm 0.001}$ $0.022_{\pm 0.001}$ $0.019_{\pm 0.000}$
+adult $0.064_{\pm 0.001}$ $0.007_{\pm 0.001}$ $0.112_{\pm 0.001}$ $0.113_{\pm 0.001}$ $0.035_{\pm 0.001}$ $0.045_{\pm 0.001}$ $0.022_{\pm 0.001}$ $0.011_{\pm 0.001}$ $0.015_{\pm 0.001}$ $0.015_{\pm 0.001}$
+bank $0.039_{\pm 0.001}$ $0.004_{\pm 0.000}$ $0.086_{\pm 0.001}$ $0.191_{\pm 0.001}$ $0.021_{\pm 0.001}$ $0.038_{\pm 0.001}$ $0.063_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.011_{\pm 0.001}$ $0.015_{\pm 0.001}$
+beijing $0.006_{\pm 0.002}$ $0.004_{\pm 0.001}$ $0.005_{\pm 0.002}$ $0.074_{\pm 0.003}$ $0.024_{\pm 0.002}$ $0.011_{\pm 0.003}$ $0.011_{\pm 0.003}$ $0.006_{\pm 0.002}$ $0.006_{\pm 0.001}$ $0.007_{\pm 0.002}$
+churn $0.012_{\pm 0.004}$ $0.011_{\pm 0.004}$ $0.095_{\pm 0.003}$ $0.048_{\pm 0.004}$ $0.015_{\pm 0.006}$ $0.043_{\pm 0.001}$ $0.031_{\pm 0.002}$ $0.011_{\pm 0.002}$ $0.010_{\pm 0.002}$ $0.012_{\pm 0.003}$
+covertype - $0.002_{\pm 0.000}$ $0.044_{\pm 0.000}$ $0.043_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.044_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.007_{\pm 0.000}$
+default $0.042_{\pm 0.001}$ $0.008_{\pm 0.001}$ $0.194_{\pm 0.001}$ $0.177_{\pm 0.001}$ $0.028_{\pm 0.001}$ $0.073_{\pm 0.002}$ $0.093_{\pm 0.001}$ $0.012_{\pm 0.002}$ $0.016_{\pm 0.001}$ $0.017_{\pm 0.001}$
+diabetes $0.067_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.093_{\pm 0.000}$ $0.187_{\pm 0.000}$ $0.098_{\pm 0.000}$ $0.024_{\pm 0.000}$ $0.025_{\pm 0.000}$ $0.031_{\pm 0.000}$
+lending $0.143_{\pm 0.001}$ $0.049_{\pm 0.002}$ $0.092_{\pm 0.001}$ $0.188_{\pm 0.001}$ $0.287_{\pm 0.002}$ $0.119_{\pm 0.001}$ $0.056_{\pm 0.001}$ $0.057_{\pm 0.001}$ $0.065_{\pm 0.002}$
+news $0.063_{\pm 0.001}$ $0.002_{\pm 0.001}$ $0.022_{\pm 0.001}$ $0.128_{\pm 0.001}$ $0.017_{\pm 0.001}$ $0.012_{\pm 0.001}$ $0.017_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$
+nmes $0.060_{\pm 0.001}$ $0.008_{\pm 0.002}$ $0.117_{\pm 0.002}$ $0.029_{\pm 0.003}$ $0.025_{\pm 0.004}$ $0.027_{\pm 0.003}$ $0.019_{\pm 0.001}$ $0.009_{\pm 0.001}$ $0.007_{\pm 0.002}$ $0.012_{\pm 0.004}$
 Table 14: Wasserstein distance (incl. standard errors in subscripts) for seven benchmark models and for CDTD with three different noise schedules.
@@
 acsincome - $0.007_{\pm 0.000}$ $0.037_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.017_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.001_{\pm 0.000}$ $0.001_{\pm 0.000}$
-adult $0.003_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.016_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$
-bank $0.002_{\pm 0.001}$ $0.012_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.040_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.030_{\pm 0.001}$ $0.005_{\pm 0.000}$ $0.006_{\pm 0.001}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$
-beijing $0.002_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.058_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.019_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.002_{\pm 0.000}$
-churn $0.006_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.027_{\pm 0.001}$ $0.032_{\pm 0.001}$ $0.008_{\pm 0.002}$ $0.048_{\pm 0.002}$ $0.013_{\pm 0.002}$ $0.008_{\pm 0.001}$ $0.007_{\pm 0.001}$ $0.006_{\pm 0.001}$
-covertype - $0.006_{\pm 0.000}$ $0.041_{\pm 0.000}$ $0.022_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.017_{\pm 0.000}$ $0.017_{\pm 0.000}$ $0.015_{\pm 0.000}$ $0.012_{\pm 0.000}$
-default $0.002_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.011_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$
-diabetes $0.004_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.020_{\pm 0.000}$ $0.038_{\pm 0.000}$ $0.011_{\pm 0.000}$ $0.038_{\pm 0.000}$ $0.020_{\pm 0.000}$ $0.042_{\pm 0.000}$
-lending $0.006_{\pm 0.000}$ $0.013_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.016_{\pm 0.000}$ $0.425_{\pm 0.001}$ $0.050_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.010_{\pm 0.000}$ $0.011_{\pm 0.000}$
-news $0.007_{\pm 0.000}$ $0.024_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.018_{\pm 0.000}$ $0.078_{\pm 0.001}$ $0.030_{\pm 0.000}$ $0.025_{\pm 0.000}$ $0.007_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.008_{\pm 0.000}$
-nmes $0.005_{\pm 0.001}$ $0.012_{\pm 0.000}$ $0.036_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.007_{\pm 0.001}$ $0.016_{\pm 0.001}$ $0.038_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.000}$
+adult $0.003_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.016_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.007_{\pm 0.000}$ $0.007_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.002_{\pm 0.000}$
+bank $0.002_{\pm 0.001}$ $0.012_{\pm 0.000}$ $0.021_{\pm 0.000}$ $0.040_{\pm 0.001}$ $0.004_{\pm 0.000}$ $0.030_{\pm 0.001}$ $0.006_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.003_{\pm 0.001}$ $0.007_{\pm 0.001}$
+beijing $0.002_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.030_{\pm 0.000}$ $0.036_{\pm 0.000}$ $0.007_{\pm 0.000}$ $0.019_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.002_{\pm 0.000}$
+churn $0.006_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.027_{\pm 0.001}$ $0.032_{\pm 0.001}$ $0.007_{\pm 0.002}$ $0.048_{\pm 0.002}$ $0.013_{\pm 0.002}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.001}$
+covertype - $0.006_{\pm 0.000}$ $0.041_{\pm 0.000}$ $0.022_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.019_{\pm 0.000}$ $0.018_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.013_{\pm 0.000}$
+default $0.002_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.011_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.003_{\pm 0.000}$
+diabetes $0.004_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.020_{\pm 0.000}$ $0.038_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.042_{\pm 0.000}$ $0.034_{\pm 0.000}$ $0.041_{\pm 0.000}$
+lending $0.006_{\pm 0.000}$ $0.013_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.016_{\pm 0.000}$ $0.410_{\pm 0.001}$ $0.053_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.011_{\pm 0.000}$
+news $0.007_{\pm 0.000}$ $0.024_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.018_{\pm 0.000}$ $0.033_{\pm 0.001}$ $0.030_{\pm 0.000}$ $0.029_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.008_{\pm 0.000}$
+nmes $0.005_{\pm 0.001}$ $0.012_{\pm 0.000}$ $0.036_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.007_{\pm 0.001}$ $0.016_{\pm 0.001}$ $0.032_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.000}$ $0.006_{\pm 0.000}$
 Table 15: Detection score (incl. standard errors in subscripts) for seven benchmark models and for CDTD with three different noise schedules.
 SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
-acsincome - $0.808_{\pm 0.001}$ $0.989_{\pm 0.001}$ $0.985_{\pm 0.000}$ $0.825_{\pm 0.002}$ $0.680_{\pm 0.002}$ $0.540_{\pm 0.003}$ $0.532_{\pm 0.004}$ $0.526_{\pm 0.002}$
-adult $0.320_{\pm 0.006}$ $0.889_{\pm 0.002}$ $0.997_{\pm 0.000}$ $0.967_{\pm 0.001}$ $0.590_{\pm 0.003}$ $0.992_{\pm 0.001}$ $0.630_{\pm 0.003}$ $0.604_{\pm 0.002}$ $0.588_{\pm 0.002}$ $0.591_{\pm 0.005}$
-bank $0.633_{\pm 0.008}$ $0.955_{\pm 0.002}$ $1.000_{\pm 0.000}$ $0.988_{\pm 0.001}$ $0.783_{\pm 0.003}$ $1.000_{\pm 0.000}$ $0.843_{\pm 0.002}$ $0.795_{\pm 0.003}$ $0.739_{\pm 0.003}$ $0.694_{\pm 0.006}$
-beijing $0.976_{\pm 0.001}$ $0.995_{\pm 0.000}$ $0.998_{\pm 0.000}$ $0.993_{\pm 0.001}$ $0.966_{\pm 0.002}$ $0.997_{\pm 0.001}$ $0.966_{\pm 0.001}$ $0.951_{\pm 0.002}$ $0.949_{\pm 0.001}$ $0.947_{\pm 0.002}$
-churn $0.339_{\pm 0.020}$ $0.853_{\pm 0.002}$ $0.945_{\pm 0.006}$ $0.843_{\pm 0.011}$ $0.561_{\pm 0.005}$ $0.730_{\pm 0.012}$ $0.865_{\pm 0.012}$ $0.621_{\pm 0.016}$ $0.533_{\pm 0.007}$ $0.544_{\pm 0.031}$
-covertype - $0.945_{\pm 0.002}$ $0.997_{\pm 0.000}$ $0.989_{\pm 0.001}$ $0.586_{\pm 0.002}$ $0.900_{\pm 0.002}$ $0.979_{\pm 0.001}$ $0.991_{\pm 0.001}$ $0.992_{\pm 0.001}$ $0.991_{\pm 0.001}$
-default $0.493_{\pm 0.009}$ $0.991_{\pm 0.001}$ $0.998_{\pm 0.001}$ $0.997_{\pm 0.001}$ $0.821_{\pm 0.002}$ $0.995_{\pm 0.000}$ $0.902_{\pm 0.001}$ $0.827_{\pm 0.004}$ $0.802_{\pm 0.003}$ $0.871_{\pm 0.001}$
-diabetes $0.367_{\pm 0.001}$ $0.854_{\pm 0.002}$ $0.935_{\pm 0.002}$ $0.997_{\pm 0.001}$ $0.940_{\pm 0.001}$ $0.858_{\pm 0.001}$ $0.780_{\pm 0.002}$ $0.866_{\pm 0.002}$
-lending $0.926_{\pm 0.004}$ $0.997_{\pm 0.001}$ $0.995_{\pm 0.002}$ $0.995_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.998_{\pm 0.001}$ $0.955_{\pm 0.006}$ $0.954_{\pm 0.009}$ $0.961_{\pm 0.004}$
-news $0.993_{\pm 0.001}$ $0.998_{\pm 0.000}$ $1.000_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.966_{\pm 0.002}$ $1.000_{\pm 0.000}$ $0.999_{\pm 0.000}$ $0.973_{\pm 0.001}$ $0.953_{\pm 0.001}$ $0.977_{\pm 0.001}$
-nmes $0.905_{\pm 0.007}$ $0.987_{\pm 0.002}$ $0.992_{\pm 0.003}$ $0.988_{\pm 0.002}$ $0.650_{\pm 0.014}$ $0.988_{\pm 0.000}$ $0.841_{\pm 0.008}$ $0.636_{\pm 0.008}$ $0.623_{\pm 0.008}$ $0.642_{\pm 0.010}$
+acsincome - $0.808_{\pm 0.001}$ $0.989_{\pm 0.001}$ $0.985_{\pm 0.000}$ $0.825_{\pm 0.002}$ $0.688_{\pm 0.003}$ $0.529_{\pm 0.005}$ $0.527_{\pm 0.002}$ $0.528_{\pm 0.002}$
+adult $0.687_{\pm 0.003}$ $0.889_{\pm 0.002}$ $0.997_{\pm 0.000}$ $0.967_{\pm 0.001}$ $0.594_{\pm 0.003}$ $0.992_{\pm 0.001}$ $0.641_{\pm 0.003}$ $0.621_{\pm 0.002}$ $0.590_{\pm 0.004}$ $0.605_{\pm 0.004}$
+bank $0.839_{\pm 0.003}$ $0.955_{\pm 0.002}$ $1.000_{\pm 0.000}$ $0.988_{\pm 0.001}$ $0.781_{\pm 0.002}$ $1.000_{\pm 0.000}$ $0.853_{\pm 0.003}$ $0.808_{\pm 0.003}$ $0.701_{\pm 0.005}$ $0.835_{\pm 0.001}$
+beijing $0.938_{\pm 0.002}$ $0.989_{\pm 0.002}$ $0.996_{\pm 0.001}$ $0.995_{\pm 0.001}$ $0.738_{\pm 0.004}$ $0.989_{\pm 0.001}$ $0.723_{\pm 0.003}$ $0.574_{\pm 0.003}$ $0.620_{\pm 0.005}$ $0.614_{\pm 0.003}$
+churn $0.567_{\pm 0.015}$ $0.853_{\pm 0.002}$ $0.945_{\pm 0.006}$ $0.843_{\pm 0.011}$ $0.556_{\pm 0.013}$ $0.730_{\pm 0.012}$ $0.859_{\pm 0.005}$ $0.614_{\pm 0.016}$ $0.541_{\pm 0.008}$ $0.639_{\pm 0.011}$
+covertype - $0.945_{\pm 0.002}$ $0.997_{\pm 0.000}$ $0.989_{\pm 0.001}$ $0.584_{\pm 0.002}$ $0.900_{\pm 0.002}$ $0.991_{\pm 0.000}$ $0.989_{\pm 0.000}$ $0.981_{\pm 0.000}$ $0.989_{\pm 0.001}$
+default $0.928_{\pm 0.004}$ $0.991_{\pm 0.001}$ $0.998_{\pm 0.001}$ $0.997_{\pm 0.001}$ $0.827_{\pm 0.005}$ $0.995_{\pm 0.000}$ $0.914_{\pm 0.002}$ $0.823_{\pm 0.001}$ $0.793_{\pm 0.005}$ $0.827_{\pm 0.003}$
+diabetes $0.726_{\pm 0.001}$ $0.854_{\pm 0.002}$ $0.935_{\pm 0.002}$ $0.997_{\pm 0.001}$ $0.946_{\pm 0.001}$ $0.864_{\pm 0.002}$ $0.837_{\pm 0.001}$ $0.862_{\pm 0.001}$
+lending $0.966_{\pm 0.003}$ $0.997_{\pm 0.001}$ $0.995_{\pm 0.002}$ $0.995_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.998_{\pm 0.000}$ $0.959_{\pm 0.009}$ $0.955_{\pm 0.003}$ $0.960_{\pm 0.005}$
+news $0.998_{\pm 0.000}$ $0.998_{\pm 0.000}$ $1.000_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.974_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.999_{\pm 0.000}$ $0.947_{\pm 0.002}$ $0.950_{\pm 0.002}$ $0.972_{\pm 0.002}$
+nmes $0.926_{\pm 0.007}$ $0.987_{\pm 0.002}$ $0.992_{\pm 0.003}$ $0.988_{\pm 0.002}$ $0.652_{\pm 0.011}$ $0.988_{\pm 0.000}$ $0.829_{\pm 0.010}$ $0.646_{\pm 0.005}$ $0.633_{\pm 0.010}$ $0.636_{\pm 0.010}$
 Table 16: Distance to closest record of the generated data (incl. standard errors in subscripts) for seven benchmark models and for CDTD with three different noise schedules.
 Test Set SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
-acsincome $7.673_{\pm 0.017}$ $8.637_{\pm 0.027}$ $10.758_{\pm 0.054}$ $6.652_{\pm 0.032}$ $10.877_{\pm 0.092}$ $10.305_{\pm 0.073}$ $8.346_{\pm 0.056}$ $8.322_{\pm 0.047}$ $8.349_{\pm 0.033}$
-adult $1.870_{\pm 0.000}$ $1.371_{\pm 0.018}$ $2.523_{\pm 0.012}$ $5.012_{\pm 0.028}$ $2.227_{\pm 0.013}$ $1.647_{\pm 0.009}$ $2.735_{\pm 0.028}$ $2.341_{\pm 0.013}$ $1.112_{\pm 0.019}$ $1.231_{\pm 0.011}$ $1.294_{\pm 0.009}$
-bank $2.369_{\pm 0.000}$ $1.369_{\pm 0.011}$ $3.025_{\pm 0.017}$ $3.840_{\pm 0.014}$ $3.136_{\pm 0.007}$ $2.327_{\pm 0.010}$ $3.062_{\pm 0.012}$ $2.973_{\pm 0.012}$ $1.828_{\pm 0.008}$ $1.943_{\pm 0.007}$ $2.062_{\pm 0.008}$
-beijing $0.385_{\pm 0.000}$ $0.139_{\pm 0.003}$ $0.735_{\pm 0.003}$ $1.004_{\pm 0.006}$ $0.926_{\pm 0.003}$ $0.739_{\pm 0.006}$ $0.610_{\pm 0.002}$ $0.626_{\pm 0.001}$ $0.490_{\pm 0.002}$ $0.489_{\pm 0.001}$ $0.477_{\pm 0.002}$
-churn $0.347_{\pm 0.000}$ $0.232_{\pm 0.028}$ $1.136_{\pm 0.015}$ $1.804_{\pm 0.036}$ $1.146_{\pm 0.039}$ $0.342_{\pm 0.031}$ $0.852_{\pm 0.016}$ $1.130_{\pm 0.018}$ $0.332_{\pm 0.021}$ $0.274_{\pm 0.021}$ $0.276_{\pm 0.012}$
-covertype $0.529_{\pm 0.001}$ $1.741_{\pm 0.011}$ $5.773_{\pm 0.017}$ $3.173_{\pm 0.013}$ $0.889_{\pm 0.007}$ $1.508_{\pm 0.020}$ $3.086_{\pm 0.009}$ $2.297_{\pm 0.026}$ $2.209_{\pm 0.022}$ $2.252_{\pm 0.013}$
-default $1.812_{\pm 0.000}$ $1.032_{\pm 0.010}$ $3.095_{\pm 0.026}$ $5.880_{\pm 0.020}$ $3.216_{\pm 0.013}$ $1.422_{\pm 0.013}$ $2.593_{\pm 0.020}$ $2.603_{\pm 0.018}$ $1.127_{\pm 0.028}$ $1.269_{\pm 0.014}$ $1.253_{\pm 0.012}$
-diabetes $15.608_{\pm 0.055}$ $13.909_{\pm 0.050}$ $17.736_{\pm 0.107}$ $21.935_{\pm 0.046}$ $8.214_{\pm 0.022}$ $28.955_{\pm 0.060}$ $15.279_{\pm 0.026}$ $15.126_{\pm 0.058}$ $15.350_{\pm 0.059}$
-lending $11.184_{\pm 0.000}$ $17.752_{\pm 0.143}$ $17.776_{\pm 0.132}$ $20.239_{\pm 0.222}$ $10.688_{\pm 0.025}$ $12.537_{\pm 0.076}$ $16.222_{\pm 0.092}$ $13.775_{\pm 0.147}$ $14.162_{\pm 0.188}$ $13.966_{\pm 0.282}$
-news $3.615_{\pm 0.000}$ $3.553_{\pm 0.134}$ $6.147_{\pm 0.010}$ $4.789_{\pm 0.005}$ $5.821_{\pm 0.003}$ $4.960_{\pm 0.006}$ $4.661_{\pm 0.023}$ $5.351_{\pm 0.008}$ $3.635_{\pm 0.004}$ $3.687_{\pm 0.006}$ $3.749_{\pm 0.048}$
-nmes $1.931_{\pm 0.000}$ $1.394_{\pm 0.019}$ $2.203_{\pm 0.028}$ $2.971_{\pm 0.008}$ $1.710_{\pm 0.019}$ $0.891_{\pm 0.033}$ $1.231_{\pm 0.024}$ $2.260_{\pm 0.034}$ $0.664_{\pm 0.029}$ $0.710_{\pm 0.032}$ $0.771_{\pm 0.023}$
+acsincome $7.673_{\pm 0.017}$ $8.637_{\pm 0.027}$ $10.758_{\pm 0.054}$ $6.652_{\pm 0.032}$ $10.877_{\pm 0.092}$ $10.797_{\pm 0.101}$ $8.337_{\pm 0.053}$ $8.397_{\pm 0.062}$ $8.344_{\pm 0.027}$
+adult $1.870_{\pm 0.000}$ $1.371_{\pm 0.018}$ $2.523_{\pm 0.012}$ $5.012_{\pm 0.028}$ $2.227_{\pm 0.013}$ $1.656_{\pm 0.008}$ $2.735_{\pm 0.028}$ $2.408_{\pm 0.031}$ $1.138_{\pm 0.015}$ $1.327_{\pm 0.015}$ $1.334_{\pm 0.012}$
+bank $2.369_{\pm 0.000}$ $1.369_{\pm 0.011}$ $3.025_{\pm 0.017}$ $3.840_{\pm 0.014}$ $3.136_{\pm 0.007}$ $2.211_{\pm 0.011}$ $3.062_{\pm 0.012}$ $3.022_{\pm 0.009}$ $1.749_{\pm 0.008}$ $1.913_{\pm 0.006}$ $1.997_{\pm 0.008}$
+beijing $0.385_{\pm 0.000}$ $0.139_{\pm 0.003}$ $0.731_{\pm 0.003}$ $0.800_{\pm 0.002}$ $0.724_{\pm 0.005}$ $0.639_{\pm 0.002}$ $0.588_{\pm 0.003}$ $0.633_{\pm 0.002}$ $0.474_{\pm 0.002}$ $0.473_{\pm 0.002}$ $0.469_{\pm 0.001}$
+churn $0.347_{\pm 0.000}$ $0.232_{\pm 0.028}$ $1.136_{\pm 0.015}$ $1.804_{\pm 0.036}$ $1.146_{\pm 0.039}$ $0.368_{\pm 0.044}$ $0.852_{\pm 0.016}$ $1.209_{\pm 0.010}$ $0.329_{\pm 0.010}$ $0.278_{\pm 0.012}$ $0.350_{\pm 0.023}$
+covertype $0.529_{\pm 0.001}$ $1.741_{\pm 0.011}$ $5.773_{\pm 0.017}$ $3.173_{\pm 0.013}$ $0.877_{\pm 0.008}$ $1.508_{\pm 0.020}$ $3.033_{\pm 0.012}$ $1.825_{\pm 0.016}$ $1.594_{\pm 0.009}$ $1.805_{\pm 0.009}$
+default $1.812_{\pm 0.000}$ $1.032_{\pm 0.010}$ $3.095_{\pm 0.026}$ $5.880_{\pm 0.020}$ $3.216_{\pm 0.013}$ $1.437_{\pm 0.020}$ $2.593_{\pm 0.020}$ $2.801_{\pm 0.032}$ $1.192_{\pm 0.022}$ $1.355_{\pm 0.017}$ $1.352_{\pm 0.016}$
+diabetes $15.608_{\pm 0.055}$ $13.909_{\pm 0.050}$ $17.736_{\pm 0.107}$ $21.935_{\pm 0.046}$ $8.214_{\pm 0.022}$ $28.794_{\pm 0.054}$ $14.356_{\pm 0.050}$ $14.468_{\pm 0.028}$ $14.866_{\pm 0.033}$
+lending $11.184_{\pm 0.000}$ $17.752_{\pm 0.143}$ $17.776_{\pm 0.132}$ $20.239_{\pm 0.222}$ $10.688_{\pm 0.025}$ $14.310_{\pm 0.093}$ $16.239_{\pm 0.052}$ $14.958_{\pm 0.292}$ $14.962_{\pm 0.090}$ $14.146_{\pm 0.240}$
+news $3.615_{\pm 0.000}$ $3.553_{\pm 0.134}$ $6.147_{\pm 0.010}$ $4.789_{\pm 0.005}$ $5.821_{\pm 0.003}$ $4.358_{\pm 0.013}$ $4.661_{\pm 0.023}$ $5.410_{\pm 0.005}$ $3.615_{\pm 0.009}$ $3.676_{\pm 0.008}$ $3.736_{\pm 0.094}$
+nmes $1.931_{\pm 0.000}$ $1.394_{\pm 0.019}$ $2.203_{\pm 0.028}$ $2.971_{\pm 0.008}$ $1.710_{\pm 0.019}$ $0.890_{\pm 0.027}$ $1.231_{\pm 0.024}$ $2.105_{\pm 0.022}$ $0.801_{\pm 0.017}$ $0.780_{\pm 0.031}$ $0.803_{\pm 0.017}$
 Table 17: Machine learning efficiency F1 score for seven benchmark models, the real training data and for CDTD with three different noise schedules.
@@
 Real Data SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
-adult $0.797_{\pm 0.000}$ $0.784_{\pm 0.001}$ $0.769_{\pm 0.002}$ $0.647_{\pm 0.015}$ $0.756_{\pm 0.002}$ $0.787_{\pm 0.001}$ $0.745_{\pm 0.004}$ $0.782_{\pm 0.001}$ $0.787_{\pm 0.001}$ $0.787_{\pm 0.001}$ $0.787_{\pm 0.001}$
-bank $0.745_{\pm 0.002}$ $0.740_{\pm 0.004}$ $0.682_{\pm 0.006}$ $0.680_{\pm 0.006}$ $0.629_{\pm 0.006}$ $0.720_{\pm 0.006}$ $0.673_{\pm 0.006}$ $0.711_{\pm 0.007}$ $0.776_{\pm 0.003}$ $0.767_{\pm 0.004}$ $0.737_{\pm 0.004}$
-churn $0.873_{\pm 0.003}$ $0.865_{\pm 0.008}$ $0.780_{\pm 0.015}$ $0.761_{\pm 0.009}$ $0.802_{\pm 0.017}$ $0.857_{\pm 0.007}$ $0.865_{\pm 0.008}$ $0.771_{\pm 0.014}$ $0.854_{\pm 0.011}$ $0.852_{\pm 0.006}$ $0.845_{\pm 0.011}$
-covertype $0.817_{\pm 0.001}$ $0.783_{\pm 0.001}$ $0.442_{\pm 0.008}$ $0.711_{\pm 0.002}$ $0.799_{\pm 0.001}$ $0.767_{\pm 0.001}$ $0.614_{\pm 0.015}$ $0.734_{\pm 0.002}$ $0.754_{\pm 0.001}$ $0.722_{\pm 0.002}$
-default $0.674_{\pm 0.001}$ $0.677_{\pm 0.001}$ $0.627_{\pm 0.003}$ $0.686_{\pm 0.002}$ $0.632_{\pm 0.007}$ $0.678_{\pm 0.002}$ $0.638_{\pm 0.008}$ $0.496_{\pm 0.009}$ $0.670_{\pm 0.002}$ $0.671_{\pm 0.001}$ $0.673_{\pm 0.003}$
-diabetes $0.621_{\pm 0.002}$ $0.615_{\pm 0.002}$ $0.572_{\pm 0.005}$ $0.557_{\pm 0.004}$ $0.553_{\pm 0.003}$ $0.560_{\pm 0.006}$ $0.617_{\pm 0.002}$ $0.617_{\pm 0.002}$ $0.611_{\pm 0.002}$
+adult $0.797_{\pm 0.000}$ $0.784_{\pm 0.001}$ $0.769_{\pm 0.002}$ $0.647_{\pm 0.015}$ $0.756_{\pm 0.002}$ $0.788_{\pm 0.001}$ $0.745_{\pm 0.004}$ $0.782_{\pm 0.002}$ $0.788_{\pm 0.001}$ $0.789_{\pm 0.002}$ $0.789_{\pm 0.002}$
+bank $0.745_{\pm 0.002}$ $0.740_{\pm 0.004}$ $0.682_{\pm 0.006}$ $0.680_{\pm 0.006}$ $0.629_{\pm 0.006}$ $0.744_{\pm 0.005}$ $0.673_{\pm 0.006}$ $0.661_{\pm 0.008}$ $0.782_{\pm 0.003}$ $0.766_{\pm 0.005}$ $0.747_{\pm 0.004}$
+churn $0.873_{\pm 0.003}$ $0.865_{\pm 0.008}$ $0.780_{\pm 0.015}$ $0.761_{\pm 0.009}$ $0.802_{\pm 0.017}$ $0.855_{\pm 0.012}$ $0.865_{\pm 0.008}$ $0.748_{\pm 0.015}$ $0.859_{\pm 0.007}$ $0.863_{\pm 0.007}$ $0.860_{\pm 0.007}$
+covertype $0.817_{\pm 0.001}$ $0.783_{\pm 0.001}$ $0.442_{\pm 0.008}$ $0.711_{\pm 0.002}$ $0.799_{\pm 0.001}$ $0.767_{\pm 0.001}$ $0.620_{\pm 0.014}$ $0.766_{\pm 0.001}$ $0.767_{\pm 0.001}$ $0.765_{\pm 0.001}$
+default $0.674_{\pm 0.001}$ $0.677_{\pm 0.001}$ $0.627_{\pm 0.003}$ $0.686_{\pm 0.002}$ $0.632_{\pm 0.007}$ $0.680_{\pm 0.002}$ $0.638_{\pm 0.008}$ $0.485_{\pm 0.016}$ $0.673_{\pm 0.003}$ $0.675_{\pm 0.002}$ $0.677_{\pm 0.002}$
+diabetes $0.621_{\pm 0.002}$ $0.615_{\pm 0.002}$ $0.572_{\pm 0.005}$ $0.557_{\pm 0.004}$ $0.553_{\pm 0.003}$ $0.566_{\pm 0.006}$ $0.614_{\pm 0.003}$ $0.619_{\pm 0.002}$ $0.614_{\pm 0.003}$
 Table 18: Machine learning efficiency AUC score for seven benchmark models, the real training data and for CDTD with three different noise schedules.
@@
 Real Data SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
-adult $0.915_{\pm 0.000}$ $0.906_{\pm 0.001}$ $0.901_{\pm 0.000}$ $0.836_{\pm 0.006}$ $0.889_{\pm 0.002}$ $0.908_{\pm 0.000}$ $0.880_{\pm 0.005}$ $0.906_{\pm 0.001}$ $0.910_{\pm 0.000}$ $0.910_{\pm 0.001}$ $0.909_{\pm 0.000}$
-bank $0.947_{\pm 0.000}$ $0.943_{\pm 0.001}$ $0.938_{\pm 0.001}$ $0.934_{\pm 0.003}$ $0.830_{\pm 0.020}$ $0.940_{\pm 0.005}$ $0.929_{\pm 0.005}$ $0.939_{\pm 0.003}$ $0.945_{\pm 0.000}$ $0.945_{\pm 0.001}$ $0.943_{\pm 0.004}$
-churn $0.964_{\pm 0.001}$ $0.961_{\pm 0.002}$ $0.939_{\pm 0.007}$ $0.882_{\pm 0.006}$ $0.948_{\pm 0.004}$ $0.957_{\pm 0.004}$ $0.961_{\pm 0.001}$ $0.919_{\pm 0.006}$ $0.962_{\pm 0.001}$ $0.962_{\pm 0.001}$ $0.959_{\pm 0.003}$
-covertype $0.892_{\pm 0.000}$ $0.860_{\pm 0.001}$ $0.677_{\pm 0.007}$ $0.777_{\pm 0.001}$ $0.876_{\pm 0.000}$ $0.845_{\pm 0.001}$ $0.671_{\pm 0.013}$ $0.816_{\pm 0.002}$ $0.828_{\pm 0.001}$ $0.802_{\pm 0.002}$
-default $0.768_{\pm 0.000}$ $0.759_{\pm 0.003}$ $0.754_{\pm 0.002}$ $0.744_{\pm 0.002}$ $0.751_{\pm 0.004}$ $0.763_{\pm 0.002}$ $0.739_{\pm 0.008}$ $0.746_{\pm 0.011}$ $0.762_{\pm 0.003}$ $0.765_{\pm 0.002}$ $0.765_{\pm 0.002}$
-diabetes $0.693_{\pm 0.001}$ $0.679_{\pm 0.001}$ $0.669_{\pm 0.002}$ $0.626_{\pm 0.003}$ $0.592_{\pm 0.002}$ $0.645_{\pm 0.002}$ $0.675_{\pm 0.001}$ $0.673_{\pm 0.001}$ $0.667_{\pm 0.001}$
+adult $0.915_{\pm 0.000}$ $0.906_{\pm 0.001}$ $0.901_{\pm 0.000}$ $0.836_{\pm 0.006}$ $0.889_{\pm 0.002}$ $0.909_{\pm 0.000}$ $0.880_{\pm 0.005}$ $0.906_{\pm 0.001}$ $0.909_{\pm 0.000}$ $0.909_{\pm 0.001}$ $0.908_{\pm 0.001}$
+bank $0.947_{\pm 0.000}$ $0.943_{\pm 0.001}$ $0.938_{\pm 0.001}$ $0.934_{\pm 0.003}$ $0.830_{\pm 0.020}$ $0.942_{\pm 0.004}$ $0.929_{\pm 0.005}$ $0.922_{\pm 0.006}$ $0.945_{\pm 0.000}$ $0.946_{\pm 0.000}$ $0.944_{\pm 0.003}$
+churn $0.964_{\pm 0.001}$ $0.961_{\pm 0.002}$ $0.939_{\pm 0.007}$ $0.882_{\pm 0.006}$ $0.948_{\pm 0.004}$ $0.957_{\pm 0.006}$ $0.961_{\pm 0.001}$ $0.911_{\pm 0.013}$ $0.960_{\pm 0.002}$ $0.957_{\pm 0.008}$ $0.960_{\pm 0.003}$
+covertype $0.892_{\pm 0.000}$ $0.860_{\pm 0.001}$ $0.677_{\pm 0.007}$ $0.777_{\pm 0.001}$ $0.876_{\pm 0.001}$ $0.845_{\pm 0.001}$ $0.675_{\pm 0.013}$ $0.840_{\pm 0.001}$ $0.844_{\pm 0.000}$ $0.840_{\pm 0.001}$
+default $0.768_{\pm 0.000}$ $0.759_{\pm 0.003}$ $0.754_{\pm 0.002}$ $0.744_{\pm 0.002}$ $0.751_{\pm 0.004}$ $0.765_{\pm 0.002}$ $0.739_{\pm 0.008}$ $0.732_{\pm 0.021}$ $0.763_{\pm 0.002}$ $0.764_{\pm 0.002}$ $0.765_{\pm 0.002}$
+diabetes $0.693_{\pm 0.001}$ $0.679_{\pm 0.001}$ $0.669_{\pm 0.002}$ $0.626_{\pm 0.003}$ $0.592_{\pm 0.002}$ $0.642_{\pm 0.002}$ $0.671_{\pm 0.002}$ $0.673_{\pm 0.002}$ $0.672_{\pm 0.002}$
 Table 19: Machine learning efficiency RMSE for seven benchmark models, the real training data and for CDTD with three different noise schedules.
@@
 Real Data SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
-acsincome $0.804_{\pm 0.012}$ $0.757_{\pm 0.007}$ $2.292_{\pm 0.013}$ $1.054_{\pm 0.011}$ $0.857_{\pm 0.010}$ $0.959_{\pm 0.012}$ $0.838_{\pm 0.015}$ $0.811_{\pm 0.014}$ $0.820_{\pm 0.011}$
-beijing $0.712_{\pm 0.001}$ $0.739_{\pm 0.002}$ $0.792_{\pm 0.007}$ $1.246_{\pm 0.010}$ $1.690_{\pm 0.016}$ $0.606_{\pm 0.006}$ $0.912_{\pm 0.005}$ $0.788_{\pm 0.011}$ $0.774_{\pm 0.005}$ $0.770_{\pm 0.005}$ $0.762_{\pm 0.005}$
-lending $0.030_{\pm 0.000}$ $0.042_{\pm 0.001}$ $0.274_{\pm 0.007}$ $0.137_{\pm 0.007}$ $0.404_{\pm 0.007}$ $0.795_{\pm 0.031}$ $0.268_{\pm 0.004}$ $0.061_{\pm 0.001}$ $0.060_{\pm 0.001}$ $0.066_{\pm 0.002}$
-news $1.001_{\pm 0.002}$ $1.180_{\pm 0.107}$ $0.923_{\pm 0.052}$ $1.906_{\pm 0.019}$ $3.999_{\pm 0.175}$ $0.083_{\pm 0.001}$ $1.302_{\pm 0.074}$ $0.374_{\pm 0.028}$ $0.819_{\pm 0.103}$ $0.776_{\pm 0.091}$ $0.755_{\pm 0.066}$
-nmes $1.001_{\pm 0.003}$ $1.112_{\pm 0.044}$ $0.972_{\pm 0.024}$ $1.331_{\pm 0.052}$ $1.127_{\pm 0.047}$ $1.154_{\pm 0.047}$ $1.137_{\pm 0.052}$ $0.535_{\pm 0.013}$ $1.108_{\pm 0.083}$ $1.184_{\pm 0.076}$ $1.203_{\pm 0.081}$
-Appendix S Ablation Study Details
+acsincome $0.804_{\pm 0.012}$ $0.757_{\pm 0.007}$ $2.292_{\pm 0.013}$ $1.054_{\pm 0.011}$ $0.857_{\pm 0.010}$ $0.955_{\pm 0.010}$ $0.827_{\pm 0.009}$ $0.807_{\pm 0.008}$ $0.812_{\pm 0.004}$
+beijing $0.711_{\pm 0.001}$ $0.742_{\pm 0.002}$ $0.779_{\pm 0.007}$ $1.050_{\pm 0.010}$ $1.295_{\pm 0.016}$ $0.799_{\pm 0.007}$ $0.849_{\pm 0.004}$ $0.789_{\pm 0.010}$ $0.772_{\pm 0.003}$ $0.766_{\pm 0.003}$ $0.762_{\pm 0.002}$
+lending $0.030_{\pm 0.000}$ $0.042_{\pm 0.001}$ $0.274_{\pm 0.007}$ $0.137_{\pm 0.007}$ $0.404_{\pm 0.007}$ $0.789_{\pm 0.033}$ $0.305_{\pm 0.006}$ $0.071_{\pm 0.001}$ $0.066_{\pm 0.002}$ $0.062_{\pm 0.002}$
+news $1.001_{\pm 0.002}$ $1.180_{\pm 0.107}$ $0.923_{\pm 0.052}$ $1.906_{\pm 0.019}$ $3.999_{\pm 0.175}$ $0.171_{\pm 0.006}$ $1.302_{\pm 0.074}$ $0.397_{\pm 0.037}$ $0.848_{\pm 0.081}$ $0.752_{\pm 0.067}$ $0.717_{\pm 0.045}$
+nmes $1.001_{\pm 0.003}$ $1.112_{\pm 0.044}$ $0.972_{\pm 0.024}$ $1.331_{\pm 0.052}$ $1.127_{\pm 0.047}$ $1.200_{\pm 0.054}$ $1.137_{\pm 0.052}$ $0.563_{\pm 0.005}$ $1.154_{\pm 0.052}$ $1.109_{\pm 0.055}$ $1.136_{\pm 0.074}$
+Appendix T Ablation Study Details
 Table 20: L2 norm (incl.
@@
 Configuration A B C D CDTD (per type)
-acsincome $0.131_{\pm 0.003}$ $0.119_{\pm 0.004}$ $0.124_{\pm 0.006}$ $0.129_{\pm 0.004}$ $0.129_{\pm 0.003}$
-adult $0.131_{\pm 0.007}$ $0.128_{\pm 0.008}$ $0.168_{\pm 0.017}$ $0.107_{\pm 0.011}$ $0.125_{\pm 0.009}$
-beijing $0.065_{\pm 0.009}$ $0.066_{\pm 0.012}$ $0.067_{\pm 0.011}$ $0.067_{\pm 0.010}$ $0.073_{\pm 0.009}$
-churn $0.244_{\pm 0.015}$ $0.272_{\pm 0.034}$ $0.299_{\pm 0.066}$ $0.264_{\pm 0.012}$ $0.289_{\pm 0.043}$
+acsincome $0.138_{\pm 0.004}$ $0.138_{\pm 0.003}$ $0.135_{\pm 0.001}$ $0.152_{\pm 0.001}$ $0.135_{\pm 0.002}$
+adult $0.135_{\pm 0.006}$ $0.113_{\pm 0.004}$ $0.159_{\pm 0.015}$ $0.100_{\pm 0.010}$ $0.125_{\pm 0.005}$
+bank $0.478_{\pm 0.019}$ $0.243_{\pm 0.015}$ $0.269_{\pm 0.012}$ $0.196_{\pm 0.012}$ $0.231_{\pm 0.009}$
+beijing $0.076_{\pm 0.009}$ $0.070_{\pm 0.007}$ $0.067_{\pm 0.004}$ $0.068_{\pm 0.006}$ $0.073_{\pm 0.007}$
+churn $0.294_{\pm 0.053}$ $0.280_{\pm 0.043}$ $0.262_{\pm 0.030}$ $0.239_{\pm 0.038}$ $0.255_{\pm 0.021}$
+covertype $1.178_{\pm 0.193}$ $2.099_{\pm 0.010}$ $1.974_{\pm 0.136}$ $1.808_{\pm 0.010}$ $1.357_{\pm 0.155}$
+default $0.905_{\pm 0.110}$ $0.799_{\pm 0.123}$ $0.727_{\pm 0.132}$ $0.521_{\pm 0.124}$ $0.641_{\pm 0.130}$
+diabetes $0.719_{\pm 0.049}$ $1.435_{\pm 0.021}$ $1.397_{\pm 0.005}$ $1.230_{\pm 0.061}$ $1.213_{\pm 0.029}$
+lending $1.480_{\pm 0.046}$ $1.127_{\pm 0.083}$ $1.178_{\pm 0.059}$ $1.295_{\pm 0.044}$ $1.239_{\pm 0.090}$
+news $2.484_{\pm 0.138}$ $2.136_{\pm 0.417}$ $1.973_{\pm 0.417}$ $2.016_{\pm 0.483}$ $1.811_{\pm 0.295}$
+nmes $0.483_{\pm 0.048}$ $0.421_{\pm 0.032}$ $0.457_{\pm 0.041}$ $0.450_{\pm 0.041}$ $0.444_{\pm 0.075}$
 Table 21: Jensen-Shannon divergence (incl. standard errors in subscripts) for five CDTD configurations with progressive addition of model components.
 Configuration A B C D CDTD (per type)
-acsincome $0.025_{\pm 0.000}$ $0.025_{\pm 0.001}$ $0.025_{\pm 0.001}$ $0.024_{\pm 0.001}$ $0.024_{\pm 0.000}$
-adult $0.012_{\pm 0.001}$ $0.013_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.014_{\pm 0.001}$ $0.013_{\pm 0.001}$
-beijing $0.004_{\pm 0.001}$ $0.006_{\pm 0.002}$ $0.005_{\pm 0.003}$ $0.004_{\pm 0.002}$ $0.005_{\pm 0.002}$
-churn $0.010_{\pm 0.002}$ $0.008_{\pm 0.002}$ $0.009_{\pm 0.004}$ $0.010_{\pm 0.002}$ $0.012_{\pm 0.002}$
+acsincome $0.016_{\pm 0.000}$ $0.022_{\pm 0.000}$ $0.026_{\pm 0.001}$ $0.032_{\pm 0.001}$ $0.022_{\pm 0.001}$
+adult $0.010_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.015_{\pm 0.001}$ $0.015_{\pm 0.001}$
+bank $0.008_{\pm 0.000}$ $0.017_{\pm 0.000}$ $0.011_{\pm 0.001}$ $0.011_{\pm 0.001}$ $0.011_{\pm 0.001}$
+beijing $0.005_{\pm 0.001}$ $0.005_{\pm 0.004}$ $0.004_{\pm 0.002}$ $0.004_{\pm 0.001}$ $0.006_{\pm 0.001}$
+churn $0.015_{\pm 0.003}$ $0.011_{\pm 0.002}$ $0.009_{\pm 0.003}$ $0.010_{\pm 0.003}$ $0.010_{\pm 0.002}$
+covertype $0.002_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$
+default $0.014_{\pm 0.001}$ $0.015_{\pm 0.001}$ $0.012_{\pm 0.002}$ $0.017_{\pm 0.002}$ $0.016_{\pm 0.001}$
+diabetes $0.024_{\pm 0.000}$ $0.030_{\pm 0.000}$ $0.022_{\pm 0.000}$ $0.024_{\pm 0.001}$ $0.025_{\pm 0.000}$
+lending $0.055_{\pm 0.001}$ $0.055_{\pm 0.002}$ $0.055_{\pm 0.001}$ $0.057_{\pm 0.001}$ $0.057_{\pm 0.001}$
+news $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.004_{\pm 0.001}$ $0.003_{\pm 0.001}$
+nmes $0.008_{\pm 0.002}$ $0.008_{\pm 0.002}$ $0.008_{\pm 0.002}$ $0.009_{\pm 0.001}$ $0.007_{\pm 0.002}$
 Table 22: Wasserstein distance (incl. standard errors in subscripts) for five CDTD configurations with progressive addition of model components.
 Configuration A B C D CDTD (per type)
-acsincome $0.002_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.001_{\pm 0.000}$ $0.001_{\pm 0.000}$
-adult $0.004_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.004_{\pm 0.000}$
-beijing $0.003_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.003_{\pm 0.000}$
-churn $0.006_{\pm 0.001}$ $0.006_{\pm 0.000}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.007_{\pm 0.001}$
+acsincome $0.005_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.001_{\pm 0.000}$ $0.001_{\pm 0.000}$
+adult $0.008_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.003_{\pm 0.000}$
+bank $0.007_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.005_{\pm 0.001}$ $0.003_{\pm 0.000}$ $0.003_{\pm 0.001}$
+beijing $0.006_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.002_{\pm 0.000}$
+churn $0.008_{\pm 0.001}$ $0.007_{\pm 0.001}$ $0.007_{\pm 0.001}$ $0.007_{\pm 0.002}$ $0.006_{\pm 0.001}$
+covertype $0.003_{\pm 0.000}$ $0.020_{\pm 0.000}$ $0.019_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.009_{\pm 0.000}$
+default $0.003_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.003_{\pm 0.000}$
+diabetes $0.021_{\pm 0.000}$ $0.042_{\pm 0.000}$ $0.041_{\pm 0.000}$ $0.033_{\pm 0.000}$ $0.034_{\pm 0.000}$
+lending $0.012_{\pm 0.000}$ $0.011_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.013_{\pm 0.000}$
+news $0.009_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.007_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.006_{\pm 0.000}$
+nmes $0.008_{\pm 0.001}$ $0.007_{\pm 0.001}$ $0.008_{\pm 0.000}$ $0.008_{\pm 0.001}$ $0.006_{\pm 0.000}$
 Table 23: Detection score (incl. standard errors in subscripts) for five CDTD configurations with progressive addition of model components.
 Configuration A B C D CDTD (per type)
-acsincome $0.534_{\pm 0.002}$ $0.534_{\pm 0.001}$ $0.538_{\pm 0.003}$ $0.532_{\pm 0.002}$ $0.532_{\pm 0.004}$
-adult $0.597_{\pm 0.002}$ $0.593_{\pm 0.001}$ $0.615_{\pm 0.003}$ $0.580_{\pm 0.003}$ $0.588_{\pm 0.002}$
-beijing $0.953_{\pm 0.002}$ $0.959_{\pm 0.001}$ $0.952_{\pm 0.003}$ $0.953_{\pm 0.001}$ $0.949_{\pm 0.001}$
-churn $0.557_{\pm 0.014}$ $0.573_{\pm 0.014}$ $0.564_{\pm 0.012}$ $0.541_{\pm 0.015}$ $0.533_{\pm 0.007}$
+acsincome $0.547_{\pm 0.001}$ $0.533_{\pm 0.003}$ $0.540_{\pm 0.002}$ $0.546_{\pm 0.003}$ $0.527_{\pm 0.002}$
+adult $0.640_{\pm 0.003}$ $0.595_{\pm 0.002}$ $0.621_{\pm 0.002}$ $0.581_{\pm 0.002}$ $0.590_{\pm 0.004}$
+bank $0.880_{\pm 0.001}$ $0.739_{\pm 0.006}$ $0.798_{\pm 0.001}$ $0.662_{\pm 0.005}$ $0.701_{\pm 0.005}$
+beijing $0.699_{\pm 0.003}$ $0.658_{\pm 0.004}$ $0.615_{\pm 0.002}$ $0.626_{\pm 0.002}$ $0.620_{\pm 0.005}$
+churn $0.710_{\pm 0.008}$ $0.610_{\pm 0.006}$ $0.580_{\pm 0.011}$ $0.560_{\pm 0.019}$ $0.541_{\pm 0.008}$
+covertype $0.887_{\pm 0.002}$ $0.991_{\pm 0.001}$ $0.989_{\pm 0.001}$ $0.984_{\pm 0.001}$ $0.981_{\pm 0.000}$
+default $0.925_{\pm 0.002}$ $0.816_{\pm 0.003}$ $0.774_{\pm 0.003}$ $0.759_{\pm 0.004}$ $0.793_{\pm 0.005}$
+diabetes $0.762_{\pm 0.001}$ $0.895_{\pm 0.001}$ $0.870_{\pm 0.002}$ $0.851_{\pm 0.001}$ $0.837_{\pm 0.001}$
+lending $0.989_{\pm 0.002}$ $0.938_{\pm 0.011}$ $0.958_{\pm 0.003}$ $0.944_{\pm 0.005}$ $0.955_{\pm 0.003}$
+news $0.993_{\pm 0.001}$ $0.956_{\pm 0.002}$ $0.965_{\pm 0.002}$ $0.946_{\pm 0.002}$ $0.950_{\pm 0.002}$
+nmes $0.663_{\pm 0.014}$ $0.651_{\pm 0.007}$ $0.667_{\pm 0.009}$ $0.648_{\pm 0.015}$ $0.633_{\pm 0.010}$
 Table 24: Distance to closest record of the generated data (incl. standard errors in subscripts) for five CDTD configurations with progressive addition of model components.
 Real Test Set A B C D CDTD (per type)
-acsincome $7.673_{\pm 0.017}$ $8.335_{\pm 0.064}$ $8.222_{\pm 0.035}$ $8.305_{\pm 0.021}$ $8.352_{\pm 0.025}$ $8.322_{\pm 0.047}$
-adult $1.870_{\pm 0.000}$ $1.221_{\pm 0.018}$ $1.294_{\pm 0.015}$ $1.252_{\pm 0.014}$ $1.427_{\pm 0.008}$ $1.231_{\pm 0.011}$
-beijing $0.385_{\pm 0.000}$ $0.545_{\pm 0.001}$ $0.559_{\pm 0.002}$ $0.539_{\pm 0.003}$ $0.541_{\pm 0.002}$ $0.489_{\pm 0.001}$
-churn $0.347_{\pm 0.000}$ $0.307_{\pm 0.016}$ $0.326_{\pm 0.009}$ $0.294_{\pm 0.022}$ $0.298_{\pm 0.013}$ $0.274_{\pm 0.021}$
+acsincome $7.673_{\pm 0.017}$ $8.550_{\pm 0.055}$ $8.342_{\pm 0.037}$ $8.284_{\pm 0.030}$ $8.311_{\pm 0.041}$ $8.397_{\pm 0.062}$
+adult $1.870_{\pm 0.000}$ $1.231_{\pm 0.016}$ $1.318_{\pm 0.009}$ $1.296_{\pm 0.012}$ $1.520_{\pm 0.011}$ $1.327_{\pm 0.015}$
+bank $2.369_{\pm 0.000}$ $1.583_{\pm 0.006}$ $2.016_{\pm 0.007}$ $2.069_{\pm 0.014}$ $2.187_{\pm 0.009}$ $1.913_{\pm 0.006}$
+beijing $0.385_{\pm 0.000}$ $0.592_{\pm 0.001}$ $0.537_{\pm 0.002}$ $0.505_{\pm 0.001}$ $0.511_{\pm 0.001}$ $0.473_{\pm 0.002}$
+churn $0.347_{\pm 0.000}$ $0.526_{\pm 0.014}$ $0.383_{\pm 0.014}$ $0.327_{\pm 0.022}$ $0.333_{\pm 0.020}$ $0.278_{\pm 0.012}$
+covertype $0.529_{\pm 0.001}$ $0.871_{\pm 0.009}$ $1.919_{\pm 0.016}$ $1.872_{\pm 0.022}$ $1.698_{\pm 0.024}$ $1.594_{\pm 0.009}$
+default $1.812_{\pm 0.000}$ $1.321_{\pm 0.015}$ $1.333_{\pm 0.021}$ $1.306_{\pm 0.020}$ $1.484_{\pm 0.012}$ $1.355_{\pm 0.017}$
+diabetes $15.608_{\pm 0.055}$ $13.299_{\pm 0.020}$ $14.958_{\pm 0.029}$ $15.007_{\pm 0.026}$ $14.755_{\pm 0.045}$ $14.468_{\pm 0.028}$
+lending $11.184_{\pm 0.000}$ $15.130_{\pm 0.320}$ $15.112_{\pm 0.286}$ $14.957_{\pm 0.149}$ $14.916_{\pm 0.196}$ $14.962_{\pm 0.090}$
+news $3.615_{\pm 0.000}$ $3.894_{\pm 0.016}$ $3.737_{\pm 0.012}$ $3.692_{\pm 0.014}$ $3.746_{\pm 0.010}$ $3.676_{\pm 0.008}$
+nmes $1.931_{\pm 0.000}$ $1.202_{\pm 0.019}$ $1.014_{\pm 0.013}$ $0.958_{\pm 0.008}$ $0.952_{\pm 0.022}$ $0.780_{\pm 0.031}$
 Table 25: Machine learning efficiency F1 score for five CDTD configurations with progressive addition of model components.
@@
 Real Data A B C D CDTD (per type)
-adult $0.797_{\pm 0.000}$ $0.788_{\pm 0.001}$ $0.788_{\pm 0.001}$ $0.787_{\pm 0.001}$ $0.788_{\pm 0.002}$ $0.787_{\pm 0.001}$
-churn $0.873_{\pm 0.003}$ $0.856_{\pm 0.008}$ $0.856_{\pm 0.014}$ $0.857_{\pm 0.007}$ $0.849_{\pm 0.006}$ $0.852_{\pm 0.006}$
+adult $0.797_{\pm 0.000}$ $0.780_{\pm 0.002}$ $0.788_{\pm 0.001}$ $0.786_{\pm 0.001}$ $0.790_{\pm 0.001}$ $0.789_{\pm 0.002}$
+bank $0.745_{\pm 0.002}$ $0.759_{\pm 0.004}$ $0.758_{\pm 0.006}$ $0.760_{\pm 0.005}$ $0.751_{\pm 0.005}$ $0.766_{\pm 0.005}$
+churn $0.873_{\pm 0.003}$ $0.850_{\pm 0.006}$ $0.861_{\pm 0.008}$ $0.863_{\pm 0.004}$ $0.860_{\pm 0.008}$ $0.863_{\pm 0.007}$
+covertype $0.817_{\pm 0.001}$ $0.791_{\pm 0.001}$ $0.745_{\pm 0.001}$ $0.761_{\pm 0.001}$ $0.768_{\pm 0.001}$ $0.767_{\pm 0.001}$
+default $0.674_{\pm 0.001}$ $0.672_{\pm 0.002}$ $0.674_{\pm 0.002}$ $0.671_{\pm 0.002}$ $0.673_{\pm 0.003}$ $0.675_{\pm 0.002}$
+diabetes $0.621_{\pm 0.002}$ $0.616_{\pm 0.002}$ $0.612_{\pm 0.003}$ $0.612_{\pm 0.003}$ $0.617_{\pm 0.002}$ $0.619_{\pm 0.002}$
 Table 26: Machine learning efficiency AUC score for five CDTD configurations with progressive addition of model components.
@@
 Real Data A B C D CDTD (per type)
-adult $0.915_{\pm 0.000}$ $0.909_{\pm 0.000}$ $0.910_{\pm 0.000}$ $0.909_{\pm 0.000}$ $0.910_{\pm 0.000}$ $0.910_{\pm 0.001}$
-churn $0.964_{\pm 0.001}$ $0.962_{\pm 0.002}$ $0.961_{\pm 0.003}$ $0.960_{\pm 0.002}$ $0.961_{\pm 0.001}$ $0.962_{\pm 0.001}$
+adult $0.915_{\pm 0.000}$ $0.906_{\pm 0.001}$ $0.909_{\pm 0.000}$ $0.909_{\pm 0.000}$ $0.909_{\pm 0.001}$ $0.909_{\pm 0.001}$
+bank $0.947_{\pm 0.000}$ $0.945_{\pm 0.003}$ $0.946_{\pm 0.002}$ $0.945_{\pm 0.002}$ $0.946_{\pm 0.000}$ $0.946_{\pm 0.000}$
+churn $0.964_{\pm 0.001}$ $0.959_{\pm 0.003}$ $0.961_{\pm 0.002}$ $0.962_{\pm 0.002}$ $0.960_{\pm 0.002}$ $0.957_{\pm 0.008}$
+covertype $0.892_{\pm 0.000}$ $0.870_{\pm 0.000}$ $0.826_{\pm 0.001}$ $0.838_{\pm 0.001}$ $0.844_{\pm 0.001}$ $0.844_{\pm 0.000}$
+default $0.768_{\pm 0.000}$ $0.763_{\pm 0.002}$ $0.764_{\pm 0.001}$ $0.764_{\pm 0.002}$ $0.764_{\pm 0.002}$ $0.764_{\pm 0.002}$
+diabetes $0.693_{\pm 0.001}$ $0.675_{\pm 0.001}$ $0.664_{\pm 0.001}$ $0.671_{\pm 0.001}$ $0.673_{\pm 0.001}$ $0.673_{\pm 0.002}$
 Table 27: Machine learning efficiency RMSE for five CDTD configurations with progressive addition of model components.
@@
 Real Data A B C D CDTD (per type)
-acsincome $0.804_{\pm 0.012}$ $0.815_{\pm 0.009}$ $0.813_{\pm 0.018}$ $0.823_{\pm 0.017}$ $0.814_{\pm 0.014}$ $0.811_{\pm 0.014}$
-beijing $0.712_{\pm 0.001}$ $0.782_{\pm 0.004}$ $0.785_{\pm 0.004}$ $0.778_{\pm 0.005}$ $0.776_{\pm 0.004}$ $0.770_{\pm 0.005}$
-Appendix T Training and Sampling Times Details
+acsincome $0.804_{\pm 0.012}$ $0.868_{\pm 0.011}$ $0.808_{\pm 0.014}$ $0.820_{\pm 0.012}$ $0.800_{\pm 0.011}$ $0.807_{\pm 0.008}$
+beijing $0.711_{\pm 0.001}$ $0.801_{\pm 0.006}$ $0.780_{\pm 0.005}$ $0.771_{\pm 0.005}$ $0.769_{\pm 0.006}$ $0.766_{\pm 0.003}$
+lending $0.030_{\pm 0.000}$ $0.124_{\pm 0.006}$ $0.059_{\pm 0.002}$ $0.072_{\pm 0.001}$ $0.067_{\pm 0.002}$ $0.066_{\pm 0.002}$
+news $1.001_{\pm 0.002}$ $0.772_{\pm 0.019}$ $0.835_{\pm 0.062}$ $0.805_{\pm 0.079}$ $0.763_{\pm 0.062}$ $0.752_{\pm 0.067}$
+nmes $1.001_{\pm 0.003}$ $0.967_{\pm 0.064}$ $1.128_{\pm 0.088}$ $1.195_{\pm 0.087}$ $1.225_{\pm 0.070}$ $1.109_{\pm 0.055}$
+Appendix U Training and Sampling Times Details
 Table 28: Training times in minutes.
@@
 Table 29: Sample times in seconds per 1000 samples.
+TabDDPM produces NaNs during training on acsincome and diabetes, and is therefore excluded for these data.
+CoDi is considered prohibitively expensive to train on diabetes and lending.
 SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (per feature)
@@
 nmes 0.69 1.54 0.31 0.17 4.17 2.91 0.82 0.55
+Figures 24 and 25 show the benefit of deep generative models over SMOTE.
+Even though SMOTE is often praised as a simple, easy-to-use oversampling tool for tabular data, it relies on identifying nearest neighbors, making sampling very inefficient for larger datasets.
+As a consequence, we deem SMOTE to be infeasible to use for the acsincome and covertype datasets.
+The figures also illustrate the performance edge of CDTD, in particular compared to other diffusion-based models.
+Figure 24: Average sample quality metrics as a function of sampling time.
+Diffusion-based models are indicated in orange.
+Only datasets for which we could retrieve results from all models are included, this excludes acsincome, covertype, diabetes, lending.
+Figure 25: Average ML efficiency metrics and DCR as a function of sampling time.
+Diffusion-based models are indicated in orange.
+The dotted line indicates the test set performance of the real data.
+Only datasets for which we could retrieve results from all models are included, this excludes acsincome, covertype, diabetes, lending.